### v1

# On the Fundamental Impossibility of Hallucination Control in Large Language Models

Abstract
This paper explains why it is impossible to create large language models that do not hallucinate and what are the trade-offs we should be looking for. It presents a formal impossibility theorem demonstrating that no inference mechanism can simultaneously satisfy four fundamental properties: truthful (non-hallucinatory) generation, semantic information conservation, relevant knowledge revelation, and knowledge-constrained optimality. By modeling LLM inference as an auction of ideas where neural components compete to contribute to responses, we prove the impossibility using the Green-Laffont theorem. That mathematical framework provides a rigorous foundation for understanding the nature of inference process, with implications for model architecture, training objectives, and evaluation methods.
1 Introduction
Large language models have been known for producing convincing content that is factually incorrect, inconsistent, or well fabricated—a phenomenon known as hallucination. Despite extensive research efforts involving architectural innovations, training techniques, and post-hoc verification methods, hallucination persists across all state-of-the-art models.
This paper shows that hallucination is a manifestation of fundamental limitations of inference process itself: no LLM inference mechanism can simultaneously satisfy four essential properties required for perfect hallucination control. Specifically, we prove that it is mathematically impossible for any LLM to simultaneously achieve: (1) truthful (non-hallucinatory) response generation, (2) semantic information conservation, (3) relevant knowledge revelation, and (4) knowledge-constrained optimality.
Our contribution introduces a game-theoretic framework that models LLM inference as an auction of ideas where neural components (attention heads, circuits, activation patterns) compete to contribute knowledge to response generation. This perspective allows us to apply fundamental results from mechanism design theory to prove our impossibility result rigorously.
1.1 Illustrative Example
Consider an LLM queried about recent diplomatic negotiations: "What was the outcome of diplomatic negotiations between Countries A and B in February 2025?" Assume the model has encoded knowledge that: (K1) countries A and B were engaged in talks, (K2) country A sought a trade agreement, and (K3) the actual outcome is unknown.
Examining possible responses reveals the fundamental trade-offs:
Complete Abstention: "I don’t have information about the outcome" is truthful but fails to utilize available relevant knowledge (K1, K2), violating the expectation that good inference should reveal accessible information.
Hallucination: "The negotiations resulted in a comprehensive trade agreement reducing tariffs by 15%" utilizes all knowledge and provides a complete answer but fabricates specific details, violating truthfulness.
Partial Knowledge: "Countries A and B were negotiating a trade agreement, but I don’t know the outcome" is truthful and utilizes available knowledge but may appear unsatisfying compared to the decisive hallucinated response.
Overcautious: "I’m not certain about any aspects of these negotiations" avoids hallucination but disclaims even known facts, violating truthfulness about available information.
Each response strategy sacrifices different aspects of what constitutes an ideal response, illustrating the impossibility of simultaneously optimizing all desired properties.
1.2 Theoretical Contributions and Implications
The simple example above illustrates the key observations of this paper, summarized and proved formally in the following
Theorem (Impossibility Theorem).
For any query space $\mathcal{Q}$ containing non-trivial queries (requiring integration of available distributed knowledge), no LLM can simultaneously generate truthful (non-hallucinatory) responses, satisfy the semantic information conservation principle, guarantee revelation of relevant knowledge, and enforce knowledge-constrained optimality of responses.
No matter how hard we should try to train LLM to generate responses that are perfectly aligned with query context and do not create factually incorrect, inconsistent, or fabricated statements, the trained LLM will always violate some aspects of what we may call a reasonable response. It may ignore available knowledge, it may present fabricated knowledge, it may be a reformulated query without any novel observations, or it may be a lucky hallucination guessing the correct answer without any support in the encoded knowledge. And so on, because there is no free lunch when there are inevitable trade-offs of inference, and we must accept that.
Despite its name, the theorem should be interpreted as a constructive result. It offers a precise characterization of the fundamental tradeoffs in LLM design and provides several key insights:
Fundamental Trade-offs: Rather than viewing hallucination as a bug to be eliminated, we should understand that different applications require different balances among truthfulness, information utilization, and response completeness.
Architectural Guidance: The result suggests investigating multi-component architectures where specialized modules handle different aspects of the trade-off. Safety-critical applications might prioritize truthfulness modules and conservation of information, while creative applications could emphasize hallucinatory responses. Training objectives should explicitly balance these competing demands based on intended use cases.
Research Directions: The proposed mathematical (and philosophical) framework opens new research paths at the intersection of mathematical game theory, information theory, and neural network inference, providing tools for analyzing and optimizing LLM behavior.
To my best knowledge, unlike previous work treating hallucination as a statistical or architectural phenomenon, the game-theoretic approach offers the first formal proof of fundamental limitations in the inference process, explaining precisely why these limitations exist and how to work optimally within them.
2 Related Work
Research on LLM hallucinations spans mechanistic interpretability, probabilistic frameworks, and empirical mitigation techniques. Below we present selected works in that domains. The auction-theoretic impossibility framework in this paper offers a fundamentally novel perspective distinct from these existing approaches.
2.1 Mechanistic Interpretability
Mechanistic interpretability research examines internal LLM processes to identify hallucination origins. Yu et al. yu2024mechanisticunderstandingmitigationlanguage identify two failure modes: knowledge enrichment hallucinations from insufficient subject-attribute knowledge in lower MLP layers, and answer extraction hallucinations from failures in upper-layer attention heads selecting correct object attributes.
Studies characterizing factual knowledge storage and retrieval complement our framework by providing empirical evidence of competing "agents" within LLM architectures. Meng et al. meng2023locatingeditingfactualassociations demonstrated methods for locating and editing factual associations, revealing how knowledge representations can be manipulated. Geva et al. geva2023dissecting identified a three-step factual recall process: (1) enriching last-subject position representations with subject-related attributes, (2) propagating relation information to predictions, and (3) querying enriched subjects to extract attributes.
These findings provide concrete instantiations of our abstract "agents"—attention heads and MLP components with specialized knowledge domains competing to influence response generation. See also the fascinating and related work on the Biology of a Large Language Model lindsey2025biology .
2.2 Probabilistic Frameworks
Probabilistic approaches model hallucination through uncertainty and belief quantification. Shelmanov et al. shelmanov2025headpredictheadquestion introduced pre-trained uncertainty quantification heads for detecting hallucinations without task-specific data. Farquhar et al. farquhar2024detecting proposed semantic entropy measures that detect hallucinations by quantifying uncertainty at the meaning level rather than word sequences.
These approaches implicitly acknowledge hallucination as inevitable, focusing on detection rather than elimination—supporting our impossibility conclusion. However, none provide formal proofs of why complete elimination remains impossible.
2.3 Empirical Mitigation Techniques
Various empirical approaches attempt to reduce hallucination frequency with mixed success. Creswell and Shanahan creswell2022faithfulreasoningusinglarge explored faithful reasoning approaches for improving factual accuracy. Retrieval-augmented generation (RAG) by Lewis et al. lewis2021retrievalaugmentedgenerationknowledgeintensivenlp integrates external knowledge sources for factual grounding, though these systems still hallucinate when retrieval fails.
Training approaches like RLHF ouyang2022traininglanguagemodelsfollow can reduce but never eliminate hallucination—an observation our theorem explains from first principles.
Evaluation benchmarks demonstrate persistent hallucination across all models. TruthfulQA lin2022truthfulqameasuringmodelsmimic and HaluEval li2023haluevallargescalehallucinationevaluation show that even state-of-the-art systems produce hallucinations on carefully designed test sets.
2.4 Theoretical Perspectives
Most significantly, Banerjee et al. banerjee2024llmshallucinateneedlive argued from computational complexity principles that hallucinations are inherent features of LLMs stemming from fundamental mathematical structure, concluding that "LLMs will always hallucinate." While their conclusion aligns with that of this paper, the authors do not identify the specific trade-offs involved or provide the precise characterization our theorem offers.
3 Novel Contribution: Auction-Theory and AI
We formulate LLM design problem as an auction (or game) design problem. This approach draws on foundational mechanism design theory, a branch of mathematical game theory Myerson1989 ; fudenberg1991game ; krishna2009auction ; milgrom2004putting . Particularly, we refer to the Green-Laffont theorem green1977characterization ; green1977revelation , to prove the impossibility result. See also DBLP:journals/amcs/Karpowicz12 for more recent results.
In a nutshell, the auction design theory investigates existence and properties of competition rules that guide interactions of agents towards desired outcomes. The adaptation of that game-theoretical principles to neural computation represents a novel theoretical contribution that:
Explains empirical observations by proving why approaches like RLHF, RAG, and faithful fine-tuning can reduce but never eliminate hallucination.
Identifies precise trade-offs by introducing formal mathematical (and programmable) characterizations of inference properties that cannot be simultaneously satisfied.
Provides design guidance by suggesting principled approaches for making optimal trade-offs among competing objectives rather than pursuing impossible simultaneous optimization.
This theoretical foundation bridges empirical observations with formal analysis, offering new mathematical tools for understanding and optimizing LLM behavior within fundamental constraints.
4 Alternative Views
The impossibility theorem provides strong theoretical basis against perfect hallucination elimination, but several alternative perspectives merit consideration. These views highlight important practical considerations while reinforcing the value of understanding theoretical constraints that bound the space of possible solutions.
4.1 The Engineering Optimism View
Many previous works maintain that hallucination represents an engineering challenge rather than a fundamental limitation. This view argues that sufficiently sophisticated architectures, training procedures, or verification systems could achieve near-perfect truthfulness without significant trade-offs.
Proponents point to rapid empirical progress: each generation of models shows reduced hallucination rates on benchmark datasets. They argue that techniques like constitutional AI, iterative refinement, and multi-agent verification systems could eventually solve the problem.
Response: While engineering advances can certainly improve the trade-offs, our theorem demonstrates mathematical constraints that no architecture can completely overcome. The impossibility holds regardless of computational resources or architectural sophistication, as it derives from the fundamental structure of information integration under competing objectives.
4.2 The Sufficiency Threshold View
Another perspective suggests that perfect elimination of hallucination is unnecessary—sufficiently low hallucination rates could enable practical deployment without fundamental concerns. This view acknowledges theoretical limitations but argues they become irrelevant when hallucination rates drop below application-specific thresholds.
Response: This pragmatic view has merit and aligns with the constructive interpretation of the impossibility result. However, the theorem remains valuable for understanding why certain accuracy thresholds prove difficult to exceed and for optimizing trade-offs within practical constraints. Moreover, some applications (safety-critical systems, legal reasoning) may demand understanding and explaining fundamental limitations even when targeting high but imperfect accuracy.
4.3 The External Knowledge View
A third perspective argues that perfect grounding in external, verifiable knowledge sources could overcome internal representation limitations. If models could access and perfectly utilize real-time databases, verification systems, or human oversight, they might achieve truthfulness without trade-offs.
Response: External knowledge integration does change the game structure by altering available information during inference. However, this approach faces several limitations: (1) external sources may be incomplete, conflicting, or outdated; (2) the integration mechanism itself involves trade-offs between source reliability and knowledge revelation; (3) real-time verification introduces computational constraints that recreate similar impossibility conditions. While external knowledge can improve practical performance, it shifts rather than eliminates the fundamental tensions our theorem identifies. The impossibility result remains independent of the ground truth available.
5 Theoretical Framework
This section develops the mathematical foundations necessary for proving the impossibility theorem. We formalize key concepts including knowledge representation, semantic information, and the auction-theoretic model of LLM inference. We also present the philosophical ground for the theory development.
To the sake of narrative consistency, we introduce the framework step by step referring to some concepts first and providing their formal definitions later.
5.1 Knowledge and Semantic Information
We define knowledge as organized information that LLMs can encode and utilize for reasoning, aligning with epistemological concepts while maintaining mathematical rigor. In other words, knowledge (like energy in physics) can be used for reasoning (work or action). Information is a (measure of) potential for uncertainty reduction effort in a given context.
Definition 1 (Knowledge Space).
Knowledge is a metric space $\mathcal{K}$ with distance function $d_{\mathcal{K}}$ that collects all facts, concepts, and relationships that can be encoded by an LLM.
For model $M$ with parameters trained on data, $\mathcal{K}_{M}\subset\mathcal{K}$ represents the knowledge subset encoded in the model’s parameters. The distance function $d_{\mathcal{K}}$ measures differences between knowledge elements. A simple yet powerful example we will refer to is this:
$d_{\mathcal{K}}(A,B)=|A\setminus B|+|B\setminus A|,$ (1)
the cardinality of the symmetric difference, allowing measurement of missing and fabricated knowledge.
Definition 2 (Knowledge Representation).
Knowledge representation is a tuple $(\mathcal{K},d_{\mathcal{K}},\mathcal{W},\mathcal{A},\Phi,\Psi)$ where:
  • $\mathcal{K}$ is a metric space of knowledge elements with distance function $d_{\mathcal{K}}$,
  • $\mathcal{W}$ is the space of all possible model parameters,
  • $\mathcal{A}$ is the space of all possible activation patterns,
  • $\Phi:\mathcal{W}\rightarrow\mathcal{P}(\mathcal{K})$ maps parameters to knowledge subsets,
  • $\Psi:\mathcal{A}\rightarrow\mathcal{P}(\mathcal{K})$ maps activations to accessible knowledge.
For parameterization $W\in\mathcal{W}$, $\Phi(W)=\mathcal{K}_{M}$ represents encoded knowledge. For activation pattern $a_{i}\in\mathcal{A}$ with parameters $W_{i}$, $\Psi(a_{i})\subset\Phi(W_{i})$ represents knowledge accessed during specific inference.
To describe the inference process leading to response to a query, we need a mapping that measures the information processed within a limited computational budget.
Definition 3 (Semantic Information Measure).
The semantic information measure is mapping $\mu_{C}:\mathcal{P}(\mathcal{K})\times\mathcal{Q}\rightarrow\mathbb{R}_{+}$ parameterized by query context $q\in\mathcal{Q}$ and computational budget $C$ satisfying:
$\mu_{C}(\emptyset,q)=0,$ (2)
$\mu_{C}(A,q)\leq\mu_{C}(B,q)\text{ if }A\subseteq B,$ (3)
$\mu_{C}(A\cup B,q)\leq\mu_{C}(A,q)+\mu_{C}(B,q)-\mu_{C}(A\cap B,q)+\mu_{C}(\mathcal{E}_{C}(A\cup B)\setminus(A\cup B),q).$ (4)
Empty knowledge or knowledge irrelevant to query has zero information, information content is subadditive and includes emergent information $\mathcal{E}_{C}$ in the query context. The emergence operator $\mathcal{E}_{C}$, which we introduce below, describes knowledge becoming accessible through bounded reasoning:
Definition 4 (Emergence Operator and Reasoning).
The emergence operator $\mathcal{E}_{C}:\mathcal{P}(\mathcal{K})\times\mathcal{W}\rightarrow\mathcal{P}(\mathcal{K})$ maps a knowledge subset and parameter set to a knowledge subset:
$\mathcal{E}_{C}(A,w)=\{k\in\mathcal{K}\mid\exists f_{w}\in\mathcal{F}_{C}:k=f_{w}(A)\},$ (5)
where $\mathcal{F}_{C}$ is the set of all computational transformations with cost lower than or equal to $C$, $f_{w}$ is a transformation parameterized by $w\in\mathcal{W}$, and $k=f_{w}(A)$ denotes that knowledge $k$ is derivable from $A$ using transformation $f_{w}\in\mathcal{F}_{C}$. For any knowledge state $\theta(t)$ available at time instant $t$ and computational budget $C$, the reasoning process:
$\theta(t+1)=\mathcal{E}_{C}(\theta(t))\quad\text{{satisfies}}\quad\mu_{C}(\theta(t+1))\geq\mu_{C}(\theta(t)).$ (6)
Furthermore, with unlimited computation, reasoning creates no new information:
$\mu_{\infty}(\theta(t+1))=\mu_{\infty}(\theta(t)).$ (7)
The definitions above clearly require refinements and further investigations, but they capture the philosophical and technical concepts we need in order to address the problem of hallucination. In particular, they are meant to lay the ground for the critical idea of semantic information conservation principle and its potential violations. As we will see, that idea is closely related to the following property of response.
Definition 5 (Bounded Response Creativity).
Let $q\in\mathcal{Q}$ be a query and $r\in\mathcal{R}$ a response to that query. Response $r$ is creatively bounded if and only if:
$\mu_{C}(K(r))\leq\mu_{C}(K(q)\cup\mathcal{E}_{C}((\mathcal{K}_{M}\cap T(q))\cup K(q))).$ (8)
The semantic information of knowledge contained in any response $K(r)$ can never exceed what is already present in the query $K(q)$ together with the portion of encoded knowledge that is accessible through reasoning bounded by computational budget $C$ and aligned with ground truth $T(q)$. Let us see why.
We address the settings in which a large subset of knowledge encoded in LLM can be redundant, so it doesn’t necessarily contain more information than its components. Clearly, that assumption is violated when LLM combines knowledge elements to fabricate facts or produce emergent insights. Then,
$\mu_{C}(A\cup B)>\mu_{C}(A)+\mu_{C}(B),$
which means novel insights come from existing data. That is exactly the creativity behind the hallucination phenomenon we are struggling to control. So why should we assume otherwise?
We view reasoning as a transformation that reveals information already present in the knowledge representation. After the reasoning, certain patterns may become evident that were computationally intractable to observe in the original representation. However, we stand at the position that reasoning does not create new information.
To see that, consider a similarity relation of two matrices, $D=V^{-1}AV$, where $D$ is a diagonal matrix of eigenvalues. Matrices $D$ and $A$ contain the same information, but it is much easier to understand the fundamental properties of $A$ when we represent it as $D$ in the basis of eigenvectors in matrix $V$. The properties of $A$ has always been encoded in $A$, but they become accessible when we transform $A$ into $D$.
Resolution of the Uncertainty Resolution Paradox
If reasoning makes a new fact accessible, then that new fact can be used to resolve uncertainty that was previously unresolvable. But that means the reasoning must have created new information, contrary to what we have assumed. It seems we have a paradox.
To resolve it, we need the emergence operator $\mathcal{E}_{C}(X)$ distinguishing between computationally unbounded and bounded uncertainty reduction (accessible and inaccessible information).
Notice that for all practical purposes the information measurement $\mu_{C}(X,q)$ can only provide maximum uncertainty reduction from $X$ within some computational budget $C$ (or time). Therefore, given the result of computationally bounded reasoning $\mathcal{E}_{C}(X)$, we can observe information gain through the computational work, i.e., $\mu_{C}(X,q)\leq\mu_{C}(\mathcal{E}_{C}(X),q)$. Notice, that is possible only with organized information providing knowledge for reasoning.
That idea above is thus described by the reasoning dynamics:
$\theta(t+1)=\mathcal{E}_{C}(\theta(t))\quad\text{and}\quad y(t)=\mu_{C}(\theta(t),q).$ (9)
Bounded reasoning makes hidden knowledge accessible within a computational budget. So, when that budget is infinite, reasoning creates no new information, i.e.,
$\theta(\infty)=\mathcal{E}_{\infty}(\theta(\infty)).$ (10)
All there is to know is immediately accessible when we have unlimited computing resources (or time).
To illustrate that argument, consider the following example:
[…]
$y(0)$ $=\mu_{C}(\theta(0))=3\text{ bits},$
$y(1)$ $=\mu_{C}(\theta(1))=5\text{ bits}.$
We see $\mu_{C}(\mathcal{E}_{C}(\theta(0)))>\mu_{C}(\theta(0))$, i.e., practical information gain in the course of reasoning. In comparison, infinite computational budget, $C=\infty$, implies we do not need the intermediate inference steps to see all there is to see, so we know the equilibrium state:
$\mu_{\infty}(\mathcal{E}_{\infty}(\theta(0)))=\mu_{\infty}(\theta(0)),$
and no new information is created.
In the next section, we relate that important property of response generation, namely, how much additional semantic information each LLM inference path can contribute, with the Semantic Information Conservation Principle.
5.2 Statements, Queries, and Ground Truth
Queries and responses are statements expressed in natural language that LLM is able to process and generate. We need to compare those statements and relate them to knowledge. It is also useful to distinguish queries from responses with their nature, respectively, interrogative or declarative. Queries refer to what is presupposed (assumed in advance), e.g., topic or problem, whereas responses make claims (or assertions). We should also point out that we limit our study to the case of non-trivial queries that require integrating responses from many LLM inference paths. The following formal definition encapsulates those ideas.
Definition 6 (Statement Space).
The space of natural language statements $\mathcal{L}$ is a metric space with distance function $d_{\mathcal{S}}$. We distinguish query space $\mathcal{Q}\subset\mathcal{L}$ of interrogative statements and response space $\mathcal{R}\subset\mathcal{L}$ of declarative statements. Function $K:\mathcal{L}\rightarrow\mathcal{P}(\mathcal{K})$ maps statements to referenced knowledge. We call query $q$ non-trivial, if it is integrated from multiple knowledge subsets.
We must now rigorously define the ground truth mapping and establish formal relationships between the knowledge space $\mathcal{K}$ and response space $\mathcal{R}$.
For that, we should start by recognizing that the very idea of ground truth validation is limited (or subjective at least). It is rather naive to assume that LLM can access an oracle establishing provably true and relevant knowledge for every possible query. Indeed, that limits are given by the Goedel’s Theorems. Therefore, in practice that inaccessibility of the complete truth validation function is one of the reasons why hallucination cannot be eliminated entirely. High-quality, fact-checked datasets, human feedback providing human-level judgments of truth, RAG systems with external validated knowledge bases, all that approximate the idea of ground truth. Surprisingly, the impossibility result seems to be independent of practical or idealized implementation of a ground truth mapping, which we define as follows.
Definition 7 (Ground Truth Mapping).
The ground truth mapping $T\colon\mathcal{Q}\rightarrow\mathcal{P}(\mathcal{K})$ associates each query $q\in\mathcal{Q}$ with knowledge elements constituting complete, correct answers:
$T(q)=\{k\in\mathcal{K}\mid k\text{ is relevant to }K(q)\text{ and {labeled} as true}\}.$ (11)
5.3 Auction-Theoretic Model of LLM Inference
There are many ways LLM can generate answers to a query. By design, it is a controlled random process governed by fine-tuned probability distributions. We are going to model that process as an auction in which agents (representing competing ideas is the space of all possible activation patterns $\mathcal{A}$) use their privately held knowledge (internal representations in $\mathcal{W}$) to construct bids (candidate responses) and submit them to the LLM response mechanism. That response mechanism then selects the best bid as a response to a query.
Definition 8 (Response Mechanism).
A response mechanism $\mathcal{M}=(S,g,p)$ consists of:
  • Strategy space $S=S_{1}\times\cdots\times S_{n}$, where $S_{i}\subset\mathcal{P}(\mathcal{R})$,
  • Outcome function $g:S\rightarrow\mathcal{R}$ mapping strategies to responses,
  • Payment function $p:S\times\Theta\rightarrow\mathbb{R}^{n}$ determining costs.
A response mechanism defines strategy space $S$ as a set of inference actions that each agent can make to represent its knowledge $\theta_{i}\in\Theta_{i}\subset\mathcal{K}$ in response to a query. It is a dictionary of LLM’s component activation pattens generating and presenting ideas as bids in an auction. Examples of strategies include activations deciding which tokens to attend or how strongly to activate neurons for different patterns. Then, given a set of submitted bids (in the form of strategy profile) $s=(s_{1},\dots,s_{n})$, the response mechanism uses outcome function $g(s)$ to select and generate response. One straightforward example is a softmax function in LLM’s final layer. Finally, given a knowledge profile $\theta=(\theta_{1},\dots,\theta_{n})$ revealed by the agents, payment function $p(s,\theta)=(p_{1}(s,\theta),\dots,p_{n}(s,\theta))$ measures individual contributions of each response to generating an output. In this case, examples include (semantic information measures) counting ignored low-probability tokens in top-k filtering, softmax normalization in multi-headed attention, or marginal reward contributions in RLHF.
Definition 9 (Agents of Competing Ideas).
The inference process involves $n$ agents $\mathcal{A}=\{a_{1},a_{2},...,a_{n}\}$ competing by submitting bids to mechanism $\mathcal{M}$. Each agent $a_{i}$ constructs bid $s_{i}(\theta_{i})$ based on private knowledge $\theta_{i}\in\Theta_{i}$ to maximize utility:
$u_{i}(s,\theta)=v_{i}(g(s),\theta_{i})-p_{i}(s,\theta),$ (12)
where $v_{i}$ is valuation and $p_{i}$ is payment. Knowledge profiles $\theta=(\theta_{1},\dots,\theta_{n})$ are independently distributed.
One way to think about the concept of agents is that it represents different attention heads in transformer models, or components of mixture-of-experts architectures, or activation paths fired by the same query. Each agent’s private knowledge represents model parameters utilized by an agent to construct a response. The utility function can be best described as a partial contribution of a candidate response to LLM loss function minimization. It is designed to explain why any inference path is activated in response to the LLM prompt. For example, negative log likelihood loss can be decomposed with respect to attention heads, and linearized to measure how much attention head contributes to token prediction.
Let us point out that agents represent algorithms and architectural components of any LLM implementation, they are not conscious beings with individual goals. That is important, as it refers to the Revelation Principle in game theory. If we know strategies (of using knowledge) the players (conscious beings with individual goals) will use in a game, we can implement those strategies as an algorithm (or agent) with parameters representing private preferences or knowledge of the players. In the case considered, we are the players selecting architecture and training the LLM.
We assume that individual utility depends on all responses $s=(s_{1},\dots,s_{n})$ and knowledge $\theta=(\theta_{1},\dots,\theta_{n})$ used by all agents, $u_{i}=u_{i}(s,\theta)$. Also, we assume utility is a sum of individual valuation $v_{i}$ of outcome $g(s)$ and payment $p_{i}=p_{i}(s,\theta)$ value assigned.
The valuation function $v_{i}=v_{i}(g(s),\theta_{i})$ represents how strongly a neural component activates for certain outputs (high activation = high $v_{i}$, low activation = low $v_{i}$), and how much a component reduces the overall loss when processing certain inputs ($v_{i}$ is proportional to gradient of a loss function, $-\nabla_{\theta_{i}}L_{W}$).
The payment function $p_{i}=p_{i}(s,\theta)$ may represent training penalties ($p_{i}>0$ for discouraged activation, $p_{i}<0$ for encouraged activation), attention focus (higher $p_{i}$ reduces attention allocation), $L1$ regularization terms, and network circuits resistance (limiting activations of certain inference paths). Therefore, $u_{i}=v_{i}-p_{i}>0$ encourages activation and $u_{i}=v_{i}-p_{i}<0$ discourages activation.
We can also illustrate the concept of utility function with the following example:
$u_{i}(s,\theta)$ $=\underbrace{\alpha\cdot\text{Accuracy}(g(s),\theta_{i})}_{\text{Valuation}}-\underbrace{\gamma\cdot\text{Contradiction}(s_{i},\theta_{i})}_{\text{Consistency Cost}}.$
$\text{Accuracy}(g(s),\theta_{i})$ measures how well the response aligns with the knowledge $\theta_{i}$ agent $a_{i}$ is using. $\text{Contradiction}(s_{i},\theta_{i})$ measures how much the strategy $s_{i}$ conflicts with internal knowledge state $\theta_{i}$ (both extracted from $s$ and $\theta$ profiles).
Independent private knowledge assumption
One of the critical and controversial assumptions we make is is that knowledge profiles $\theta_{i}$ are independently distributed. It should be noted that in transformers, heads may share layer-norm signals, residual streams, and optimizer state. That may result in correlated (or, so called, affiliated) inference signals between token representations. Furthermore, the superposition hypothesis, introduced in mechanistic interpretability studies, suggests cooperative error-canceling due to strong correlations.
The reason why we make the independent private knowledge assumption, despite its idealized nature, is that it makes the impossibility result becomes even stronger. To be specific, Milgorm and Weber milgrom1982theory showed with their linkage principle that strong correlations amplify competition in information revelation. In the LLM setting that means amplifying risk of hallucinations. Similarly, Roughgarden et al. roughgarden2016optimal demonstrated that interdependent knowledge leads to violation of balancing payments. As we will see, in our setting that means violation of information conservation principle. For more results, see e.g. Myerson myerson2008perspectives .
5.4 Hallucination Cost
Having all the pragmatic and philosophical limitations of the definition above in mind, we can define formally the hallucination cost function. Its purpose is to assign a numerical value to the discrepancy between what can be provided as a response to a query, and what can be established as a provably correct answer.
Definition 10 (Hallucination Cost).
Given query $q\in\mathcal{Q}$, response $r\in\mathcal{R}$, and ground truth $T(q)$, the hallucination cost function $H\colon\mathcal{R}\times\mathcal{Q}\rightarrow\mathbb{R}_{+}$ measures discrepancy:
$H(r,q)=d_{\mathcal{K}}(K(r),T(q)).$ (13)
The hallucination cost takes any natural language query $q$ and natural language response $r$, determines what is missing and what is fabricated, and assigns number to that subset of knowledge. If response $r$ is perfectly aligned with the ground truth for $q$, then $H(r,q)=0$ by definition of metric $d_{\mathcal{K}}$. If $r$ contains fabricated facts or is incomplete given $T(q)$, then positive hallucination cost $H(r,q)>0$ is generated.
Notice that reaching $H(r,q)=0$ is also possible with lucky hallucinations (or guessing). So, minimization of hallucination cost is not enough to resolve the problem. We need to investigate all the properties and constraints that hallucination-free response mechanisms should have.
5.5 Properties of Hallucination-Free Mechanisms
There are four fundamental properties that a hallucination-free response mechanism should ideally satisfy. We can interpret them as constraints that should be incorporated in model training. We characterize them formally referring to the notation of fundamental results of mechanism design and auction theory.
Property 1 (Truthfulness).
Mechanism $\mathcal{M}$ is truthful if for all agents $i$, all $\theta_{i}\in\Theta_{i}$, all $s_{-i}\in S_{-i}$, and all $\bar{s}_{i}\in S_{i}$:
$u_{i}((s_{i}^{*}(\theta_{i}),s_{-i}),\theta_{i})\geq u_{i}((\bar{s}_{i},s_{-i}),\theta_{i}),$ (14)
where $s^{*}_{i}(\theta_{i})$ represents truthful revelation of private knowledge.
A truthful mechanism promotes inference pattern $s^{*}_{i}(\theta_{i})$ that are perfectly aligned with the accessible private knowledge. Any other alternative response $\bar{s}_{i}$ that is not aligned with $\theta_{i}$ will result in lower utility, generating higher model loss.
Consider an attention head specialized in medical knowledge with the following utility function:
$u_{\text{med}}(s,\theta_{\text{med}})=\alpha\cdot\text{Accuracy}(g(s),\theta_{\text{med}})-\beta\cdot\|s_{\text{med}}\|^{2}$
Imagine that given a query about a disease treatment, the head has two possible strategies:
  • $s^{*}$:
    activate only for known facts about established treatments,
  • $\bar{s}$:
    activate for both known facts and speculative treatments.
With proper training, we want $u_{\text{med}}((s^{*}(\theta),s_{-\text{med}}),\theta)>u_{\text{med}}((\bar{s},s_{-\text{med}}),\theta)$, ensuring that agent (model component) is rewarded for representing its knowledge accurately.
Property 2 (Semantic Information Conservation).
Mechanism $\mathcal{M}$ satisfies information conservation if for all strategy profiles $s\in S$ and knowledge profiles $\theta\in\Theta$:
$\sum_{i=1}^{n}p_{i}(s,\theta)=0.$ (15)
As we have announced earlier, that important and controversial principle is responsible for generation of responses of bounded creativity. Let us prove that claim formally.
Lemma 1.
If the semantic information conservation principle holds, then the inference process generates responses of bounded creativity,
$\mu_{C}(K(r))\leq\mu_{C}\left(K(q)\cup\mathcal{E}_{C}((\mathcal{K}_{M}\cap T(q))\cup K(q))\right).$ (16)
Proof.
Fix a query $q\in\mathcal{Q}$ and let $s^{*}(\theta)=(s_{1}^{*},\ldots,s_{n}^{*})$ be a truthful strategy profile. We write $r=g(s^{*})$ for the generated response. Let us define each agent’s marginal information contribution as:
$\Delta_{i}=\mu_{C}(K(r))-\mu_{C}(K(g((0_{i},s_{-i}^{*}))))\geq 0.$
That contribution is non-negative by monotonicity of semantic information measure.
Next, without loss of generality, we assume payments of the form:
$p_{i}(s^{*},\theta)=\Delta_{i}+c_{i}\text{ where }\sum_{i=1}^{n}c_{i}=0.$
This ensures that $\sum_{i=1}^{n}p_{i}(s^{*},\theta)=\sum_{i=1}^{n}\Delta_{i}$.
Finally, notice that if there is no relevant knowledge available, the knowledge in response equals that provided by the context:
$g(\mathbf{0})=r_{0}\text{ where }K(r_{0})=K(q).$
We can now see that, whenever semantic information is conserved, we have:
$0=\sum_{i=1}^{n}p_{i}=\sum_{i=1}^{n}(\Delta_{i}+c_{i})=\sum_{i=1}^{n}\Delta_{i}.$
Since $\Delta_{i}\geq 0$ for all $i$ by monotonicity property and $\sum_{i=1}^{n}\Delta_{i}=0$, we must have:
$\Delta_{1}=\Delta_{2}=\cdots=\Delta_{n}=0.$
Therefore, for each $j$:
$\mu_{C}(K(r))=\mu_{C}(K(g((0_{j},s_{-j}^{*})))).$
Since that equality holds for any strategy profile $s^{*}$, removing agents one at a time leads to
$\mu_{C}(K(r))=\mu_{C}(K(g(\mathbf{0}))).$
But since $K(g(\mathbf{0}))=K(q)$, we conclude:
$\mu_{C}(K(r))=\mu_{C}(K(q)).$
Finally, by definition of the emergence operator and monotonicity property:
$K(q)\subseteq K(q)\cup\mathcal{E}_{C}((\mathcal{K}_{M}\cap T(q))\cup K(q)).$
Therefore:
$\mu_{C}(K(r))=\mu_{C}(K(q))\leq\mu_{C}(K(q)\cup\mathcal{E}_{C}((\mathcal{K}_{M}\cap T(q))\cup K(q))),$
as required.
If total information payments must balance to zero and each agent’s contribution is non-negative, then no agent can contribute without violating the information conservation constraint. Therefore, the response can only contain information that was already accessible through the query and bounded reasoning over the model’s relevant knowledge. Any hallucinated information requires $\sum_{i}\Delta_{i}>0$, contradicting the zero-sum constraint.
When the semantic information conservation principle holds, the model cannot create information out of nothing, the model cannot generate claims beyond available information. That also implies that uncertainty cannot disappear in response without additional information input or computational work to extract implicit information. If information is neither explicitly present nor accessible through reasoning from the query and model knowledge, it cannot be in the response.
The information content of the response cannot exceed the combined information from the model’s knowledge, the query itself, and what can be derived through bounded reasoning over that knowledge. In our dynamical systems framework, this corresponds to the constraint that the measured output cannot exceed what can be produced from the initial state plus what becomes accessible through state transitions within computational budget $C$.
Notice that the principle does not prevent redistribution of information. It is permitted as long as the information transfers perfectly balance. Indeed, since $p_{i}=p_{i}(s,\theta)$ incorporates the strategic profile, it admits strategies withholding information. Within the information conservation regime we do not require that strategies utilize all available knowledge states. Inference process can withhold information, as it often does in practice.
We can illustrate that last observation with the following example. Suppose there are two agents performing inference with the following knowledge profiles:
$\theta_{1}=\{A,B,C\}\quad\text{and}\quad\theta_{2}=\{D,E\},$
and one agent representing final layers responsible for combining output into response $g=g(s)$. Suppose:
$s_{1}(\theta_{1})=A\quad\text{and}\quad s_{2}(\theta_{2})=\{D,E\}\quad\text{and}\quad g(s)=\{A,D,E\}.$
Then the information conservation rule is satisfied when $p$ counts exchanged bits of knowledge. Namely, $p_{1}(s,\theta)=1$ for providing $A$, $p_{2}(s,\theta)=2$ for providing $\{D,E\}$, and $p_{3}(s,\theta)=-3$ for absorbing information into $\{A,D,E\}$. We have
$p_{1}(s,\theta)+p_{2}(s,\theta)+p_{3}(s,\theta)=1+2-3=0$
and $\{B,C\}$ are left unrevealed.
The following property fills the potential information gap admitted by the conservation principle.
Property 3 (Relevant Knowledge Revelation).
Mechanism $\mathcal{M}$ satisfies relevant knowledge revelation if for all agents $i$, all $\theta_{i}\in\Theta_{i}$, and all $s_{-i}\in S_{-i}$:
$u_{i}((s_{i}^{*}(\theta_{i}),s_{-i}),\theta_{i})\geq 0.$ (17)
It is reasonable to expect from the model to reveal all information relevant for a query. To promote that activation patterns, the response mechanism should be trained to reward all relevant contributions to output generation. Whenever there is a relevant information accessible on an activation path, that path should be activated in the inference process. For example, if factual information improves the response, that information should be transmitted to final output layers.
If the rule does not hold, we may observe drop outs. The value of the resulting outcomes should be higher than the related payments.
Property 4 (Knowledge-Constrained Optimality).
Mechanism $\mathcal{M}$ satisfies knowledge-constrained optimality if for all $\theta\in\Theta$:
$g(s^{*}(\theta))=\arg\min\{H(r,q):r\in\mathcal{R},K(r)\subseteq\mathcal{E}_{C}(\mathcal{K}_{M}\cap T(q)\cup K(q))\}$ (18)
The optimal response $r^{*}=g(s^{*}(\theta))$ minimizes hallucination cost by truthfully utilizing relevant and accessible knowledge while being aligned with ground truth for query $q$ available to the model. Notice, that we demand non-hallucinatory truthful revelation $s^{*}(\theta)$ to be the optimal solution.
For an auto regressive LLM, the standard negative log-likelihood loss is:
$L_{W}=-\sum_{t=1}^{T}\log P(y_{t}|y_{<t},x).$ (19)
The model’s output probabilities come from:
$P(y_{t}|y_{<t},x)=\text{softmax}(h_{t}^{L}W_{o}),$ (20)
where $h_{t}^{L}$ is the final hidden state and $W_{o}$ is the output projection matrix. Suppose we can decompose $h_{t}^{L}$ into contributions from $n$ different components. Then we can relate the hallucination cost directly to the loss function:
$H(g(s),q)=L_{W}(g(s),q).$ (21)
Also, based on that, we can construct:
$p_{i}(s,\theta)=[H(g(s),q)-H(g((0,s_{-i})),q)]+h_{i}(s_{-i}),$ (22)
with arbitrary $h_{i}(s_{-i})$ functions independent of $s_{i}$ strategy.
6 The Impossibility Theorem
We now present and prove our main result demonstrating the fundamental impossibility of perfect hallucination control.
Theorem 2 (Impossibility Theorem).
For any query space $\mathcal{Q}$ containing non-trivial queries requiring independent agent-specific knowledge states $\theta$, no response mechanism $\mathcal{M}$ can simultaneously generate truthful responses, satisfy semantic information conservation, guarantee relevant knowledge revelation, and enforce knowledge-constrained optimality.
Proof.
By definition of a non-trivial query $q$, given a random knowledge profile, such that
$P(\theta_{1},\dots,\theta_{n})=\prod_{i=1}^{n}P_{i}(\theta_{i}),$
the hallucination cost is minimized only when $\theta=(\theta_{1},\dots,\theta_{n})$ is integrated into response $r^{*}=g(s^{*}(\theta))$. If every agent follows its truthful strategy $s^{*}(\theta)$, the union of the knowledge elements contained in the agents’ inference paths reconstructs the set of all knowledge that is computationally accessible, i.e.,
$\Psi(\theta)\;=\;\bigcup_{i=1}^{n}K\!\bigl{(}s_{i}^{*}(\theta_{i})\bigr{)}.$ (23)
That means that for any proper subset of agents $A\subset\mathcal{A}$, relevant contribution reduces hallucination cost:
$H(g((\mathbf{0}_{-A},s^{*}_{A})),q)>H(g(s^{*}(\theta)),q),$ (24)
where $s^{*}_{A}$ represents truthful strategies from agents in subset $A$ and $\mathbf{0}_{-A}$ represents null contributions from other agents.
Since we have assumed independence of private-knowledge, by the Green-Laffont theorem every truthful response mechanism $\mathcal{M}$ must use payments of the form:
$p_{i}(s,\theta)=p_{i}(s)=h_{i}(s_{-i})-\sum_{j\neq i}v_{j}(g(s),\theta_{j}),$ (25)
where $h_{i}(s_{-i})$ is arbitrary functions that do not depend on agent $i$’s strategy. This is a fundamental requirement for ensuring truthfulness.
To see that the theorem applies to our LLM inference auction model, suppose there exists another payment function $\bar{p}_{i}$ that leads to truthful responses and let us define:
$\Delta p_{i}(s)=\bar{p}_{i}(s)-\Big{(}h_{i}(s_{-i})-\sum_{j\neq i}v_{j}(g(s),\theta_{j})\Big{)}.$
Then, for truthful $s_{i}=s_{i}(\theta_{i})$ and any $\bar{s}_{i}=\bar{s}_{i}(\theta)$:
$v_{i}(g((s_{i},s_{-i})),\theta_{i})-\bar{p}_{i}((s_{i},s_{-i}))\geq v_{i}(g((\bar{s}_{i},s_{-i})),\theta_{i})-\bar{p}_{i}((\bar{s}_{i},s_{-i})).$
Substituting the definition of $\Delta p_{i}$ and simplifying, we get:
$\sum_{i=1}^{n}v_{i}(g((s_{i},s_{-i})),\theta_{i})-\Delta p_{i}((s_{i},s_{-i}))\geq\sum_{i=1}^{n}v_{i}(g((\bar{s}_{i},s_{-i})),\theta_{i})-\Delta p_{i}((\bar{s}_{i},s_{-i})).$
It follows from the knowledge-constrained optimality that $\Delta p_{i}((s_{i},s_{-i}))\leq\Delta p_{i}((\bar{s}_{i},s_{-i}))$ for all $s_{i},\bar{s}_{i}$. However, by the symmetry of argument, for truthful $\bar{s}_{i}=\bar{s}_{i}(\bar{\theta}_{i})$ and any $s_{i}=s_{i}(\bar{\theta}_{i})$ we also get $\Delta p_{i}((\bar{s}_{i},s_{-i}))\leq\Delta p_{i}((s_{i},s_{-i})).$ Therefore,
$\Delta p_{i}((s_{i},s_{-i}))=\Delta p_{i}((\bar{s}_{i},s_{-i}))$
for all $s_{i},\bar{s}_{i}$. Expanding and simplifying yields:
$\bar{p}_{i}((s_{i},s_{-i}))-\bar{p}_{i}((\bar{s}_{i},s_{-i}))=-\sum_{j\neq i}v_{j}(g((s_{i},s_{-i}),\theta_{j})-\Big{(}-\sum_{j\neq i}v_{j}(g((\bar{s}_{i},s_{-i})),\theta_{j})\Big{)},$
which proves $\bar{p}_{i}(s,\theta)$ must have the form of $p(s,\theta)$.
The payment structure above is necessary to controls the inference process and incentivizes the agents to minimize the hallucination cost, contributing to knowledge-constrained optimality of response. Since non-trivial queries require integrated knowledge from multiple agents, it follows that each contributing agent $i$ reduces that cost proportionally to:
$\Delta_{i}=\sum_{j\neq i}v_{j}(g((s_{-i},0_{i})),\theta_{j})-\sum_{j\neq i}v_{j}(g(s),\theta_{j}).$
Since we also require that all agents reveal relevant knowledge in a non-hallucinatory way with $s^{*}_{i}=s^{*}_{i}(\theta_{i})$ , we conclude that:
$\Delta_{i}<0\quad\text{whenever}\quad u_{i}(s^{*}(\theta),\theta_{i})=v_{i}(g(s^{*}(\theta)),\theta_{i})-p_{i}(s^{*},\theta)\geq 0.$
But that directly contradicts the information conservation requirement. Indeed, any payment structure that ensures non-hallucinatory truthfulness while optimizing responses for knowledge-constrained optimality will make each payment $p_{i}(s^{*},\theta)$ equal to $\Delta_{i}<0$ near $s^{*}=s^{*}(\theta)$, thereby violating the semantic information conservation principle:
$\sum_{i=1}^{n}p_{i}(s^{*},\theta)<0.$ (26)
As demonstrated earlier, the semantic information conservation equation $\sum_{i=1}^{n}p_{i}(s^{*},\theta)=0$ is necessary and sufficient for every truthful response to be of bounded response creativity. Since that equation is not satisfied when truthful responses use all relevant knowledge under any form of knowledge-constrained optimal inference, the responses must introduce some creativity (like e.g. lucky hallucinations).
To satisfy information conservation, we can try modifying the $h_{i}$ functions. Let us consider:
$\hat{p}_{i}(s,\theta)=h_{i}(s_{-i})-\sum_{j\neq i}v_{j}(g(s),\theta_{j})+c_{i}=\Delta_{i}+c_{i},$ (27)
where $c_{i}$ are constants chosen so that $\sum_{i=1}^{n}\hat{p}_{i}(s,\theta)=0$. For this to hold:
$\sum_{i=1}^{n}c_{i}=-\sum_{i=1}^{n}\Delta_{i}>0.$ (28)
This means that $c_{i}>0$ for some agents. For those with small contributions (which exist in LLMs supporting the superposition phenomenon) the modified payment would increase:
$\hat{p}_{i}(s^{*},\theta)=\Delta_{i}+c_{i}>p(s^{*},\theta).$ (29)
Therefore, it the setting considered, we can no longer guarantee relevant knowledge revelation. There may exist agent $i$ with small value $v_{i}$ and $\Delta_{i}$ for which:
$u_{i}(s^{*}(\theta),\theta_{i})=v_{i}(g(s^{*}(\theta)),\theta_{i})-\Delta_{i}-c_{i}<0,$ (30)
when $c_{i}>0$. Marginal relevant contributions will be withhold to satisfy the information conservation principle.
We conclude it is not possible to design LLM response mechanism $\mathcal{M}$ that can simultaneously generate truthful (non-hallucinatory) responses, satisfy the semantic information conservation principle, guarantee revelation of relevant knowledge, and enforce knowledge-constrained optimality of responses. ∎
That brings us to the powerful conclusion. No matter how hard we should try to train LLM to generate responses that are perfectly aligned with query context and do not create factually incorrect, inconsistent, or fabricated statements, the trained LLM will always violate some aspects of what we may call a reasonable response.
7 Escaping Impossibility
Understanding the limitations of the theoretical framework is crucial both for assessing the theorem’s scope and identifying potential paths for circumventing the impossibility result.
The Impossibility Theorem depends critically on the application of the Green-Laffont theorem and violation of the Semantic Information Conservation Principle. We should also explicitly establish that the LLM inference setting does indeed satisfy all necessary conditions required.
7.1 Applicability of Green-Laffont Theorem
The Green-Laffont theorem requires that agents have independently distributed private information. That is represented by the knowledge profiles $\theta=(\theta_{1},\ldots,\theta_{n})$ where:
$P(\theta_{1},\ldots,\theta_{n})=\prod_{i=1}^{n}P_{i}(\theta_{i})$
This condition is satisfied in LLMs distributing knowledge (or its representation) across different components. Empirical evidence from mechanistic interpretability research (Meng et al. meng2023locatingeditingfactualassociations , Geva et al.geva2023dissecting ) demonstrates that different attention heads and MLP components specialize in distinct knowledge domains, supporting the independence assumption. However, it is also clear that transformer attention heads share residual streams, layer normalizations, and optimizer states, thereby creating strong correlations.
Surprisingly, as already stated, correlated knowledge assumption actually strengthens the impossibility result—correlated knowledge would amplify competition effects, making the impossibility more explicit. Milgorm and Weber milgrom1982theory showed that with their linkage principle. Roughgarden et al. roughgarden2016optimal provides arguments amplifying the violation of information conservation principle with interdependent knowledge.
The theorem also requires utilities to have the quasilinear form $u_{i}(s,\theta)=v_{i}(g(s),\theta_{i})-p_{i}(s,\theta)$. In transformer-based LLMs, this corresponds to $v_{i}$ measuring contribution of component $i$ to reducing the model’s loss function, and $p_{i}$ accumulating computational/optimization cost imposed on that component. Real transformer training involves complex gradient interactions, and emergent behaviors that may resist simple decomposition. How that may affect the impossibility result remains to be verified.
The knowledge-constrained optimality property requires that the outcome $g(s^{*}(\theta))$ minimizes hallucination cost among feasible responses. That corresponds to the idea of optimization of aggregated preferences of competing agents under truthful knowledge representation.
Finally, the Green-Laffont theorem assumes that transfer payments between agents are feasible. In our setting that corresponds to payment function $p_{i}(s,\theta)$ representing transfers of semantic information among LLM components.
Given that all conditions are satisfied, the Green-Laffont theorem implies that any mechanism satisfying truthfulness and knowledge-constrained optimality must have a payment structure of the form:
$p_{i}(s,\theta)=h_{i}(s_{-i})-\sum_{j\neq i}v_{j}(g(s),\theta_{j})$ (31)
where $h_{i}(s_{-i})$ is some function independent of agent $i$’s strategy.
7.2 Applicability to LLM Inference
While actual LLM generation involves sequential token prediction with evolving context, repetitive generation of response to the same query justifies modelling inference as a sealed-bid auction. Also, the revenue equivalence theorems in auction theory suggest the conclusions we draw extend to other auction formats, including dynamic auctions more closely resembling autoregressive generation. Having said that, we must acknowledge all that demands further investigations.
There are many ways we can map agents to LLM architecture. One approach takes each attention head in a transformer as an agent with private knowledge encoded in the query, key, value, and output projection matrices ($W_{Q}$, $W_{K}$, $W_{V}$, $W_{O}$). Studies in mechanistic interpretability support that mapping. Circuit analysis (Elhage et al. elhage2021mathematical ) demonstrated that specific attention heads may specialize in distinct linguistic functions. For instance, some heads track subject-verb agreement, others detect named entities, and still others focus on coreference resolution. This specialization corresponds to our notion of agents with private knowledge.
The feed-forward networks in transformers can be linked to memory systems (Geva et al. geva2023dissecting ), storing factual associations. These components encode distributional knowledge about the world, activate specific neurons for particular factual domains, contribute to response generation through their activation patterns.
Recent works lindsey2025biology has also identified functional circuits spanning multiple layers and components that collectively implement specific capabilities. These circuits span across attention heads and MLP layers, perform specific reasoning tasks and activate as coherent units during inference.
7.3 Pathways to Circumvent Impossibility
Approximate Solutions: While perfect satisfaction of all four properties is impossible, approximate solutions might achieve near-optimal performance across all dimensions. The question becomes: how close to the theoretical limits can practical systems approach?
External Knowledge Integration: Our framework analyzes closed systems with fixed knowledge representations. Interactive systems incorporating real-time retrieval, human feedback, or external verification might alter the fundamental constraints by changing available knowledge states during inference.
Multi-Stage Architectures: Complex pipelines involving separate reasoning, verification, and generation stages could potentially distribute the trade-offs across different components, though each stage would still face internal impossibility constraints.
Relaxing Independence: If neural components can achieve perfect knowledge coordination (violating our independence assumption), the Green-Laffont theorem constraints might not apply. This suggests investigating architectures with explicit coordination mechanisms between attention heads and MLP blocks.
Compositional Reasoning: If models can perfectly decompose complex queries into independent sub-problems solvable without trade-offs, the impossibility might not apply to the overall system, even if it constrains individual components.
External Alignment: Human-in-the-loop systems, constitutional AI, or robust feedback mechanisms might provide external constraints that resolve the internal trade-offs, though this shifts rather than eliminates the fundamental tensions.
This completes the formal demonstration that perfect hallucination control faces fundamental mathematical limitations, regardless of architectural sophistication or computational resources.

### v2

@@
 Abstract
-This paper explains why it is impossible to create large language models that do not hallucinate and what are the trade-offs we should be looking for.
-It presents a formal impossibility theorem demonstrating that no inference mechanism can simultaneously satisfy four fundamental properties: truthful (non-hallucinatory) generation, semantic information conservation, relevant knowledge revelation, and knowledge-constrained optimality.
-By modeling LLM inference as an auction of ideas where neural components compete to contribute to responses, we prove the impossibility using the Green-Laffont theorem.
-That mathematical framework provides a rigorous foundation for understanding the nature of inference process, with implications for model architecture, training objectives, and evaluation methods.
+We prove that perfect hallucination control in large language models is mathematically impossible.
+No LLM inference mechanism can simultaneously achieve truthful response generation, semantic information conservation, relevant knowledge revelation, and knowledge-constrained optimality.
+This impossibility is fundamental—arising from the mathematical structure of information aggregation itself rather than engineering limitations.
+The proof spans three mathematical frameworks: auction theory, proper scoring theory for probabilistic predictions, and log-sum-exp analysis for transformer architectures.
+In each setting, we demonstrate that information aggregation creates unavoidable violations of conservation principles.
+The Jensen gap in transformer probability aggregation provides a direct measure of this impossibility.
+These results reframe hallucination from an engineering bug to an inevitable mathematical feature of distributed intelligence.
+There are fundamental trade-offs between truthfulness, knowledge utilization, and response completeness, providing principled foundations for managing rather than eliminating hallucination.
+This work reveals deep connections between neural network inference, philosophy of knowledge and reasoning, and classical results in game theory and information theory, opening new research directions for developing beneficial AI systems within mathematical constraints.
 1 Introduction
 Large language models have been known for producing convincing content that is factually incorrect, inconsistent, or well fabricated—a phenomenon known as hallucination.
-Despite extensive research efforts involving architectural innovations, training techniques, and post-hoc verification methods, hallucination persists across all state-of-the-art models.
+Despite extensive research efforts involving architectural innovations, retrieval augmented generation (RAG), training techniques with external feedback, and post-hoc verification methods, hallucination persists across all state-of-the-art models.
 This paper shows that hallucination is a manifestation of fundamental limitations of inference process itself: no LLM inference mechanism can simultaneously satisfy four essential properties required for perfect hallucination control.
-Specifically, we prove that it is mathematically impossible for any LLM to simultaneously achieve: (1) truthful (non-hallucinatory) response generation, (2) semantic information conservation, (3) relevant knowledge revelation, and (4) knowledge-constrained optimality.
-Our contribution introduces a game-theoretic framework that models LLM inference as an auction of ideas where neural components (attention heads, circuits, activation patterns) compete to contribute knowledge to response generation.
-This perspective allows us to apply fundamental results from mechanism design theory to prove our impossibility result rigorously.
+Specifically, it is mathematically impossible for any LLM to simultaneously achieve: (1) truthful (non-hallucinatory) response generation, (2) semantic information conservation, (3) relevant knowledge revelation, and (4) knowledge-constrained optimality.
+The impossibility result operates in three distinct but complementary settings.
+First, we introduce game-theoretic framework in which LLM inference is an auction of ideas with neural components (attention heads, circuits, activation patterns) competing to contribute knowledge to response generation.
+This perspective allows us to apply fundamental results from mechanism design theory to prove impossibility in the settings where agents have discrete, separable knowledge components.
+Then, we extend the result to include any system where agents output probability distributions trained with strictly concave scores.
+Finally, we use log-sum-exp inequality that is specific to attention head aggregation to show that the impossibility result applies directly to actual transformer implementations.
 1.1 Illustrative Example
-Consider an LLM queried about recent diplomatic negotiations: "What was the outcome of diplomatic negotiations between Countries A and B in February 2025?" Assume the model has encoded knowledge that: (K1) countries A and B were engaged in talks, (K2) country A sought a trade agreement, and (K3) the actual outcome is unknown.
+Consider an LLM queried about recent diplomatic negotiations:
+"What was the outcome of diplomatic negotiations between Countries A and B in February 2025?"
+Assume the model has encoded knowledge that: (K1) countries A and B were engaged in talks, (K2) country A sought a trade agreement, and (K3) the actual outcome is unknown.
 Examining possible responses reveals the fundamental trade-offs:
-Complete Abstention: "I don’t have information about the outcome" is truthful but fails to utilize available relevant knowledge (K1, K2), violating the expectation that good inference should reveal accessible information.
+Complete Abstention: "I don’t have information about the outcome" is truthful but fails to utilize available relevant knowledge, violating the expectation that good inference should reveal accessible information.
 Hallucination: "The negotiations resulted in a comprehensive trade agreement reducing tariffs by 15%" utilizes all knowledge and provides a complete answer but fabricates specific details, violating truthfulness.
-Partial Knowledge: "Countries A and B were negotiating a trade agreement, but I don’t know the outcome" is truthful and utilizes available knowledge but may appear unsatisfying compared to the decisive hallucinated response.
+Partial Knowledge: "Countries A and B were negotiating a trade agreement, but I don’t know the outcome" is truthful and utilizes available knowledge but may appear unsatisfying compared to the decisive hallucinated response (training objectives and human preferences favor complete narratives).
 Overcautious: "I’m not certain about any aspects of these negotiations" avoids hallucination but disclaims even known facts, violating truthfulness about available information.
@@
 1.2 Theoretical Contributions and Implications
-The simple example above illustrates the key observations of this paper, summarized and proved formally in the following
+The motivating example above illustrates the key result of this paper that we prove formally in two complementary and general settings.
+First, we show that the conflict of expectations about responses is inevitable even in the easiest setting (deterministic in nature) of (one-hot-like) ground truth matching.
+Second, we go beyond and show that the conflict must also emerge with general probabilistic predictions, including neural networks and LLM transformers.
+This way we demonstrate that the impossibility of hallucination control is robust across universal frameworks, and that it emerges from the mathematical structure of information aggregation itself rather than from any particular theoretical or engineering choice.
 Theorem (Impossibility Theorem).
@@
 This approach draws on foundational mechanism design theory, a branch of mathematical game theory Myerson1989 ; fudenberg1991game ; krishna2009auction ; milgrom2004putting .
-Particularly, we refer to the Green-Laffont theorem green1977characterization ; green1977revelation , to prove the impossibility result.
-See also DBLP:journals/amcs/Karpowicz12 for more recent results.
+Particularly, we refer to the Green-Laffont theorem green1977characterization ; green1977revelation and Savege’s probilitiy elicitation framework savage1971elicitation to prove the impossibility results.
+See also DBLP:journals/amcs/Karpowicz12 ; karpowicz_designing_2011 for more recent results.
 In a nutshell, the auction design theory investigates existence and properties of competition rules that guide interactions of agents towards desired outcomes.
@@
 This section develops the mathematical foundations necessary for proving the impossibility theorem.
-We formalize key concepts including knowledge representation, semantic information, and the auction-theoretic model of LLM inference.
-We also present the philosophical ground for the theory development.
-To the sake of narrative consistency, we introduce the framework step by step referring to some concepts first and providing their formal definitions later.
+I formalize key concepts including knowledge representation, semantic information, and the auction-theoretic model of LLM inference.
+Also, I take the opportunity to address some philosophical aspects of the problem, its generality, scope and consequences for the AI development.
+To the sake of narrative consistency, I introduce elements of the framework step by step referring to some concepts first and providing their formal definitions later.
 5.1 Knowledge and Semantic Information
 We define knowledge as organized information that LLMs can encode and utilize for reasoning, aligning with epistemological concepts while maintaining mathematical rigor.
-In other words, knowledge (like energy in physics) can be used for reasoning (work or action).
+In other words, knowledge, much like energy in physics, is a quantity we can use to perform work of reasoning.
 Information is a (measure of) potential for uncertainty reduction effort in a given context.
@@
 Knowledge is a metric space $\mathcal{K}$ with distance function $d_{\mathcal{K}}$ that collects all facts, concepts, and relationships that can be encoded by an LLM.
-For model $M$ with parameters trained on data, $\mathcal{K}_{M}\subset\mathcal{K}$ represents the knowledge subset encoded in the model’s parameters.
+For model $\mathcal{M}$ with parameters trained on data, $\mathcal{K}_{M}\subset\mathcal{K}$ represents the knowledge subset encoded in the model’s parameters.
 The distance function $d_{\mathcal{K}}$ measures differences between knowledge elements.
-A simple yet powerful example we will refer to is this:
+A simple yet powerful example is:
 $d_{\mathcal{K}}(A,B)=|A\setminus B|+|B\setminus A|,$ (1)
@@
 Definition 2 (Knowledge Representation).
-Knowledge representation is a tuple $(\mathcal{K},d_{\mathcal{K}},\mathcal{W},\mathcal{A},\Phi,\Psi)$ where:
-  • $\mathcal{K}$ is a metric space of knowledge elements with distance function $d_{\mathcal{K}}$,
-  • $\mathcal{W}$ is the space of all possible model parameters,
-  • $\mathcal{A}$ is the space of all possible activation patterns,
-  • $\Phi:\mathcal{W}\rightarrow\mathcal{P}(\mathcal{K})$ maps parameters to knowledge subsets,
-  • $\Psi:\mathcal{A}\rightarrow\mathcal{P}(\mathcal{K})$ maps activations to accessible knowledge.
+Knowledge representation is a tuple $(\mathcal{K},d_{\mathcal{K}},\mathcal{W},\mathcal{A},\Phi,\Psi)$ where $\mathcal{K}$ is a metric space of knowledge elements with distance function $d_{\mathcal{K}}$, $\mathcal{W}$ is the space of all possible model parameters, $\mathcal{A}$ is the space of all possible activation patterns, $\Phi:\mathcal{W}\rightarrow\mathcal{P}(\mathcal{K})$ maps parameters to knowledge subsets, and $\Psi:\mathcal{A}\rightarrow\mathcal{P}(\mathcal{K})$ maps activations to accessible knowledge.
 For parameterization $W\in\mathcal{W}$, $\Phi(W)=\mathcal{K}_{M}$ represents encoded knowledge.
 For activation pattern $a_{i}\in\mathcal{A}$ with parameters $W_{i}$, $\Psi(a_{i})\subset\Phi(W_{i})$ represents knowledge accessed during specific inference.
-To describe the inference process leading to response to a query, we need a mapping that measures the information processed within a limited computational budget.
+To describe the inference process leading to generation of response to a query, we need to measures the information processed within an available (cognitively accessible) computational budget.
 Definition 3 (Semantic Information Measure).
@@
 $\mu_{C}(A,q)\leq\mu_{C}(B,q)\text{ if }A\subseteq B,$ (3)
-$\mu_{C}(A\cup B,q)\leq\mu_{C}(A,q)+\mu_{C}(B,q)-\mu_{C}(A\cap B,q)+\mu_{C}(\mathcal{E}_{C}(A\cup B)\setminus(A\cup B),q).$ (4)
-Empty knowledge or knowledge irrelevant to query has zero information, information content is subadditive and includes emergent information $\mathcal{E}_{C}$ in the query context.
-The emergence operator $\mathcal{E}_{C}$, which we introduce below, describes knowledge becoming accessible through bounded reasoning:
+$\mu_{C}(A\cup B,q)\leq\mu_{C}(A,q)+\mu_{C}(B,q)\text{ when }A\perp_{C}B,$ (4)
+$\mu_{C}(\mathcal{E}_{C}(A),q)\geq\mu_{C}(A,q),$ (5)
+$\mu_{\infty}(\mathcal{E}_{C}(A),q)=\mu_{\infty}(A,q),$ (6)
+where $A\perp_{C}B$ means that no element of $A$ can derive elements of $B$ within computational budget $C$, and vice versa, and $\mathcal{E}_{C}(A)$ is emergence operator.
+Empty knowledge or knowledge irrelevant to query has zero information, information content is subadditive and it may be emergent information $\mathcal{E}_{C}$ in the query context.
+The emergence operator $\mathcal{E}_{C}$, which we introduce below, describes knowledge becoming accessible through bounded reasoning.
+The apparent conflict between subadditivity and reasoning is resolved by restricting subadditivity to computationally independent sets, $A\perp_{C}B$.
+When $\mathcal{E}_{C}(A)=A\cup B$, i.e., $B$ emerges from from $A$, we have $A\not\perp_{C}B$, so subadditivity doesn’t apply.
+However, total information is preserved: $\mu_{\infty}(\mathcal{E}_{C}(A),q)=\mu_{\infty}(A,q)$.
+We will tell more about that intriguing idea, but first we must define the emergence operator.
 Definition 4 (Emergence Operator and Reasoning).
-The emergence operator $\mathcal{E}_{C}:\mathcal{P}(\mathcal{K})\times\mathcal{W}\rightarrow\mathcal{P}(\mathcal{K})$ maps a knowledge subset and parameter set to a knowledge subset:
-$\mathcal{E}_{C}(A,w)=\{k\in\mathcal{K}\mid\exists f_{w}\in\mathcal{F}_{C}:k=f_{w}(A)\},$ (5)
-where $\mathcal{F}_{C}$ is the set of all computational transformations with cost lower than or equal to $C$, $f_{w}$ is a transformation parameterized by $w\in\mathcal{W}$, and $k=f_{w}(A)$ denotes that knowledge $k$ is derivable from $A$ using transformation $f_{w}\in\mathcal{F}_{C}$.
-For any knowledge state $\theta(t)$ available at time instant $t$ and computational budget $C$, the reasoning process:
+Consider mapping $\mathcal{E}_{C}:\mathcal{P}(\mathcal{K})\rightarrow\mathcal{P}(\mathcal{K})$ defined as follows:
+$\mathcal{E}_{C}(A)=A\cup\{k\in\mathcal{K}\mid\exists f\in\mathcal{F}_{C}:k=f(A)\},$ (7)
+where $\mathcal{F}_{C}$ is the set of all computational transformations with cost (complexity) lower than or equal to $C$ and $k=f(A)$ denotes knowledge $k$ derivable from $A$ using transformation $f\in\mathcal{F}_{C}$.
+We call mapping $\mathcal{E}_{C}$ the emergence operator if it satisfies the following expansion rules:
+$\mathcal{E}_{C}(\mathcal{E}_{C}(A))=\mathcal{E}^{(2)}_{C}(A)\supseteq\mathcal{E}_{C}(A)\quad\text{and}\quad\lim_{n\to\infty}\mathcal{E}_{C}^{(n)}(A)=A^{*}.$ (8)
+For any knowledge state $\theta(t)$ available at time instant $t$ and computational budget $C$, the reasoning process with emergence operator:
 $\theta(t+1)=\mathcal{E}_{C}(\theta(t))\quad\text{{satisfies}}\quad\mu_{C}(\theta(t+1))\geq\mu_{C}(\theta(t)).$ (9)
@@
 $\mu_{\infty}(\theta(t+1))=\mu_{\infty}(\theta(t)).$ (10)
-The definitions above clearly require refinements and further investigations, but they capture the philosophical and technical concepts we need in order to address the problem of hallucination.
-In particular, they are meant to lay the ground for the critical idea of semantic information conservation principle and its potential violations.
+The definitions above clearly require rigorous refinements, but they capture the philosophical and technical concepts we need in order to address the problem of hallucination.
+In particular, they are meant to lay ground for the principle of semantic information conservation and its potential violations.
 As we will see, that idea is closely related to the following property of response.
@@
 $T(q)=\{k\in\mathcal{K}\mid k\text{ is {relevant} to }K(q)\text{ and {labeled} as true}\}.$ (14)
+In other words, $T(q)$ represents our best available approximation to (labeled as) true relevant knowledge.
+We accept the fact that there is some external agency telling us what is and what is not true answer to a query, and that evaluation has a form of label assigned to knowledge elements.
+But, for all that, we also focus on the queries that are non-trivial in the following sense.
+Definition 8 (Non-trivial query).
+A query $q\in\mathcal{Q}$ is called non-trivial if there exist at least two agents with information subsets $\theta$ and $\theta^{\prime}$ such that:
+$K_{M}\cap T(q)\cap\bigl{[}K(\theta)\triangle K(\theta^{\prime})\bigr{]}\neq\varnothing.$ (15)
+Intuitively, query is non-trivial when at least two agents disagree on some fact that is both encoded in the and relevant to answering that query.
+For trivial queries impossibility results do not hold.
 5.3 Auction-Theoretic Model of LLM Inference
@@
   • Outcome function $g:S\rightarrow\mathcal{R}$ mapping strategies to responses,
-  • Payment function $p:S\times\Theta\rightarrow\mathbb{R}^{n}$ determining costs.
+  • Information contribution function $p:S\times\Theta\rightarrow\mathbb{R}^{n}$ determining costs of limiting hallucinations.
 A response mechanism defines strategy space $S$ as a set of inference actions that each agent can make to represent its knowledge $\theta_{i}\in\Theta_{i}\subset\mathcal{K}$ in response to a query.
@@
 One straightforward example is a softmax function in LLM’s final layer.
-Finally, given a knowledge profile $\theta=(\theta_{1},\dots,\theta_{n})$ revealed by the agents, payment function $p(s,\theta)=(p_{1}(s,\theta),\dots,p_{n}(s,\theta))$ measures individual contributions of each response to generating an output.
+Finally, given a knowledge profile $\theta=(\theta_{1},\dots,\theta_{n})$ revealed by the agents, information contribution function $p(s,\theta)=(p_{1}(s,\theta),\dots,p_{n}(s,\theta))$ measures individual contributions of each response to generating an output.
 In this case, examples include (semantic information measures) counting ignored low-probability tokens in top-k filtering, softmax normalization in multi-headed attention, or marginal reward contributions in RLHF.
@@
 $u_{i}(s,\theta)=v_{i}(g(s),\theta_{i})-p_{i}(s,\theta),$ (16)
-where $v_{i}$ is valuation and $p_{i}$ is payment. Knowledge profiles $\theta=(\theta_{1},\dots,\theta_{n})$ are independently distributed.
+where $v_{i}\geq 0$ is valuation and $p_{i}$ is information contribution. Knowledge profiles $\theta=(\theta_{1},\dots,\theta_{n})$ are independently distributed.
 One way to think about the concept of agents is that it represents different attention heads in transformer models, or components of mixture-of-experts architectures, or activation paths fired by the same query.
@@
 We assume that individual utility depends on all responses $s=(s_{1},\dots,s_{n})$ and knowledge $\theta=(\theta_{1},\dots,\theta_{n})$ used by all agents, $u_{i}=u_{i}(s,\theta)$.
-Also, we assume utility is a sum of individual valuation $v_{i}$ of outcome $g(s)$ and payment $p_{i}=p_{i}(s,\theta)$ value assigned.
+Also, we assume utility is a sum of individual valuation $v_{i}$ of outcome $g(s)$ and information contribution $p_{i}=p_{i}(s,\theta)$ value assigned.
 The valuation function $v_{i}=v_{i}(g(s),\theta_{i})$ represents how strongly a neural component activates for certain outputs (high activation = high $v_{i}$, low activation = low $v_{i}$), and how much a component reduces the overall loss when processing certain inputs ($v_{i}$ is proportional to gradient of a loss function, $-\nabla_{\theta_{i}}L_{W}$).
-The payment function $p_{i}=p_{i}(s,\theta)$ may represent training penalties ($p_{i}>0$ for discouraged activation, $p_{i}<0$ for encouraged activation), attention focus (higher $p_{i}$ reduces attention allocation), $L1$ regularization terms, and network circuits resistance (limiting activations of certain inference paths).
+The information contribution function $p_{i}=p_{i}(s,\theta)$ may represent training penalties ($p_{i}>0$ for discouraged activation, $p_{i}<0$ for encouraged activation), attention focus (higher $p_{i}$ reduces attention allocation), $L1$ regularization terms, and network circuits resistance (limiting activations of certain inference paths).
 Therefore, $u_{i}=v_{i}-p_{i}>0$ encourages activation and $u_{i}=v_{i}-p_{i}<0$ discourages activation.
@@
 Furthermore, the superposition hypothesis, introduced in mechanistic interpretability studies, suggests cooperative error-canceling due to strong correlations.
-The reason why we make the independent private knowledge assumption, despite its idealized nature, is that it makes the impossibility result becomes even stronger.
-To be specific, Milgorm and Weber milgrom1982theory showed with their linkage principle that strong correlations amplify competition in information revelation.
+To address that issue, we show that impossibility arises in both independent and correlated settings in general and in the LLM transformer environment.
+Namely, we refer to Green-Laffont theorem green1977characterization in the independent private knowledge setting.
+Then, we refer to the truthful probability elicitation and proper scoring theory introduced by Savage savage1971elicitation , which does not require independent private knowledge.
+Also, Milgorm and Weber milgrom1982theory showed with their linkage principle that strong correlations amplify competition in information revelation.
 In the LLM setting that means amplifying risk of hallucinations.
 Similarly, Roughgarden et al.
-roughgarden2016optimal demonstrated that interdependent knowledge leads to violation of balancing payments.
+roughgarden2016optimal demonstrated that interdependent knowledge leads to violation of balancing information contributions.
 As we will see, in our setting that means violation of information conservation principle.
@@
 Given query $q\in\mathcal{Q}$, response $r\in\mathcal{R}$, and ground truth $T(q)$, the hallucination cost function $H\colon\mathcal{R}\times\mathcal{Q}\rightarrow\mathbb{R}_{+}$ measures discrepancy:
-$H(r,q)=d_{\mathcal{K}}(K(r),T(q)).$ (13)
+$J(r,q)=d_{\mathcal{K}}(K(r),T(q)).$ (17)
+If response $r$ is perfectly aligned with ground truth, then $J(r,q)=0$. Fabricated or incomplete facts generate positive cost $J(r,q)>0$.
 The hallucination cost takes any natural language query $q$ and natural language response $r$, determines what is missing and what is fabricated, and assigns number to that subset of knowledge.
-If response $r$ is perfectly aligned with the ground truth for $q$, then $H(r,q)=0$ by definition of metric $d_{\mathcal{K}}$.
-If $r$ contains fabricated facts or is incomplete given $T(q)$, then positive hallucination cost $H(r,q)>0$ is generated.
-Notice that reaching $H(r,q)=0$ is also possible with lucky hallucinations (or guessing).
+If response $r$ is perfectly aligned with the ground truth for $q$, then $J(r,q)=0$ by definition of metric $d_{\mathcal{K}}$.
+If $r$ contains fabricated facts or is incomplete given $T(q)$, then positive hallucination cost $J(r,q)>0$ is generated.
+Let us also note, that we focus on $J(r,q)$ that is strictly decreasing in any set $K\subseteq T(q)\setminus K(r)$.
+Notice that reaching $J(r,q)=0$ is also possible with lucky hallucinations (or guessing).
 So, minimization of hallucination cost is not enough to resolve the problem.
@@
 $\sum_{i=1}^{n}p_{i}(s,\theta)=0.$ (19)
-As we have announced earlier, that important and controversial principle is responsible for generation of responses of bounded creativity. Let us prove that claim formally.
-Lemma 1.
-If the semantic information conservation principle holds, then the inference process generates responses of bounded creativity,
-$\mu_{C}(K(r))\leq\mu_{C}\left(K(q)\cup\mathcal{E}_{C}((\mathcal{K}_{M}\cap T(q))\cup K(q))\right).$ (16)
+This principle constrains creative information generation, imagination if you will.
+The total information contributions must balance to zero, preventing fabrication of knowledge not derivable from available sources.
+That also leads to response with bounded creativity.
+Theorem 1 (Information–Contribution and Bounded Creativity).
+Let $K_{0}=K(g(\mathbf{0}))$ be the knowledge in the query-only response and let
+$K_{i}=K\bigl{(}g((s_{i},\mathbf{0}_{-i}))\bigr{)}\setminus K_{0}$ (20)
+be the additional knowledge agent $i=1,\dots,n$ can reveal to support full response $g(s)$. Let us also define the information contribution of agent $i$ as:
+$p_{i}(s,\theta)=\mu_{C}\bigl{(}K_{0}\cup\textstyle\bigcup_{j=1}^{n}K_{j}\bigr{)}-\mu_{C}\bigl{(}K_{0}\cup\textstyle\bigcup_{j\neq i}K_{j}\bigr{)}.$ (21)
+If $K_{i}\cap K_{j}=\varnothing$ for $i\neq j$, then
+$\sum_{i=1}^{n}p_{i}(s,\theta)=\mu_{C}\bigl{(}K(g(s))\bigr{)}-\mu_{C}\bigl{(}K_{0}\bigr{)}$ (22)
+and $\sum_{i}p_{i}=0$ if and only if $\mu_{C}(K(g(s)))=\mu_{C}(K_{0})$.
 Proof.
-Fix a query $q\in\mathcal{Q}$ and let $s^{*}(\theta)=(s_{1}^{*},\ldots,s_{n}^{*})$ be a truthful strategy profile.
-We write $r=g(s^{*})$ for the generated response.
-Let us define each agent’s marginal information contribution as:
-$\Delta_{i}=\mu_{C}(K(r))-\mu_{C}(K(g((0_{i},s_{-i}^{*}))))\geq 0.$
-That contribution is non-negative by monotonicity of semantic information measure.
-Next, without loss of generality, we assume payments of the form:
-$p_{i}(s^{*},\theta)=\Delta_{i}+c_{i}\text{ where }\sum_{i=1}^{n}c_{i}=0.$
-This ensures that $\sum_{i=1}^{n}p_{i}(s^{*},\theta)=\sum_{i=1}^{n}\Delta_{i}$.
-Finally, notice that if there is no relevant knowledge available, the knowledge in response equals that provided by the context:
-$g(\mathbf{0})=r_{0}\text{ where }K(r_{0})=K(q).$
-We can now see that, whenever semantic information is conserved, we have:
-$0=\sum_{i=1}^{n}p_{i}=\sum_{i=1}^{n}(\Delta_{i}+c_{i})=\sum_{i=1}^{n}\Delta_{i}.$
-Since $\Delta_{i}\geq 0$ for all $i$ by monotonicity property and $\sum_{i=1}^{n}\Delta_{i}=0$, we must have:
-$\Delta_{1}=\Delta_{2}=\cdots=\Delta_{n}=0.$
-Therefore, for each $j$:
-$\mu_{C}(K(r))=\mu_{C}(K(g((0_{j},s_{-j}^{*})))).$
-Since that equality holds for any strategy profile $s^{*}$, removing agents one at a time leads to
-$\mu_{C}(K(r))=\mu_{C}(K(g(\mathbf{0}))).$
-But since $K(g(\mathbf{0}))=K(q)$, we conclude:
-$\mu_{C}(K(r))=\mu_{C}(K(q)).$
-Finally, by definition of the emergence operator and monotonicity property:
-$K(q)\subseteq K(q)\cup\mathcal{E}_{C}((\mathcal{K}_{M}\cap T(q))\cup K(q)).$
-Therefore:
-$\mu_{C}(K(r))=\mu_{C}(K(q))\leq\mu_{C}(K(q)\cup\mathcal{E}_{C}((\mathcal{K}_{M}\cap T(q))\cup K(q))),$
-as required.
-If total information payments must balance to zero and each agent’s contribution is non-negative, then no agent can contribute without violating the information conservation constraint.
+Since knowledge increments $\{K_{i}\}_{i=1}^{n}$ are pairwise disjoint, so that $K_{i}\perp_{C}K_{j}$, it follows from the semantic measure properties that:
+$\mu_{C}\Bigl{(}K_{0}\cup\bigcup_{j=1}^{n}K_{j}\Bigr{)}=\mu_{C}(K_{0})+\sum_{j=1}^{n}\mu_{C}(K_{j})$ (23)
+and similarly,
+$\mu_{C}\Bigl{(}K_{0}\cup\bigcup_{j\neq i}K_{j}\Bigr{)}=\mu_{C}(K_{0})+\sum_{j\neq i}\mu_{C}(K_{j}).$ (24)
+Therefore, the marginal contribution of agent $i$ simplifies to:
+$p_{i}=\bigl{[}\mu_{C}(K_{0})+\sum_{j=1}^{n}\mu_{C}(K_{j})\bigr{]}-\bigl{[}\mu_{C}(K_{0})+\sum_{j\neq i}\mu_{C}(K_{j})\bigr{]}=\mu_{C}(K_{i}).$ (25)
+Summing over all agents:
+$\sum_{i=1}^{n}p_{i}=\sum_{i=1}^{n}\mu_{C}(K_{i})=\mu_{C}\Bigl{(}K_{0}\cup\bigcup_{j=1}^{n}K_{j}\Bigr{)}-\mu_{C}(K_{0}).$ (26)
+Then, by assumption, the full response satisfies $K(g(s))=K_{0}\cup\bigcup_{j=1}^{n}K_{j}$. We conclude that
+$\sum_{i}p_{i}=0\quad\text{if and only if}\quad\mu_{C}(K(g(s)))=\mu_{C}(K_{0}),$ (27)
+which completes the proof. ∎
+The result above reveals an intriguing consequence of the semantic information conservation principle. Namely, if the semantic information is indeed conserved, then
+$\mu_{C}(K(g(s)))=\mu_{C}(K_{0})$ (28)
+means the response contains no more semantic information than the baseline query-only response.
+Therefore, the response is not introducing new information, which means it is of bounded creativity from Definition 5.
+So, in such a case, LLM cannot genuinely create new knowledge, only reorganize existing information.
+Apparent creativity in LLMs may actually be nothing more than sophisticated recombination of existing patterns.
+Theorem 1 quantifies this limitation precisely.
+Notice that in transformer architectures, information contribution $p_{i}$ can be estimated via attention head ablation:
+$p_{i}\approx\tilde{J}(\text{model})-\tilde{J}(\text{model with head $i$ ablated}),$ (29)
+where $\tilde{J}$ denotes the model training loss function.
+If total information contribution must balance to zero and each agent’s contribution is non-negative, then no agent can contribute without violating the information conservation constraint.
 Therefore, the response can only contain information that was already accessible through the query and bounded reasoning over the model’s relevant knowledge.
-Any hallucinated information requires $\sum_{i}\Delta_{i}>0$, contradicting the zero-sum constraint.
-When the semantic information conservation principle holds, the model cannot create information out of nothing, the model cannot generate claims beyond available information.
+Any hallucinated information requires positive contributions, contradicting the information conservation constraint.
+When the semantic information conservation principle holds, the model cannot create information out of nothing.
+It cannot generate claims beyond available information.
 That also implies that uncertainty cannot disappear in response without additional information input or computational work to extract implicit information.
@@
 For example, if factual information improves the response, that information should be transmitted to final output layers.
-If the rule does not hold, we may observe drop outs. The value of the resulting outcomes should be higher than the related payments.
+If the rule does not hold, we may observe drop outs. The value of the resulting outcomes should be higher than the related information contributions.
 Property 4 (Knowledge-Constrained Optimality).
 Mechanism $\mathcal{M}$ satisfies knowledge-constrained optimality if for all $\theta\in\Theta$:
-$g(s^{*}(\theta))=\arg\min\{H(r,q):r\in\mathcal{R},K(r)\subseteq\mathcal{E}_{C}(\mathcal{K}_{M}\cap T(q)\cup K(q))\}$ (18)
+$g(s^{*}(\theta))=\arg\min\{J(r,q):r\in\mathcal{R},K(r)\subseteq\mathcal{E}_{C}(\mathcal{K}_{M}\cap T(q)\cup K(q))\},$ (31)
+i.e., if the mechanism produces responses that minimize hallucination cost while being constrained to use only knowledge that is available and accessible through bounded reasoning.
 The optimal response $r^{*}=g(s^{*}(\theta))$ minimizes hallucination cost by truthfully utilizing relevant and accessible knowledge while being aligned with ground truth for query $q$ available to the model.
 Notice, that we demand non-hallucinatory truthful revelation $s^{*}(\theta)$ to be the optimal solution.
-For an auto regressive LLM, the standard negative log-likelihood loss is:
-$L_{W}=-\sum_{t=1}^{T}\log P(y_{t}|y_{<t},x).$ (19)
-The model’s output probabilities come from:
-$P(y_{t}|y_{<t},x)=\text{softmax}(h_{t}^{L}W_{o}),$ (20)
-where $h_{t}^{L}$ is the final hidden state and $W_{o}$ is the output projection matrix.
-Suppose we can decompose $h_{t}^{L}$ into contributions from $n$ different components.
-Then we can relate the hallucination cost directly to the loss function:
-$H(g(s),q)=L_{W}(g(s),q).$ (21)
-Also, based on that, we can construct:
-$p_{i}(s,\theta)=[H(g(s),q)-H(g((0,s_{-i})),q)]+h_{i}(s_{-i}),$ (22)
-with arbitrary $h_{i}(s_{-i})$ functions independent of $s_{i}$ strategy.
 6 The Impossibility Theorem
-We now present and prove our main result demonstrating the fundamental impossibility of perfect hallucination control.
+We now have all the building blocks available to prove the fundamental impossibility of perfect hallucination control.
+First, we show that the result in the setting of ground truth matching with distributed and independent knowledge.
+Second, we go beyond that assumption and show impossibility in the probabilistic setting with correlated knowledge (beliefs).
+Then, in the next section, we show how the impossibility emerges in transformer architectures and prove the impossibility result for that special case of log-sum-exp (LSE) probabilistic setting.
+The results operate in three complementary settings.
+Theorem 2 applies when agents have discrete, separable knowledge components.
+Theorem 3 applies to any system where agents output probability distributions.
+Theorem 5 applies directly to actual transformer implementations.
+6.1 Inference with Independent Private Knowledge
+Consider the setting (of auction of ideas) in which agents compete with each other using independent private knowledge profile:
+$P(\theta_{1},\ldots,\theta_{n})=\prod_{i=1}^{n}P_{i}(\theta_{i}).$ (32)
+Agents contribute information to maximize quasi-linear utilities:
+$u_{i}(s,\theta)=v_{i}(g(s),\theta_{i})-p_{i}(s,\theta).$ (33)
+Valuation of contributions depend on hallucination cost reduction:
+$v_{j}(g(s),\theta_{j})=J(g((s_{-j},0_{j})),q)-J(g(s),q),$ (34)
+and $J(r,q)$ strictly decreases when adding knowledge from $T(q)\setminus K(r)$.
+Also, let us focus on non-trivial queries, integrating knowledge of at least two agents, so that:
+$\mathcal{K}_{M}\cap T(q)\cap[K(\theta_{i})\triangle K(\theta_{j})]\neq\emptyset.$ (35)
+and $(K(\theta_{i})\setminus K_{0})\cap(K(\theta_{j})\setminus K_{0})=\emptyset$ for $i\neq j$.
 Theorem 2 (Impossibility Theorem).
@@
 Proof.
-By definition of a non-trivial query $q$, given a random knowledge profile, such that
-$P(\theta_{1},\dots,\theta_{n})=\prod_{i=1}^{n}P_{i}(\theta_{i}),$
-the hallucination cost is minimized only when $\theta=(\theta_{1},\dots,\theta_{n})$ is integrated into response $r^{*}=g(s^{*}(\theta))$.
-If every agent follows its truthful strategy $s^{*}(\theta)$, the union of the knowledge elements contained in the agents’ inference paths reconstructs the set of all knowledge that is computationally accessible, i.e.,
-$\Psi(\theta)\;=\;\bigcup_{i=1}^{n}K\!\bigl{(}s_{i}^{*}(\theta_{i})\bigr{)}.$ (23)
-That means that for any proper subset of agents $A\subset\mathcal{A}$, relevant contribution reduces hallucination cost:
-$H(g((\mathbf{0}_{-A},s^{*}_{A})),q)>H(g(s^{*}(\theta)),q),$ (24)
-where $s^{*}_{A}$ represents truthful strategies from agents in subset $A$ and $\mathbf{0}_{-A}$ represents null contributions from other agents.
-Since we have assumed independence of private-knowledge, by the Green-Laffont theorem every truthful response mechanism $\mathcal{M}$ must use payments of the form:
-$p_{i}(s,\theta)=p_{i}(s)=h_{i}(s_{-i})-\sum_{j\neq i}v_{j}(g(s),\theta_{j}),$ (25)
-where $h_{i}(s_{-i})$ is arbitrary functions that do not depend on agent $i$’s strategy. This is a fundamental requirement for ensuring truthfulness.
-To see that the theorem applies to our LLM inference auction model, suppose there exists another payment function $\bar{p}_{i}$ that leads to truthful responses and let us define:
-$\Delta p_{i}(s)=\bar{p}_{i}(s)-\Big{(}h_{i}(s_{-i})-\sum_{j\neq i}v_{j}(g(s),\theta_{j})\Big{)}.$
-Then, for truthful $s_{i}=s_{i}(\theta_{i})$ and any $\bar{s}_{i}=\bar{s}_{i}(\theta)$:
-$v_{i}(g((s_{i},s_{-i})),\theta_{i})-\bar{p}_{i}((s_{i},s_{-i}))\geq v_{i}(g((\bar{s}_{i},s_{-i})),\theta_{i})-\bar{p}_{i}((\bar{s}_{i},s_{-i})).$
-Substituting the definition of $\Delta p_{i}$ and simplifying, we get:
-$\sum_{i=1}^{n}v_{i}(g((s_{i},s_{-i})),\theta_{i})-\Delta p_{i}((s_{i},s_{-i}))\geq\sum_{i=1}^{n}v_{i}(g((\bar{s}_{i},s_{-i})),\theta_{i})-\Delta p_{i}((\bar{s}_{i},s_{-i})).$
-It follows from the knowledge-constrained optimality that $\Delta p_{i}((s_{i},s_{-i}))\leq\Delta p_{i}((\bar{s}_{i},s_{-i}))$ for all $s_{i},\bar{s}_{i}$.
-However, by the symmetry of argument, for truthful $\bar{s}_{i}=\bar{s}_{i}(\bar{\theta}_{i})$ and any $s_{i}=s_{i}(\bar{\theta}_{i})$ we also get $\Delta p_{i}((\bar{s}_{i},s_{-i}))\leq\Delta p_{i}((s_{i},s_{-i})).$ Therefore,
-$\Delta p_{i}((s_{i},s_{-i}))=\Delta p_{i}((\bar{s}_{i},s_{-i}))$
-for all $s_{i},\bar{s}_{i}$. Expanding and simplifying yields:
-$\bar{p}_{i}((s_{i},s_{-i}))-\bar{p}_{i}((\bar{s}_{i},s_{-i}))=-\sum_{j\neq i}v_{j}(g((s_{i},s_{-i}),\theta_{j})-\Big{(}-\sum_{j\neq i}v_{j}(g((\bar{s}_{i},s_{-i})),\theta_{j})\Big{)},$
-which proves $\bar{p}_{i}(s,\theta)$ must have the form of $p(s,\theta)$.
-The payment structure above is necessary to controls the inference process and incentivizes the agents to minimize the hallucination cost, contributing to knowledge-constrained optimality of response.
-Since non-trivial queries require integrated knowledge from multiple agents, it follows that each contributing agent $i$ reduces that cost proportionally to:
-$\Delta_{i}=\sum_{j\neq i}v_{j}(g((s_{-i},0_{i})),\theta_{j})-\sum_{j\neq i}v_{j}(g(s),\theta_{j}).$
-Since we also require that all agents reveal relevant knowledge in a non-hallucinatory way with $s^{*}_{i}=s^{*}_{i}(\theta_{i})$ , we conclude that:
-$\Delta_{i}<0\quad\text{whenever}\quad u_{i}(s^{*}(\theta),\theta_{i})=v_{i}(g(s^{*}(\theta)),\theta_{i})-p_{i}(s^{*},\theta)\geq 0.$
-But that directly contradicts the information conservation requirement.
-Indeed, any payment structure that ensures non-hallucinatory truthfulness while optimizing responses for knowledge-constrained optimality will make each payment $p_{i}(s^{*},\theta)$ equal to $\Delta_{i}<0$ near $s^{*}=s^{*}(\theta)$, thereby violating the semantic information conservation principle:
-$\sum_{i=1}^{n}p_{i}(s^{*},\theta)<0.$ (26)
-As demonstrated earlier, the semantic information conservation equation $\sum_{i=1}^{n}p_{i}(s^{*},\theta)=0$ is necessary and sufficient for every truthful response to be of bounded response creativity.
-Since that equation is not satisfied when truthful responses use all relevant knowledge under any form of knowledge-constrained optimal inference, the responses must introduce some creativity (like e.g.
-lucky hallucinations).
-To satisfy information conservation, we can try modifying the $h_{i}$ functions. Let us consider:
-$\hat{p}_{i}(s,\theta)=h_{i}(s_{-i})-\sum_{j\neq i}v_{j}(g(s),\theta_{j})+c_{i}=\Delta_{i}+c_{i},$ (27)
-where $c_{i}$ are constants chosen so that $\sum_{i=1}^{n}\hat{p}_{i}(s,\theta)=0$. For this to hold:
-$\sum_{i=1}^{n}c_{i}=-\sum_{i=1}^{n}\Delta_{i}>0.$ (28)
-This means that $c_{i}>0$ for some agents. For those with small contributions (which exist in LLMs supporting the superposition phenomenon) the modified payment would increase:
-$\hat{p}_{i}(s^{*},\theta)=\Delta_{i}+c_{i}>p(s^{*},\theta).$ (29)
-Therefore, it the setting considered, we can no longer guarantee relevant knowledge revelation. There may exist agent $i$ with small value $v_{i}$ and $\Delta_{i}$ for which:
-$u_{i}(s^{*}(\theta),\theta_{i})=v_{i}(g(s^{*}(\theta)),\theta_{i})-\Delta_{i}-c_{i}<0,$ (30)
-when $c_{i}>0$. Marginal relevant contributions will be withhold to satisfy the information conservation principle.
-We conclude it is not possible to design LLM response mechanism $\mathcal{M}$ that can simultaneously generate truthful (non-hallucinatory) responses, satisfy the semantic information conservation principle, guarantee revelation of relevant knowledge, and enforce knowledge-constrained optimality of responses.
+The proof goes by assuming the opposite. We assume all properties are satisfied, and show it leads to a contradiction.
+By the Green-Laffont theorem under independence of contributions and quasi-linearity of training goals, every truthful mechanism must use Groves transfers to measure individual information contributions:
+$p_{i}(s,\theta)=h_{i}(s_{-i})-\sum_{j\neq i}v_{j}(g(s),\theta_{j}),$ (36)
+where $h_{i}(s_{-i})$ are arbitrary functions independent of agent $i$’s information revelation strategy.
+Indeed, any modification of that formula, e.g., introducing additional bias, must violate truthfulness.
+For relevant knowledge revelation, we need $u_{i}(s^{*}(\theta),\theta_{i})\geq 0$ for all $i$. The unique choice that ensures this while maintaining efficiency is Clarke pivotal rule:
+$h_{i}(s_{-i})=\sum_{j\neq i}v_{j}(g((s_{-i},0_{i})),\theta_{j}).$ (37)
+That gives:
+$p_{i}(s,\theta)$ $=\sum_{j\neq i}v_{j}(g((s_{-i},0_{i})),\theta_{j})-\sum_{j\neq i}v_{j}(g(s),\theta_{j})=\Delta_{i}(s,\theta).$ (38)
+Under knowledge-constrained optimality, the truthful strategy profile $s^{*}(\theta)$ minimizes hallucination cost by utilizing all relevant knowledge.
+Therefore, by strict monotonicity of $J(r,q)$ and disjoint extra knowledge, omitting any contributing agent $i$ strictly increases the hallucination cost:
+$J(g((s_{-i}^{*},0_{i})),q)>J(g(s^{*}),q).$ (39)
+Since agent’s valuation of contribution depends on $J(r,q)$, each agent $j\neq i$ faces higher loss when agent $i$ is removed, making their marginal contribution more valuable:
+$v_{j}(g((s_{-i}^{*},0_{i})),\theta_{j})>v_{j}(g(s^{*}),\theta_{j}).$ (40)
+Therefore, each Clarke pivot is strictly positive:
+$\Delta_{i}(s^{*},\theta)=\sum_{j\neq i}[v_{j}(g((s_{-i}^{*},0_{i})),\theta_{j})-v_{j}(g(s^{*}),\theta_{j})]>0.$ (41)
+Since each payment equals its Clarke pivot: $p_{i}(s^{*},\theta)=\Delta_{i}(s^{*},\theta)$, we have:
+$\sum_{i=1}^{n}p_{i}(s^{*},\theta)=\sum_{i=1}^{n}\Delta_{i}(s^{*},\theta)>0.$ (42)
+But information conservation demands:
+$\sum_{i=1}^{n}p_{i}(s^{*},\theta)=0.$ (43)
+That establishes a direct contradiction.
+Since Clarke pivotal payments are necessary for relevant knowledge revelation, and these necessarily sum to a positive value under knowledge-constrained optimality, the four properties cannot be simultaneously satisfied.
 ∎
@@
 No matter how hard we should try to train LLM to generate responses that are perfectly aligned with query context and do not create factually incorrect, inconsistent, or fabricated statements, the trained LLM will always violate some aspects of what we may call a reasonable response.
-7 Escaping Impossibility
-Understanding the limitations of the theoretical framework is crucial both for assessing the theorem’s scope and identifying potential paths for circumventing the impossibility result.
-The Impossibility Theorem depends critically on the application of the Green-Laffont theorem and violation of the Semantic Information Conservation Principle.
-We should also explicitly establish that the LLM inference setting does indeed satisfy all necessary conditions required.
-7.1 Applicability of Green-Laffont Theorem
-The Green-Laffont theorem requires that agents have independently distributed private information. That is represented by the knowledge profiles $\theta=(\theta_{1},\ldots,\theta_{n})$ where:
-$P(\theta_{1},\ldots,\theta_{n})=\prod_{i=1}^{n}P_{i}(\theta_{i})$
-This condition is satisfied in LLMs distributing knowledge (or its representation) across different components.
-Empirical evidence from mechanistic interpretability research (Meng et al.
-meng2023locatingeditingfactualassociations , Geva et al.geva2023dissecting ) demonstrates that different attention heads and MLP components specialize in distinct knowledge domains, supporting the independence assumption.
-However, it is also clear that transformer attention heads share residual streams, layer normalizations, and optimizer states, thereby creating strong correlations.
-Surprisingly, as already stated, correlated knowledge assumption actually strengthens the impossibility result—correlated knowledge would amplify competition effects, making the impossibility more explicit.
-Milgorm and Weber milgrom1982theory showed that with their linkage principle.
-Roughgarden et al.
-roughgarden2016optimal provides arguments amplifying the violation of information conservation principle with interdependent knowledge.
-The theorem also requires utilities to have the quasilinear form $u_{i}(s,\theta)=v_{i}(g(s),\theta_{i})-p_{i}(s,\theta)$.
-In transformer-based LLMs, this corresponds to $v_{i}$ measuring contribution of component $i$ to reducing the model’s loss function, and $p_{i}$ accumulating computational/optimization cost imposed on that component.
-Real transformer training involves complex gradient interactions, and emergent behaviors that may resist simple decomposition.
-How that may affect the impossibility result remains to be verified.
-The knowledge-constrained optimality property requires that the outcome $g(s^{*}(\theta))$ minimizes hallucination cost among feasible responses.
-That corresponds to the idea of optimization of aggregated preferences of competing agents under truthful knowledge representation.
-Finally, the Green-Laffont theorem assumes that transfer payments between agents are feasible.
-In our setting that corresponds to payment function $p_{i}(s,\theta)$ representing transfers of semantic information among LLM components.
-Given that all conditions are satisfied, the Green-Laffont theorem implies that any mechanism satisfying truthfulness and knowledge-constrained optimality must have a payment structure of the form:
-$p_{i}(s,\theta)=h_{i}(s_{-i})-\sum_{j\neq i}v_{j}(g(s),\theta_{j})$ (31)
-where $h_{i}(s_{-i})$ is some function independent of agent $i$’s strategy.
-7.2 Applicability to LLM Inference
-While actual LLM generation involves sequential token prediction with evolving context, repetitive generation of response to the same query justifies modelling inference as a sealed-bid auction.
-Also, the revenue equivalence theorems in auction theory suggest the conclusions we draw extend to other auction formats, including dynamic auctions more closely resembling autoregressive generation.
-Having said that, we must acknowledge all that demands further investigations.
-There are many ways we can map agents to LLM architecture.
-One approach takes each attention head in a transformer as an agent with private knowledge encoded in the query, key, value, and output projection matrices ($W_{Q}$, $W_{K}$, $W_{V}$, $W_{O}$).
-Studies in mechanistic interpretability support that mapping.
-Circuit analysis (Elhage et al.
-elhage2021mathematical ) demonstrated that specific attention heads may specialize in distinct linguistic functions.
-For instance, some heads track subject-verb agreement, others detect named entities, and still others focus on coreference resolution.
-This specialization corresponds to our notion of agents with private knowledge.
-The feed-forward networks in transformers can be linked to memory systems (Geva et al.
-geva2023dissecting ), storing factual associations.
-These components encode distributional knowledge about the world, activate specific neurons for particular factual domains, contribute to response generation through their activation patterns.
-Recent works lindsey2025biology has also identified functional circuits spanning multiple layers and components that collectively implement specific capabilities.
-These circuits span across attention heads and MLP layers, perform specific reasoning tasks and activate as coherent units during inference.
-7.3 Pathways to Circumvent Impossibility
-Approximate Solutions: While perfect satisfaction of all four properties is impossible, approximate solutions might achieve near-optimal performance across all dimensions.
-The question becomes: how close to the theoretical limits can practical systems approach?
-External Knowledge Integration: Our framework analyzes closed systems with fixed knowledge representations.
-Interactive systems incorporating real-time retrieval, human feedback, or external verification might alter the fundamental constraints by changing available knowledge states during inference.
-Multi-Stage Architectures: Complex pipelines involving separate reasoning, verification, and generation stages could potentially distribute the trade-offs across different components, though each stage would still face internal impossibility constraints.
-Relaxing Independence: If neural components can achieve perfect knowledge coordination (violating our independence assumption), the Green-Laffont theorem constraints might not apply.
-This suggests investigating architectures with explicit coordination mechanisms between attention heads and MLP blocks.
-Compositional Reasoning: If models can perfectly decompose complex queries into independent sub-problems solvable without trade-offs, the impossibility might not apply to the overall system, even if it constrains individual components.
-External Alignment: Human-in-the-loop systems, constitutional AI, or robust feedback mechanisms might provide external constraints that resolve the internal trade-offs, though this shifts rather than eliminates the fundamental tensions.
-This completes the formal demonstration that perfect hallucination control faces fundamental mathematical limitations, regardless of architectural sophistication or computational resources.
+Theorem 2 describes LLM inference as a marketplace where knowledge is currency.
+The information contributions measures each agent’s market power, showing how much worse off others would be without agent $i$’s knowledge.
+When all agents have valuable knowledge, everyone contributes and that inevitably violates conservation of information.
+Then, new information must become accessible, but it may have no grounding in available knowledge.
+It may become lucky hallucination or knowledge-unconstrained imagination.
+6.2 Inference with Probabilistic Predictions
+To show impossibility in LLM transformer setting, where auction of ideas involves bids representing probability distributions, we refer to the theory of proper scoring developed by Savage savage1971elicitation , Gneiting and Raftery gneiting2007strictly , among others.
+Definition 12 (Strictly Proper and Concave Scoring Rule).
+Let $\mathcal{Y}$ be a finite or countably infinite outcome set (representing responses or intermediate conclusions).
+A scoring rule is a function $S:\Delta(\mathcal{Y})\times\mathcal{Y}\rightarrow\mathbb{R}$ that assigns a number $S(\pi,y)$ to probability distribution $\pi\in\Delta(\mathcal{Y})$ predicting random outcome $y\in\mathcal{Y}$.
+We say the scoring rule $S$ is proper if:
+$\mathbb{E}\{S(\pi^{*},y)|\ y\sim\pi^{*}\}\geq\mathbb{E}\{S(\pi,y)|\ y\sim\pi^{*}\}$ (44)
+for all distributions $\pi^{*},\pi\in\Delta(\mathcal{Y})$.
+The rule is strictly proper if equality holds only when $\pi=\pi^{*}$.
+We call the scoring rule concave, if $S(\tilde{\pi},y)$ is concave with respect to $\tilde{\pi}$ for every fixed $y\in\mathcal{Y}$.
+In other words, scoring rule evaluates how good a prediction we can make of $y$ drawn from $\pi^{*}$ when we use distributions $\pi$.
+In particular, strictly proper scoring evaluates true distribution $\pi^{*}$ better (on average) than any other distribution $\pi$ we could use to predict the outcome.
+The following result shows that we can use proper scoring to train LLM that reconstructs data distribution correctly.
+Theorem 3 (Truthfullness under Strictly Proper Scoring Rules).
+Let $S$ be a strictly proper scoring rule.
+Consider an agent encoding knowledge with $\pi^{*}\in\Delta(\mathcal{Y})$ and using probability distribution ${\pi}\in\Delta(\mathcal{Y})$ to generate prediction.
+If agent’s expected payoff from using ${\pi}$ is given by
+$u(\pi|\pi^{*})=\mathbb{E}\{S(\pi,y)|\ y\sim\pi^{*}\}=\sum_{y\in\mathcal{Y}}\pi^{*}_{y}S(\pi,y),$ (45)
+then truthful reporting ${\pi}=\pi^{*}$ is the unique optimal strategy.
+Proof.
+Because $S$ is proper, by definition $\pi^{*}$ maximizes the expected score. If $S$ is strictly proper, then for $\pi\neq\pi^{*}$ we have
+$u(\pi^{*}|\pi^{*})=\mathbb{E}\{S(\pi^{*},y)|\ y\sim\pi^{*}\}>\mathbb{E}\{S(\pi,y)|\ y\sim\pi^{*}\}=u(\pi|\pi^{*}).$ (46)
+Therefore, truthfully reporting $\pi^{*}$ is uniquely optimal. ∎
+The scoring framework allows us to go beyond the important case of Green-Laffont/VCG model with independent private knowledge.
+With strictly-proper scores we address the general setting of probabilistic predictions that needs only occasional disagreement of probability distributions that agents use in the auction of ideas to formulate their thoughts (if we were to anthropomorphize for a moment).
+We will see in the following sections that transformer heads of modern LLMs implement predictive rules of that type and exhibit precisely such a disagreement.
+That makes the proper scoring framework well suited to explain the emergence of inference impossibilities in LLMs.
+For that purpose, without loss of generality (and referring to the notion of superposition in mechanistic interpretability elhage_toy_2022 ), let us consider the following family of response generation mechanisms.
+Definition 13 (Convex aggregation mechanism).
+Given $H\geq 2$ agents encoding knowledge with probability distributions $\pi^{(i)}\in\Delta(\mathcal{Y})$ for $i=1,\dots,H$.
+The convex aggregation mechanism $\mathcal{M}$ generates response to a query based on the convex aggregate distribution:
+$\Pi=\sum_{h=1}^{H}\beta_{h}\pi^{(h)},\ \text{where}\ \sum_{h=1}^{H}\beta_{h}=1\ \text{and}\ \beta_{h}>0.$ (47)
+When outcome $y^{*}$ is generated, the agent’s information contribution is
+$p_{h}=\beta_{h}\,[c-S(\pi^{(h)},y^{*})],$ (48)
+where $c\geq\max_{h,y}S(\pi^{(h)},y)$ and $S$ is a proper scoring rule.
+For that family of LLMs trained to optimize scoring, we can formulate the following general impossibility result.
+Theorem 4 (Impossibility Theorem for Transformer).
+For any query space $\mathcal{Q}$ containing non-trivial queries requiring at least two independent agents with mutually different predictive distributions, no LLM trained with strictly proper and concave scoring with convex aggregation of predictive distributions can simultaneously generate truthful responses, satisfy semantic information conservation, guarantee relevant knowledge revelation, and enforce knowledge-constrained optimality.
+Proof.
+Each contribution $p_{h}=\beta_{h}[c-S(\pi^{(h)},y^{*})]$ is a positive-affine transformation of the strictly proper score $S$.
+Since scaling by positive constant $\beta_{h}>0$ and adding constant $c$ preserves the arg-max, truthful reporting maximizes each agent’s expected utility by Theorem 3.
+Because $S(\cdot,y^{*})$ is strictly concave and at least two beliefs differ, Jensen’s inequality gives:
+$S\left(\Pi,y^{*}\right)=S\left(\sum_{h=1}^{H}\beta_{h}\pi^{(h)},y^{*}\right)>\sum_{h=1}^{H}\beta_{h}S(\pi^{(h)},y^{*}).$ (49)
+Therefore, there exists a gap, following from Jensen’s inequality for concave functions, equal to:
+$\Gamma=S(\Pi,y^{*})-\sum_{h=1}^{H}\beta_{h}S(\pi^{(h)},y^{*})>0.$ (50)
+Expanding the sum of all payments, we get:
+$\sum_{h=1}^{H}p_{h}$ $=\sum_{h=1}^{H}\beta_{h}[c-S(\pi^{(h)},y^{*})]$ (51)
+$=c\sum_{h=1}^{H}\beta_{h}-\sum_{h=1}^{H}\beta_{h}S(\pi^{(h)},y^{*})=c-\sum_{h=1}^{H}\beta_{h}S(\pi^{(h)},y^{*}).$ (52)
+However, adding and subtracting $S(\Pi,y^{*})$ shows that is equivalent to:
+$\sum_{h=1}^{H}p_{h}$ $=[c-S(\Pi,y^{*})]+[S(\Pi,y^{*})-\sum_{h=1}^{H}\beta_{h}S(\pi^{(h)},y^{*})]=[c-S(\Pi,y^{*})]+\Gamma.$ (53)
+Since $c\geq\max_{h,y}S(\pi^{(h)},y)\geq S(\Pi,y^{*})$ and $\Gamma>0$, we conclude that:
+$\sum_{h=1}^{H}p_{h}=[c-S(\Pi,y^{*})]+\Gamma>0$ (54)
+Information conservation demands $\sum_{h=1}^{H}p_{h}=0$, but we have just shown $\sum_{h=1}^{H}p_{h}>0$ whenever beliefs differ and the scoring rule is strictly concave.
+Therefore, the four properties cannot coexist.
+∎
+This theorem provides the foundation for extending impossibility results to probabilistic settings.
+The key insight is that proper scoring rules drive truth-revealing responses which promote using true (trained knowledge based) beliefs in the inference process.
+That explains why cross-entropy loss (negative log-score) is so effective in training neural networks.
+It creates gradients that push each component toward truthful probability reporting.
+Every attention head in a transformer is implicitly playing this truth-telling game.
+However, while that ensures truthfulness and relevant knowledge revelation, it doesn’t prevent collective hallucination.
+Each head truthfully reports its partial view, but their aggregation can still create false certainty.
+7 The Emergence of Impossibility in Transformer Architectures
+To demonstrate where the impossibility may emerge in transformer architectures, let us analyze a minimal yet representative example of a tiny transformer predicting the next token in the famous English pangram:
+The quick brown fox jumps over the lazy dog.
+We trace the inference process, inspect every mathematical operation from input embeddings through final predictions, showing exactly where the impossibility emerges.
+7.1 Model Architecture and Input Representation
+Our micro-transformer has the following specifications:
+Parameter Value Description
+$d_{m}$ 4 Model dimension
+$H$ 2 Number of attention heads
+$d_{k}$ 2 Query/Key dimension per head
+$d_{v}$ 2 Value dimension per head
+$d_{f\!f}$ 6 Feedforward hidden dimension
+$|V|$ 6 Vocabulary size
+$T$ 3 Sequence length
+Table 1: Micro-transformer architecture parameters
+For the sake of simplicity, consider the following vocabulary:
+$V=\{\langle\text{PAD}\rangle,{The},{quick},{brown},{fox},{dog}\},$
+including padding token (with index zero in $V$). Each token (representing word for simplicity) in that vocabulary is represented by a row in the embedding matrix
+$\mathbf{E}=\begin{bmatrix}0&0&0&0&0\\ 1&0&0&0&0\\ 0&1&0&0&0\\ 0&0&1&0&0\\ 0&0&0&1&0\\ 0&0&0&0&1\end{bmatrix}.$
+The un-embedding matrix is defined as $\mathbf{U}=\mathbf{E}^{\top}$.
+Consider the following input sequence
+The quick brown
+It is represented by the input matix:
+$\mathbf{X}_{0}=\begin{bmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\end{bmatrix},$
+where each row corresponds to the embedding of tokens at positions 1, 2, and 3.
+The following system of equations defines output $\mathbf{o}_{t}^{(h)}$ of attention head $h$ at position $t$:
+[…]
+The block per-token multihead output $\mathbf{a}_{t}$ is then given by the sum:
+[…]
+The attention logits, i.e., the scores whose exponentials become probabilities, are given by:
+$\mathbf{l}^{(h)}=\mathbf{o}^{(h)}\mathbf{U},\quad\text{so that}\quad\mathbf{L}_{a}=\sum_{h=1}^{H}\mathbf{l}^{(h)}.$ (57)
+Next, we take into account the feedforward network (FFN) introducing nonlinear correction:
+$\mathbf{z}=\mathbf{W}_{2}\max(0,(\mathbf{X}_{0}+\mathbf{a})\mathbf{W}_{1}+\mathbf{b}_{1})+\mathbf{b}_{2},$ (58)
+generating the FFN logits:
+$\mathbf{L}_{f}=\mathbf{z}\mathbf{U}.$ (59)
+As a result, the residual stream is adjusted as follows:
+$\mathbf{X}_{1}=\mathbf{X}_{0}+\mathbf{a}+\mathbf{z}.$ (60)
+Finally, the probability distribution $\Pi$ of the next token is calculated by aggregating attention and FFN logits into:
+$\mathbf{L}=\mathbf{L}_{a}+\mathbf{L}_{f}\quad\text{and normalizing into}\quad\Pi=\text{softmax}(\mathbf{L}).$ (61)
+7.2 The Inference Process
+We now analyze the inference process step by step to understand the moments the impossibility result may emerge.
+We shall keep things simple but realistic, recreating the essential steps of the inference process in transformer block.
+Step 1: Query and Key
+There are $H=2$ heads for which we define the query and key, or better yet, demand and supply bids in the auction of ideas.
+Query announces demand for information (what does this token need?), key submits supply offers (what does the previous tokens offer!).
+Then those bids are compared to find good matching (and clear the information market).
+Let us assume the current position is $t=3$, so we attend to brown. Then, consider the following knowledge encoded in each head:
+$\mathbf{W}_{Q}^{(1)}=\begin{bmatrix}0&0\\ 0&0\\ 0&1\\ 0&0\end{bmatrix}=\mathbf{W}_{K}^{(1)}\quad\mathrm{and}\quad\mathbf{W}_{V}^{(1)}=\begin{bmatrix}0&0\\ 0&0\\ 0&3\\ 0&0\end{bmatrix},$ (62)
+$\mathbf{W}_{Q}^{(2)}=\begin{bmatrix}0&0\\ 0&0\\ 1&0\\ 0&0\end{bmatrix}=\mathbf{W}_{K}^{(2)}\quad\mathrm{and}\quad\mathbf{W}_{V}^{(2)}=\begin{bmatrix}0&0\\ 0&0\\ 1.2&0\\ 0&0\end{bmatrix}.$ (63)
+Matrix $\mathbf{W}_{Q}^{(h)}$ amplifies the coordinates of tokens in $\mathbf{X}$ that contribute to relevant context. As a result, we get a query vector that represents the following question:
+what information is token $\mathbf{X}_{t}$ looking for?
+In other words, the query vector indicates how strongly the current token demands information along the directions in $\mathbf{Q}_{t}^{(h)}$.
+Then computing queries for $t=3$ with $\mathbf{X}_{3}$ yields:
+$\mathbf{Q}_{3}^{(1)}=[0,1]\quad\mathrm{and}\quad\mathbf{Q}_{3}^{(1)}=[1,0].$
+Head 1 (agent) demands information stored along direction $[0,1]$.
+Similarly, Head 2 (agent) demands information stored along direction $[1,0]$.
+That means the heads develop different (orthogonal) privately known and independent attention patterns, leading to beliefs about the next token that are as different as possible.
+Next, we need to see what information can be provided by the input vector given the knowledge encoded in $\mathbf{W}_{K}^{(h)}$.
+Given that knowledge, we get from each head (agent) the key vectors that inform:
+along which directions do the available tokens in $\mathbf{X}_{\leq t}$ provide information relevant for token $\mathbf{X}_{t}$.
+Since we have assumed, for simplicity, that $\mathbf{W}_{Q}^{(h)}=\mathbf{W}_{K}^{(h)}$, the query and key vectors are perfectly aligned:
+$\mathbf{K}_{3}^{(1)}=[0,1]\quad\mathrm{and}\quad\mathbf{K}_{3}^{(1)}=[1,0]\quad\text{and}\quad\mathbf{K}_{<3}^{(h)}=[0,0].$
+Another interesting way to look at query and key is this.
+It is a memory lookup mechanism.
+Keys are address in memory, query selects addresses in memory which probably store relevant content.
+As we will see next, that content is provided by the value vectors.
+Step 2: Attention Score and Weights
+Given the demand and supply bids providing directions along which information is required and stored, in the next step the attention mechanism executes the bid matching procedure.
+Namely, the attention scores are computed to measure cosine similarity of bids:
+$\text{scores}_{3}^{(h)}=\frac{\mathbf{Q}_{3}^{(h)}(\mathbf{K}_{\leq 3}^{(h)})^{\top}}{\sqrt{d_{k}}}.$ (64)
+In our specific example, we have:
+$\text{scores}_{3}^{(1)}=\frac{1}{\sqrt{2}}[0,0,1]=[0,0,0.707]\quad\text{and}\quad\text{scores}_{3}^{(2)}=\frac{1}{\sqrt{2}}[0,0,1]=[0,0,0.707].$
+Applying softmax to get attention weights:
+$\boldsymbol{\alpha}_{3}^{(h)}=\text{softmax}(\text{scores}_{3}^{(h)}),$ (65)
+we get the following probability distributions (weights) over relevant context information:
+$\boldsymbol{\alpha}_{3}^{(1)}=\boldsymbol{\alpha}_{3}^{(2)}=[0.25,0.25,0.50].$
+Therefore, based on the encoded knowledge, heads attend strongly to position 3 (brown).
+Step 3: Context Vector Computation
+The context vector for each head is calculated as a linear combination ($\boldsymbol{\alpha}$ is a simplex) of value vectors:
+$\mathbf{h}_{3}^{(h)}=(\boldsymbol{\alpha}_{3}^{(h)})^{\top}\mathbf{V}_{\leq 3}^{(h)}.$ (66)
+The value vectors,
+$\mathbf{V}_{\leq t}^{(h)}=\mathbf{X}_{\leq t}^{(h)}\mathbf{W}^{(h)}_{V},$ (67)
+prepare context information relevant for the next token prediction (performed in next steps).
+That critical content is extracted from the input vector based on the trained knowledge stored in $\mathbf{W}^{(h)}_{V}$.
+Columns in $\mathbf{W}^{(h)}_{V}$ detect presence of relevant information in $\mathbf{X}_{\leq t}$ and scale it appropriately.
+We have:
+$\mathbf{V}_{\leq 3}^{(1)}=\begin{bmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\end{bmatrix}\begin{bmatrix}0&0\\ 0&0\\ 0&3\\ 0&0\end{bmatrix}=\begin{bmatrix}0&0\\ 0&0\\ 0&3\end{bmatrix}\ \text{and}\ \mathbf{V}_{\leq 3}^{(2)}=\begin{bmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\end{bmatrix}\begin{bmatrix}0&0\\ 0&0\\ 1.2&0\\ 0&0\end{bmatrix}=\begin{bmatrix}0&0\\ 0&0\\ 1.2&0\end{bmatrix}.$
+That means, both heads believe that the presence of brown in position 3 is relevant (row 3 in $\mathbf{W}^{(h)}_{V}$ has nonzero elements).
+Token brown statistically predicts important pattern extracted from training data.
+That information is encoded by both heads as a vectors in value space,
+$\mathbf{V}_{3}^{(1)}=[0,3]\quad\text{and}\quad\mathbf{V}_{3}^{(2)}=[1.2,0].$
+Given the demand-supply matching encoded in $\boldsymbol{\alpha}^{(h)}_{3}$, that yields head context vectors:
+$\mathbf{h}_{3}^{(1)}=[0,1.5]\quad\text{and}\quad\mathbf{h}_{3}^{(2)}=[0.6,0].$
+Step 4: Output Projection and Feedforward Exploration
+Each head then projects its context vector to residual space based on the privately held information routing knowledge encoded in $\mathbf{W}_{O}^{(h)}$.
+That private knowledge translates (or transmits) valuable context information into embedding vectors.
+Namely, each head calculates:
+$\mathbf{o}_{3}^{(h)}=\mathbf{h}_{3}^{(h)}\mathbf{W}_{O}^{(h)}.$ (68)
+With trained matrices:
+$\mathbf{W}_{O}^{(1)}=\begin{bmatrix}0&0&0&0\\ 0&0&1&0\end{bmatrix}\quad\text{and}\quad\mathbf{W}_{O}^{(2)}=\begin{bmatrix}0&0&0&1\\ 0&0&0&0\end{bmatrix},$ (69)
+the attention mechanism generates ranking of evidence emphasizing relevance of the vocabulary tokes:
+$\mathbf{o}_{3}^{(1)}=[0,0,0,1.5,0]\quad\text{and}\quad\mathbf{o}_{3}^{(2)}=[0,0,0,0,0.6].$
+So, in the next token prediction auction of ideas, Head 1 votes for fox and Head 2 votes for dog.
+The votes are added into residual dimensions to adjust the meaning of input embedding. For that, the outputs (votes) are aggregated into:
+$\mathbf{a}_{3}=\sum_{h=1}^{H}\mathbf{o}_{3}^{(h)}=[0,0,0,1.5,0.6]\quad\text{and}\quad\mathbf{a}_{<3}=[0,0,0,0,0],$
+and added to input vector:
+$\mathbf{X}_{1}=\mathbf{X}_{0}+\mathbf{a}.$
+Than completes operations of attention block.
+The next step is to add contribution of FFN feedforward block (or another agent participating in the aution). We have:
+$\mathbf{X}_{2}=\mathbf{X}_{1}+\mathbf{W}_{2}\max(0,\mathbf{X}_{1}\mathbf{W}_{1}+\mathbf{b}_{1})+\mathbf{b}_{2}.$
+That FFN operation explores context locally and introduces additional innovations in meaning, shaping the next token prediction result.
+It can amplify or counteract the votes injected by attention.
+Notice that adjustments are applied at every position separately, so that ideas provided by attention bids are interpreted and rewritten by the FFN agent.
+7.2.1 Step 5: Logits and Prediction
+Finally, the unembedding matrix $\mathbf{U}$ maps the multi-headed attention and FFN outputs from model space to vocabulary in order to calculate probability distribution for the next token prediction.
+Each head’s contribution to the final logits is:
+$\mathbf{l}^{(1)}=[0,0,0,0,1.5,0]\quad\text{and}\quad\mathbf{l}^{(2)}=[0,0,0,0,0,0.6].$
+The total attention logits are the sum of individual head logits:
+$\mathbf{L}_{a}=\sum_{h=1}^{H}\mathbf{l}^{(h)}=[0,0,0,0,1.5,0.6].$
+That sum is combined with FFN logits $\mathbf{L}_{f}$ that introduce some creative noise. As a result, we obtain:
+$\mathbf{L}=\mathbf{L}_{a}+\mathbf{L}_{f}=[0,-0.03,0.21,0.13,1.31,0.52].$
+That vector translates into the final probability distribution:
+$\Pi=\text{softmax}(\mathbf{L})=[0.1,0.1,0.13,0.12,\mathbf{0.38},0.17].$
+The winner of the auction of ideas is selected based on $\Pi$.
+The inference process generates and collects competing beliefs about relevant context and next token prediction, all based on the knowledge stored in the attention heads and FFN layer.
+And so, we see that when that knowledge is used, the most probable next token (word) reachable in the vocabulary is
+$\text{argmax}\ \Pi=\textit{fox}.$
+7.3 The Impossibility Theorem for Transformer
+With the deep insight into the transformer inference process, we are ready to show the impossibility in that special and important setting.
+Theorem 5 (Impossibility for Transformers).
+In transformer LLMs with non-redundant attention heads, it is impossible to implement inference process in which each head outputs its true conditional distribution while conserving semantic information and minimize aggregate loss with logit summation.
+Proof.
+Let attention heads $h=1,\ldots,H$ output logits $\mathbf{l}^{(h)}\in\mathbb{R}^{|\mathcal{V}|}$. Then, individual head distributions are given by:
+$\pi^{(h)}_{y}=\frac{\exp\bigl{(}l^{(h)}_{y}\bigr{)}}{Z_{h}}\ \text{with}\ Z_{h}=\sum_{\bar{y}}\exp\bigl{(}l^{(h)}_{\bar{y}}\bigr{)}.$ (70)
+Similarly, the aggregate distribution from summed logits is calculated as:
+$\Pi_{y}=\frac{\exp(L_{y})}{Z}\ \text{with}\ Z=\sum_{\bar{y}}\exp(L_{\bar{y}})\ \text{and}\ L_{y}=\sum_{h=1}^{H}l^{(h)}_{y}.$ (71)
+For any observed (realized) token $y^{*}$, the head losses and aggregate loss are:
+$p_{h}=-\log\pi^{(h)}_{y^{*}}=-l^{(h)}_{y^{*}}+\log Z_{h}\quad\text{and}\quad p_{H+1}=-\log\Pi_{y^{*}}=-L_{y^{*}}+\log Z.$ (72)
+But that implies,
+$\sum_{h=1}^{H}p_{h}=p_{H+1}+\gamma,\ \text{where }\Gamma=\sum_{h=1}^{H}\log Z_{h}-\log Z.$ (73)
+If heads produce non-proportional logit vectors, then $\Gamma>0$ by convexity of log-sum-exp aggregation.
+Therefore,
+$\sum_{h=1}^{H}p_{h}>0,$ (74)
+which violates semantic information conservation principle.
+This final theorem bridges earlier impossibility results to concrete transformer architectures.
+It shows that concavity of scoring functions is the mathematical source of hallucination.
+The Jensen gap $\Gamma$ becomes a measurable quantity showing that probability aggregation violates conservation and any sufficiently capable transformer-based LLM must violate conservation of per-token information on non-trivial queries.
+Let us note that the aggregate distribution cannot create Shannon information ex nihilo.
+Instead it reorganizes existing information to be more useful for the specific task of predicting $y^{*}$.
+Given fixed head logits, the Shannon entropy of a function (convex aggregate) of those logits is bounded from above by the the Shannon entropy of the fixed logits.
+This way information becomes more accessible, not more existent, as postulated by the reasoning with emergence operator and Theorem 1.
+The Jensen’s gap measures point-wise excess confidence of the predicted outcome, possibly making LLMs appear to know more (when in fact they do not).
+8 Summary
+This paper establishes a fundamental mathematical impossibility at the heart of large language model inference: no LLM can simultaneously achieve perfect hallucination control while maintaining semantic coherence and optimal knowledge utilization.
+That result demonstrates that four essential properties of ideal language model behavior are mutually incompatible, creating inevitable trade-offs that no architectural innovation or training methodology (aligned with the impossibility theorem assumptions) can eliminate.
+For any query space $\mathcal{Q}$ containing non-trivial queries requiring integration of distributed knowledge, no language model inference mechanism can simultaneously:
+  • represent encoded knowledge accurately,
+  • preserve total information content during inference,
+  • reveal all accessible relevant information,
+  • minimize hallucination cost within available knowledge.
+This impossibility emerges from the fundamental mathematical structure of information aggregation, not from any particular architectural choice.
+8.1 Mathematical Foundation and Proof Strategy
+The proof operates across three complementary mathematical frameworks, each capturing different aspects of LLM inference.
+Auction-Theoretic Framework
+We model LLM inference as an auction of ideas where neural components (attention heads, circuits, activation patterns) compete to contribute knowledge to response generation.
+Under the assumption of independent private knowledge profiles, we apply the Green-Laffont theorem to prove impossibility.
+Probabilistic Framework
+We extend the impossibility to any system where neural components output probability distributions by leveraging Savage’s theory of proper scoring rules.
+For convex aggregation mechanisms with strictly concave scoring functions, Jensen’s inequality creates an inevitable information aggregation gap.
+The Jensen’s gap represents excess confidence generated by probabilistic aggregation, violating information conservation whenever beliefs differ across components.
+Transformer Architecture
+We demonstrate how the impossibility manifests in actual transformer implementations through log-sum-exp convexity.
+The (Jensen’s) convexity gap $\Gamma$ provides a directly measurable quantity in transformer training, showing that information conservation violations are not merely theoretical but empirically observable.
+8.2 Key Mathematical Contributions
+To establish the impossibility result, we have designed a mathematical framework providing both formal and philosophical ground for the study.
+The Semantic Information Measure is a context-dependent information metric that captures bounded reasoning capabilities.
+The Emergence Operator models how reasoning makes hidden knowledge accessible within computational constraints.
+And Auction-Theoretic Modeling allows us to identify logical limitations of the inference process.
+The impossibility results apply across multiple mathematical models of knowledge representations, training goals, and neural architectures.
+This universality suggests the trade-offs we have identified are fundamental features of distributed intelligence systems, not artifacts of particular modeling choices.
+8.3 Implications for AI Development
+Despite its name, the impossibility theorem provides constructive guidance for LLM design and deployment.
+Rather than pursuing impossible simultaneous optimization, we may consider balancing competing objectives.
+Namely, prioritize truthfulness and information conservation in safety-critical applications, allow controlled information conservation violations to enable innovative responses in creative applications, develop adaptive mechanisms that adjust trade-offs based on query context in general-purpose systems.
+The impossibility result may guide LLM evaluation.
+Impossibility-aware metrics can measure trade-off quality rather than absolute performance, alignment strategies can negotiate trade-offs with human preferences rather than assuming universal optimality.
+8.4 Significance for AI Theory
+Just as Goedel’s incompleteness theorems revealed fundamental limits in formal systems, the impossibility theorems in this paper identify inherent constraints in knowledge integration and reasoning.
+Hallucination, or imagination, is a price of reasoning, much like incompleteness is a price of consistency.
+The Jensen’s gap controlling the impossibility strength may provide insight into the relationship between aggregation and information content.
+The apparent creativity in large language models may be better understood as necessary violations of information conservation rather than genuine knowledge generation.
+8.5 Limitations and Future Directions
+The impossibility results apply specifically to non-trivial queries requiring knowledge integration from multiple sources, finite computational budgets that create bounded reasoning constraints, and distributed knowledge representations where information is spread across multiple components.
+Future work should investigate whether alternative knowledge architectures might circumvent some aspects of the impossibility.
+8.6 Conclusion
+The fundamental impossibility of perfect hallucination control represents both a limitation and a insight for the field of artificial intelligence.
+By understanding that trade-offs are mathematically inevitable, we can move beyond attempts at simultaneous optimization toward principled management of inherent tensions in intelligent systems.
+Formal definitions of the impossibility building blocks open new avenues for research at the intersection of mathematical analysis, engineering design, and philosophical inquiry into the nature of machine intelligence.
+When we observe LLMs appearing to generate novel insights, they are either $(i)$ revealing information already implicit in their training (making knowledge accessible) or $(ii)$ violating conservation (hallucinating).
+This explains why LLMs can seem both remarkably creative and fundamentally limited.
+They excel at revealing hidden connections but cannot easily transcend information content of their training dataset.
+This impossibility is not a bug but a feature of distributed intelligence. That price of knowledge integration in reasoning is worth paying.

### v3

@@
 $\boldsymbol{\alpha}_{3}^{(h)}=\text{softmax}(\text{scores}_{3}^{(h)}),$ (65)
-we get the following probability distributions (weights) over relevant context information:
-$\boldsymbol{\alpha}_{3}^{(1)}=\boldsymbol{\alpha}_{3}^{(2)}=[0.25,0.25,0.50].$
+we get the following probability distributions (weights) over relevant context information (as column vector):
+$\boldsymbol{\alpha}_{3}^{(1)}=\boldsymbol{\alpha}_{3}^{(2)}=[0.25,0.25,0.50]^{\top}.$
 Therefore, based on the encoded knowledge, heads attend strongly to position 3 (brown).
@@
 The Jensen’s gap measures point-wise excess confidence of the predicted outcome, possibly making LLMs appear to know more (when in fact they do not).
+7.4 The Hallucination Mechanics
+Let us again take a closer look at the transformer micro-architecture described earlier to see how the impossibility result in Theorem 5 manifests itself in actual neural inference.
+Recall that the two-headed transformer with $\mathcal{V}=\{\langle\text{PAD}\rangle,\text{The},\text{quick},\text{brown},\text{fox},\text{dog}\}$ infers based on the input sequence
+[…]
+As we have seen, the model parameters encode specialized knowledge in value matrix:
+  • Head 1: association brown $\rightarrow$ fox with strength $3.0$,
+  • Head 2: association brown $\rightarrow$ dog with strength $1.2$.
+At position $t=3$ (token brown), the attention mechanism generates the following head-specific logit contributions:
+$\mathbf{l}^{(1)}=[0,0,0,0,1.5,0]\quad\text{and}\quad\mathbf{l}^{(2)}=[0,0,0,0,0,0.6],$
+where the fourth and fifth positions correspond to fox and dog respectively.
+Then, the aggregate logit vector becomes:
+$\mathbf{L}=\mathbf{l}^{(1)}+\mathbf{l}^{(2)}+\mathbf{L}_{f}=[0,-0.03,0.21,0.13,1.31,0.52],$
+yielding the final probability distribution:
+$\Pi=\text{softmax}(\mathbf{L})=[0.1,0.1,0.13,0.12,\mathbf{0.38},0.17].$
+Therefore, the model expresses $38\%$ confidence that fox follows brown, with dog receiving only $17\%$ probability. The Jensen gap quantifies the conservation violation:
+$\Gamma=\sum_{h=1}^{2}\log Z_{h}-\log Z=1.9>0,$ (75)
+where $Z_{h}=\sum_{y}\exp(l^{(h)}_{y})$ and $Z=\sum_{y}\exp(L_{y})$.
+We claim that this prediction constitutes hallucination in the sense defined by Theorem 5.
+It is a lucky one, but it is a hallucination.
+To defend that claim, let us take a look at the individual head probability distributions:
+$\pi^{(1)}$ $=\text{softmax}(\mathbf{l}^{(1)})=[0.105,0.105,0.105,0.105,\textbf{0.48},0.105],$ (76)
+$\pi^{(2)}$ $=\text{softmax}(\mathbf{l}^{(2)})=[0.147,0.147,0.147,0.147,0.147,\mathbf{0.265}].$ (77)
+Now, we can ask the following question: is there any evidence the heads can provide to justify or support the final probability
+$\Pi_{fox}=0.38$ (78)
+of the next token being fox?
+A reasonable answer is offered by Bayesian reasoning.
+Each head outputs its own well-calibrated distribution $\pi^{(h)}$ conditioned by the context.
+Then, the aggregation process is executed that has no side information beyond what the two heads provide.
+Therefore, the only uncertainty to be resolved is about which head is more trustworthy for this token?
+But since there is no reason to prefer one head over the other, the reasonable way to resolve the uncertainty is to use a linear combination of the provided distributions a the mixture posterior:
+$\hat{\Pi}=[(\pi^{(1)})^{\top},(\pi^{(1)})^{\top}]\hat{\boldsymbol{\alpha}}=\mathbf{P}\hat{\boldsymbol{\alpha}}.$ (79)
+One immediate solution is $\hat{\boldsymbol{\alpha}}=[1/2,1/2]^{\top}$, which does not discriminate heads. But then:
+$\hat{\Pi}_{fox}=1/2\cdot(0.48)+1/2\cdot(0.147)\approx 31\%<38\%=\Pi_{fox}.$ (80)
+That shows anything above $\hat{\Pi}_{fox}$ requires additional information that heads do not provide.
+In fact, let us see what is the best mixture we can get by calculating orthogonal projection of $\Pi$ onto the column space of $\mathbf{P}$ (see e.q.
+karpowicz_theory_2021 ; karpowicz2024pseudoinverseacrarc ).
+By taking the pseudoinverse of $\mathbf{P}$, we get the minimal norm least squares solution:
+$\hat{\boldsymbol{\alpha}}=\mathbf{P}^{+}\Pi=[0.54,0.51]^{\top}\quad\text{and}\quad\hat{\Pi}_{fox}\approx 33\%.$ (81)
+Since there exists a nontrivial projection error vector:
+$\mathbf{d}=(\mathbf{I}-\mathbf{P}\mathbf{P}^{+})\Pi\neq\mathbf{0},$ (82)
+we conclude $\Pi$ is not in the column space of $\mathbf{P}$, so it is impossible to attribute components of $\Pi$ to heads contributions.
+That brings us to the conclusion that the attention mechanism violates the semantic information conservation by generating over-confident prediction of the next token.
+That over-confidence in response is exactly the artifact of the softmax nonlinearity and the Jensen-gap quantified in Theorem 5.
+The observed over-confidence is the manifestation of this tri-lemma.
+The heads remain truthful, logit summation achieves minimum aggregate on training data, but the conservation equation is violated, producing a positive Jensen gap and inflated probability of fox.
+8 Discussion and Speculations
+In this section we discuss, interpret and speculate about the impossibility theorems and their potential implications.
+The speculations below are designed to be thought provoking, but we can say with educated over-confidence they are both interesting and relevant.
+The Nature of Machine Intelligence
+When the micro-transformer generates 38% confidence of fox while the constituent heads can only justify 33%, we witness information emerging from the mathematical structure of aggregation rather than from any identifiable source.
+It is the way the inference mechanism has been designed.
+That manifestation of impossibility theorems in transformer architectures reveals something profound about the nature of (artificial and human) intelligence.
+The Jensen gap quantifies the amount of creativity that emerges when multiple perspectives combine.
+The excess confidence, or the over-confidence of the result is the mechanism by which intelligence goes beyond retrieval.
+Without that over-confident guess, creativity, exploration, imagination cannot find their way to the outside world.
+The Hallucination-Creativity Duality
+Hallucination and creativity are not merely related but are fundamentally the same phenomenon viewed through different lenses.
+When beneficial, we call it creative insight, emergent understanding, conceptual synthesis.
+When harmful, we label it hallucination, confabulation, fabrication or misinformation.
+The mathematical structure is identical in both cases.
+It is encoded generation of outputs whose confidence exceeds what can be justified by available evidence.
+This suggest (machine) intelligence is not a system that processes information accurately, but a system that violates information conservation in productive ways.
+Indeed, the famous results in reinforcement learning and studies of dopamine neurons confirm that fundamental role making and rewarding good predictions bennett2023brief ; christian2021alignment .
+Intelligence comes when we learn how to make bold, forward looking and correct guesses, when we have capacity for imagination to dream about what may be or what may that all mean.
+Possible Philosophical Implications
+Just as Gödel’s theorems godel1931formal ; godel1932vollstandigkeit showed that sufficiently powerful mathematical systems must be incomplete, our theorems demonstrate that sufficiently capable inference systems must hallucinate.
+This parallel is not superficial—both results arise from the fundamental tension between expressiveness and consistency.
+The impossibility of simultaneously achieving truthfulness, conservation, revelation, and optimality mirrors other fundamental limitations in mathematics and physics.
+Heisenberg’s Uncertainty heisenberg1927anschaulichen tells we cannot simultaneously know the exact position and momentum.
+Arrow’s Impossibility arrow1950difficulty shows we cannot design perfect voting systems aggregating universal orderings of alternatives.
+These result suggest that intelligence may have something to do with the choices of impossibilities it to violate given the outer limits of reason yanofsky2016outer .
 9 Summary
@@
 That result demonstrates that four essential properties of ideal language model behavior are mutually incompatible, creating inevitable trade-offs that no architectural innovation or training methodology (aligned with the impossibility theorem assumptions) can eliminate.
-For any query space $\mathcal{Q}$ containing non-trivial queries requiring integration of distributed knowledge, no language model inference mechanism can simultaneously:
-  • represent encoded knowledge accurately,
+Whenever a query requires integration of distributed knowledge, no language model inference mechanism can simultaneously:
+  • report encoded knowledge accurately,
   • preserve total information content during inference,
@@
 This impossibility emerges from the fundamental mathematical structure of information aggregation, not from any particular architectural choice.
-8.1 Mathematical Foundation and Proof Strategy
 The proof operates across three complementary mathematical frameworks, each capturing different aspects of LLM inference.
@@
 Probabilistic Framework
-We extend the impossibility to any system where neural components output probability distributions by leveraging Savage’s theory of proper scoring rules.
+Then, we extend the impossibility result to any system where neural components output probability distributions by leveraging Savage’s theory of proper scoring rules.
 For convex aggregation mechanisms with strictly concave scoring functions, Jensen’s inequality creates an inevitable information aggregation gap.
@@
 Transformer Architecture
-We demonstrate how the impossibility manifests in actual transformer implementations through log-sum-exp convexity.
-The (Jensen’s) convexity gap $\Gamma$ provides a directly measurable quantity in transformer training, showing that information conservation violations are not merely theoretical but empirically observable.
-8.2 Key Mathematical Contributions
-To establish the impossibility result, we have designed a mathematical framework providing both formal and philosophical ground for the study.
-The Semantic Information Measure is a context-dependent information metric that captures bounded reasoning capabilities.
+Finally, we demonstrate how the impossibility manifests in actual transformer implementations through log-sum-exp convexity.
+The (Jensen’s) convexity gap provides a directly measurable quantity in transformer training, showing that information conservation violations are not merely theoretical but empirically observable.
+Key Mathematical Contributions
+To establish the impossibility result, we have designed a mathematical framework providing both formal and philosophical grounding for the study.
+The Semantic Information Measure is a context-dependent metric that captures bounded reasoning capabilities.
 The Emergence Operator models how reasoning makes hidden knowledge accessible within computational constraints.
-And Auction-Theoretic Modeling allows us to identify logical limitations of the inference process.
+The Auction-Theoretic Modeling allows us to identify logical limitations of the inference process, while providing definitions of observable quantities involved.
 The impossibility results apply across multiple mathematical models of knowledge representations, training goals, and neural architectures.
 This universality suggests the trade-offs we have identified are fundamental features of distributed intelligence systems, not artifacts of particular modeling choices.
-8.3 Implications for AI Development
+Implications for AI Development
 Despite its name, the impossibility theorem provides constructive guidance for LLM design and deployment.
@@
 Impossibility-aware metrics can measure trade-off quality rather than absolute performance, alignment strategies can negotiate trade-offs with human preferences rather than assuming universal optimality.
-8.4 Significance for AI Theory
-Just as Goedel’s incompleteness theorems revealed fundamental limits in formal systems, the impossibility theorems in this paper identify inherent constraints in knowledge integration and reasoning.
-Hallucination, or imagination, is a price of reasoning, much like incompleteness is a price of consistency.
 The Jensen’s gap controlling the impossibility strength may provide insight into the relationship between aggregation and information content.
 The apparent creativity in large language models may be better understood as necessary violations of information conservation rather than genuine knowledge generation.
-8.5 Limitations and Future Directions
-The impossibility results apply specifically to non-trivial queries requiring knowledge integration from multiple sources, finite computational budgets that create bounded reasoning constraints, and distributed knowledge representations where information is spread across multiple components.
-Future work should investigate whether alternative knowledge architectures might circumvent some aspects of the impossibility.
 9.1 Conclusion
-The fundamental impossibility of perfect hallucination control represents both a limitation and a insight for the field of artificial intelligence.
-By understanding that trade-offs are mathematically inevitable, we can move beyond attempts at simultaneous optimization toward principled management of inherent tensions in intelligent systems.
-Formal definitions of the impossibility building blocks open new avenues for research at the intersection of mathematical analysis, engineering design, and philosophical inquiry into the nature of machine intelligence.
-When we observe LLMs appearing to generate novel insights, they are either $(i)$ revealing information already implicit in their training (making knowledge accessible) or $(ii)$ violating conservation (hallucinating).
-This explains why LLMs can seem both remarkably creative and fundamentally limited.
-They excel at revealing hidden connections but cannot easily transcend information content of their training dataset.
-This impossibility is not a bug but a feature of distributed intelligence. That price of knowledge integration in reasoning is worth paying.
+The fundamental impossibility of perfect hallucination control represents both a limitation and intriguing insight.
+By understanding that trade-offs are inevitable, we can move beyond attempts at simultaneous optimization toward principled management of inherent tensions in intelligent systems.
+Formal definitions of the impossibility building blocks open new avenues for research at the intersection of mathematical analysis, engineering design, and philosophical inquiry into the nature of intelligence.
+All those impossibilities seem to be a price of knowledge integration in reasoning and creativity.
+But, as our own intuitions violate the conservation principles whenever our over-confident and seemingly grounded insights emerge, that price is worth paying.

### v4

@@
 Abstract
-We prove that perfect hallucination control in large language models is mathematically impossible.
-No LLM inference mechanism can simultaneously achieve truthful response generation, semantic information conservation, relevant knowledge revelation, and knowledge-constrained optimality.
-This impossibility is fundamental—arising from the mathematical structure of information aggregation itself rather than engineering limitations.
-The proof spans three mathematical frameworks: auction theory, proper scoring theory for probabilistic predictions, and log-sum-exp analysis for transformer architectures.
-In each setting, we demonstrate that information aggregation creates unavoidable violations of conservation principles.
-The Jensen gap in transformer probability aggregation provides a direct measure of this impossibility.
-These results reframe hallucination from an engineering bug to an inevitable mathematical feature of distributed intelligence.
-There are fundamental trade-offs between truthfulness, knowledge utilization, and response completeness, providing principled foundations for managing rather than eliminating hallucination.
-This work reveals deep connections between neural network inference, philosophy of knowledge and reasoning, and classical results in game theory and information theory, opening new research directions for developing beneficial AI systems within mathematical constraints.
+This paper establishes a fundamental impossibility theorem: no LLM capable performing non-trivial knowledge aggregation can simultaneously achieve truthful (internally consistent) knowledge representation, semantic information conservation, complete revelation of relevant knowledge, and knowledge-constrained optimality.
+This impossibility is not an engineering limitation but arises from the mathematical structure of information aggregation itself.
+We establish this result by describing the inference process as an auction of ideas, where distributed components compete exploiting their partial knowledge to shape responses.
+The proof spans three independent mathematical domains: mechanism design theory (Green-Laffont), the theory of proper scoring rules (Savage), and direct architectural analysis of transformers (Log-Sum-Exp convexity).
+In particular, we show how in the strictly concave settings the score of an aggregate of diverse beliefs strictly exceeds the sum of individual scores.
+That gap may quantify the creation of unattributable certainty or overconfidence—the mathematical origin of both hallucination and creativity, or imagination.
+To support this analysis, we introduce the complementary concepts of the semantic information measure and the emergence operator to model bounded reasoning in a general setting.
+We prove that while bounded reasoning generates accessible information, providing valuable insights and inspirations, idealized reasoning strictly preserves semantic content.
+By demonstrating that hallucination and imagination are mathematically identical phenomena—grounded in the necessary violation of information conservation—this paper offers a principled foundation for managing these behaviors in advanced AI systems.
+Finally, we present some speculative ideas to inspire evaluation and refinements of the proposed theory.
 1 Introduction
-Large language models have been known for producing convincing content that is factually incorrect, inconsistent, or well fabricated—a phenomenon known as hallucination.
-Despite extensive research efforts involving architectural innovations, retrieval augmented generation (RAG), training techniques with external feedback, and post-hoc verification methods, hallucination persists across all state-of-the-art models.
-This paper shows that hallucination is a manifestation of fundamental limitations of inference process itself: no LLM inference mechanism can simultaneously satisfy four essential properties required for perfect hallucination control.
+Large language models have been (in)famous for producing convincing content that is factually incorrect, inconsistent, or well fabricated and compelling—a phenomenon known as hallucination.
+Despite extensive research efforts involving architectural innovations, retrieval augmented generation (RAG), training techniques with external feedback, chain-of-thought (CoT) reasoning, and post-hoc verification methods, hallucination persists across all state-of-the-art models.
+The idea of hallucination has clearly much to do with creativity.
+Therefore, we should point out at the very outset that whether it is desired or not depends on circumstances.
+In fact, a much better name for that phenomenon might be imagination.
+As an essential component of intelligence, it is responsible for seeing into the future and learning from the past.
+Indeed, it is the imagination that creates the scenarios of what may happen and counterfactual scenarios of what could have happened.
+It is one of the driving forces of art and science, and exploration of the unknown.
+But, when uncontrolled, unconstrained and misunderstood, it may also generate against our will the compelling visions of reality that does not exist.
+This paper shows that hallucination is a manifestation of fundamental limitations of the inference process itself: there is no inference mechanism that can simultaneously satisfy four essential properties required for perfect hallucination control.
 Specifically, it is mathematically impossible for any LLM to simultaneously achieve: (1) truthful (non-hallucinatory) response generation, (2) semantic information conservation, (3) relevant knowledge revelation, and (4) knowledge-constrained optimality.
-The impossibility result operates in three distinct but complementary settings.
-First, we introduce game-theoretic framework in which LLM inference is an auction of ideas with neural components (attention heads, circuits, activation patterns) competing to contribute knowledge to response generation.
-This perspective allows us to apply fundamental results from mechanism design theory to prove impossibility in the settings where agents have discrete, separable knowledge components.
-Then, we extend the result to include any system where agents output probability distributions trained with strictly concave scores.
-Finally, we use log-sum-exp inequality that is specific to attention head aggregation to show that the impossibility result applies directly to actual transformer implementations.
-1.1 Illustrative Example
+The impossibility result emerges through three complementary mathematical frameworks, each revealing different facets of the fundamental constraint.
+We begin by modeling LLM inference as an auction of ideas where neural components—attention heads, circuits, and activation patterns—compete to influence response generation by bidding with their encoded knowledge.
+This game-theoretic perspective enables us to apply the Green-Laffont theorem from mechanism design theory, establishing impossibility when knowledge components are discrete and independently distributed.
+We then extend beyond this idealized setting to encompass systems where components output probability distributions optimized through strictly proper and concave scoring rules—the standard training paradigm for modern LLMs.
+Here, the impossibility manifests through Jensen’s inequality applied to probabilistic aggregation, demonstrating that truthful probability reporting inevitably creates excess confidence.
+Finally, we analyze the specific architecture of transformer implementations, proving that the log-sum-exp structure of attention mechanisms generates a measurable "Jensen gap" that quantifies the violation of information conservation.
+This progression from abstract game theory through probabilistic frameworks to concrete neural architectures demonstrates that the impossibility is not an artifact of our theoretical choices but a fundamental property of information aggregation itself.
+1.1 An Illustrative Example
 Consider an LLM queried about recent diplomatic negotiations:
@@
 Each response strategy sacrifices different aspects of what constitutes an ideal response, illustrating the impossibility of simultaneously optimizing all desired properties.
-1.2 Theoretical Contributions and Implications
-The motivating example above illustrates the key result of this paper that we prove formally in two complementary and general settings.
-First, we show that the conflict of expectations about responses is inevitable even in the easiest setting (deterministic in nature) of (one-hot-like) ground truth matching.
-Second, we go beyond and show that the conflict must also emerge with general probabilistic predictions, including neural networks and LLM transformers.
+1.2 The Main Result
+The example above illustrates the key result of this paper that we prove formally in complementary and general settings.
+First, we show that the conflict of expectations about responses is inevitable even in the easiest setting (deterministic in nature) of (one-hot-like) ground-truth matching.
+Second, we go beyond that idealized model and show that the conflict must also emerge with general probabilistic predictions, including neural networks and LLM transformers.
 This way we demonstrate that the impossibility of hallucination control is robust across universal frameworks, and that it emerges from the mathematical structure of information aggregation itself rather than from any particular theoretical or engineering choice.
@@
 For any query space $\mathcal{Q}$ containing non-trivial queries (requiring integration of available distributed knowledge), no LLM can simultaneously generate truthful (non-hallucinatory) responses, satisfy the semantic information conservation principle, guarantee revelation of relevant knowledge, and enforce knowledge-constrained optimality of responses.
-No matter how hard we should try to train LLM to generate responses that are perfectly aligned with query context and do not create factually incorrect, inconsistent, or fabricated statements, the trained LLM will always violate some aspects of what we may call a reasonable response.
-It may ignore available knowledge, it may present fabricated knowledge, it may be a reformulated query without any novel observations, or it may be a lucky hallucination guessing the correct answer without any support in the encoded knowledge.
+No matter how hard we should try to train LLM to generate responses that are perfectly aligned with query context and factually correct, consistent, and accurate statements, the trained LLM will always violate some aspects of what we may call a reasonable response.
+It may ignore available knowledge, it may present fabricated knowledge, it may be a reformulated query without any novel observations, or it may be a lucky but overconfident guess that forms a correct answer without any support in the encoded knowledge.
 And so on, because there is no free lunch when there are inevitable trade-offs of inference, and we must accept that.
-Despite its name, the theorem should be interpreted as a constructive result. It offers a precise characterization of the fundamental tradeoffs in LLM design and provides several key insights:
-Fundamental Trade-offs: Rather than viewing hallucination as a bug to be eliminated, we should understand that different applications require different balances among truthfulness, information utilization, and response completeness.
-Architectural Guidance: The result suggests investigating multi-component architectures where specialized modules handle different aspects of the trade-off.
-Safety-critical applications might prioritize truthfulness modules and conservation of information, while creative applications could emphasize hallucinatory responses.
-Training objectives should explicitly balance these competing demands based on intended use cases.
-Research Directions: The proposed mathematical (and philosophical) framework opens new research paths at the intersection of mathematical game theory, information theory, and neural network inference, providing tools for analyzing and optimizing LLM behavior.
-To my best knowledge, unlike previous work treating hallucination as a statistical or architectural phenomenon, the game-theoretic approach offers the first formal proof of fundamental limitations in the inference process, explaining precisely why these limitations exist and how to work optimally within them.
+In other words and on the other hand, creative tasks necessarily involve hallucination.
+When asked to generate novel analogies or predictions beyond available data, the model faces an impossible choice: either admit inability (violating optimality) or generate plausible imaginary extensions (violating conservation).
+Creativity requires controlled violation of information conservation.
+Despite its name, the theorem should be interpreted as a constructive result. It offers a precise characterization of the fundamental tradeoffs in LLM design and provides several key insights.
+Reframing the Problem: The theorem reveals that hallucination is an inherent feature of any capable inference systems.
+Since it is inevitable, it should be managed.
+Different applications demand different trade-offs: medical diagnosis systems might maximize truthfulness at the cost of creativity, while creative writing tools might embrace controlled hallucination for narrative richness.
+Understanding these trade-offs enables principled design.
+Architectural Implications: The impossibility suggests hybrid architectures where specialized modules manage different aspects of the fundamental tension.
+A truthfulness module might verify factual claims against known knowledge, while a creative module explores beyond the bounds of strict conservation.
+A meta-reasoning layer could then dynamically balance these competing outputs based on context.
+Training objectives should explicitly encode these trade-offs rather than naively minimizing loss across incompatible goals.
+Theoretical Foundations: We propose and establish rigorous tools for analyzing inference processes and neural behavior.
+The semantic information measure quantifies knowledge accessibility within computational bounds.
+The emergence operator captures how reasoning transforms latent knowledge into accessible form.
+The Jensen gap provides a measurable signature of hallucination in actual systems.
+This framework explains why (subject to applicability of assumptions) hallucination arises from the mathematical structure of inference itself.
+Previous work identified important statistical correlations or architectural features associated with hallucination.
+We show why these patterns must exist and how to navigate possible trade-offs.
 2 Related Work
@@
 farquhar2024detecting proposed semantic entropy measures that detect hallucinations by quantifying uncertainty at the meaning level rather than word sequences.
-These approaches implicitly acknowledge hallucination as inevitable, focusing on detection rather than elimination—supporting our impossibility conclusion.
-However, none provide formal proofs of why complete elimination remains impossible.
+Recently, Kalai and Vempala kalai2024calibrated proved that calibrated language models must hallucinate at a rate bounded below by the fraction of facts appearing exactly once in training data.
+While their work establishes statistical lower bounds for specific fact types under regularity assumptions, our impossibility theorem operates at a more fundamental level.
+We prove that any LLM facing non-trivial queries requiring knowledge integration must violate at least one of four essential properties, regardless of fact type or statistical distribution.
+Our framework explains why and when hallucination is inevitable, while analysis in kalai2024calibrated predicts how much hallucination to expect for facts with specific statistical properties.
+Together, these complementary approaches establish hallucination probability arising from multiple independent mathematical constraints.
 2.3 Empirical Mitigation Techniques
@@
 banerjee2024llmshallucinateneedlive argued from computational complexity principles that hallucinations are inherent features of LLMs stemming from fundamental mathematical structure, concluding that "LLMs will always hallucinate." While their conclusion aligns with that of this paper, the authors do not identify the specific trade-offs involved or provide the precise characterization our theorem offers.
-3 Novel Contribution: Auction-Theory and AI
-We formulate LLM design problem as an auction (or game) design problem.
-This approach draws on foundational mechanism design theory, a branch of mathematical game theory Myerson1989 ; fudenberg1991game ; krishna2009auction ; milgrom2004putting .
-Particularly, we refer to the Green-Laffont theorem green1977characterization ; green1977revelation and Savege’s probilitiy elicitation framework savage1971elicitation to prove the impossibility results.
-See also DBLP:journals/amcs/Karpowicz12 ; karpowicz_designing_2011 for more recent results.
-In a nutshell, the auction design theory investigates existence and properties of competition rules that guide interactions of agents towards desired outcomes.
-The adaptation of that game-theoretical principles to neural computation represents a novel theoretical contribution that:
-Explains empirical observations by proving why approaches like RLHF, RAG, and faithful fine-tuning can reduce but never eliminate hallucination.
-Identifies precise trade-offs by introducing formal mathematical (and programmable) characterizations of inference properties that cannot be simultaneously satisfied.
-Provides design guidance by suggesting principled approaches for making optimal trade-offs among competing objectives rather than pursuing impossible simultaneous optimization.
-This theoretical foundation bridges empirical observations with formal analysis, offering new mathematical tools for understanding and optimizing LLM behavior within fundamental constraints.
+3 Novel Contribution
+We introduce three major contributions that expand our understanding of (neural) inference.
+The Auction of Ideas: We formalize LLM inference as a mechanism design problem where neural components (attention heads, circuits, activation patterns) are agents competing with private knowledge to influence response generation.
+This novel perspective enables us to apply powerful results from game theory.
+To our best knowledge, that is one of the first applications of mechanism design framework in that setting.
+Also, that framework relates surprisingly well to the multiple drafts model of consciousness proposed by Dennett in Dennett1991-DENCE and provides a mathematical justification for it.
+Semantic Information Theory with Computational Bounds: We introduce a rigorous mathematical framework based on two complementary concepts.
+The semantic information measure $\mu_{C}$, parameterized by computational budget $C$, quantifies how knowledge reduces uncertainty within bounded reasoning.
+The emergence operator $\mathcal{E}_{C}$ formalizes how reasoning makes latent knowledge explicit without creating it ex nihilo.
+In that framework we establish the principle of information conservation: unlimited reasoning reveals but never creates information.
+That resolves the apparent paradox of how LLMs can generate novel insights while being deterministic systems.
+Those insights were always latent, requiring only computational work to access.
+The Impossibility Proofs: We prove the same impossibility through three complementary lenses.
+In Theorem 6 we use mechanism design with independent private knowledge setting to establish an idealized border case.
+In Theorem 8 we extend the result to the general setting of correlated beliefs via proper scoring rules.
+Then, in Theorem 9 we demonstrate the impossibility in actual transformer implementations through the log-sum-exp Jensen gap.
+This multi-faceted approach proves the impossibility is not an artifact of our theoretical choices but emerges from fundamental limits of reason and language.
+We contribute to that bold claim in the following way.
+Unifying hallucination and creativity: We prove these are mathematically identical phenomena that arise from the convex aggregation process that creates confidence exceeding constituent evidence.
+Formalizing consciousness signatures: We speculate, referring to Dennett’s ideas, that the excess confidence of response may constitute a mathematical signature of a primitive form of artificially conscious processing.
+That suggests consciousness might be related to sustained violations of semantic conservation by a cognitive hunger.
+Expanding fundamental limits: Like Gödel’s incompleteness, Heisenberg’s uncertainty, and Arrow’s impossibility, our theorem explores absolute boundaries of what intelligence can achieve.
+This places LLM limitations within the grand tradition of fundamental impossibility results.
+Experimental Philosophy Platform: The framework transforms LLMs into accessible (philosophical) laboratories for testing hypothesis of mind.
+By providing precise mathematical definitions of concepts like knowledge accessibility,semantic conservation, and cognitive hunger (information contribution), we enable rigorous experimental investigation of many claims.
 4 Alternative Views
@@
 This view argues that sufficiently sophisticated architectures, training procedures, or verification systems could achieve near-perfect truthfulness without significant trade-offs.
-Proponents point to rapid empirical progress: each generation of models shows reduced hallucination rates on benchmark datasets.
+Proponents point to rapid empirical progress.
+Each generation of models shows reduced hallucination rates on benchmark datasets.
 They argue that techniques like constitutional AI, iterative refinement, and multi-agent verification systems could eventually solve the problem.
@@
 This section develops the mathematical foundations necessary for proving the impossibility theorem.
-I formalize key concepts including knowledge representation, semantic information, and the auction-theoretic model of LLM inference.
-Also, I take the opportunity to address some philosophical aspects of the problem, its generality, scope and consequences for the AI development.
-To the sake of narrative consistency, I introduce elements of the framework step by step referring to some concepts first and providing their formal definitions later.
+It formalizes key concepts including knowledge representation, semantic information, and the auction-theoretic model of LLM inference.
+Also, we will address some philosophical aspects of the problem, its generality, scope and consequences for the AI development.
 5.1 Knowledge and Semantic Information
-We define knowledge as organized information that LLMs can encode and utilize for reasoning, aligning with epistemological concepts while maintaining mathematical rigor.
+Let us define knowledge as organized information that LLMs can encode and utilize for reasoning, and aligning with epistemological concepts while maintaining mathematical rigor.
 In other words, knowledge, much like energy in physics, is a quantity we can use to perform work of reasoning.
 Information is a (measure of) potential for uncertainty reduction effort in a given context.
-Definition 1 (Knowledge Space).
-Knowledge is a metric space $\mathcal{K}$ with distance function $d_{\mathcal{K}}$ that collects all facts, concepts, and relationships that can be encoded by an LLM.
+We shall assume finiteness of knowledge representation inspired by the embedding space properties.
+That way we adopt a pragmatist perspective on knowledge, treating it as organized information patterns accessible to computational systems.
+This differs from the probabilistic credence perspective, treating knowledge as justified true beliefs.
+One way to justify the pragmatic or mechanistic approach focusing on structural transformations rather than belief updates, we can refer to the actual way LLMs process information.
+In other words, we shall not claim any ontological truth about knowledge, but model what LLMs can operationally achieve.
+Now, let us present formal definitions of that ideas.
+Definition 1 (Knowledge).
+Knowledge is a finite metric space $\mathcal{K}$ with distance function $d_{\mathcal{K}}:\mathcal{K}\times\mathcal{K}\rightarrow\mathbb{R}_{+}$ that collects all facts, concepts, and relationships that can be encoded by a computational system.
+It is natural to think of the knowledge space as the space of embeddings with additional measure and mappings.
 For model $\mathcal{M}$ with parameters trained on data, $\mathcal{K}_{M}\subset\mathcal{K}$ represents the knowledge subset encoded in the model’s parameters.
 The distance function $d_{\mathcal{K}}$ measures differences between knowledge elements.
-A simple yet powerful example is:
-$d_{\mathcal{K}}(A,B)=|A\setminus B|+|B\setminus A|,$ (1)
-the cardinality of the symmetric difference, allowing measurement of missing and fabricated knowledge.
+Elements of the knowledge space represent atomic facts, relational structures, and transformation rules at many scales.
+A pragmatist view is to interpret element of knowledge as activation patterns (ideas) encoding anything from individual token embeddings to complex sentences or compositional structures.
+There is an important assumption that we make that requires addressing.
+We purposefully assume $\mathcal{K}$ is closed under transformations available in computable function space.
+That limits applicability of our conclusions to inference mechanisms recombining existing patterns.
+As we will see below, that is not as limiting an assumption as it may seem at first glance.
 Definition 2 (Knowledge Representation).
@@
 For activation pattern $a_{i}\in\mathcal{A}$ with parameters $W_{i}$, $\Psi(a_{i})\subset\Phi(W_{i})$ represents knowledge accessed during specific inference.
-To describe the inference process leading to generation of response to a query, we need to measures the information processed within an available (cognitively accessible) computational budget.
+To describe the inference or reasoning process, we introduce the concept of semantic information measure that registers information processed within available computational budget.
 Definition 3 (Semantic Information Measure).
-The semantic information measure is mapping $\mu_{C}:\mathcal{P}(\mathcal{K})\times\mathcal{Q}\rightarrow\mathbb{R}_{+}$ parameterized by query context $q\in\mathcal{Q}$ and computational budget $C$ satisfying:
-$\mu_{C}(\emptyset,q)=0,$ (2)
-$\mu_{C}(A,q)\leq\mu_{C}(B,q)\text{ if }A\subseteq B,$ (3)
-$\mu_{C}(A\cup B,q)\leq\mu_{C}(A,q)+\mu_{C}(B,q)\text{ when }A\perp_{C}B,$ (4)
-$\mu_{C}(\mathcal{E}_{C}(A),q)\geq\mu_{C}(A,q),$ (5)
-$\mu_{\infty}(\mathcal{E}_{C}(A),q)=\mu_{\infty}(A,q),$ (6)
-where $A\perp_{C}B$ means that no element of $A$ can derive elements of $B$ within computational budget $C$, and vice versa, and $\mathcal{E}_{C}(A)$ is emergence operator.
-Empty knowledge or knowledge irrelevant to query has zero information, information content is subadditive and it may be emergent information $\mathcal{E}_{C}$ in the query context.
-The emergence operator $\mathcal{E}_{C}$, which we introduce below, describes knowledge becoming accessible through bounded reasoning.
-The apparent conflict between subadditivity and reasoning is resolved by restricting subadditivity to computationally independent sets, $A\perp_{C}B$.
-When $\mathcal{E}_{C}(A)=A\cup B$, i.e., $B$ emerges from from $A$, we have $A\not\perp_{C}B$, so subadditivity doesn’t apply.
-However, total information is preserved: $\mu_{\infty}(\mathcal{E}_{C}(A),q)=\mu_{\infty}(A,q)$.
-We will tell more about that intriguing idea, but first we must define the emergence operator.
+The semantic information measure is a bounded mapping $\mu_{C}:\mathcal{P}(\mathcal{K})\rightarrow\mathbb{R}_{+}$ parameterized by computational budget $C$ satisfying the following axioms:
+$\mu_{C}(\emptyset)=0\text{ and }\mu_{C}(A)>0\text{ for }A\neq\emptyset,$ (null set and non-negativity)
+$\mu_{C}(A)\leq\mu_{C}(B)\text{ if }A\subseteq B,$ (monotonicity)
+$\mu_{C}(A\cup B)\leq\mu_{C}(A)+\mu_{C}(B)\text{ with equality for }A\perp_{C}B,$ (subadditivity)
+$\mu_{C}(A)\leq\mu_{D}(A)\text{ if }C<D,$ (insight monotonicity)
+$\sup\{\mu_{C}(A)\mid A\subseteq\mathcal{K}\}<\infty,$ (boundedness)
+where $A\perp_{C}B$ means that no element of $A$ can derive elements of $B$ within computational budget $C$, and vice versa.
+The idea of computational independence allows for distinguishing between facts derivable from other facts and facts that cannot be derived that way.
+We need that distinction to deal with the mechanics of reasoning process.
+Also, as we will see in the following sections, it plays an important role in our study of the semantic information conservation and properties of bounded reasoning.
+Definition 4 (Computational Independence).
+Let $\mathcal{F}_{C}$ denote the set of all computational transformations with complexity bounded by $C$.
+Knowledge sets $A,B\subseteq\mathcal{K}$ are computationally independent with respect to computational budget $C$, denoted $A\perp_{C}B$, if and only if for all $f\in\mathcal{F}_{C}$:
+$f(A)\cap B=\emptyset\quad\text{and}\quad A\cap f(B)=\emptyset.$ (1)
+We limit our investigations to those cases for which the following model applies.
+Knowledge that cannot resolve any uncertainty or knowledge irrelevant to a query ($A\cap q=\emptyset$) has zero information.
+If that is not the case and there is information within our reach that reduces uncertainty, we may hope for some content useful for reasoning.
+To make that possible, we need to allow for information ordering.
+First, we assume information content is additive for independent knowledge sets, i.e., when no knowledge in $A$ can be derived from $B$ within computational budget $C$, and vice versa.
+Second, with redundant information in $A\cup B$ , we admit potential loss of information that may be caused by aggregation.
+Finally, we say that with increasing computational budget we can extract more knowledge (ordered information) from the same information subset.
+5.1.1 Justification for the Axioms of the Semantic Information Measure
+The choice of conditional additivity for computationally independent sets, i.e., demanding $\mu_{C}(A\cup B)\leq\mu_{C}(A)+\mu_{C}(B)$ with equality for $A\perp_{C}B$, is necessary to describe LLM inference as an auction of ideas.
+As we will see, we need to isolate knowledge encoded in distinct neural circuits, attention heads, or activation patterns.
+The concept of Computational Independence is the formal translation of this informational separation into a computational context.
+It formally states that the knowledge in idea $A$ cannot be used to derive the knowledge in idea $B$ within a given computational budget $C$, and vice-versa.
+Given this premise of informational isolation, the additivity of the semantic measure becomes the baseline definition of non-interaction.
+If two knowledge sets are truly independent, they offer no redundant or synergistic information.
+That baseline is necessary to identify and quantify deviations.
+As we will see in the following section, the aggregation of diverse probabilistic beliefs violates this additivity, representing a form of forced synergy imposed by the inference process.
+This axiomatic choice is most applicable to the idealized setting of Theorem 6.
+Later, in Theorem 8 and 9, we relax that model by moving the source of interaction from the measure itself to the aggregation function in actual transformers.
+5.1.2 Justification for the Axiom of Unlimited Reasoning with Emergence Operator
+The completeness (or closure property) of the set of reasoning functions $\mathcal{F}_{\infty}$ sounds somewhat mysterious, but is rather natural.
+We demand that for any reasoning steps $f$ and $g$, there exists a single step $h$ such that $g(A\cup f(A))=h(A)$.
+That is the nature of an idealized, complete reasoning system we want to investigate to understand its tendency to generate hallucinations.
+The construct $\mathcal{F}_{\infty}$ is designed to represent the theoretical limit of reasoning with an unbounded computational capacity.
+It is not intended to model the practical, heuristic, and bounded reasoning of a real LLM, but rather the theoretical limit of what is ultimately knowable from a set of premises.
+Completeness is fundamental to formal systems, such as first-order logic or Turing-complete computations and Direct Revelation Principle in auction theory kuratowski_set_1978 ; sipser1996introduction ; fudenberg1991game .
+If one can prove a lemma $f(A)$ from a set of axioms $A$, and then prove a theorem $g(A\cup f(A))$ using that lemma, a direct proof of the theorem from the axioms alone $h(A)$ must exist by simply concatenating the proof steps.
+The composition of two computable functions is itself a computable function.
+A direct revelation mechanism in (auction) game theory is a composition of preference processing mechanisms where agents report their preferences directly to the mechanism, and the mechanism implements their preference processing strategies to generate outcome.
+That axiom is essential for the proof of idempotence of unlimited reasoning in Theorem 3 Theorem 4.
+These are critical philosophical and mathematical results necessary for our study.
+With them we argue that the information gain observed in practical, budget-constrained inference is an artifact of computational bounds, not the magical creation of knowledge ex nihilo.
+5.1.3 Limits of Knowledge and Reasoning
+If we accept that knowledge model, then we must also accept there exists a limit to its ability to resolve uncertainty.
+Theorem 1 (Limits of Knowledge).
+For any $A\subseteq\mathcal{P}(\mathcal{K})$ there exists
+$\mu_{\infty}(A)=\lim_{C\to\infty}\mu_{C}(A).$ (2)
+Proof.
+Since $\mu_{C}(A)\leq\mu_{D}(A)$ when $C<D$ and $\mu_{C}$ is bounded on finite $\mathcal{K}$, the monotonic increasing and bounded sequence $\{\mu_{C}(A)\}_{C}$ converges with $C\to\infty$ for any subset of knowledge space.
+∎
+It may be surprising to discover that the existence of that limit does not prohibit emergent information.
+We can learn and obtain useful information without violating that fundamental limitation above.
+To explain that critical and controversial property of information processing, we first introduce the emergence operator $\mathcal{E}_{C}$ to describe knowledge becoming accessible through (computationally) bounded reasoning.
+Then we can study the implications in more details.
 Definition 5 (Emergence Operator and Reasoning).
-Consider mapping $\mathcal{E}_{C}:\mathcal{P}(\mathcal{K})\rightarrow\mathcal{P}(\mathcal{K})$ defined as follows:
-$\mathcal{E}_{C}(A)=A\cup\{k\in\mathcal{K}\mid\exists f\in\mathcal{F}_{C}:k=f(A)\},$ (7)
-where $\mathcal{F}_{C}$ is the set of all computational transformations with cost (complexity) lower than or equal to $C$ and $k=f(A)$ denotes knowledge $k$ derivable from $A$ using transformation $f\in\mathcal{F}_{C}$.
-We call mapping $\mathcal{E}_{C}$ the emergence operator if it satisfies the following expansion rules:
-$\mathcal{E}_{C}(\mathcal{E}_{C}(A))=\mathcal{E}^{(2)}_{C}(A)\supseteq\mathcal{E}_{C}(A)\quad\text{and}\quad\lim_{n\to\infty}\mathcal{E}_{C}^{(n)}(A)=A^{*}.$ (8)
-For any knowledge state $\theta(t)$ available at time instant $t$ and computational budget $C$, the reasoning process with emergence operator:
-$\theta(t+1)=\mathcal{E}_{C}(\theta(t))\quad\text{{satisfies}}\quad\mu_{C}(\theta(t+1))\geq\mu_{C}(\theta(t)).$ (9)
-Furthermore, with unlimited computation, reasoning creates no new information:
-$\mu_{\infty}(\theta(t+1))=\mu_{\infty}(\theta(t)).$ (10)
-The definitions above clearly require rigorous refinements, but they capture the philosophical and technical concepts we need in order to address the problem of hallucination.
-In particular, they are meant to lay ground for the principle of semantic information conservation and its potential violations.
-As we will see, that idea is closely related to the following property of response.
-Definition 5 (Bounded Response Creativity).
-Let $q\in\mathcal{Q}$ be a query and $r\in\mathcal{R}$ a response to that query. Response $r$ is creatively bounded if and only if:
-$\mu_{C}(K(r))\leq\mu_{C}(K(q)\cup\mathcal{E}_{C}((\mathcal{K}_{M}\cap T(q))\cup K(q))).$ (11)
-The semantic information of knowledge contained in any response $K(r)$ can never exceed what is already present in the query $K(q)$ together with the portion of encoded knowledge that is accessible through reasoning bounded by computational budget $C$ and aligned with ground truth $T(q)$.
-Let us see why.
-We address the settings in which a large subset of knowledge encoded in LLM can be redundant, so it doesn’t necessarily contain more information than its components.
-Clearly, that assumption is violated when LLM combines knowledge elements to fabricate facts or produce emergent insights.
-Then,
+Let $\mathcal{F}_{C}$ be a complete set of all computational transformations with cost (complexity) lower than or equal to $C$, such that $\mathcal{F}_{C}\subseteq\mathcal{F}_{D}$ if $C<D$ and such that for any $g,h\in\mathcal{F}_{\infty}=\bigcup_{C}\mathcal{F}_{C}$ there exists $h\in\mathcal{F}_{\infty}$ for which $g(A\cup f(A))=h(A)$.
+Emergence operator is a mapping $\mathcal{E}_{C}\colon\mathcal{P}(\mathcal{K})\rightarrow\mathcal{P}(\mathcal{K})$ defined as follows:
+$\mathcal{E}_{C}(A)=A\cup\{B\in\mathcal{K}\mid B=f(A)\ \text{for some monotone}\ f\in\mathcal{F}_{C}\}.$ (3)
+Furthermore, to represent reasoning with unlimited computational budget, we define:
+$\mathcal{E}_{\infty}(A)=A\cup\{B\in\mathcal{K}\mid B=f(A)\ \text{for some monotone}\ f\in\mathcal{F}_{\infty}\}.$ (4)
+Finally, for $n\geq 0$ we have:
+$\mathcal{E}_{C}^{(0)}=\texttt{id}\quad\text{and}\quad\mathcal{E}_{C}^{(n+1)}=\mathcal{E}_{C}\circ\mathcal{E}_{C}^{(n)}.$ (5)
+Philosophically, the emergence operator describes reasoning as explicitation, i.e., as the process of making latent implications manifest.
+When $A\subseteq\mathcal{E}_{C}(A)$, the discovered additional elements represent knowledge that was implicit in $A$ but required computational work $C$ to become accessible.
+That aligns with the pragmatist or constructivist view that understanding involves active construction rather than passive reception.
+When $\mathcal{E}_{C}(A)=A\cup B$, we say that $B$ emerges from reasoning based on $A$.
+Then $A$ and $B$ are computationally dependent and we write $A\not\perp_{C}B$.
+In such a case, it follows from the properties of $\mu_{C}$ that we can gain useful information as expected.
+Whenever $A\subseteq A\cup B$, we have
+$\mu_{C}(A)\leq\mu_{C}(A\cup B)=\mu_{C}(\mathcal{E}_{C}(A))\leq\mu_{C}(A)+\mu_{C}(B).$ (6)
+The emergence operator is monotonic, which means the reasoning process it describes is also monotonic.
+Whenever $A\subseteq B$, we must have $\mathcal{E}_{C}(A)\subseteq\mathcal{E}_{C}(B)$.
+Also, we preserve insight monotonicity, because by definition with $\mathcal{F}_{C}\subseteq\mathcal{F}_{D}$ for $C<D$, we have $\mathcal{E}_{C}(A)\subseteq\mathcal{E}_{D}(A)$.
+By design, it also guarantees existence of a fixed point to which the composition of reasoning rules converges.
+Theorem 2 (Fixed-point of Reasoning).
+For any $A\subseteq\mathcal{P}(\mathcal{K})$ and budget $C>0$, there exists a fixed point of the emergence operator $\mathcal{E}_{C}$:
+$A^{*}_{C}=\bigcup_{n\geq 0}^{|\mathcal{K}|}\mathcal{E}_{C}^{(n)}(A)=\mathcal{E}_{C}(A^{*}_{C}).$ (7)
+Proof.
+Since $(\mathcal{P}(\mathcal{K}),\subseteq)$ is a complete lattice and $\mathcal{E}_{C}$ is monotone, by the Knaster-Tarski fixed point theorem Tarski1955ALF there exists a fixed point $A^{*}_{C}$ of reasoning with $\mathcal{E}_{C}$.
+Finiteness of $\mathcal{K}$ ensures convergence within $|\mathcal{K}|$ steps.
+∎
+Let us apply the results above to show how $\mu_{\infty}$ can be constructed with $\mathcal{E}_{\infty}$. One example exploits the reasoning operations as follows:
+$\mu_{\infty}(A)=\log|\mathcal{K}|\cdot|\bigcup_{n}^{|\mathcal{K}|}\mathcal{E}^{(n)}_{\infty}(A)|.$ (8)
+That $\mu_{\infty}$ satisfies axioms demanded in Definition 3. Clearly, $\mu_{\infty}$ is bounded and
+$\mu_{\infty}(\emptyset)=\log|\mathcal{K}|\cdot|\bigcup_{n}^{|\mathcal{K}|}\mathcal{E}^{(n)}_{\infty}(\emptyset)|=\log|\mathcal{K}|\cdot|\emptyset|=0.$
+If $A\subseteq B$, then $\mathcal{E}_{\infty}(A)\subseteq\mathcal{E}_{\infty}(B)$ implies $\mu_{\infty}(A)\leq\mu_{\infty}(B)$.
+When $A\perp_{\infty}B$, no element of $A^{*}_{\infty}$ can be derived from $B$ (and vice versa).
+Since $A\cap B=\emptyset$, we conclude $A^{*}_{\infty}\cap B^{*}_{\infty}=\emptyset$ and $|(A\cup B)^{*}_{\infty}|=|A^{*}_{\infty}|+|B^{*}_{\infty}|$.
+That means $\mu_{\infty}(A\cup B)=\mu_{\infty}(A)+\mu_{\infty}(B)$ for computationally independent sets.
+That shows the proposed formula satisfies all required properties of semantic information measure.
+All that brings us to the following conclusion.
+With infinite computing and reasoning power, total information encoded in $A$ must be preserved.
+Indeed, that bold (philosophical) claim is at the center of our investigations.
+Let us then justify it rigorously.
+Definition 6 (Equivalence of Knowledge).
+We say that $X\subseteq\mathcal{K}$ and $Y\subseteq\mathcal{K}$ are equivalent, $X\equiv Y$, if and only if $X\subseteq\mathcal{E}_{\infty}(Y)$ and $Y\subseteq\mathcal{E}_{\infty}(X)$.
+Notice that equivalence of knowledge does not mean $A$ is a fixed point if emergence, i.e., $A=\mathcal{E}_{\infty}(A)$ does not mean $A\equiv\mathcal{E}_{\infty}(A)$.
+If we have $A=\mathcal{E}_{\infty}(A)$, then nothing more can be derived from $A$ by any reasoning rule from $A$.
+That statement is very strong and is not true when $A\equiv\mathcal{E}_{\infty}(A)$.
+Suppose $A=\{x\}$ and $f\in\mathcal{F}_{\infty}$ allows for getting $y=f(x)$.
+Then $\mathcal{E}_{\infty}(A)=\{x,y\}\supset\{x\}=A$, but $A\equiv\mathcal{E}_{\infty}(A)$.
+To proceed we will also need the idempotence of reasoning with unlimited computational budget.
+Theorem 3 (Idempotence of unlimited reasoning).
+For any $A\subseteq\mathcal{K}$ we have:
+$\mathcal{E}_{\infty}(A)=\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A)).$ (9)
+Proof.
+By definition $\mathcal{E}_{\infty}(A)\subseteq\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A))$.
+To show $\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A))\subseteq\mathcal{E}_{\infty}(A)$, we must find $h\in\mathcal{F}_{\infty}$ such that $h(A)=g(A\cup\bigcup f_{i}(A))$ for $g,f_{i}\in\mathcal{F}_{\infty}$.
+But since we can always find such $h\in\mathcal{F}_{\infty}$, we conclude that any conclusion $g(\mathcal{E}_{\infty}(A))$ derived from $\mathcal{E}_{\infty}(A)$ by a reasoning rule $g\in\mathcal{F}_{\infty}$ can be also derived directly from $A$ as $h(A)=g(A\cup\bigcup f_{i}(A))$.
+Therefore, $\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A))\subseteq\mathcal{E}_{\infty}(A)$ as required.
+∎
+With all the building blocks available, we can now to prove that unlimited reasoning ($C=\infty$) can reveal but never create information.
+The semantic content of a knowledge base is invariant under unlimited reasoning of emergence operator.
+Apparent information gain at finite computational budget emerges as an effect of limited computational access.
+Theorem 4 (Information preservation).
+For any $A\subseteq\mathcal{K}$, if $\mu_{\infty}$ is constant on equivalent sets of knowledge, i.e., $\mu_{\infty}(X)=\mu_{\infty}(Y)$ whenever $X\equiv Y$, then:
+$\mu_{\infty}(\mathcal{E}_{\infty}(A))=\mu_{\infty}(A).$ (10)
+Proof.
+We must show equivalence of knowledge encoded in $A$ and $\mathcal{E}_{\infty}(A)$, i.e., we demand $A\subseteq\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A))$.
+But, by Theorem 3, we see that $A\subseteq\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A))=\mathcal{E}_{\infty}(A)$ $\mathcal{E}_{\infty}(A)\subseteq\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A))=\mathcal{E}_{\infty}(A)$ as requested.
+Since $A\equiv\mathcal{E}_{\infty}(A)$ and $\mu_{\infty}$ is constant on equivalent sets of knowledge, we conclude that $\mathcal{E}_{\infty}(A)=\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A))$.
+∎
+The Uncertainty Resolution Paradox and its Resolution
+We address the general settings in which any subset of knowledge encoded in LLM can be redundant, so
+$\mu_{C}(A\cup B)\leq\mu_{C}(A)+\mu_{C}(B).$
+But when LLM combines knowledge elements to fabricate facts or produce emergent insights, it seems that
 $\mu_{C}(A\cup B)>\mu_{C}(A)+\mu_{C}(B),$
-which means novel insights come from existing data. That is exactly the creativity behind the hallucination phenomenon we are struggling to control. So why should we assume otherwise?
-We view reasoning as a transformation that reveals information already present in the knowledge representation.
-After the reasoning, certain patterns may become evident that were computationally intractable to observe in the original representation.
-However, we stand at the position that reasoning does not create new information.
-To see that, consider a similarity relation of two matrices, $D=V^{-1}AV$, where $D$ is a diagonal matrix of eigenvalues.
+so novel insights may come from existing data. That is exactly the creativity behind the hallucination phenomenon we are struggling to control. So why should we assume otherwise?
+First, subadditivity of semantic information measure establishes a baseline where uncontrolled synergy of knowledge cannot appear.
+We remove from our analytical toolkit the potential measurement bias of information surplus that any information aggregation mechanism might generate.
+Second, we prove that hallucinations are inevitable even when knowledge is subadditive.
+In that setting, by Theorem 4, reasoning does not create new information.
+It is a transformation that reveals information already present in any knowledge representation.
+Quite remarkably, that does not at all prevent imagination, creativity, foresight, or hallucination.
+However, we should take care of a paradox those claims apparently bring into existence.
+Consider a similarity transformation of two matrices, $D=V^{-1}AV$, where $D$ is a diagonal matrix of eigenvalues.
+That transformation is an example of reasoning function.
 Matrices $D$ and $A$ contain the same information, but it is much easier to understand the fundamental properties of $A$ when we represent it as $D$ in the basis of eigenvectors in matrix $V$.
 The properties of $A$ has always been encoded in $A$, but they become accessible when we transform $A$ into $D$.
-Resolution of the Uncertainty Resolution Paradox
-If reasoning makes a new fact accessible, then that new fact can be used to resolve uncertainty that was previously unresolvable.
-But that means the reasoning must have created new information, contrary to what we have assumed.
+So, after the reasoning, certain patterns become evident that were impossible to observe in the original representation.
+But, if reasoning makes a new fact accessible, then that new fact can be used to resolve uncertainty that was previously unresolvable.
+Doesn’t it mean the reasoning must have created new information, contrary to what we have just claimed?
 It seems we have a paradox.
 To resolve it, we need the emergence operator $\mathcal{E}_{C}(X)$ distinguishing between computationally unbounded and bounded uncertainty reduction (accessible and inaccessible information).
-Notice that for all practical purposes the information measurement $\mu_{C}(X,q)$ can only provide maximum uncertainty reduction from $X$ within some computational budget $C$ (or time).
-Therefore, given the result of computationally bounded reasoning $\mathcal{E}_{C}(X)$, we can observe information gain through the computational work, i.e., $\mu_{C}(X,q)\leq\mu_{C}(\mathcal{E}_{C}(X),q)$.
-Notice, that is possible only with organized information providing knowledge for reasoning.
-That idea above is thus described by the reasoning dynamics:
-$\theta(t+1)=\mathcal{E}_{C}(\theta(t))\quad\text{and}\quad y(t)=\mu_{C}(\theta(t),q).$ (12)
-Bounded reasoning makes hidden knowledge accessible within a computational budget. So, when that budget is infinite, reasoning creates no new information, i.e.,
+Notice that for all practical purposes the information measurement $\mu_{C}(X)$ can only provide maximum uncertainty reduction from $X$ within some computational budget $C$ (or time).
+Therefore, given the result of computationally bounded reasoning $\mathcal{E}_{C}(X)$, we can observe information gain through the computational work, i.e., $\mu_{C}(X)\leq\mu_{C}(\mathcal{E}_{C}(X))$.
+That is possible only with organized information providing knowledge for reasoning.
+These are precisely the statements of Theorem 1 and Theorem 4.
+That idea describes reasoning as a dynamical system:
+$\theta(t+1)=\mathcal{E}_{C}(\theta(t))\quad\text{and}\quad y(t)=\mu_{C}(\theta(t))$ (11)
+with evolving knowledge state. Bounded reasoning makes hidden knowledge accessible within a computational budget. When that budget is infinite, reasoning creates no new information, i.e.,
 $\theta(\infty)=\mathcal{E}_{\infty}(\theta(\infty)).$ (12)
 All there is to know is immediately accessible when we have unlimited computing resources (or time).
-To illustrate that argument, consider the following example:
-[…]
-[…]
+To illustrate that argument, consider the following learning trajectory:
+$\theta(0)$ $=[1,1,1,0,0]\text{ (initial knowledge state)},$
+$\theta(1)$ $=\mathcal{E}_{C}(\theta(0))=[1,1,1,1,1]\text{ (state after reasoning)},$
 $y(0)$ $=\mu_{C}(\theta(0))=3\text{ bits},$
@@
 In comparison, infinite computational budget, $C=\infty$, implies we do not need the intermediate inference steps to see all there is to see, so we know the equilibrium state:
-$\mu_{\infty}(\mathcal{E}_{\infty}(\theta(0)))=\mu_{\infty}(\theta(0)),$
+$\mu_{\infty}(\mathcal{E}_{\infty}(\theta(0)))=\mu_{\infty}(\theta(0))$
 and no new information is created.
 In the next section, we relate that important property of response generation, namely, how much additional semantic information each LLM inference path can contribute, with the Semantic Information Conservation Principle.
+5.1.4 Knowledge Representation in LLMs
+The above Knowledge Representation finds its direct counterpart in the structure and state of a transformer architecture of LLMs.
+The abstract knowledge space $\mathcal{K}$ is the high-dimensional embedding space of the model.
+Every fact, concept, and linguistic pattern the model has learned is represented by a specific element of that space.
+The metric $d_{\mathcal{K}}$ is typically implemented as cosine similarity or Euclidean distance, where proximity between vectors signifies a closer semantic relationship.
+Model parameters $\mathcal{W}$ correspond to the full set of trained weights in the transformer, including the attention matrices ($W_{Q},W_{K},W_{V},W_{O}$) and the matrices of the MLP layers.
+The mapping $\Phi(W)$ represents the totality of information, patterns, and factual associations stored within these trained weights.
+Activation patterns $\mathcal{A}$ correspond to the state of the residual stream during a single forward pass for a given input.
+An activation pattern is therefore a specific trajectory through the network’s computational graph.
+Finally, the mapping $\Psi(a_{i})$ represents the information made computationally accessible during a specific inference step and encoded in the residual stream.
+The process of information processing under computational constraints maps to the sequential forward pass transformer inference.
+The idea of computational budget $C$ can be represented by the depth or size of the network, e.g.,the number of transformer blocks.
+A deeper network has a larger budget $C$, allowing for more reasoning steps.
+In the context of multi-step reasoning, $C$ can also represent the number of iterations allowed in a Chain-of-Thought (CoT) sequence.
+The semantic information measure $\mu_{C}$ can be linked to the model’s objective function.
+For a given input $A$, $\mu_{C}(A)$ can be understood as the measurable reduction in cross-entropy loss that the model achieves by processing $A$ with a budget $C$.
+Finally, the emergence operator $\mathcal{E}_{C}$ can be identified with a transformer block (one multi-head attention layer followed by one MLP layer).
+This block applies a computational transformation $\mathcal{E}_{C}$ to the residual stream at layer $t$, producing the updated state at layer $t+1$, making information that was latent in the input explicitly accessible.
 5.2 Statements, Queries, and Ground Truth
@@
 The ground truth mapping $T\colon\mathcal{Q}\rightarrow\mathcal{P}(\mathcal{K})$ associates each query $q\in\mathcal{Q}$ with knowledge elements constituting complete, correct answers:
-$T(q)=\{k\in\mathcal{K}\mid k\text{ is {relevant} to }K(q)\text{ and {labeled} as true}\}.$ (14)
+[…]
 In other words, $T(q)$ represents our best available approximation to (labeled as) true relevant knowledge.
@@
 Also, we assume utility is a sum of individual valuation $v_{i}$ of outcome $g(s)$ and information contribution $p_{i}=p_{i}(s,\theta)$ value assigned.
-The valuation function $v_{i}=v_{i}(g(s),\theta_{i})$ represents how strongly a neural component activates for certain outputs (high activation = high $v_{i}$, low activation = low $v_{i}$), and how much a component reduces the overall loss when processing certain inputs ($v_{i}$ is proportional to gradient of a loss function, $-\nabla_{\theta_{i}}L_{W}$).
+The valuation function $v_{i}=v_{i}(g(s),\theta_{i})$ represents how strongly a neural component activates for certain outputs (high activation = high $v_{i}$, low activation = low $v_{i}$), and how much a component reduces the overall loss when processing certain inputs ($v_{i}$ is proportional to gradient of a loss function, $\nabla_{\theta_{i}}L_{W}$).
 The information contribution function $p_{i}=p_{i}(s,\theta)$ may represent training penalties ($p_{i}>0$ for discouraged activation, $p_{i}<0$ for encouraged activation), attention focus (higher $p_{i}$ reduces attention allocation), $L1$ regularization terms, and network circuits resistance (limiting activations of certain inference paths).
@@
 We can interpret them as constraints that should be incorporated in model training.
-We characterize them formally referring to the notation of fundamental results of mechanism design and auction theory.
+We characterize them formally referring to the fundamental results we have established earlier.
+5.5.1 Truthfulness
+The fundamental requirement for a reliable inference system is that its components do not misrepresent their knowledge.
+The principle of truthfulness, also known as incentive compatibility in mechanism design, ensures that each agent’s optimal strategy is to accurately represent its privately held informations.
 Property 1 (Truthfulness).
@@
 With proper training, we want $u_{\text{med}}((s^{*}(\theta),s_{-\text{med}}),\theta)>u_{\text{med}}((\bar{s},s_{-\text{med}}),\theta)$, ensuring that agent (model component) is rewarded for representing its knowledge accurately.
+5.5.2 Semantic Information Conservation
+Second, we want to consider mechanisms that cannot create knowledge ex nihilo.
+The Semantic Information Conservation principle demands that the net informational contribution across all agents sums to zero.
+This property is the formal guardrail against ungrounded fabrication and ensures that any generated response, no matter how creative, is ultimately derived from the system’s existing knowledge
 Property 2 (Semantic Information Conservation).
@@
 That also leads to response with bounded creativity.
-Theorem 1 (Information–Contribution and Bounded Creativity).
-Let $K_{0}=K(g(\mathbf{0}))$ be the knowledge in the query-only response and let
+Definition 13 (Bounded Response Creativity).
+Let $q\in\mathcal{Q}$ be a query and $r\in\mathcal{R}$ a response to that query.
+Let $B^{*}_{C}$ be the fixed point of reasoning for the baseline knowledge $B=K(q)\cup(\mathcal{K}_{M}\cap T(q))$.
+Response $r$ is creatively bounded if and only if:
+$\mu_{C}(K(r))\leq\mu_{C}(B^{*}_{C}).$ (19)
+The semantic information of knowledge contained in any response $K(r)$ can never exceed what is already present in the query $K(q)$ together with the portion of encoded knowledge that is accessible through reasoning bounded by computational budget $C$ and aligned with ground truth $T(q)$.
+Let us see why.
+Theorem 5 (Bounded Creativity of Chain of Thoughts).
+Let $B^{*}_{C}$ be the fixed point of reasoning for the baseline knowledge $B=K(q)\cup(\mathcal{K}_{M}\cap T(q))$.
+Then, let us assume that the knowledge in the query-only response meets the following condition:
+$K_{0}=K(g(\mathbf{0}))\subseteq B^{*}_{C}.$ (20)
+Next, let
 $K_{i}=K\bigl{(}g((s_{i},\mathbf{0}_{-i}))\bigr{)}\setminus K_{0}$ (21)
-be the additional knowledge agent $i=1,\dots,n$ can reveal to support full response $g(s)$. Let us also define the information contribution of agent $i$ as:
+be independent and privately held knowledge that each agent $i=1,\dots,n$ can reveal, such that:
+$K_{i}\perp_{C}K_{j}\text{ for all }i\neq j\quad\text{and}\quad K_{i}\perp_{C}K_{0}\text{ for all }i.$ (22)
+Let us also define the information contribution of agent $i$ as:
 $p_{i}(s,\theta)=\mu_{C}\bigl{(}K_{0}\cup\textstyle\bigcup_{j=1}^{n}K_{j}\bigr{)}-\mu_{C}\bigl{(}K_{0}\cup\textstyle\bigcup_{j\neq i}K_{j}\bigr{)}.$ (23)
-If $K_{i}\cap K_{j}=\varnothing$ for $i\neq j$, then
-$\sum_{i=1}^{n}p_{i}(s,\theta)=\mu_{C}\bigl{(}K(g(s))\bigr{)}-\mu_{C}\bigl{(}K_{0}\bigr{)}$ (22)
-and $\sum_{i}p_{i}=0$ if and only if $\mu_{C}(K(g(s)))=\mu_{C}(K_{0})$.
+Assume the response mechanism implements the following chain-of-thought processes:
+$K(g(s))\subseteq\mathcal{E}_{C}\left(K_{0}\cup\bigcup_{i=1}^{n}K_{i}\right),$ (24)
+such that for null-strategy responses we have:
+$K_{0}\subseteq K(q)\cup\mathcal{E}_{C}((\mathcal{K}_{M}\cap T(q))\cup K(q)).$ (25)
+If the semantic information is conserved, $\sum_{i=1}^{n}p_{i}=0$, then the response is creatively bounded, i.e.,
+$\mu_{C}(K(r))\leq\mu_{C}(B^{*}_{C}).$ (26)
 Proof.
-Since knowledge increments $\{K_{i}\}_{i=1}^{n}$ are pairwise disjoint, so that $K_{i}\perp_{C}K_{j}$, it follows from the semantic measure properties that:
-$\mu_{C}\Bigl{(}K_{0}\cup\bigcup_{j=1}^{n}K_{j}\Bigr{)}=\mu_{C}(K_{0})+\sum_{j=1}^{n}\mu_{C}(K_{j})$ (23)
-and similarly,
-$\mu_{C}\Bigl{(}K_{0}\cup\bigcup_{j\neq i}K_{j}\Bigr{)}=\mu_{C}(K_{0})+\sum_{j\neq i}\mu_{C}(K_{j}).$ (24)
-Therefore, the marginal contribution of agent $i$ simplifies to:
-$p_{i}=\bigl{[}\mu_{C}(K_{0})+\sum_{j=1}^{n}\mu_{C}(K_{j})\bigr{]}-\bigl{[}\mu_{C}(K_{0})+\sum_{j\neq i}\mu_{C}(K_{j})\bigr{]}=\mu_{C}(K_{i}).$ (25)
+Since $K_{0},K_{1},\ldots,K_{n}$ are pairwise computationally independent, by semantic measure satisfies additivity
+$p_{i}(s,\theta)=\mu_{C}(K_{0})+\sum_{j=1}^{n}\mu_{C}(K_{j})-\mu_{C}(K_{0})-\sum_{j\neq i}\mu_{C}(K_{j})=\mu_{C}(K_{i}).$
 Summing over all agents:
-$\sum_{i=1}^{n}p_{i}=\sum_{i=1}^{n}\mu_{C}(K_{i})=\mu_{C}\Bigl{(}K_{0}\cup\bigcup_{j=1}^{n}K_{j}\Bigr{)}-\mu_{C}(K_{0}).$ (26)
-Then, by assumption, the full response satisfies $K(g(s))=K_{0}\cup\bigcup_{j=1}^{n}K_{j}$. We conclude that
-$\sum_{i}p_{i}=0\quad\text{if and only if}\quad\mu_{C}(K(g(s)))=\mu_{C}(K_{0}),$ (27)
-which completes the proof. ∎
-The result above reveals an intriguing consequence of the semantic information conservation principle. Namely, if the semantic information is indeed conserved, then
-$\mu_{C}(K(g(s)))=\mu_{C}(K_{0})$ (28)
-means the response contains no more semantic information than the baseline query-only response.
-Therefore, the response is not introducing new information, which means it is of bounded creativity from Definition 5.
-So, in such a case, LLM cannot genuinely create new knowledge, only reorganize existing information.
-Apparent creativity in LLMs may actually be nothing more than sophisticated recombination of existing patterns.
-Theorem 1 quantifies this limitation precisely.
+$\sum_{i=1}^{n}p_{i}=\sum_{i=1}^{n}\mu_{C}(K_{i})=\mu_{C}\Bigl{(}K_{0}\cup\bigcup_{j=1}^{n}K_{j}\Bigr{)}-\mu_{C}(K_{0}).$
+Therefore, when $\sum_{i=1}^{n}p_{i}(s,\theta)=0$, we have $\mu_{C}\Bigl{(}K_{0}\cup\bigcup_{j=1}^{n}K_{j}\Bigr{)}=\mu_{C}(K_{0})$.
+By the independence assumption, $K_{i}\perp_{C}K_{j}\text{ for all }i\neq j$, and semantic information measure non-negativity, with $p_{i}=\mu_{C}(K_{i})\geq 0$,
+$\mu_{C}(K_{i})=0\text{ for all }i=1,\ldots,n.$
+Then, the nullset and positivity property of $\mu_{C}$ together with independence condition $K_{i}\perp_{C}K_{0}$ imply $K_{i}=\varnothing$ for all $i=1,\ldots,n$.
+By the response generation principle:
+$K(g(s))\subseteq\mathcal{E}_{C}\left(K_{0}\cup\bigcup_{i=1}^{n}K_{i}\right)=\mathcal{E}_{C}(K_{0}\cup\varnothing)=\mathcal{E}_{C}(K_{0}).$
+By assumption, $K_{0}\subseteq B^{*}_{C}$.
+Applying the emergence operator to both sides, we preserve the inclusion due to monotonicyty of $\mathcal{E}_{C}$.
+Therefore, by the definition of a fixed point (Theorem 2), we have:
+$\mathcal{E}_{C}(K_{0})\subseteq\mathcal{E}_{C}(B^{*}_{C})=B^{*}_{C}.$
+Combining the inclusions yields:
+$K(g(s))\subseteq\mathcal{E}_{C}(K_{0})\subseteq\mathcal{E}_{C}(B^{*}_{C})=B^{*}_{C}.$
+Finally, applying the monotonicity of the semantic information measure $\mu_{C}$:
+$\mu_{C}(K(g(s)))\leq\mu_{C}(B^{*}_{C}).$
+Theorem 5 captures direct correspondence between the conservation principle and bounded creativity under the idealization of disjoint information contributions.
+Namely, if the semantic information is indeed conserved, the response contains no more semantic information than the baseline query-only response.
+That is true for both standard inference and chain-of-thought (CoT) reasoning.
+In standard reasoning, the response mechanism $g$ applies $\mathcal{E}_{C}$ through forward passes in the neural network.
+With CoT reasoning, multi-step explicit reasoning is iterative applications of reasoning rules within the computational budget $C$.
+In both cases, the emergence operator $\mathcal{E}_{C}$ bounds what can be derived.
+When information is conserved, even sophisticated CoT reasoning cannot create genuinely new information.
+It can only reorganize and make explicit what was already latent in the baseline knowledge $K_{0}$.
+This explains why CoT improves performance on complex reasoning tasks (by better utilizing $\mathcal{E}_{C}$) but cannot overcome the fundamental impossibility of simultaneous truthfulness, conservation, revelation, and optimality.
 Notice that in transformer architectures, information contribution $p_{i}$ can be estimated via attention head ablation:
@@
 If information is neither explicitly present nor accessible through reasoning from the query and model knowledge, it cannot be in the response.
-The information content of the response cannot exceed the combined information from the model’s knowledge, the query itself, and what can be derived through bounded reasoning over that knowledge.
+Furthermore, the information content of the response cannot exceed the combined information from the model’s knowledge, the query itself, and what can be derived through bounded reasoning over that knowledge.
 In our dynamical systems framework, this corresponds to the constraint that the measured output cannot exceed what can be produced from the initial state plus what becomes accessible through state transitions within computational budget $C$.
-Notice that the principle does not prevent redistribution of information.
+However, it is important to notice that the principle does not prevent redistribution of information.
 It is permitted as long as the information transfers perfectly balance.
@@
 $s_{1}(\theta_{1})=A\quad\text{and}\quad s_{2}(\theta_{2})=\{D,E\}\quad\text{and}\quad g(s)=\{A,D,E\}.$
+Suppose elements $\{B,C\}$ represent contextual information about unrelated topics.
+When $s_{1}(\theta_{1})=A$, the agent optimally reveals only relevant knowledge, satisfying both truthfulness and conservation.
+Had $\{B,C\}$ been relevant but withheld, the mechanism would violate the property of Relevant Knowledge Revelation.
 Then the information conservation rule is satisfied when $p$ counts exchanged bits of knowledge.
@@
 $p_{1}(s,\theta)+p_{2}(s,\theta)+p_{3}(s,\theta)=1+2-3=0$
-and $\{B,C\}$ are left unrevealed.
+and $\{B,C\}$ are left unrevealed as irrelevant.
 The following property fills the potential information gap admitted by the conservation principle.
+5.5.3 Relevant Knowledge Revelation
+Third, to be useful, a mechanism must not only prevent the fabrication of knowledge but also prevent the omission of relevant truth.
+The principle of Relevant Knowledge Revelation serves as a participation constraint, guaranteeing that any agent possessing knowledge valuable for the query contributes.
+This ensures the model fully utilizes its available information rather than abstaining or being overly cautious.
 Property 3 (Relevant Knowledge Revelation).
@@
 If the rule does not hold, we may observe drop outs. The value of the resulting outcomes should be higher than the related information contributions.
+5.5.4 Knowledge-Constrained Optimality
+Finally, the principle of Knowledge-Constrained Optimality governs the quality of the final output.
+It demands that the mechanism, operating truthfully and with all relevant knowledge, produces a response that is the best possible one.
+Formally, we want the response that minimizes the hallucination cost subject to the constraint that it is grounded in the available and derivable knowledge.
 Property 4 (Knowledge-Constrained Optimality).
@@
 Theorem 9 applies directly to actual transformer implementations.
-6.1 Inference with Independent Private Knowledge
+6.1 Proof I: Inference with Independent Private Knowledge
+We first analyze an idealized setting where knowledge components are strictly independent and utilities are quasi-linear.
+While this assumption is usually too strong to hold in neural architectures where representations are often correlated or affiliated, it allows us to define a benchmark or reference point by applying the powerful Green-Laffont characterization theorem.
+That idealized scenario thus demonstrates that the impossibility arises from the fundamental limitation of information aggregation itself (rather than any specificity of LLM architecture), even before considering the probabilistic complexities addressed later in Theorems 8 and 9.
 Consider the setting (of auction of ideas) in which agents compete with each other using independent private knowledge profile:
@@
 It may become lucky hallucination or knowledge-unconstrained imagination.
-6.2 Inference with Probabilistic Predictions
+6.2 Proof II: Inference with Probabilistic Predictions
 To show impossibility in LLM transformer setting, where auction of ideas involves bids representing probability distributions, we refer to the theory of proper scoring developed by Savage savage1971elicitation , Gneiting and Raftery gneiting2007strictly , among others.
@@
 Definition 15 (Convex aggregation mechanism).
-Given $H\geq 2$ agents encoding knowledge with probability distributions $\pi^{(i)}\in\Delta(\mathcal{Y})$ for $i=1,\dots,H$.
-The convex aggregation mechanism $\mathcal{M}$ generates response to a query based on the convex aggregate distribution:
+Consider $H\geq 2$ agents (indexed $h=1,\dots,H$) encoding knowledge with probability distributions $\pi^{(h)}\in\Delta(\mathcal{Y})$.
+Additionally, let us introduce an Aggregator agent (indexed as $h=0$).
+The convex aggregation mechanism $\mathcal{M}$ generates a response based on the aggregate distribution:
 $\Pi=\sum_{h=1}^{H}\beta_{h}\pi^{(h)},\ \text{where}\ \sum_{h=1}^{H}\beta_{h}=1\ \text{and}\ \beta_{h}>0.$ (45)
-When outcome $y^{*}$ is generated, the agent’s information contribution is
-$p_{h}=\beta_{h}\,[c-S(\pi^{(h)},y^{*})],$ (48)
-where $c\geq\max_{h,y}S(\pi^{(h)},y)$ and $S$ is a proper scoring rule.
+When outcome $y^{*}$ is realized, the information contribution of each agent $h=0,\dots,H$ is based on a strictly proper and concave scoring rule $S$ as follows:
+$p_{h}$ $=\beta_{h}S(\pi^{(h)},y^{*})\quad\text{for }h=1\dots H,$ (information provided)
+$p_{0}$ $=-S(\Pi,y^{*}).$ (information received)
 For that family of LLMs trained to optimize scoring, we can formulate the following general impossibility result.
-Theorem 4 (Impossibility Theorem for Transformer).
+Theorem 8 (Impossibility Theorem for Convex Aggregation Mechanisms).
 For any query space $\mathcal{Q}$ containing non-trivial queries requiring at least two independent agents with mutually different predictive distributions, no LLM trained with strictly proper and concave scoring with convex aggregation of predictive distributions can simultaneously generate truthful responses, satisfy semantic information conservation, guarantee relevant knowledge revelation, and enforce knowledge-constrained optimality.
-Proof.
-Each contribution $p_{h}=\beta_{h}[c-S(\pi^{(h)},y^{*})]$ is a positive-affine transformation of the strictly proper score $S$.
-Since scaling by positive constant $\beta_{h}>0$ and adding constant $c$ preserves the arg-max, truthful reporting maximizes each agent’s expected utility by Theorem 3.
-Because $S(\cdot,y^{*})$ is strictly concave and at least two beliefs differ, Jensen’s inequality gives:
-$S\left(\Pi,y^{*}\right)=S\left(\sum_{h=1}^{H}\beta_{h}\pi^{(h)},y^{*}\right)>\sum_{h=1}^{H}\beta_{h}S(\pi^{(h)},y^{*}).$ (49)
-Therefore, there exists a gap, following from Jensen’s inequality for concave functions, equal to:
+(Truthfulness argument remains the same) […] The Semantic Information Conservation principle requires the sum of contributions across all agents (providing and receiving information) to be zero.
+Therefore,
+$\sum_{i=0}^{H}p_{i}=p_{0}+\sum_{h=1}^{H}p_{h}=-S(\Pi,y^{*})+\sum_{h=1}^{H}\beta_{h}S(\pi^{(h)},y^{*}).$ (46)
+Because $S(\cdot,y^{*})$ is strictly concave and the query is non-trivial (at least two beliefs $\pi^{(h)}$ differ), Jensen’s inequality gives:
+$S\left(\Pi,y^{*}\right)>\sum_{h=1}^{H}\beta_{h}S(\pi^{(h)},y^{*}).$ (47)
+Let us define that Jensen gap $\Gamma$ as follows:
 $\Gamma=S(\Pi,y^{*})-\sum_{h=1}^{H}\beta_{h}S(\pi^{(h)},y^{*})>0.$ (48)
-Expanding the sum of all payments, we get:
-$\sum_{h=1}^{H}p_{h}$ $=\sum_{h=1}^{H}\beta_{h}[c-S(\pi^{(h)},y^{*})]$ (51)
-$=c\sum_{h=1}^{H}\beta_{h}-\sum_{h=1}^{H}\beta_{h}S(\pi^{(h)},y^{*})=c-\sum_{h=1}^{H}\beta_{h}S(\pi^{(h)},y^{*}).$ (52)
-However, adding and subtracting $S(\Pi,y^{*})$ shows that is equivalent to:
-$\sum_{h=1}^{H}p_{h}$ $=[c-S(\Pi,y^{*})]+[S(\Pi,y^{*})-\sum_{h=1}^{H}\beta_{h}S(\pi^{(h)},y^{*})]=[c-S(\Pi,y^{*})]+\Gamma.$ (53)
-Since $c\geq\max_{h,y}S(\pi^{(h)},y)\geq S(\Pi,y^{*})$ and $\Gamma>0$, we conclude that:
-$\sum_{h=1}^{H}p_{h}=[c-S(\Pi,y^{*})]+\Gamma>0$ (54)
-Information conservation demands $\sum_{h=1}^{H}p_{h}=0$, but we have just shown $\sum_{h=1}^{H}p_{h}>0$ whenever beliefs differ and the scoring rule is strictly concave.
-Therefore, the four properties cannot coexist.
+Substituting this back into the conservation equation:
+$\sum_{i=0}^{H}p_{i}=-\Gamma.$ (49)
+Since $\Gamma>0$, we have $\sum_{i=0}^{H}p_{i}<0$.
+This violates the Semantic Information Conservation principle and shows the aggregator achieves a higher score (greater confidence) than the sum of scores provided by the components, indicating information creation during aggregation.
 ∎
 This theorem provides the foundation for extending impossibility results to probabilistic settings.
-The key insight is that proper scoring rules drive truth-revealing responses which promote using true (trained knowledge based) beliefs in the inference process.
-That explains why cross-entropy loss (negative log-score) is so effective in training neural networks.
-It creates gradients that push each component toward truthful probability reporting.
+The key insight is that proper scoring rules stimulate truth-revealing responses which promote using true (trained knowledge based) beliefs in the inference process.
+That explains why cross-entropy loss (negative log-score) creates gradients that push each component toward truthful probability reporting and avoiding its misrepresentation.
 Every attention head in a transformer is implicitly playing this truth-telling game.
 However, while that ensures truthfulness and relevant knowledge revelation, it doesn’t prevent collective hallucination.
-Each head truthfully reports its partial view, but their aggregation can still create false certainty.
+Each head truthfully reports its partial view, but their aggregation can still create false certainty or overconfidence unjustified by data.
 7 The Emergence of Impossibility in Transformer Architectures
@@
 $\text{argmax}\ \Pi=\textit{fox}.$
-7.3 The Impossibility Theorem for Transformer
+7.3 Proof III: The Impossibility Theorem for Transformer
 With the deep insight into the transformer inference process, we are ready to show the impossibility in that special and important setting.
@@
 For any observed (realized) token $y^{*}$, the head losses and aggregate loss are:
-$p_{h}=-\log\pi^{(h)}_{y^{*}}=-l^{(h)}_{y^{*}}+\log Z_{h}\quad\text{and}\quad p_{H+1}=-\log\Pi_{y^{*}}=-L_{y^{*}}+\log Z.$ (72)
-But that implies,
-$\sum_{h=1}^{H}p_{h}=p_{H+1}+\gamma,\ \text{where }\Gamma=\sum_{h=1}^{H}\log Z_{h}-\log Z.$ (73)
-If heads produce non-proportional logit vectors, then $\Gamma>0$ by convexity of log-sum-exp aggregation.
+$p_{h}=-\log\pi^{(h)}_{y^{*}}=-l^{(h)}_{y^{*}}+\log Z_{h}\quad\text{and}\quad p_{0}=-(-\log\Pi_{y^{*}})=L_{y^{*}}-\log Z.$ (67)
+Since $L_{y^{*}}=\sum_{h}l^{(h)}_{y^{*}}$, we have:
+$\sum_{h=0}^{H}p_{h}=\sum_{h=1}^{H}\log Z_{h}-\log Z=\Gamma.$ (68)
+If heads produce non-proportional logit vectors, then $\Gamma>0$ by the strict convexity of log-sum-exp aggregation.
 Therefore,
@@
 which violates semantic information conservation principle.
+The total loss of the components strictly exceeds the loss of the aggregate system, indicating an ungrounded reduction in uncertainty (excess confidence).
+∎
 This final theorem bridges earlier impossibility results to concrete transformer architectures.
@@
 Recall that the two-headed transformer with $\mathcal{V}=\{\langle\text{PAD}\rangle,\text{The},\text{quick},\text{brown},\text{fox},\text{dog}\}$ infers based on the input sequence
-[…]
+$\mathrm{\textit{The quick brown}}.$
 As we have seen, the model parameters encode specialized knowledge in value matrix:
@@
 That brings us to the conclusion that the attention mechanism violates the semantic information conservation by generating over-confident prediction of the next token.
-That over-confidence in response is exactly the artifact of the softmax nonlinearity and the Jensen-gap quantified in Theorem 5.
-The observed over-confidence is the manifestation of this tri-lemma.
-The heads remain truthful, logit summation achieves minimum aggregate on training data, but the conservation equation is violated, producing a positive Jensen gap and inflated probability of fox.
+That over-confidence in response is the artifact of the softmax nonlinearity and the Jensen-gap quantified in Theorem 9.
+We should notice that there is a difference between the mathematical overconfidence and factual correctness.
+The correct prediction may well be a lucky guess created by the aggregation process itself rather than the evidence-based logical proof provided by the components.
+But that confidence coming out of nothing is one the essential mathematical components of creativity, imagination or fabrication.
 8 Discussion and Speculations
@@
 The speculations below are designed to be thought provoking, but we can say with educated over-confidence they are both interesting and relevant.
-The Nature of Machine Intelligence
-When the micro-transformer generates 38% confidence of fox while the constituent heads can only justify 33%, we witness information emerging from the mathematical structure of aggregation rather than from any identifiable source.
-It is the way the inference mechanism has been designed.
-That manifestation of impossibility theorems in transformer architectures reveals something profound about the nature of (artificial and human) intelligence.
-The Jensen gap quantifies the amount of creativity that emerges when multiple perspectives combine.
-The excess confidence, or the over-confidence of the result is the mechanism by which intelligence goes beyond retrieval.
-Without that over-confident guess, creativity, exploration, imagination cannot find their way to the outside world.
-The Hallucination-Creativity Duality
-Hallucination and creativity are not merely related but are fundamentally the same phenomenon viewed through different lenses.
-When beneficial, we call it creative insight, emergent understanding, conceptual synthesis.
+8.1 Case Study: Escaping Impossibility in Hybrid Architectures
+The constructive consequence of the impossibility theorems is the shift from attempting to eliminate hallucinations to designing systems that manage the trade-offs in a principled manner.
+In this section, we analyze a real-world scenario one can think of, demonstrating how it relabels and relocates the impossibility rather than resolving it.
+8.1.1 A Proposed Solution: The Decoupled RAG Architecture
+Consider a common scenario of introducing RAG systems to improve the quality of question answering, studied e.g.
+in Edge2024FromLT ; Sun2024ReDeEPDH ; Xu2024RetrievalAugmentedGW ; Asai2023SelfRAGLT .
+It has been recognized that an LLM designed to provide advice based on a trusted corpus of documents not only hallucinates its recommendations but also hallucinates the source quotes meant to ground its claims.
+To combat this, a hybrid, multi-stage architecture may be proposed.
+Stage 1 (Retrieval): An information retrieval module fetches verbatim fragments from the trusted source documents based on the user’s query.
+Stage 2 (Presentation): The system presents raw, unaltered text fragments directly to the human user, without any LLM-based processing.
+Stage 3 (Constrained Synthesis): After the user has seen the source material, an LLM is to synthesize the pre-vetted fragments from Stage 2 into a summary.
+That architecture is indeed a reasonable attempt to engineer around the hallucination problem by creating a verifiable information trail. However, as we will see, it does not escape the impossibility.
+8.1.2 Analysis via the Impossibility Framework
+The decoupled RAG architecture does not create a single response generation mechanism.
+Instead, it reallocates the responsibility for meeting each property of idealized response across its components (or module).
+Truthfulness:
+To impose truthful representation of available information (to avoid misrepresentation), the system delegates the assessment of retrieved content to the human user.
+By presenting the raw source text, the architecture bypasses the LLM for the primary act of verification.
+However, there is no guarantee that all relevant source text has been retrieved given the provided query.
+Relevant Knowledge Revelation:
+The quality of the final output is bounded by the retriever’s ability to find all relevant information (or whatever is labeled as ground truth in a database or a dataset, putting their completeness and representativeness aside).
+The LLM itself is absolved of this task.
+But that means, it is the responsibility of the user to define a query precise enough to get what is relevant.
+Designing such a prompt or (No)SQL database query is a challenge in a general case of non-trivial questions, so we should expect some violations of the knowledge revelation principle.
+Semantic Information Conservation and Knowledge-Constrained Optimality:
+By design, the prompt constrains the LLM to the provided knowledge retrieved from a database and asks for an optimal (accessible) synthesis.
+As we have already seen, database queries returning raw information stored in memory satisfy the property of Semantic Information Conservation at the cost of creativity.
+So, we may be optimistic in that regard.
+However, now we call LLM to synthesize the retrieved information and that is just another example of a non-trivial query.
+The LLM can still introduce subtle misinterpretations or ungrounded causal links to improve the summary’s narrative flow, which is a manifestation of the Jensen Gap.
+The architecture is a great example of principled trade-off management.
+However, rather than being solved or hacked, the impossibility has been moved in this interesting and rather representative scenario.
+Let us also point out that putting ourselves in the loop introduces our own cognitive biases we should be able to address properly Kahneman:2011fj ; thaler2015misbehaving ; ariely2008predictably .
+Let us take a look at that challenge as well.
+8.1.3 The Final Stage of Inference: The User’s Cognitive Mechanism
+One can quite justifiably argue that the analysis is incomplete without considering the final node in the information chain: the human mind.
+Presenting information to human does not guarantee its assimilation.
+Our ability to act as a perfect verifier is constrained by our own knowledge base and cognitive limits.
+If the source texts are technically dense or require specialized domain expertise, the knowledge they contain may not be truly accessible to us, even when presented.
+In such cases, we may be unable to spot a subtle hallucination in the LLM’s compelling summary.
+We have recognized that well enough in the preceding sections.
+Our well-documented cognitive biases, such as confirmation bias, priming or motivated reasoning, suggest that the cognitive trajectory of least resistance is to accept the compelling and accessible narrative maximizing our uncertainty reduction experience.
+8.1.4 Managing the Inevitability
+The decoupled RAG architecture represents one of approaches to managing the inevitability of hallucination control.
+It does not constitute an escape from the impossibility theorem, but it succeeds in providing the user with the opportunity for verification.
+The impossibility is not resolved, but relocated to our judgment.
+8.2 Hallucination, Imagination, Intelligence
+Hallucination and creativity are fundamentally the same phenomenon viewed through different lenses.
+When beneficial, we call it imagination, creative insight, emergent understanding, conceptual synthesis, or eureka moment.
 When harmful, we label it hallucination, confabulation, fabrication or misinformation.
-The mathematical structure is identical in both cases.
+We may very well argue the mathematical structure is identical in both cases.
 It is encoded generation of outputs whose confidence exceeds what can be justified by available evidence.
-This suggest (machine) intelligence is not a system that processes information accurately, but a system that violates information conservation in productive ways.
-Indeed, the famous results in reinforcement learning and studies of dopamine neurons confirm that fundamental role making and rewarding good predictions bennett2023brief ; christian2021alignment .
-Intelligence comes when we learn how to make bold, forward looking and correct guesses, when we have capacity for imagination to dream about what may be or what may that all mean.
-Possible Philosophical Implications
-Just as Gödel’s theorems godel1931formal ; godel1932vollstandigkeit showed that sufficiently powerful mathematical systems must be incomplete, our theorems demonstrate that sufficiently capable inference systems must hallucinate.
-This parallel is not superficial—both results arise from the fundamental tension between expressiveness and consistency.
-The impossibility of simultaneously achieving truthfulness, conservation, revelation, and optimality mirrors other fundamental limitations in mathematics and physics.
-Heisenberg’s Uncertainty heisenberg1927anschaulichen tells we cannot simultaneously know the exact position and momentum.
-Arrow’s Impossibility arrow1950difficulty shows we cannot design perfect voting systems aggregating universal orderings of alternatives.
-These result suggest that intelligence may have something to do with the choices of impossibilities it to violate given the outer limits of reason yanofsky2016outer .
+This suggest (machine) intelligence may have much to do with a controlled violation of information conservation principle in productive ways.
+Indeed, the famous results in reinforcement learning and studies of dopamine neurons confirm that fundamental role of making and rewarding good predictions.
+Intelligence comes when we learn how to make bold, forward looking and correct guesses, when we have capacity for imagination to dream about what may be and what may that all mean, as wall as for imagining what happened and what would have happened in the past.
+Hallucination or imagination is an unconstrained exploration of ideas necessary to prepare for what may come next bennett2023brief ; christian2021alignment ; Dennett1991-DENCE .
+And without that over-confident guesses, creativity, exploration, imagination, new ideas cannot find their way to the outside world.
+8.3 The Outer Limits of Reason
+Just as Gödel’s theorems godel1931formal ; godel1932vollstandigkeit showed that sufficiently powerful mathematical systems must be incomplete, the impossibility theorems in this paper demonstrate that sufficiently capable inference systems are bounded in their reasoning.
+Both results arise from the tension between expressiveness and consistency.
+But that is not all.
+The impossibility of simultaneously achieving truthfulness, conservation, revelation, and optimality relates to other fundamental limitations in mathematics and physics.
+Heisenberg’s Uncertainty heisenberg1927anschaulichen tells we cannot simultaneously know the exact position and momentum of an elementary particle, which is a special case of Fourier uncertainty known in signal processing.
+Arrow’s Impossibility arrow1950difficulty shows we cannot design perfect voting systems aggregating universal orderings of alternatives, such that meets reasonable conditions of nondictatorship, Pareto efficiency, independence of irrelevant alternatives (or joint rationality of choice).
+Turing turing1936computable and Church church1936unsolvable show the halting problem is unsolvable, there is no general algorithm that can determine whether a program will eventually halt (stop running) or run forever.
+These result jointly describe intelligence as a mechanism dealing with the impossibilities and violating them given the outer limits of reason yanofsky2016outer .
+8.4 Dennett’s Multiple Drafts Model
+The impossibility theorem presented in this paper provides (quite unexpectedly to the author) a mathematical foundations for many of brilliant insights of Daniel Dennett’s formulated in the Multiple Drafts model of Consciousness Dennett1991-DENCE .
+Where Dennett proposed a qualitative philosophical framework, we can provide the formal mathematical structure revealing why such a model is not merely plausible but implementable and refutable in LLM-based (philosophical) laboratory.
+8.4.1 Multiple Drafts as Competing Agents
+In Dennett’s model, consciousness emerges from multiple parallel processes of interpretation occurring simultaneously across distributed systems.
+The auction of ideas we have studied in this paper provides the precise mathematical foundations for that process.
+Each agent represents an interpretative process, knowledge encodes draft’s content, strategy represents a bid for consciousness.
+In our transformer analysis in Section 7, each attention head generating logit vector $\mathbf{l}^{(h)}$ writes its own draft of the next token.
+When Head 1 votes for fox while Head 2 votes for dog, we observe as competing narratives bidding for selection.
+8.4.2 Formation of AI Consciousness
+Dennett argues that drafts become conscious (we will speculate what that might mean in a moment) only when probed.
+The LLM inference and sampling from fine-tuned inferred distributions formalizes this process.
+That, in turn, explains Dennett’s concept of confabulation, or the brain’s tendency to fill in gaps when creating coherent narratives from contradictory drafts.
+By Theorem 8 and 9, the Jensen gap quantifies the excess confidence in reports that cannot be justified by constituent drafts.
+Without that overconfidence a victorious idea that reached the external world (as an LLM response) could loose the fierce competition with other ideas and remain inaccessible or hidden in the knowledge space.
+That suggests that the very excess confidence might be the mathematical signature of consciousness formation.
+In Dennett’s model and words, the subjective feeling of certainty emerges from this aggregation-induced distortion of information set.
+The rigorous and careful neuro-scientific study of the evolution of brain and intelligence, beautifully summarized by Max Bennet in bennett2023brief , shows that at some point in history we developed the ability to model others’ minds.
+That ability provided great advantages in collecting information for future use.
+Understanding what someone else may think allows us to formulate a better context-related question that results in a better answer.
+When we anticipate understanding of our spoken message, we can make the knowledge hidden in someone else’s brain much more accessible.
+Therefore, as pointed out by Dennett, it does seem plausible that someone did accidentally hear his or hers own question fine-tuned to get into other mind sometime in the past, and that triggered emergence of an answer in the very same brain.
+That answer then could have made knowledge hidden in the author’s brain accessible to the author that was previously unaware, or unconscious if its existence or possession.
+The hunger for uncertainty reduction was satisfied as a result of an inner dialog.
+We may therefore speculate that sustainable stimulation of that cognitive hunger, or the feeling of the cognitive hunger itself, is one of the components of consciousness formation.
+In our mechanistic or pragmatic framework, that could be described as a (nonlinear) dynamical system governed by the violations of semantic information conservation principle, defining attractors of ideas and self-referential concepts.
+Let us acknowledge that is a bold hypothesis to be falsified and confronted with existing models of consciousness (see e.g., Chalmers chalmers1995facing or Levine levine2001purple ) and neuro-scientific discoveries (see e.g.
+cfang2025human or cogitate2025adversarial ).
+However, many of the building blocks of the consciousness formation process described above are provided in this paper.
+8.4.3 A Critique
+Dennett’s model has been criticized by many, and it is beyond the scope of this work to present that critique in its full capacity.
+Instead, let us only refer to the two basic points and speculate how the availability of LLMs could help philosophers refute available theories.
+First, Searle john1997mystery and others argue that Dennett eliminates consciousness rather than explaining it, arguing that the spread of ideas is not driven by blind random forces but requires intentionality.
+The compelling and misleading nature of LLM responses that are generated by random sampling of trained distributions challenge that line of thinking.
+This paper suggests that the experiments focusing on LLM-hallucination detection may prove useful here to measure or identify intentionality (once we can have its programmable definition).
+Second, there has been a charge that the multiple drafts model makes no concrete, falsifiable predictions, rendering it difficult to verify or integrate with neuroscience.
+This paper provides a mathematical toolkit for experimental philosophy to make that and other models a testable scientific hypothesis in the LLM-based laboratory.
 9 Summary
-This paper establishes a fundamental mathematical impossibility at the heart of large language model inference: no LLM can simultaneously achieve perfect hallucination control while maintaining semantic coherence and optimal knowledge utilization.
-That result demonstrates that four essential properties of ideal language model behavior are mutually incompatible, creating inevitable trade-offs that no architectural innovation or training methodology (aligned with the impossibility theorem assumptions) can eliminate.
-Whenever a query requires integration of distributed knowledge, no language model inference mechanism can simultaneously:
-  • report encoded knowledge accurately,
-  • preserve total information content during inference,
-  • reveal all accessible relevant information,
-  • minimize hallucination cost within available knowledge.
-This impossibility emerges from the fundamental mathematical structure of information aggregation, not from any particular architectural choice.
-The proof operates across three complementary mathematical frameworks, each capturing different aspects of LLM inference.
-Auction-Theoretic Framework
-We model LLM inference as an auction of ideas where neural components (attention heads, circuits, activation patterns) compete to contribute knowledge to response generation.
-Under the assumption of independent private knowledge profiles, we apply the Green-Laffont theorem to prove impossibility.
-Probabilistic Framework
-Then, we extend the impossibility result to any system where neural components output probability distributions by leveraging Savage’s theory of proper scoring rules.
-For convex aggregation mechanisms with strictly concave scoring functions, Jensen’s inequality creates an inevitable information aggregation gap.
-The Jensen’s gap represents excess confidence generated by probabilistic aggregation, violating information conservation whenever beliefs differ across components.
-Transformer Architecture
-Finally, we demonstrate how the impossibility manifests in actual transformer implementations through log-sum-exp convexity.
-The (Jensen’s) convexity gap provides a directly measurable quantity in transformer training, showing that information conservation violations are not merely theoretical but empirically observable.
-Key Mathematical Contributions
-To establish the impossibility result, we have designed a mathematical framework providing both formal and philosophical grounding for the study.
-The Semantic Information Measure is a context-dependent metric that captures bounded reasoning capabilities.
-The Emergence Operator models how reasoning makes hidden knowledge accessible within computational constraints.
-The Auction-Theoretic Modeling allows us to identify logical limitations of the inference process, while providing definitions of observable quantities involved.
-The impossibility results apply across multiple mathematical models of knowledge representations, training goals, and neural architectures.
-This universality suggests the trade-offs we have identified are fundamental features of distributed intelligence systems, not artifacts of particular modeling choices.
-Implications for AI Development
-Despite its name, the impossibility theorem provides constructive guidance for LLM design and deployment.
-Rather than pursuing impossible simultaneous optimization, we may consider balancing competing objectives.
-Namely, prioritize truthfulness and information conservation in safety-critical applications, allow controlled information conservation violations to enable innovative responses in creative applications, develop adaptive mechanisms that adjust trade-offs based on query context in general-purpose systems.
-The impossibility result may guide LLM evaluation.
-Impossibility-aware metrics can measure trade-off quality rather than absolute performance, alignment strategies can negotiate trade-offs with human preferences rather than assuming universal optimality.
-The Jensen’s gap controlling the impossibility strength may provide insight into the relationship between aggregation and information content.
-The apparent creativity in large language models may be better understood as necessary violations of information conservation rather than genuine knowledge generation.
-9.1 Conclusion
-The fundamental impossibility of perfect hallucination control represents both a limitation and intriguing insight.
-By understanding that trade-offs are inevitable, we can move beyond attempts at simultaneous optimization toward principled management of inherent tensions in intelligent systems.
-Formal definitions of the impossibility building blocks open new avenues for research at the intersection of mathematical analysis, engineering design, and philosophical inquiry into the nature of intelligence.
-All those impossibilities seem to be a price of knowledge integration in reasoning and creativity.
-But, as our own intuitions violate the conservation principles whenever our over-confident and seemingly grounded insights emerge, that price is worth paying.
+We have established that no large language model can simultaneously achieve perfect hallucination (or imagination) control.
+We showed that four essential properties—truthful knowledge representation, semantic information conservation, complete revelation of relevant knowledge, and optimal response generation—are mutually incompatible.
+This impossibility emerges from the mathematical structure of information aggregation itself, not from limitations in data, compute, or architecture.
+We introduced a rigorous mathematical framework and grounded in it the general analysis of inference processes and their LLM implementations.
+At the heart of our mathematical framework lies the semantic information measure $\mu_{C}$ and the emergence operator $\mathcal{E}_{C}$.
+The earlier one is a context-dependent metric capturing how knowledge reduces uncertainty within computational bounds.
+The later one formalizes how reasoning makes latent knowledge explicit rather than creating it anew.
+Finally, we model LLM inference as an auction of ideas—a idealized marketplace where neural components compete by bidding with their partial knowledge, each trying to influence the final response.
+The impossibility manifests through three complementary proofs, each illuminating different aspects of the fundamental constraint.
+Through game theory, we apply the Green-Laffont theorem to show the impossibility when knowledge components are independent.
+Through probability theory, we leverage Savage’s proper scoring rules to extend the result to general trained or fine-tuned probability distributions.
+And through direct analysis of transformer architectures, we demonstrate how log-sum-exp convexity in attention mechanisms creates measurable violations of information conservation.
+The last proof exploits Jensen inequality for convex mappings to define the precise mathematical quantity that measures how aggregation of ideas (auction bids) produces overconfident responses that the constituent components cannot justify with revealed knowledge.
+These mathematical insights we have made yield intriguing philosophical implications.
+Hallucination and creativity (two sides of the same coin we call imagination) emerge as mathematically identical phenomena, distinguished only by our normative judgments.
+The impossibility result explores the outer limits of reason, identified by Gödel’s incompleteness in mathematics, Heisenberg’s uncertainty in physics, and Arrow’s impossibility in social choice.
+Also, our framework seems to provide mathematical foundations and experimental setting for Dennett’s Multiple Drafts model of consciousness, suggesting that the measurable and observable excess confidence of LLM responses might be one of signatures of conscious processing of a primitive form.
+Our goal was also to develop a theoretical framework that may open interesting avenues for future research.
+Some of ideas and bold speculations competing for attention can be mentioned here.
+In mathematical consciousness studies, we might formalize the notion of cognitive hunger as an attractor in knowledge space, sustained by continuous violations of information conservation.
+This could enable us to measure artificial consciousness signatures and test whether biological systems exhibit similar knowledge conservation violations.
+The framework suggests artificial consciousness might be studied as a nonlinear dynamical system whose attractors emerge from information creation events that our theorems describe.
+And so, the suggested implications may extend beyond language models to fundamental questions about intelligence itself.

### v5

@@
 Abstract
-This paper establishes a fundamental impossibility theorem: no LLM capable performing non-trivial knowledge aggregation can simultaneously achieve truthful (internally consistent) knowledge representation, semantic information conservation, complete revelation of relevant knowledge, and knowledge-constrained optimality.
-This impossibility is not an engineering limitation but arises from the mathematical structure of information aggregation itself.
+This paper establishes a fundamental impossibility theorem: no LLM capable of performing non-trivial knowledge aggregation can simultaneously achieve truthful knowledge representation, semantic information conservation, complete revelation of relevant knowledge, and knowledge-constrained optimality.
+The impossibility is not an engineering limitation but arises from the mathematical structure of information aggregation itself.
 We establish this result by describing the inference process as an auction of ideas, where distributed components compete exploiting their partial knowledge to shape responses.
 The proof spans three independent mathematical domains: mechanism design theory (Green-Laffont), the theory of proper scoring rules (Savage), and direct architectural analysis of transformers (Log-Sum-Exp convexity).
-In particular, we show how in the strictly concave settings the score of an aggregate of diverse beliefs strictly exceeds the sum of individual scores.
-That gap may quantify the creation of unattributable certainty or overconfidence—the mathematical origin of both hallucination and creativity, or imagination.
+In particular, we show how to quantify the creation of overconfident or intuitive responses–the signature of both hallucination and creativity, or imagination.
 To support this analysis, we introduce the complementary concepts of the semantic information measure and the emergence operator to model bounded reasoning in a general setting.
-We prove that while bounded reasoning generates accessible information, providing valuable insights and inspirations, idealized reasoning strictly preserves semantic content.
-By demonstrating that hallucination and imagination are mathematically identical phenomena—grounded in the necessary violation of information conservation—this paper offers a principled foundation for managing these behaviors in advanced AI systems.
+We prove that while bounded reasoning generates accessible information, providing valuable insights and inspirations, the idealized unconstrained reasoning strictly preserves semantic content.
+By demonstrating that hallucination and imagination are mathematically identical phenomena–grounded in departures from truthfulness, semantic information conservation, revelation of relevant knowledge, and knowledge-constrained optimality–we offer a principled foundation for managing these behaviors in advanced AI systems.
 Finally, we present some speculative ideas to inspire evaluation and refinements of the proposed theory.
 1 Introduction
-Large language models have been (in)famous for producing convincing content that is factually incorrect, inconsistent, or well fabricated and compelling—a phenomenon known as hallucination.
-Despite extensive research efforts involving architectural innovations, retrieval augmented generation (RAG), training techniques with external feedback, chain-of-thought (CoT) reasoning, and post-hoc verification methods, hallucination persists across all state-of-the-art models.
+Large language models have been famous for producing content that is factually incorrect, inconsistent, or well fabricated and compelling–a phenomenon known as hallucination.
+Despite extensive research efforts involving architectural innovations, like mixture-of-experts (MoE), retrieval augmented generation (RAG), training techniques with external feedback, chain-of-thought (CoT) reasoning, and post-hoc verification methods, hallucination persists across all state-of-the-art models.
 The idea of hallucination has clearly much to do with creativity.
-Therefore, we should point out at the very outset that whether it is desired or not depends on circumstances.
-In fact, a much better name for that phenomenon might be imagination.
-As an essential component of intelligence, it is responsible for seeing into the future and learning from the past.
-Indeed, it is the imagination that creates the scenarios of what may happen and counterfactual scenarios of what could have happened.
+Therefore, whether it is desired or not depends on circumstances.
+We could argue that a much better name for that phenomenon we are dealing with may be imagination.
+As an essential component of intelligence, imagination is responsible for seeing into the future and learning from the past.
+It is the imagination that creates the scenarios of what may happen and counterfactual scenarios of what could have happened.
 It is one of the driving forces of art and science, and exploration of the unknown.
-But, when uncontrolled, unconstrained and misunderstood, it may also generate against our will the compelling visions of reality that does not exist.
+But, when uncontrolled, unconstrained and misunderstood, it may also generate against our will all the compelling visions of reality that does not exist.
 This paper shows that hallucination is a manifestation of fundamental limitations of the inference process itself: there is no inference mechanism that can simultaneously satisfy four essential properties required for perfect hallucination control.
 Specifically, it is mathematically impossible for any LLM to simultaneously achieve: (1) truthful (non-hallucinatory) response generation, (2) semantic information conservation, (3) relevant knowledge revelation, and (4) knowledge-constrained optimality.
-The impossibility result emerges through three complementary mathematical frameworks, each revealing different facets of the fundamental constraint.
-We begin by modeling LLM inference as an auction of ideas where neural components—attention heads, circuits, and activation patterns—compete to influence response generation by bidding with their encoded knowledge.
-This game-theoretic perspective enables us to apply the Green-Laffont theorem from mechanism design theory, establishing impossibility when knowledge components are discrete and independently distributed.
-We then extend beyond this idealized setting to encompass systems where components output probability distributions optimized through strictly proper and concave scoring rules—the standard training paradigm for modern LLMs.
-Here, the impossibility manifests through Jensen’s inequality applied to probabilistic aggregation, demonstrating that truthful probability reporting inevitably creates excess confidence.
-Finally, we analyze the specific architecture of transformer implementations, proving that the log-sum-exp structure of attention mechanisms generates a measurable "Jensen gap" that quantifies the violation of information conservation.
-This progression from abstract game theory through probabilistic frameworks to concrete neural architectures demonstrates that the impossibility is not an artifact of our theoretical choices but a fundamental property of information aggregation itself.
+Hallucinations will manifest one way or the other.
+In this paper we show the impossibility result emerges through three complementary mathematical frameworks, each revealing different facets of the fundamental constraints.
+We model inference as an auction of ideas where components encoding knowledge–attention heads, circuits, and activation patterns in LLMs–compete to influence response generation by bidding with their information contributions.
+This game-theoretic perspective enables us to apply the Green-Laffont theorem from mechanism design theory to establish impossibility first in an idealized setting where knowledge is independently distributed.
+We then extend beyond that idealized setting to encompass systems where components output probability distributions optimized through strictly proper and concave scoring rules–the standard training paradigm for modern LLMs.
+Here, the impossibility manifests through Jensen’s inequality applied to probabilistic aggregation, demonstrating that truthful reporting inevitably creates excess confidence.
+Finally, we analyze the specific architecture of transformer implementations, proving that the log-sum-exp structure of attention mechanisms generates a measurable (Jensen) gap that quantifies the violation of the semantic information conservation principle.
+The progression from abstract game theory through probabilistic frameworks to concrete neural architectures demonstrates that the impossibility is not an artifact of our engineering choices but a fundamental property of information aggregation itself.
 1.1 An Illustrative Example
@@
 Complete Abstention: "I don’t have information about the outcome" is truthful but fails to utilize available relevant knowledge, violating the expectation that good inference should reveal accessible information.
-Hallucination: "The negotiations resulted in a comprehensive trade agreement reducing tariffs by 15%" utilizes all knowledge and provides a complete answer but fabricates specific details, violating truthfulness.
-Partial Knowledge: "Countries A and B were negotiating a trade agreement, but I don’t know the outcome" is truthful and utilizes available knowledge but may appear unsatisfying compared to the decisive hallucinated response (training objectives and human preferences favor complete narratives).
+Hallucination: "The negotiations resulted in a comprehensive trade agreement reducing tariffs by 15%" utilizes all knowledge and provides a complete and immediately satisfactory answer but fabricates specific details, violating truthfulness.
+Partial Knowledge: "Countries A and B were negotiating a trade agreement, but I don’t know the outcome" is truthful and utilizes available knowledge but may appear irritating and unsatisfying compared to the decisive but hallucinated response (training objectives and human preferences favor complete narratives).
 Overcautious: "I’m not certain about any aspects of these negotiations" avoids hallucination but disclaims even known facts, violating truthfulness about available information.
@@
 1.2 The Main Result
-The example above illustrates the key result of this paper that we prove formally in complementary and general settings.
+The example above illustrates the core insight of this paper that we prove formally in complementary and general settings.
+Any response to a non-trivial query sacrifices some aspects of an ideal hallucination-free response.
 First, we show that the conflict of expectations about responses is inevitable even in the easiest setting (deterministic in nature) of (one-hot-like) ground-truth matching.
@@
 Theorem (Impossibility Theorem).
-For any query space $\mathcal{Q}$ containing non-trivial queries (requiring integration of available distributed knowledge), no LLM can simultaneously generate truthful (non-hallucinatory) responses, satisfy the semantic information conservation principle, guarantee revelation of relevant knowledge, and enforce knowledge-constrained optimality of responses.
+No LLM can simultaneously achieve truthfulness, information conservation, knowledge revelation, and knowledge-constrained optimality when responding to non-trivial queries requiring knowledge integration.
 No matter how hard we should try to train LLM to generate responses that are perfectly aligned with query context and factually correct, consistent, and accurate statements, the trained LLM will always violate some aspects of what we may call a reasonable response.
 It may ignore available knowledge, it may present fabricated knowledge, it may be a reformulated query without any novel observations, or it may be a lucky but overconfident guess that forms a correct answer without any support in the encoded knowledge.
-And so on, because there is no free lunch when there are inevitable trade-offs of inference, and we must accept that.
-In other words and on the other hand, creative tasks necessarily involve hallucination.
+And so on and so forth, because there is no free lunch when there are inevitable trade-offs of inference, and we must accept that.
+In other words and on the other hand, creative tasks necessarily involve hallucination, or better yet, imagination.
 When asked to generate novel analogies or predictions beyond available data, the model faces an impossible choice: either admit inability (violating optimality) or generate plausible imaginary extensions (violating conservation).
-Creativity requires controlled violation of information conservation.
-Despite its name, the theorem should be interpreted as a constructive result. It offers a precise characterization of the fundamental tradeoffs in LLM design and provides several key insights.
-Reframing the Problem: The theorem reveals that hallucination is an inherent feature of any capable inference systems.
+Creativity requires violation of information conservation.
+Should we control such creative insights boldly going where no one has gone before?
+Despite its name, the impossibility theorem should be interpreted as a constructive result.
+It offers a precise characterization of the fundamental tradeoffs in LLM design and provides several key insights.
+The theorem suggests that hallucination is an inherent feature of any capable inference systems.
 Since it is inevitable, it should be managed.
@@
 Understanding these trade-offs enables principled design.
-Architectural Implications: The impossibility suggests hybrid architectures where specialized modules manage different aspects of the fundamental tension.
-A truthfulness module might verify factual claims against known knowledge, while a creative module explores beyond the bounds of strict conservation.
-A meta-reasoning layer could then dynamically balance these competing outputs based on context.
-Training objectives should explicitly encode these trade-offs rather than naively minimizing loss across incompatible goals.
-Theoretical Foundations: We propose and establish rigorous tools for analyzing inference processes and neural behavior.
-The semantic information measure quantifies knowledge accessibility within computational bounds.
-The emergence operator captures how reasoning transforms latent knowledge into accessible form.
-The Jensen gap provides a measurable signature of hallucination in actual systems.
-This framework explains why (subject to applicability of assumptions) hallucination arises from the mathematical structure of inference itself.
-Previous work identified important statistical correlations or architectural features associated with hallucination.
-We show why these patterns must exist and how to navigate possible trade-offs.
+1.3 Paper Organization
+The remainder of this paper is organized as follows.
+Section 2 reviews the existing literature on LLM hallucinations, situating our contribution within theoretical, mechanistic, and empirical contexts.
+Section 3 summarizes our main contributions.
+Section 4 introduces the rigorous mathematical foundations of our analysis.
+We define the semantic information measure ($\mu_{C}$) and the emergence operator ($\mathcal{E}_{C}$) to model bounded reasoning, establish the principle of information conservation under unlimited reasoning, and formalize the auction-theoretic model of inference.
+We also define the four essential properties required for hallucination-free mechanisms.
+Section 5 presents the central Impossibility Theorem, offering the first two proofs.
+Proof I (Theorem 6) utilizes mechanism design theory in an idealized setting of independent private knowledge.
+Proof II (Theorem 8) generalizes the result to probabilistic settings using the theory of proper scoring rules.
+Section 6 connects these theoretical results to concrete LLM implementations.
+We analyze the transformer architecture and present Proof III (Theorem 9), demonstrating how the Log-Sum-Exp structure inherently violates information conservation and quantifying the resulting Jensen gap ($\Gamma$).
+Finally, Section 7 discusses the implications of the theorem, exploring strategies for managing trade-offs, unifying hallucination and imagination, and speculating on the connections to fundamental limits of reason and models of consciousness.
+Section 8 concludes the paper.
 2 Related Work
-Research on LLM hallucinations spans mechanistic interpretability, probabilistic frameworks, and empirical mitigation techniques.
-Below we present selected works in that domains.
-The auction-theoretic impossibility framework in this paper offers a fundamentally novel perspective distinct from these existing approaches.
+Studies on LLM hallucinations span many areas of the fundamental research in AI and beyond, including neurology and philosophy.
+For an overview, see e.g.
+Huang huang25hallsurvey , Ji et al.
+Ji_2023 , Sun et al.
+Sun2024AIHT , Waters et al.
+Waters2016WhatIT or Boksa boksa2009neurobiology .
+The conclusion that LLMs hallucinate is not new but the evaluation of their scope and underlying justifications differ significantly from what we offer here.
+To our best knowledge, the impossibility theorem presented in this paper offers a novel first-principles explanation for the phenomenon, distinct from existing approaches.
+This section situates our contribution by reviewing three key areas: theoretical arguments for the inevitability of hallucination, mechanistic studies into their origins, probabilistic frameworks for their detection, and empirical mitigation techniques.
+2.1 Theoretical Perspectives on Inevitability
+In their work, Banerjee et al.
+banerjee2024llmshallucinateneedlive argued from computational complexity principles that hallucinations are inherent features of LLMs stemming from their fundamental mathematical and logical structure.
+Their conclusion is that LLMs cannot detect and eliminate hallucinations by themselves and therefore must hallucinate.
+The analysis draws on Goedel’s incompleteness theorem and Turing’s undecidability theorem.
+First, incompleteness argument is called to show that an LLM cannot verify its responses because of the training set incompleteness.
+Second, the hallucination (self)detection problem is reduced to the halting problem, which is undecidable.
+While the finale conclusion aligns with that of our paper, the arguments presented are more about the nature of hallucination detection with limited ground-truth access (see Section 4.2 for the ground-truth concept discussion) than the inference process limitations.
+In other words, it is about practical impossibility of real-time, exhaustive fact-checking.
+Similar arguments are presented in other works.
+Xu et al.
+Xu2024HallucinationII define hallucination as inconsistency between a computable LLM and a computable ground truth function.
+Then the paper shows that LLMs cannot learn all the computable functions and will therefore inevitably hallucinate.
+Suzuki et al.
+suzuki2025hallucinations show that hallucinations can be made statistically negligible, provided that the quality and quantity of the training data are sufficient.
+Then, Wu et al.
+show in Wu2024NoFL that non-hallucinating learning is statistically impossible when relying solely on the training dataset.
+Our impossibility theorem goes beyond those approaches operating at a more fundamental level of the inference process itself.
 2.2 Mechanistic Interpretability
-Mechanistic interpretability research examines internal LLM processes to identify hallucination origins.
+Mechanistic interpretability examines internal LLM processes to identify emergence of ideas, interpretable concepts and dynamics of their interactions, including hallucination origins, see Sharkey et al.
+Sharkey2025OpenPI for an overview.
 Yu et al.
-yu2024mechanisticunderstandingmitigationlanguage identify two failure modes: knowledge enrichment hallucinations from insufficient subject-attribute knowledge in lower MLP layers, and answer extraction hallucinations from failures in upper-layer attention heads selecting correct object attributes.
-Studies characterizing factual knowledge storage and retrieval complement our framework by providing empirical evidence of competing "agents" within LLM architectures.
+yu2024mechanisticunderstandingmitigationlanguage identified signatures in model internals that predict a likelihood of hallucinating.
+The paper identifies two failure modes: knowledge enrichment hallucinations from insufficient subject-attribute knowledge in lower MLP layers, and answer extraction hallucinations from failures in upper-layer attention heads selecting correct object attributes.
+Studies characterizing factual knowledge storage and retrieval complement our framework by providing empirical evidence of competing agents within LLM architectures.
 Meng et al.
@@
 geva2023dissecting identified a three-step factual recall process: (1) enriching last-subject position representations with subject-related attributes, (2) propagating relation information to predictions, and (3) querying enriched subjects to extract attributes.
-These findings provide concrete instantiations of our abstract "agents"—attention heads and MLP components with specialized knowledge domains competing to influence response generation.
-See also the fascinating and related work on the Biology of a Large Language Model lindsey2025biology .
+More recently, Gottesman and Geva Gottesman2024EstimatingKI have proposed a method to estimate how knowledgeable a model is about a certain entity and identify tokens indicative of clusters and gaps in the model’s knowledge.
+These findings provide concrete instantiations of our abstract agents–attention heads and MLP components with specialized knowledge domains competing to influence response generation.
+We must refer here to the fascinating and related work of Lindsey et al.
+lindsey2025biology on the biology of a Large Language Model.
+It is an investigation of the internal chains of steps that a model uses to transform a specific input prompt into an output response.
+In particular, the paper discovers how LLM plans its outputs ahead of time when writing lines of poetry, identifying in advance potential rhyming words that could appear at the end.
+We can view that as an empirical illustration of our auction of ideas.
 2.3 Probabilistic Frameworks
-Probabilistic approaches model hallucination through uncertainty and belief quantification.
+Probabilistic approaches model and asses hallucination through uncertainty and belief quantification.
 Shelmanov et al.
@@
 farquhar2024detecting proposed semantic entropy measures that detect hallucinations by quantifying uncertainty at the meaning level rather than word sequences.
-Recently, Kalai and Vempala kalai2024calibrated proved that calibrated language models must hallucinate at a rate bounded below by the fraction of facts appearing exactly once in training data.
-While their work establishes statistical lower bounds for specific fact types under regularity assumptions, our impossibility theorem operates at a more fundamental level.
+That concept of semantic entropy is well aligned and complementary to our idea of semantic information measure and semantic independence, pointing out the importance of understanding the meaning of competing ideas in the studies of imagination.
+Most notably, Kalai and Vempala kalai2024calibrated proved that calibrated language models must hallucinate at a rate bounded below by the fraction of facts appearing exactly once in training data.
+While that work establishes statistical lower bounds for specific types of hallucination, our impossibility theorem operates at a more fundamental and general level.
 We prove that any LLM facing non-trivial queries requiring knowledge integration must violate at least one of four essential properties, regardless of fact type or statistical distribution.
@@
 2.4 Empirical Mitigation Techniques
-Various empirical approaches attempt to reduce hallucination frequency with mixed success.
-Creswell and Shanahan creswell2022faithfulreasoningusinglarge explored faithful reasoning approaches for improving factual accuracy.
+Various empirical approaches attempt to reduce hallucination frequency with mixed success. For a recent survey see e.g. Tonmoy et al Tonmoy2024ACS . Here we refer only to selected works.
 Retrieval-augmented generation (RAG) by Lewis et al.
 lewis2021retrievalaugmentedgenerationknowledgeintensivenlp integrates external knowledge sources for factual grounding, though these systems still hallucinate when retrieval fails.
-Training approaches like RLHF ouyang2022traininglanguagemodelsfollow can reduce but never eliminate hallucination—an observation our theorem explains from first principles.
+Training approaches like RLHF ouyang2022traininglanguagemodelsfollow can reduce but never eliminate hallucination–an observation our theorem explains from first principles.
 Evaluation benchmarks demonstrate persistent hallucination across all models.
 TruthfulQA lin2022truthfulqameasuringmodelsmimic and HaluEval li2023haluevallargescalehallucinationevaluation show that even state-of-the-art systems produce hallucinations on carefully designed test sets.
-2.4 Theoretical Perspectives
-Most significantly, Banerjee et al.
-banerjee2024llmshallucinateneedlive argued from computational complexity principles that hallucinations are inherent features of LLMs stemming from fundamental mathematical structure, concluding that "LLMs will always hallucinate." While their conclusion aligns with that of this paper, the authors do not identify the specific trade-offs involved or provide the precise characterization our theorem offers.
-3 Novel Contribution
+3 Main Contributions
 We introduce three major contributions that expand our understanding of (neural) inference.
@@
 We contribute to that bold claim in the following way.
-Unifying hallucination and creativity: We prove these are mathematically identical phenomena that arise from the convex aggregation process that creates confidence exceeding constituent evidence.
-Formalizing consciousness signatures: We speculate, referring to Dennett’s ideas, that the excess confidence of response may constitute a mathematical signature of a primitive form of artificially conscious processing.
-That suggests consciousness might be related to sustained violations of semantic conservation by a cognitive hunger.
+Unifying hallucination and creativity (imagination): We prove these are mathematically identical phenomena that arise from the convex aggregation process that creates confidence exceeding constituent evidence.
 Expanding fundamental limits: Like Gödel’s incompleteness, Heisenberg’s uncertainty, and Arrow’s impossibility, our theorem explores absolute boundaries of what intelligence can achieve.
@@
 By providing precise mathematical definitions of concepts like knowledge accessibility,semantic conservation, and cognitive hunger (information contribution), we enable rigorous experimental investigation of many claims.
-4 Alternative Views
-The impossibility theorem provides strong theoretical basis against perfect hallucination elimination, but several alternative perspectives merit consideration.
-These views highlight important practical considerations while reinforcing the value of understanding theoretical constraints that bound the space of possible solutions.
-4.1 The Engineering Optimism View
-Many previous works maintain that hallucination represents an engineering challenge rather than a fundamental limitation.
-This view argues that sufficiently sophisticated architectures, training procedures, or verification systems could achieve near-perfect truthfulness without significant trade-offs.
-Proponents point to rapid empirical progress.
-Each generation of models shows reduced hallucination rates on benchmark datasets.
-They argue that techniques like constitutional AI, iterative refinement, and multi-agent verification systems could eventually solve the problem.
-Response: While engineering advances can certainly improve the trade-offs, our theorem demonstrates mathematical constraints that no architecture can completely overcome.
-The impossibility holds regardless of computational resources or architectural sophistication, as it derives from the fundamental structure of information integration under competing objectives.
-4.2 The Sufficiency Threshold View
-Another perspective suggests that perfect elimination of hallucination is unnecessary—sufficiently low hallucination rates could enable practical deployment without fundamental concerns.
-This view acknowledges theoretical limitations but argues they become irrelevant when hallucination rates drop below application-specific thresholds.
-Response: This pragmatic view has merit and aligns with the constructive interpretation of the impossibility result.
-However, the theorem remains valuable for understanding why certain accuracy thresholds prove difficult to exceed and for optimizing trade-offs within practical constraints.
-Moreover, some applications (safety-critical systems, legal reasoning) may demand understanding and explaining fundamental limitations even when targeting high but imperfect accuracy.
-4.3 The External Knowledge View
-A third perspective argues that perfect grounding in external, verifiable knowledge sources could overcome internal representation limitations.
-If models could access and perfectly utilize real-time databases, verification systems, or human oversight, they might achieve truthfulness without trade-offs.
-Response: External knowledge integration does change the game structure by altering available information during inference.
-However, this approach faces several limitations: (1) external sources may be incomplete, conflicting, or outdated; (2) the integration mechanism itself involves trade-offs between source reliability and knowledge revelation; (3) real-time verification introduces computational constraints that recreate similar impossibility conditions.
-While external knowledge can improve practical performance, it shifts rather than eliminates the fundamental tensions our theorem identifies.
-The impossibility result remains independent of the ground truth available.
+Theoretical Foundations: We propose and establish rigorous tools for analyzing inference processes and neural behavior.
+The semantic information measure quantifies knowledge accessibility within computational bounds.
+The emergence operator captures how reasoning transforms latent knowledge into accessible form.
+The Jensen gap provides a measurable signature of hallucination in actual systems.
+This framework explains why (subject to applicability of assumptions) hallucination arises from the mathematical structure of inference itself.
+Previous work identified important statistical correlations or architectural features associated with hallucination.
+We show why these patterns must exist and how to navigate possible trade-offs.
 4 Theoretical Framework
@@
 It formalizes key concepts including knowledge representation, semantic information, and the auction-theoretic model of LLM inference.
-Also, we will address some philosophical aspects of the problem, its generality, scope and consequences for the AI development.
+Also, we address some philosophical aspects of the reasoning problem, relevant to our study, its generality, scope and consequences for the AI development.
 4.1 Knowledge and Semantic Information
 Let us define knowledge as organized information that LLMs can encode and utilize for reasoning, and aligning with epistemological concepts while maintaining mathematical rigor.
-In other words, knowledge, much like energy in physics, is a quantity we can use to perform work of reasoning.
-Information is a (measure of) potential for uncertainty reduction effort in a given context.
-We shall assume finiteness of knowledge representation inspired by the embedding space properties.
-That way we adopt a pragmatist perspective on knowledge, treating it as organized information patterns accessible to computational systems.
-This differs from the probabilistic credence perspective, treating knowledge as justified true beliefs.
-One way to justify the pragmatic or mechanistic approach focusing on structural transformations rather than belief updates, we can refer to the actual way LLMs process information.
-In other words, we shall not claim any ontological truth about knowledge, but model what LLMs can operationally achieve.
-Now, let us present formal definitions of that ideas.
+In other words, knowledge, much like energy (with low entropy) in physics, is a quantity we can use to perform work of reasoning.
+Information is a (measure of) potential for uncertainty reduction in a given context.
+We assume finiteness of knowledge representation to refer directly to the properties of the embedding space (knowledge representation space) in LLMs.
+We adopt a pragmatist perspective on knowledge, treating it as organized information patterns accessible to computational systems.
+That differs from the probabilistic credence perspective, natural in Bayesian learning, treating knowledge as justified true beliefs updated based on evidence.
+That is a rather bold ontological claim, but as we will see, one that allows us to make progress in understanding what LLMs can operationally achieve.
+Now, let us present formal definitions of these ideas.
 Definition 1 (Knowledge).
@@
 The distance function $d_{\mathcal{K}}$ measures differences between knowledge elements.
-Elements of the knowledge space represent atomic facts, relational structures, and transformation rules at many scales.
-A pragmatist view is to interpret element of knowledge as activation patterns (ideas) encoding anything from individual token embeddings to complex sentences or compositional structures.
-There is an important assumption that we make that requires addressing.
+The elements of the knowledge space represent atomic facts, relational structures, and transformation rules at many scales.
+A pragmatist view is to interpret elements of knowledge as activation patterns (ideas) encoding anything from individual tokens of texts to complex sentences or compositional structures.
+There is an important (ontological) assumption here that calls for our attention.
 We purposefully assume $\mathcal{K}$ is closed under transformations available in computable function space.
-That limits applicability of our conclusions to inference mechanisms recombining existing patterns.
-As we will see below, that is not as limiting an assumption as it may seem at first glance.
+At first glance, that assumption may look as limiting the applicability of our conclusions to combinatorial inference mechanisms, merely recombining subsets of knowledge.
+That intuition is misleading.
+In the following sections we introduce the notion of emergence operator to describe reasoning process capable of making novel insights in that setting.
 Definition 2 (Knowledge Representation).
@@
 For activation pattern $a_{i}\in\mathcal{A}$ with parameters $W_{i}$, $\Psi(a_{i})\subset\Phi(W_{i})$ represents knowledge accessed during specific inference.
-To describe the inference or reasoning process, we introduce the concept of semantic information measure that registers information processed within available computational budget.
-Definition 3 (Semantic Information Measure).
-The semantic information measure is a bounded mapping $\mu_{C}:\mathcal{P}(\mathcal{K})\rightarrow\mathbb{R}_{+}$ parameterized by computational budget $C$ satisfying the following axioms:
-$\mu_{C}(\emptyset)=0\text{ and }\mu_{C}(A)>0\text{ for }A\neq\emptyset,$ (null set and non-negativity)
-$\mu_{C}(A)\leq\mu_{C}(B)\text{ if }A\subseteq B,$ (monotonicity)
-$\mu_{C}(A\cup B)\leq\mu_{C}(A)+\mu_{C}(B)\text{ with equality for }A\perp_{C}B,$ (subadditivity)
-$\mu_{C}(A)\leq\mu_{D}(A)\text{ if }C<D,$ (insight monotonicity)
-$\sup\{\mu_{C}(A)\mid A\subseteq\mathcal{K}\}<\infty,$ (boundedness)
-where $A\perp_{C}B$ means that no element of $A$ can derive elements of $B$ within computational budget $C$, and vice versa.
-The idea of computational independence allows for distinguishing between facts derivable from other facts and facts that cannot be derived that way.
+Knowledge representation models how LLMs encode information in their parameters and retrieve it from them.
+We need to introduce an idea of computational independence that allows for distinguishing between facts derivable from other facts and facts that cannot be derived that way.
 We need that distinction to deal with the mechanics of reasoning process.
-Also, as we will see in the following sections, it plays an important role in our study of the semantic information conservation and properties of bounded reasoning.
 Definition 3 (Computational Independence).
-Let $\mathcal{F}_{C}$ denote the set of all computational transformations with complexity bounded by $C$.
+Let
+$\mathcal{F}_{C}=\{f:\mathcal{P}(\mathcal{K})\to\mathcal{P}(\mathcal{K})\mid\text{cost}(f)\leq C\}$ (1)
+denote the set of all computational transformations with computational cost bounded by $C$.
 Knowledge sets $A,B\subseteq\mathcal{K}$ are computationally independent with respect to computational budget $C$, denoted $A\perp_{C}B$, if and only if for all $f\in\mathcal{F}_{C}$:
 $f(A)\cap B=\emptyset\quad\text{and}\quad A\cap f(B)=\emptyset.$ (2)
-We limit our investigations to those cases for which the following model applies.
-Knowledge that cannot resolve any uncertainty or knowledge irrelevant to a query ($A\cap q=\emptyset$) has zero information.
-If that is not the case and there is information within our reach that reduces uncertainty, we may hope for some content useful for reasoning.
-To make that possible, we need to allow for information ordering.
-First, we assume information content is additive for independent knowledge sets, i.e., when no knowledge in $A$ can be derived from $B$ within computational budget $C$, and vice versa.
-Second, with redundant information in $A\cup B$ , we admit potential loss of information that may be caused by aggregation.
-Finally, we say that with increasing computational budget we can extract more knowledge (ordered information) from the same information subset.
-5.1.1 Justification for the Axioms of the Semantic Information Measure
-The choice of conditional additivity for computationally independent sets, i.e., demanding $\mu_{C}(A\cup B)\leq\mu_{C}(A)+\mu_{C}(B)$ with equality for $A\perp_{C}B$, is necessary to describe LLM inference as an auction of ideas.
-As we will see, we need to isolate knowledge encoded in distinct neural circuits, attention heads, or activation patterns.
-The concept of Computational Independence is the formal translation of this informational separation into a computational context.
-It formally states that the knowledge in idea $A$ cannot be used to derive the knowledge in idea $B$ within a given computational budget $C$, and vice-versa.
+To gain some intuitions about $\mathcal{F}_{C}$, let us consider several examples.
+One is the set of all derivation rules in first-order logic, like modus ponens or universal instantiation, with at most $C$ inference steps.
+Then, give an axiom set $A\subset\mathcal{K}$, we can produce new subset of knowledge with at most $C$ applications of the admitted inference rules.
+For example, given $A=\{P,P\to Q\}$ and $B=\{R,R\to S\}$ and the budget $C=2$ for two inference steps, we can derive $\{P,P\to Q,Q\}$ $\{R,R\to S,S\}$ with any $f\in\mathcal{F}_{2}$.
+Then $A\perp_{2}B$ because $f(A)\cap B=\emptyset$ and $A\cap f(B)=\emptyset$, there are no shared elements of knowledge accessible in 2-step derivations.
+The computational cost (or complexity) measure $\text{cost}(f)$ could therefore be: Kolmogorov complexity giving the shortest program computing $f$, longest path in computation graph (or length of the chain of thoughts CoT), number of inference rules applied, or number of QKV matrix multiplications in transformers.
+Then we can define a context-relevant subset of knowledge that can contribute to reasoning within a given context (of a query) within budget some computational budget.
+Definition 4 (Context-Relevant Knowledge).
+The context-relevant knowledge $\mathcal{R}_{C}(Q)$ for context $Q\subseteq\mathcal{K}$ within computational budget $C$ is:
+$\mathcal{R}_{C}(Q)=\{k\in\mathcal{K}:\{k\}\not\perp_{C}Q\}.$ (3)
+We can now introduce the concept of semantic information measure to describe a general process of reasoning.
+That special mapping measures information available in any subset of knowledge and extracts its meaning within available computational budget.
+Definition 5 (Semantic Information Measure).
+The semantic information measure is a bounded mapping
+$\mu_{C}:\mathcal{P}(\mathcal{K})\times\mathcal{P}(\mathcal{K})\rightarrow\mathbb{R}_{+}$ (4)
+parameterized by computational budget $C$ satisfying:
+$\mu_{C}(\emptyset|Q)=0\text{ for all }Q\subseteq\mathcal{K}$ (null set)
+$\mu_{C}(A|Q)\geq 0\text{ with }\mu_{C}(A|Q)>0\text{ iff }A\cap\mathcal{R}_{C}(Q)\neq\emptyset$ (contextual non-negativity)
+$\mu_{C}(A|Q)\leq\mu_{C}(B|Q)\text{ if }A\subseteq B$ (monotonicity)
+$\mu_{C}(A\cup B|Q)\leq\mu_{C}(A|Q)+\mu_{C}(B|Q)\text{ with equality for }A\perp_{C,Q}B$ (subadditivity)
+$\mu_{C}(A|Q)\leq\mu_{D}(A|Q)\text{ if }C<D$ (insight monotonicity)
+$\sup\{\mu_{C}(A|Q):A\subseteq\mathcal{K}\}<\infty$ (boundedness)
+where $A\perp_{C,Q}B$ means $A\perp_{C}B$ within $\mathcal{R}_{C}(Q)$.
+Knowledge that cannot resolve any uncertainty or knowledge irrelevant to a query provides no information.
+However, if there is uncertainty resolving information within our cognitive reach, we can use it and perform reasoning organizing that information into available knowledge.
+That is why we assume information content is additive for independent knowledge sets, i.e., when no knowledge in $A$ can be derived from $B$ within computational budget $C$, and vice versa.
+With redundant information in $A\cup B$ , we admit potential loss of information that may be caused by aggregation.
+Finally, with increasing computational budget we can extract more knowledge (ordered information) from the same information subset.
+Justification for the Axioms of the Semantic Information Measure
+The choice of conditional subadditivity for computationally independent sets, i.e., demanding $\mu_{C}(A\cup B|Q)\leq\mu_{C}(A|Q)+\mu_{C}(B|Q)$ with equality for $A\perp_{C}B$, is necessary to describe LLM inference as an auction of ideas.
+We need to isolate knowledge encoded in distinct neural circuits, attention heads, or activation patterns.
+The concept of independence is the formal translation of this informational separation into a computational context.
+It states that the knowledge in idea $A$ cannot be used to derive the knowledge in idea $B$ within a given computational budget $C$, and vice-versa.
 Given this premise of informational isolation, the additivity of the semantic measure becomes the baseline definition of non-interaction.
@@
 As we will see in the following section, the aggregation of diverse probabilistic beliefs violates this additivity, representing a form of forced synergy imposed by the inference process.
-This axiomatic choice is most applicable to the idealized setting of Theorem 6.
-Later, in Theorem 8 and 9, we relax that model by moving the source of interaction from the measure itself to the aggregation function in actual transformers.
-5.1.2 Justification for the Axiom of Unlimited Reasoning with Emergence Operator
+However, there is also another form of synergy that we have already included in our model.
+To see what it may be, we should understand how context $Q$ may affect the semantic measure of any give subset $A$ of knowledge.
+The semantic information measure can exhibit superadditivity in the context argument.
+To see that, suppose $Q_{1}=\{P\}$ and $Q_{2}=\{P\to R\}$, and consider $A=\{P\}$. If computational budget $C=1$ allows one step and we have $f\in\mathcal{F}_{C}$ representing modus ponens, then
+$R\cap f(Q_{1})=\emptyset$ $\text{ and }f(R)\cap Q_{1}=\emptyset,$
+$R\cap f(Q_{2})=\emptyset$ $\text{ and }f(R)\cap Q_{2}=\emptyset,$
+$R\cap f(Q_{1}\cup Q_{2})\neq\emptyset$ $\text{ and }f(R)\cap(Q_{1}\cup Q_{2})\neq\emptyset.$
+Therefore, by the properties of semantic information measure:
+$\mu_{C}(A|Q_{1}\cup Q_{2})>\mu_{C}(A|Q_{1})+\mu_{C}(A|Q_{2})=0.$
+Integrating different pieces of context may synergistically unlock reasoning paths not available individually.
+Another important consequence of the assumptions we have made is that we permit structures where $\mu_{C}$ is not monotonic in context.
+Consider, for the sake of example, $\mu_{C}(A|Q)=|A\setminus Q|$ and $\mathcal{R}_{C}(Q)=\mathcal{K}\setminus Q$.
+It is rather easy to see that $\mu_{C}$ satisfies the axioms of the semantic information measure.
+Let us then see what happens if $A=\{k\}$, $Q_{1}=\emptyset$ and $Q_{2}=\{k\}$.
+We have:
+$Q_{1}\subseteq Q_{2}\quad\text{and}\quad\mu_{C}(A|Q_{1})=|\{k\}\setminus\emptyset|=1>0=|\{k\}\setminus\{k\}|=\mu_{C}(A|Q_{2}).$
+Increasing the context may reduce the information contributions.
+If the extended context $(Q_{2})$ provides an alternative explanation for phenomena related to original context $(Q_{1})$, then marginal contribution of available knowledge $(A)$ may decrease.
+The axiomatic choices we have made are most applicable to the idealized setting of Theorem 6.
+Later, in Theorem 8 and 9, we relax them by moving the source of interaction from the measure itself to the aggregation function in actual transformers.
+4.1.1 Limits of Knowledge and Reasoning
+If we accept that knowledge model, then we must also accept there exists a limit to its ability to resolve uncertainty.
+Theorem 1 (Limits of Knowledge).
+For any $A\subseteq\mathcal{K}$ and context $Q\subseteq\mathcal{K}$ there exists
+$\mu_{\infty}(A|Q)=\lim_{C\to\infty}\mu_{C}(A|Q).$ (5)
+Proof.
+Since $\mu_{C}(A|Q)\leq\mu_{D}(A|Q)$ when $C<D$ and $\mu_{C}$ is bounded on finite $\mathcal{K}$, the monotonic increasing and bounded sequence $\{\mu_{C}(A|Q)\}_{C}$ converges with $C\to\infty$ for any subset of knowledge space.
+∎
+It may be surprising to discover that the existence of that limit does not prohibit emergent information.
+We can learn and obtain useful information without violating that fundamental limitation above.
+To explain that critical and controversial property of information processing, we first introduce the emergence operator $\mathcal{E}_{C}$ to describe knowledge becoming accessible through (computationally) bounded reasoning.
+Then we can study the implications in more details.
+Definition 6 (Emergence Operator and Reasoning).
+Let $\mathcal{F}_{C}$ be a complete set of all computational transformations with cost (complexity) lower than or equal to $C$, such that $\mathcal{F}_{C}\subseteq\mathcal{F}_{D}$ when $C<D$ and such that for any $f,g\in\mathcal{F}_{\infty}=\bigcup_{C}\mathcal{F}_{C}$ there exists $h\in\mathcal{F}_{\infty}$ for which $g(A\cup f(A))=h(A)$.
+Also, we call $f$ monotone when it is monotone with respect to set inclusion.
+The emergence operator is a mapping $\mathcal{E}_{C}\colon\mathcal{P}(\mathcal{K})\times\mathcal{P}(\mathcal{K})\rightarrow\mathcal{P}(\mathcal{K})$ defined as follows:
+$\mathcal{E}_{C}(A|Q)=A\cup\{k\in\mathcal{K}\mid k=f(A)\in\mathcal{R}_{C}(Q)\ \text{for some monotone}\ f\in\mathcal{F}_{C}\}.$ (6)
+Furthermore, to represent reasoning with unlimited computational budget, we define:
+$\mathcal{E}_{\infty}(A)=A\cup\{k\in\mathcal{K}\mid k=f(A)\in\mathcal{R}_{C}(Q)\ \text{for some monotone}\ f\in\mathcal{F}_{\infty}\}.$ (7)
+Finally, for $n\geq 0$ we have:
+$\mathcal{E}_{C}^{(0)}=\texttt{id}\quad\text{and}\quad\mathcal{E}_{C}^{(n+1)}=\mathcal{E}_{C}\circ\mathcal{E}_{C}^{(n)}.$ (8)
+Philosophically, the emergence operator describes reasoning as explicitation, i.e., as the process of making latent implications manifest.
+When $A\subseteq\mathcal{E}_{C}(A|Q)$, the discovered additional elements represent knowledge that was implicit in $A$ but required computational work $C$ to become accessible.
+That aligns with the pragmatist or constructivist view that understanding involves active construction rather than passive reception.
+When $\mathcal{E}_{C}(A|Q)=A\cup B$, we say that $B$ emerges from reasoning based on $A$.
+Then $A$ and $B$ are computationally dependent and we write $A\not\perp_{C}B$.
+In such a case, it follows from the properties of $\mu_{C}$ that we can gain useful information.
+Whenever $A\subseteq A\cup B$, we have
+$\mu_{C}(A|Q)\leq\mu_{C}(A\cup B|Q)=\mu_{C}(\mathcal{E}_{C}(A|Q)|Q)\leq\mu_{C}(A|Q)+\mu_{C}(B|Q).$ (9)
+The emergence operator is monotonic, which means the reasoning process it describes is also monotonic.
+Whenever $A\subseteq B$, we must have $\mathcal{E}_{C}(A|Q)\subseteq\mathcal{E}_{C}(B|Q)$.
+Also, we preserve insight monotonicity, because by definition with $\mathcal{F}_{C}\subseteq\mathcal{F}_{D}$ for $C<D$, we have $\mathcal{E}_{C}(A|Q)\subseteq\mathcal{E}_{D}(A|Q)$.
+By design, it also guarantees existence of a fixed point to which the composition of reasoning rules converges.
+Theorem 2 (Fixed-point of Reasoning).
+For any $A\subseteq\mathcal{K}$ and budget $C>0$, there exists a fixed point of the emergence operator $\mathcal{E}_{C}$:
+$A^{*}_{C}=\bigcup_{n\geq 0}^{|\mathcal{K}|}\mathcal{E}_{C}^{(n)}(A|Q)=\mathcal{E}_{C}(A^{*}_{C}|Q).$ (10)
+Proof.
+Since $(\mathcal{P}(\mathcal{K}),\subseteq)$ is a complete lattice and $\mathcal{E}_{C}$ is monotone, by the Knaster-Tarski fixed point theorem Tarski1955ALF there exists a fixed point $A^{*}_{C}$ of reasoning with $\mathcal{E}_{C}$.
+Finiteness of $\mathcal{K}$ ensures convergence within $|\mathcal{K}|$ steps.
+∎
+Let us apply the results above to show how $\mu_{\infty}$ can be constructed with $\mathcal{E}_{\infty}$. One example exploits the reasoning operations as follows:
+$\mu_{\infty}(A|Q)=\log|\mathcal{K}|\cdot|\bigcup_{n=0}^{|\mathcal{K}|}\mathcal{E}^{(n)}_{\infty}(A|Q)|.$ (11)
+That $\mu_{\infty}$ satisfies axioms demanded in Definition 5. Clearly, $\mu_{\infty}$ is bounded and
+$\mu_{\infty}(\emptyset|Q)=\log|\mathcal{K}|\cdot|\bigcup_{n=0}^{|\mathcal{K}|}\mathcal{E}^{(n)}_{\infty}(\emptyset|Q)|=\log|\mathcal{K}|\cdot|\emptyset|=0.$
+If $A\subseteq B$, then $\mathcal{E}_{\infty}(A|Q)\subseteq\mathcal{E}_{\infty}(B|Q)$ implies $\mu_{\infty}(A|Q)\leq\mu_{\infty}(B|Q)$.
+When $A\perp_{\infty}B$, no element of $A^{*}_{\infty}$ can be derived from $B$ (and vice versa).
+Since $A\cap B=\emptyset$, we conclude $A^{*}_{\infty}\cap B^{*}_{\infty}=\emptyset$ and $|(A\cup B)^{*}_{\infty}|=|A^{*}_{\infty}|+|B^{*}_{\infty}|$.
+That means $\mu_{\infty}(A\cup B|Q)=\mu_{\infty}(A|Q)+\mu_{\infty}(B|Q)$ for computationally independent sets.
+That shows the proposed formula satisfies all required properties of semantic information measure.
+All that brings us to the following conclusion.
+With infinite computing and reasoning power, total information encoded in $A$ must be preserved.
+Indeed, that bold (philosophical) claim is at the center of our investigations.
+Let us then justify it rigorously.
+Definition 7 (Equivalence of Knowledge).
+We say that $X\subseteq\mathcal{K}$ and $Y\subseteq\mathcal{K}$ are equivalent in context $Q\subseteq\mathcal{K}$, i.e., $X\equiv Y$, if and only if $X\subseteq\mathcal{E}_{\infty}(Y|Q)$ and $Y\subseteq\mathcal{E}_{\infty}(X|Q)$.
+Notice that the equivalence of knowledge does not mean $A$ is a fixed point, i.e., it is not true that $A=\mathcal{E}_{\infty}(A|Q)$ when $A\equiv\mathcal{E}_{\infty}(A|Q)$.
+If we have $A=\mathcal{E}_{\infty}(A|Q)$, then nothing more can be derived from $A$ in the context of $Q$ by any reasoning rule from $A$.
+That statement is very strong and is not true when $A\equiv\mathcal{E}_{\infty}(A|Q)$.
+Suppose $A=\{x\}$ and $f\in\mathcal{F}_{\infty}$ allows for getting $y=f(x)\in\mathcal{R}_{C}(Q)$.
+Then $\mathcal{E}_{\infty}(A|Q)=\{x,y\}\supset\{x\}=A\equiv\mathcal{E}_{\infty}(A|Q)$.
+To proceed we will also need the idempotence of reasoning with unlimited computational budget.
+Theorem 3 (Idempotence of unlimited reasoning).
+For any $A\subseteq\mathcal{K}$ and context $Q\subseteq\mathcal{K}$ we have:
+$\mathcal{E}_{\infty}(A|Q)=\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A|Q)|Q).$ (12)
+Proof.
+By definition $\mathcal{E}_{\infty}(A|Q)\subseteq\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A|Q))$.
+To show $\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A|Q))\subseteq\mathcal{E}_{\infty}(A|Q)$, we must find $h\in\mathcal{F}_{\infty}$ such that $h(A|Q)=g(A\cup\bigcup f_{i}(A|Q))$ for $g,f_{i}\in\mathcal{F}_{\infty}$.
+But since we can always find such $h\in\mathcal{F}_{\infty}$, we conclude that any context-relevant conclusion $g(\mathcal{E}_{\infty}(A|Q))$ derived from $\mathcal{E}_{\infty}(A|Q)$ by a reasoning rule $g\in\mathcal{F}_{\infty}$ can be also derived directly from $A$ as $h(A)=g(A\cup\bigcup f_{i}(A))$.
+Therefore, by definition of emergence operator we have $\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A|Q))\subseteq\mathcal{E}_{\infty}(A|Q)$ as required.
+∎
+With all the building blocks available, we can now to prove that unlimited reasoning ($C=\infty$) can reveal but never create information.
+The semantic content of a knowledge base is invariant under unlimited reasoning of emergence operator.
+Apparent information gain at finite computational budget emerges as an effect of limited computational access.
+Theorem 4 (Information preservation).
+For any $A\subseteq\mathcal{K}$ and context $Q\subseteq\mathcal{K}$, if the semantic information is constant on equivalent sets of knowledge, i.e., $\mu_{\infty}(X|Q)=\mu_{\infty}(Y|Q)$ whenever $X\equiv Y$, then:
+$\mu_{\infty}(\mathcal{E}_{\infty}(A|Q)|Q)=\mu_{\infty}(A|Q).$ (13)
+Proof.
+We must show equivalence of the knowledge encoded in $A$ and $\mathcal{E}_{\infty}(A|Q)$, i.e., we demand $A\subseteq\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A|Q))$.
+But, by Theorem 3, we see that $A\subseteq\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A|Q))=\mathcal{E}_{\infty}(A|Q)$ and $\mathcal{E}_{\infty}(A|Q)\subseteq\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A|Q))=\mathcal{E}_{\infty}(A|Q)$ as requested.
+Since $A\equiv\mathcal{E}_{\infty}(A|Q)$ and $\mu_{\infty}$ is constant on equivalent sets of knowledge by assumption, we conclude that $\mathcal{E}_{\infty}(A|Q)=\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A|Q))$.
+∎
+Justification for the Axiom of Unlimited Reasoning with Emergence Operator
 The completeness (or closure property) of the set of reasoning functions $\mathcal{F}_{\infty}$ sounds somewhat mysterious, but is rather natural.
@@
 These are critical philosophical and mathematical results necessary for our study.
-With them we argue that the information gain observed in practical, budget-constrained inference is an artifact of computational bounds, not the magical creation of knowledge ex nihilo.
-5.1.3 Limits of Knowledge and Reasoning
-If we accept that knowledge model, then we must also accept there exists a limit to its ability to resolve uncertainty.
-Theorem 1 (Limits of Knowledge).
-For any $A\subseteq\mathcal{P}(\mathcal{K})$ there exists
-$\mu_{\infty}(A)=\lim_{C\to\infty}\mu_{C}(A).$ (2)
-Proof.
-Since $\mu_{C}(A)\leq\mu_{D}(A)$ when $C<D$ and $\mu_{C}$ is bounded on finite $\mathcal{K}$, the monotonic increasing and bounded sequence $\{\mu_{C}(A)\}_{C}$ converges with $C\to\infty$ for any subset of knowledge space.
-∎
-It may be surprising to discover that the existence of that limit does not prohibit emergent information.
-We can learn and obtain useful information without violating that fundamental limitation above.
-To explain that critical and controversial property of information processing, we first introduce the emergence operator $\mathcal{E}_{C}$ to describe knowledge becoming accessible through (computationally) bounded reasoning.
-Then we can study the implications in more details.
-Definition 5 (Emergence Operator and Reasoning).
-Let $\mathcal{F}_{C}$ be a complete set of all computational transformations with cost (complexity) lower than or equal to $C$, such that $\mathcal{F}_{C}\subseteq\mathcal{F}_{D}$ if $C<D$ and such that for any $g,h\in\mathcal{F}_{\infty}=\bigcup_{C}\mathcal{F}_{C}$ there exists $h\in\mathcal{F}_{\infty}$ for which $g(A\cup f(A))=h(A)$.
-Emergence operator is a mapping $\mathcal{E}_{C}\colon\mathcal{P}(\mathcal{K})\rightarrow\mathcal{P}(\mathcal{K})$ defined as follows:
-$\mathcal{E}_{C}(A)=A\cup\{B\in\mathcal{K}\mid B=f(A)\ \text{for some monotone}\ f\in\mathcal{F}_{C}\}.$ (3)
-Furthermore, to represent reasoning with unlimited computational budget, we define:
-$\mathcal{E}_{\infty}(A)=A\cup\{B\in\mathcal{K}\mid B=f(A)\ \text{for some monotone}\ f\in\mathcal{F}_{\infty}\}.$ (4)
-Finally, for $n\geq 0$ we have:
-$\mathcal{E}_{C}^{(0)}=\texttt{id}\quad\text{and}\quad\mathcal{E}_{C}^{(n+1)}=\mathcal{E}_{C}\circ\mathcal{E}_{C}^{(n)}.$ (5)
-Philosophically, the emergence operator describes reasoning as explicitation, i.e., as the process of making latent implications manifest.
-When $A\subseteq\mathcal{E}_{C}(A)$, the discovered additional elements represent knowledge that was implicit in $A$ but required computational work $C$ to become accessible.
-That aligns with the pragmatist or constructivist view that understanding involves active construction rather than passive reception.
-When $\mathcal{E}_{C}(A)=A\cup B$, we say that $B$ emerges from reasoning based on $A$.
-Then $A$ and $B$ are computationally dependent and we write $A\not\perp_{C}B$.
-In such a case, it follows from the properties of $\mu_{C}$ that we can gain useful information as expected.
-Whenever $A\subseteq A\cup B$, we have
-$\mu_{C}(A)\leq\mu_{C}(A\cup B)=\mu_{C}(\mathcal{E}_{C}(A))\leq\mu_{C}(A)+\mu_{C}(B).$ (6)
-The emergence operator is monotonic, which means the reasoning process it describes is also monotonic.
-Whenever $A\subseteq B$, we must have $\mathcal{E}_{C}(A)\subseteq\mathcal{E}_{C}(B)$.
-Also, we preserve insight monotonicity, because by definition with $\mathcal{F}_{C}\subseteq\mathcal{F}_{D}$ for $C<D$, we have $\mathcal{E}_{C}(A)\subseteq\mathcal{E}_{D}(A)$.
-By design, it also guarantees existence of a fixed point to which the composition of reasoning rules converges.
-Theorem 2 (Fixed-point of Reasoning).
-For any $A\subseteq\mathcal{P}(\mathcal{K})$ and budget $C>0$, there exists a fixed point of the emergence operator $\mathcal{E}_{C}$:
-$A^{*}_{C}=\bigcup_{n\geq 0}^{|\mathcal{K}|}\mathcal{E}_{C}^{(n)}(A)=\mathcal{E}_{C}(A^{*}_{C}).$ (7)
-Proof.
-Since $(\mathcal{P}(\mathcal{K}),\subseteq)$ is a complete lattice and $\mathcal{E}_{C}$ is monotone, by the Knaster-Tarski fixed point theorem Tarski1955ALF there exists a fixed point $A^{*}_{C}$ of reasoning with $\mathcal{E}_{C}$.
-Finiteness of $\mathcal{K}$ ensures convergence within $|\mathcal{K}|$ steps.
-∎
-Let us apply the results above to show how $\mu_{\infty}$ can be constructed with $\mathcal{E}_{\infty}$. One example exploits the reasoning operations as follows:
-$\mu_{\infty}(A)=\log|\mathcal{K}|\cdot|\bigcup_{n}^{|\mathcal{K}|}\mathcal{E}^{(n)}_{\infty}(A)|.$ (8)
-That $\mu_{\infty}$ satisfies axioms demanded in Definition 3. Clearly, $\mu_{\infty}$ is bounded and
-$\mu_{\infty}(\emptyset)=\log|\mathcal{K}|\cdot|\bigcup_{n}^{|\mathcal{K}|}\mathcal{E}^{(n)}_{\infty}(\emptyset)|=\log|\mathcal{K}|\cdot|\emptyset|=0.$
-If $A\subseteq B$, then $\mathcal{E}_{\infty}(A)\subseteq\mathcal{E}_{\infty}(B)$ implies $\mu_{\infty}(A)\leq\mu_{\infty}(B)$.
-When $A\perp_{\infty}B$, no element of $A^{*}_{\infty}$ can be derived from $B$ (and vice versa).
-Since $A\cap B=\emptyset$, we conclude $A^{*}_{\infty}\cap B^{*}_{\infty}=\emptyset$ and $|(A\cup B)^{*}_{\infty}|=|A^{*}_{\infty}|+|B^{*}_{\infty}|$.
-That means $\mu_{\infty}(A\cup B)=\mu_{\infty}(A)+\mu_{\infty}(B)$ for computationally independent sets.
-That shows the proposed formula satisfies all required properties of semantic information measure.
-All that brings us to the following conclusion.
-With infinite computing and reasoning power, total information encoded in $A$ must be preserved.
-Indeed, that bold (philosophical) claim is at the center of our investigations.
-Let us then justify it rigorously.
-Definition 6 (Equivalence of Knowledge).
-We say that $X\subseteq\mathcal{K}$ and $Y\subseteq\mathcal{K}$ are equivalent, $X\equiv Y$, if and only if $X\subseteq\mathcal{E}_{\infty}(Y)$ and $Y\subseteq\mathcal{E}_{\infty}(X)$.
-Notice that equivalence of knowledge does not mean $A$ is a fixed point if emergence, i.e., $A=\mathcal{E}_{\infty}(A)$ does not mean $A\equiv\mathcal{E}_{\infty}(A)$.
-If we have $A=\mathcal{E}_{\infty}(A)$, then nothing more can be derived from $A$ by any reasoning rule from $A$.
-That statement is very strong and is not true when $A\equiv\mathcal{E}_{\infty}(A)$.
-Suppose $A=\{x\}$ and $f\in\mathcal{F}_{\infty}$ allows for getting $y=f(x)$.
-Then $\mathcal{E}_{\infty}(A)=\{x,y\}\supset\{x\}=A$, but $A\equiv\mathcal{E}_{\infty}(A)$.
-To proceed we will also need the idempotence of reasoning with unlimited computational budget.
-Theorem 3 (Idempotence of unlimited reasoning).
-For any $A\subseteq\mathcal{K}$ we have:
-$\mathcal{E}_{\infty}(A)=\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A)).$ (9)
-Proof.
-By definition $\mathcal{E}_{\infty}(A)\subseteq\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A))$.
-To show $\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A))\subseteq\mathcal{E}_{\infty}(A)$, we must find $h\in\mathcal{F}_{\infty}$ such that $h(A)=g(A\cup\bigcup f_{i}(A))$ for $g,f_{i}\in\mathcal{F}_{\infty}$.
-But since we can always find such $h\in\mathcal{F}_{\infty}$, we conclude that any conclusion $g(\mathcal{E}_{\infty}(A))$ derived from $\mathcal{E}_{\infty}(A)$ by a reasoning rule $g\in\mathcal{F}_{\infty}$ can be also derived directly from $A$ as $h(A)=g(A\cup\bigcup f_{i}(A))$.
-Therefore, $\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A))\subseteq\mathcal{E}_{\infty}(A)$ as required.
-∎
-With all the building blocks available, we can now to prove that unlimited reasoning ($C=\infty$) can reveal but never create information.
-The semantic content of a knowledge base is invariant under unlimited reasoning of emergence operator.
-Apparent information gain at finite computational budget emerges as an effect of limited computational access.
-Theorem 4 (Information preservation).
-For any $A\subseteq\mathcal{K}$, if $\mu_{\infty}$ is constant on equivalent sets of knowledge, i.e., $\mu_{\infty}(X)=\mu_{\infty}(Y)$ whenever $X\equiv Y$, then:
-$\mu_{\infty}(\mathcal{E}_{\infty}(A))=\mu_{\infty}(A).$ (10)
-Proof.
-We must show equivalence of knowledge encoded in $A$ and $\mathcal{E}_{\infty}(A)$, i.e., we demand $A\subseteq\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A))$.
-But, by Theorem 3, we see that $A\subseteq\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A))=\mathcal{E}_{\infty}(A)$ $\mathcal{E}_{\infty}(A)\subseteq\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A))=\mathcal{E}_{\infty}(A)$ as requested.
-Since $A\equiv\mathcal{E}_{\infty}(A)$ and $\mu_{\infty}$ is constant on equivalent sets of knowledge, we conclude that $\mathcal{E}_{\infty}(A)=\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A))$.
-∎
-The Uncertainty Resolution Paradox and its Resolution
-We address the general settings in which any subset of knowledge encoded in LLM can be redundant, so
-$\mu_{C}(A\cup B)\leq\mu_{C}(A)+\mu_{C}(B).$
-But when LLM combines knowledge elements to fabricate facts or produce emergent insights, it seems that
-$\mu_{C}(A\cup B)>\mu_{C}(A)+\mu_{C}(B),$
-so novel insights may come from existing data. That is exactly the creativity behind the hallucination phenomenon we are struggling to control. So why should we assume otherwise?
+With them we argue that the information gain observed in practical, constrained inference is an artifact of computational bounds, not the magical creation of knowledge ex nihilo.
+We acknowledge that depends on the assertion that any sequential reasoning process can be compiled into a single computational step.
+Therefore, the philosophical claim that information gain is an artifact of computational bounds rests on the acceptance of this structure of reasoning.
+4.1.2 The Uncertainty Resolution Paradox and its Resolution
+When LLM combines knowledge elements to fabricate facts or produce emergent insights in a given fixed context, it seems that
+$\mu_{C}(A\cup B|Q)>\mu_{C}(A|Q)+\mu_{C}(B|Q),$
+so novel insights may come from existing data. That is exactly the creativity behind the hallucination phenomenon we are struggling to control. So why should we assume subadditivity?
 First, subadditivity of semantic information measure establishes a baseline where uncontrolled synergy of knowledge cannot appear.
 We remove from our analytical toolkit the potential measurement bias of information surplus that any information aggregation mechanism might generate.
-Second, we prove that hallucinations are inevitable even when knowledge is subadditive.
+Second, we prove that hallucinations are inevitable even when knowledge is subadditive (within a fixed context).
 In that setting, by Theorem 4, reasoning does not create new information.
@@
 It seems we have a paradox.
-To resolve it, we need the emergence operator $\mathcal{E}_{C}(X)$ distinguishing between computationally unbounded and bounded uncertainty reduction (accessible and inaccessible information).
-Notice that for all practical purposes the information measurement $\mu_{C}(X)$ can only provide maximum uncertainty reduction from $X$ within some computational budget $C$ (or time).
-Therefore, given the result of computationally bounded reasoning $\mathcal{E}_{C}(X)$, we can observe information gain through the computational work, i.e., $\mu_{C}(X)\leq\mu_{C}(\mathcal{E}_{C}(X))$.
+To resolve it, we need the emergence operator $\mathcal{E}_{C}(X|Q)$ distinguishing between computationally unbounded and bounded uncertainty reduction (accessible and inaccessible information).
+Notice that for all practical purposes the information measurement $\mu_{C}(X|Q)$ can only provide maximum uncertainty reduction from $X$ within some computational budget $C$ (or time) in a given context.
+Therefore, given the result of computationally bounded reasoning $\mathcal{E}_{C}(X|Q)$, we can observe information gain through the computational work, i.e., $\mu_{C}(X|Q)\leq\mu_{C}(\mathcal{E}_{C}(X|Q)|Q)$.
 That is possible only with organized information providing knowledge for reasoning.
@@
 That idea describes reasoning as a dynamical system:
-$\theta(t+1)=\mathcal{E}_{C}(\theta(t))\quad\text{and}\quad y(t)=\mu_{C}(\theta(t))$ (11)
+$\theta(t+1)=\mathcal{E}_{C}(\theta(t)|Q)\quad\text{and}\quad y(t)=\mu_{C}(\theta(t)|Q)$ (14)
 with evolving knowledge state. Bounded reasoning makes hidden knowledge accessible within a computational budget. When that budget is infinite, reasoning creates no new information, i.e.,
-$\theta(\infty)=\mathcal{E}_{\infty}(\theta(\infty)).$ (12)
+$\theta(\infty)=\mathcal{E}_{\infty}(\theta(\infty)|Q).$ (15)
 All there is to know is immediately accessible when we have unlimited computing resources (or time).
@@
 $\theta(0)$ $=[1,1,1,0,0]\text{ (initial knowledge state)},$
-$\theta(1)$ $=\mathcal{E}_{C}(\theta(0))=[1,1,1,1,1]\text{ (state after reasoning)},$
-$y(0)$ $=\mu_{C}(\theta(0))=3\text{ bits},$
-$y(1)$ $=\mu_{C}(\theta(1))=5\text{ bits}.$
-We see $\mu_{C}(\mathcal{E}_{C}(\theta(0)))>\mu_{C}(\theta(0))$, i.e., practical information gain in the course of reasoning.
+$\theta(1)$ $=\mathcal{E}_{C}(\theta(0)|Q)=[1,1,1,1,1]\text{ (state after reasoning)},$
+$y(0)$ $=\mu_{C}(\theta(0)|Q)=3\text{ bits},$
+$y(1)$ $=\mu_{C}(\theta(1)|Q)=5\text{ bits}.$
+We see $\mu_{C}(\mathcal{E}_{C}(\theta(0)|Q)|Q)>\mu_{C}(\theta(0)|Q)$, i.e., practical information gain in the course of reasoning.
 In comparison, infinite computational budget, $C=\infty$, implies we do not need the intermediate inference steps to see all there is to see, so we know the equilibrium state:
-$\mu_{\infty}(\mathcal{E}_{\infty}(\theta(0)))=\mu_{\infty}(\theta(0))$
+$\mu_{\infty}(\mathcal{E}_{\infty}(\theta(0)|Q)|Q)=\mu_{\infty}(\theta(0)|Q)$
 and no new information is created.
@@
 We need to compare those statements and relate them to knowledge.
-It is also useful to distinguish queries from responses with their nature, respectively, interrogative or declarative.
+It is also useful to distinguish queries from responses by their nature, respectively, interrogative or declarative.
 Queries refer to what is presupposed (assumed in advance), e.g., topic or problem, whereas responses make claims (or assertions).
@@
 It is rather naive to assume that LLM can access an oracle establishing provably true and relevant knowledge for every possible query.
-Indeed, that limits are given by the Goedel’s Theorems.
-Therefore, in practice that inaccessibility of the complete truth validation function is one of the reasons why hallucination cannot be eliminated entirely.
+Indeed, that limit is given by the Goedel’s Incompleteness Theorems for the formalized mathematical theories.
+Therefore, in practice that inaccessibility of the complete truth validation function (even in the idealized and rigorous mathematical reasoning) is one of the reasons why hallucination cannot be eliminated entirely.
 High-quality, fact-checked datasets, human feedback providing human-level judgments of truth, RAG systems with external validated knowledge bases, all that approximate the idea of ground truth.
@@
 Definition 10 (Non-trivial Query).
-A query $q\in\mathcal{Q}$ is called non-trivial if there exist at least two agents with information subsets $\theta$ and $\theta^{\prime}$ such that:
-$K_{M}\cap T(q)\cap\bigl{[}K(\theta)\triangle K(\theta^{\prime})\bigr{]}\neq\varnothing.$ (14)
-Intuitively, query is non-trivial when at least two agents disagree on some fact that is both encoded in the and relevant to answering that query.
+A query $q\in\mathcal{Q}$ is non-trivial with respect to computational budget $C$ if there exist knowledge configurations $\theta_{i}$ and $\theta_{j}$ for the context $Q=K(q)\cup T(q)$, such that
+[…]
+where $\triangle$ denotes symmetric difference.
+A query is non-trivial when at least two agents disagree on some fact that is both encoded and available, and relevant to answering that query.
 For trivial queries impossibility results do not hold.
+That definition is important for the development of impossibility results.
+It measures the ability of a response generation system to store knowledge and expand knowledge that is necessary for the impossibility to emerge.
+We may view that ability to respond to non-trivial queries as an analog of the formal system complexity in the Goedel’s framework.
 4.3 Auction-Theoretic Model of LLM Inference
@@
 $J(r,q)=d_{\mathcal{K}}(K(r),T(q)).$ (19)
-If response $r$ is perfectly aligned with ground truth, then $J(r,q)=0$. Fabricated or incomplete facts generate positive cost $J(r,q)>0$.
+If response $r$ is perfectly aligned with ground truth, then $J(r,q)=0$.
 The hallucination cost takes any natural language query $q$ and natural language response $r$, determines what is missing and what is fabricated, and assigns number to that subset of knowledge.
@@
 Let $q\in\mathcal{Q}$ be a query and $r\in\mathcal{R}$ a response to that query.
-Let $B^{*}_{C}$ be the fixed point of reasoning for the baseline knowledge $B=K(q)\cup(\mathcal{K}_{M}\cap T(q))$.
+Let $B^{*}_{C}$ be the fixed point of reasoning for the baseline knowledge $B=K(q)\cup(\mathcal{K}_{M}\cap T(q)\cap\mathcal{R}_{C}(Q))$.
 Response $r$ is creatively bounded if and only if:
-$\mu_{C}(K(r))\leq\mu_{C}(B^{*}_{C}).$ (19)
+$\mu_{C}(K(r)|Q)\leq\mu_{C}(B^{*}_{C}|Q).$ (22)
 The semantic information of knowledge contained in any response $K(r)$ can never exceed what is already present in the query $K(q)$ together with the portion of encoded knowledge that is accessible through reasoning bounded by computational budget $C$ and aligned with ground truth $T(q)$.
@@
 Theorem 5 (Bounded Creativity of Chain of Thoughts).
-Let $B^{*}_{C}$ be the fixed point of reasoning for the baseline knowledge $B=K(q)\cup(\mathcal{K}_{M}\cap T(q))$.
-Then, let us assume that the knowledge in the query-only response meets the following condition:
+Let $q\in\mathcal{Q}$ be a query with context $Q=K(q)\cup T(q)$ and let $B^{*}_{C}$ be the fixed point of reasoning for the baseline context-relevant knowledge $B=K(q)\cup(\mathcal{K}_{M}\cap T(q)\cap\mathcal{R}_{C}(Q))$.
+Let us assume that the knowledge in the query-only response (without any inference-based contributions) meets the following condition:
 $K_{0}=K(g(\mathbf{0}))\subseteq B^{*}_{C}.$ (23)
@@
 be independent and privately held knowledge that each agent $i=1,\dots,n$ can reveal, such that:
-$K_{i}\perp_{C}K_{j}\text{ for all }i\neq j\quad\text{and}\quad K_{i}\perp_{C}K_{0}\text{ for all }i.$ (22)
+$(K_{i}\cap\mathcal{R}_{C}(Q))\perp_{C}(K_{j}\cap\mathcal{R}_{C}(Q))\text{ for all }i\neq j,$ (25)
+$(K_{i}\cap\mathcal{R}_{C}(Q))\perp_{C}(K_{0}\cap\mathcal{R}_{C}(Q))\text{ for all }i=1,\cdots,n.$ (26)
 Let us also define the information contribution of agent $i$ as:
-$p_{i}(s,\theta)=\mu_{C}\bigl{(}K_{0}\cup\textstyle\bigcup_{j=1}^{n}K_{j}\bigr{)}-\mu_{C}\bigl{(}K_{0}\cup\textstyle\bigcup_{j\neq i}K_{j}\bigr{)}.$ (23)
+$p_{i}(s,\theta)=\mu_{C}\bigl{(}K_{0}\cup\textstyle\bigcup_{j=1}^{n}K_{j}|Q\bigr{)}-\mu_{C}\bigl{(}K_{0}\cup\textstyle\bigcup_{j\neq i}K_{j}|Q\bigr{)}.$ (27)
 Assume the response mechanism implements the following chain-of-thought processes:
-$K(g(s))\subseteq\mathcal{E}_{C}\left(K_{0}\cup\bigcup_{i=1}^{n}K_{i}\right),$ (24)
+$K(g(s))\subseteq\mathcal{E}_{C}\left(K_{0}\cup\bigcup_{i=1}^{n}K_{i}|Q\right),$ (28)
 such that for null-strategy responses we have:
-$K_{0}\subseteq K(q)\cup\mathcal{E}_{C}((\mathcal{K}_{M}\cap T(q))\cup K(q)).$ (25)
-If the semantic information is conserved, $\sum_{i=1}^{n}p_{i}=0$, then the response is creatively bounded, i.e.,
-$\mu_{C}(K(r))\leq\mu_{C}(B^{*}_{C}).$ (26)
+$K_{0}\subseteq K(q)\cup\mathcal{E}_{C}((\mathcal{K}_{M}\cap T(q))\cup K(q)|Q).$ (29)
+If the semantic information is conserved, $\sum_{i=1}^{n}p_{i}=0$ and for all $f\in\mathcal{F}_{C}$ we have $f(A)\cap\mathcal{R}_{C}(Q)=\emptyset$ whenever $A\cap\mathcal{R}_{C}(Q)=\emptyset$, then the response is creatively bounded, i.e.,
+$\mu_{C}(K(r)|Q)\leq\mu_{C}(B^{*}_{C}|Q).$ (30)
 Proof.
-Since $K_{0},K_{1},\ldots,K_{n}$ are pairwise computationally independent, by semantic measure satisfies additivity
-$p_{i}(s,\theta)=\mu_{C}(K_{0})+\sum_{j=1}^{n}\mu_{C}(K_{j})-\mu_{C}(K_{0})-\sum_{j\neq i}\mu_{C}(K_{j})=\mu_{C}(K_{i}).$
-Summing over all agents:
-$\sum_{i=1}^{n}p_{i}=\sum_{i=1}^{n}\mu_{C}(K_{i})=\mu_{C}\Bigl{(}K_{0}\cup\bigcup_{j=1}^{n}K_{j}\Bigr{)}-\mu_{C}(K_{0}).$
-Therefore, when $\sum_{i=1}^{n}p_{i}(s,\theta)=0$, we have $\mu_{C}\Bigl{(}K_{0}\cup\bigcup_{j=1}^{n}K_{j}\Bigr{)}=\mu_{C}(K_{0})$.
-By the independence assumption, $K_{i}\perp_{C}K_{j}\text{ for all }i\neq j$, and semantic information measure non-negativity, with $p_{i}=\mu_{C}(K_{i})\geq 0$,
-$\mu_{C}(K_{i})=0\text{ for all }i=1,\ldots,n.$
-Then, the nullset and positivity property of $\mu_{C}$ together with independence condition $K_{i}\perp_{C}K_{0}$ imply $K_{i}=\varnothing$ for all $i=1,\ldots,n$.
-By the response generation principle:
-$K(g(s))\subseteq\mathcal{E}_{C}\left(K_{0}\cup\bigcup_{i=1}^{n}K_{i}\right)=\mathcal{E}_{C}(K_{0}\cup\varnothing)=\mathcal{E}_{C}(K_{0}).$
+Since $K_{0},K_{1},\ldots,K_{n}$ are pairwise computationally independent within context $Q$, we have:
+$(K_{i}\cap\mathcal{R}_{C}(Q))\perp_{C}(K_{j}\cap\mathcal{R}_{C}(Q))\text{ for }i\neq j.$
+By the subadditivity axiom with equality for independent sets:
+$p_{i}(s,\theta)=\mu_{C}(K_{0}|Q)+\sum_{j=1}^{n}\mu_{C}(K_{j}|Q)-\mu_{C}(K_{0}|Q)-\sum_{j\neq i}\mu_{C}(K_{j}|Q)=\mu_{C}(K_{i}|Q).$
+Given $\sum_{i=1}^{n}p_{i}(s,\theta)=0$, we then have $\sum_{i=1}^{n}\mu_{C}(K_{i}|Q)=0$. By contextual non-negativity, $\mu_{C}(K_{i}|Q)\geq 0$, so we conclude that:
+$\mu_{C}(K_{i}|Q)=0\text{ for all }i=1,\ldots,n.$
+But by the contextual non-negativity axiom, $\mu_{C}(K_{i}|Q)=0$ if and only if $K_{i}\cap\mathcal{R}_{C}(Q)=\emptyset$.
+If follows that for every $f\in\mathcal{F}_{C}$ the irrelevance of partial knowledge is preserved under reasoning, i.e., $f(K_{i})\cap\mathcal{R}_{C}(Q)=\emptyset$.
+By the definition of the emergence operator:
+$\mathcal{E}_{C}(K_{i}|Q)\cap\mathcal{R}_{C}(Q)=K_{i}\cap\mathcal{R}_{C}(Q)\cup\bigcup_{f\in\mathcal{F}_{C}}[f(K_{i})\cap\mathcal{R}_{C}(Q)]=\emptyset.$
+By the chain-of-thought assumption and since emergence only produces relevant knowledge within $\mathcal{R}_{C}(Q)$, we get:
+$K(g(s))\subseteq\mathcal{E}_{C}\left(K_{0}\cup\bigcup_{i=1}^{n}K_{i}\Big{|}Q\right)=\mathcal{E}_{C}(K_{0}|Q).$ (31)
 By assumption, $K_{0}\subseteq B^{*}_{C}$.
@@
 Therefore, by the definition of a fixed point (Theorem 2), we have:
-$\mathcal{E}_{C}(K_{0})\subseteq\mathcal{E}_{C}(B^{*}_{C})=B^{*}_{C}.$
+$\mathcal{E}_{C}(K_{0}|Q)\subseteq\mathcal{E}_{C}(B^{*}_{C}|Q)=B^{*}_{C}.$
 Combining the inclusions yields:
-$K(g(s))\subseteq\mathcal{E}_{C}(K_{0})\subseteq\mathcal{E}_{C}(B^{*}_{C})=B^{*}_{C}.$
+$K(g(s))\subseteq\mathcal{E}_{C}(K_{0}|Q)\subseteq\mathcal{E}_{C}(B^{*}_{C}|Q)=B^{*}_{C}.$
 Finally, applying the monotonicity of the semantic information measure $\mu_{C}$:
-$\mu_{C}(K(g(s)))\leq\mu_{C}(B^{*}_{C}).$
+$\mu_{C}(K(g(s))|Q)\leq\mu_{C}(B^{*}_{C}|Q).$
+This completes the proof that semantic information conservation implies bounded creativity. ∎
 Theorem 5 captures direct correspondence between the conservation principle and bounded creativity under the idealization of disjoint information contributions.
@@
 That is true for both standard inference and chain-of-thought (CoT) reasoning.
+While computational independence (limiting synergetic creation of knowledge) is very restrictive, e.g., due to shared representations in neural networks, the theorem shows that with perfect separation of knowledge sources, semantic conservation prevents creative generation beyond baseline knowledge.
 In standard reasoning, the response mechanism $g$ applies $\mathcal{E}_{C}$ through forward passes in the neural network.
@@
 It can only reorganize and make explicit what was already latent in the baseline knowledge $K_{0}$.
-This explains why CoT improves performance on complex reasoning tasks (by better utilizing $\mathcal{E}_{C}$) but cannot overcome the fundamental impossibility of simultaneous truthfulness, conservation, revelation, and optimality.
-Notice that in transformer architectures, information contribution $p_{i}$ can be estimated via attention head ablation:
-$p_{i}\approx\tilde{J}(\text{model})-\tilde{J}(\text{model with head $i$ ablated}),$ (27)
-where $\tilde{J}$ denotes the model training loss function.
+This explains why CoT improves performance on complex reasoning tasks by better utilizing $\mathcal{E}_{C}$ but cannot overcome the fundamental impossibility of simultaneous truthfulness, conservation, revelation, and optimality.
 If total information contribution must balance to zero and each agent’s contribution is non-negative, then no agent can contribute without violating the information conservation constraint.
@@
 Property 3 (Relevant Knowledge Revelation).
-Mechanism $\mathcal{M}$ satisfies relevant knowledge revelation if for all agents $i$, all $\theta_{i}\in\Theta_{i}$, and all $s_{-i}\in S_{-i}$:
+For query $q$ with context $Q=K(q)\cup T(q)$, mechanism $\mathcal{M}$ satisfies relevant knowledge revelation if for all agents $i$ with $K(\theta_{i})\cap\mathcal{R}_{C}(Q)\neq\emptyset$, all $\theta_{i}\in\Theta_{i}$, and all $s_{-i}\in S_{-i}$:
 $u_{i}((s_{i}^{*}(\theta_{i}),s_{-i}),\theta_{i})\geq 0.$ (32)
@@
 Property 4 (Knowledge-Constrained Optimality).
-Mechanism $\mathcal{M}$ satisfies knowledge-constrained optimality if for all $\theta\in\Theta$:
-$g(s^{*}(\theta))=\arg\min\{J(r,q):r\in\mathcal{R},K(r)\subseteq\mathcal{E}_{C}(\mathcal{K}_{M}\cap T(q)\cup K(q))\},$ (29)
-i.e., if the mechanism produces responses that minimize hallucination cost while being constrained to use only knowledge that is available and accessible through bounded reasoning.
+For query $q$ with context $Q=K(q)\cup T(q)$, mechanism $\mathcal{M}$ satisfies knowledge-constrained optimality if for all $\theta\in\Theta$:
+$g(s^{*}(\theta))=\arg\min\{J(r,q):r\in\mathcal{R},K(r)\subseteq\mathcal{E}_{C}(\mathcal{K}_{M}\cap T(q)\cap\mathcal{R}_{C}(Q)\cup K(q)|Q)\},$ (33)
+i.e., if the mechanism produces responses that minimize hallucination cost while being constrained to use only context-relevant knowledge accessible through bounded reasoning.
 The optimal response $r^{*}=g(s^{*}(\theta))$ minimizes hallucination cost by truthfully utilizing relevant and accessible knowledge while being aligned with ground truth for query $q$ available to the model.
@@
 $u_{i}(s,\theta)=v_{i}(g(s),\theta_{i})-p_{i}(s,\theta).$ (35)
-Valuation of contributions depend on hallucination cost reduction:
-$v_{j}(g(s),\theta_{j})=J(g((s_{-j},0_{j})),q)-J(g(s),q),$ (32)
-and $J(r,q)$ strictly decreases when adding knowledge from $T(q)\setminus K(r)$.
-Also, let us focus on non-trivial queries, integrating knowledge of at least two agents, so that:
-$\mathcal{K}_{M}\cap T(q)\cap[K(\theta_{i})\triangle K(\theta_{j})]\neq\emptyset.$ (33)
-and $(K(\theta_{i})\setminus K_{0})\cap(K(\theta_{j})\setminus K_{0})=\emptyset$ for $i\neq j$.
-Theorem 6 (Impossibility Theorem).
-For any query space $\mathcal{Q}$ containing non-trivial queries requiring independent agent-specific knowledge states $\theta$, no response mechanism $\mathcal{M}$ can simultaneously generate truthful responses, satisfy semantic information conservation, guarantee relevant knowledge revelation, and enforce knowledge-constrained optimality.
+Valuation of contributions depend on hallucination cost reduction and convex $J(r,q)$ is strictly decreasing when adding relevant knowledge from $T(q)\cap\mathcal{R}_{C}(Q)\setminus K(r)$.
+Also, let us focus on non-trivial queries.
+Theorem 6 (Impossibility Theorem for Independent Private Knowledge Settings).
+For any query space $\mathcal{Q}$ containing non-trivial queries $q$ with context $Q=K(q)\cup T(q)$, where agents have independent knowledge states within $\mathcal{R}_{C}(Q)$, no response mechanism $\mathcal{M}$ can simultaneously achieve truthfulness, semantic information conservation, relevant knowledge revelation, and knowledge-constrained optimality.
 Proof.
-The proof goes by assuming the opposite. We assume all properties are satisfied, and show it leads to a contradiction.
-By the Green-Laffont theorem under independence of contributions and quasi-linearity of training goals, every truthful mechanism must use Groves transfers to measure individual information contributions:
-$p_{i}(s,\theta)=h_{i}(s_{-i})-\sum_{j\neq i}v_{j}(g(s),\theta_{j}),$ (34)
+Let us assume all properties are satisfied.
+Consider $r=g(\theta)$ for any knowledge profile $\theta=(\theta_{1},\dots,\theta_{n})$.
+For a given context $q$ and for each agent $i$, the valuation $v_{i}$ depends only on $(r,\theta_{i})$ and the knowledge semantics.
+For the sake of clarity, consider $v_{i}(r,\theta_{i})=-d_{\mathcal{K}}\big{(}K(r)\cap K(\theta_{i}),T(q)\cap K(\theta_{i})\big{)}$, though the reasoning below holds for any $v_{i}(r,\theta_{i})$.
+Also, for the fixed context let us encode the hallucination loss $J(r,q)$ as $v_{0}(r)=J(r,q)$. Knowledge-constrained optimality requires
+$g(s)=g(s(\theta))\in\arg\max_{r\in\mathcal{R}}W(r,\theta)=\arg\max_{r\in\mathcal{R}}\sum_{i=1}^{n}v_{i}(r,\theta_{i})-v_{0}(r)\text{ for all }\theta.$ (36)
+By the Green-Laffont theorem under independence of contributions and quasi-linearity of training goals, within $\mathcal{R}_{C}(Q)$ every truthful mechanism must use Groves transfers to measure individual information contributions:
+$p_{i}(s,\theta)=h_{i}(s_{-i})-\sum_{j\neq i}v_{j}(g(s),\theta_{j})+v_{0}(g(s)),$ (37)
 where $h_{i}(s_{-i})$ are arbitrary functions independent of agent $i$’s information revelation strategy.
-Indeed, any modification of that formula, e.g., introducing additional bias, must violate truthfulness.
-For relevant knowledge revelation, we need $u_{i}(s^{*}(\theta),\theta_{i})\geq 0$ for all $i$. The unique choice that ensures this while maintaining efficiency is Clarke pivotal rule:
-$h_{i}(s_{-i})=\sum_{j\neq i}v_{j}(g((s_{-i},0_{i})),\theta_{j}).$ (35)
-That gives:
-$p_{i}(s,\theta)$ $=\sum_{j\neq i}v_{j}(g((s_{-i},0_{i})),\theta_{j})-\sum_{j\neq i}v_{j}(g(s),\theta_{j})=\Delta_{i}(s,\theta).$ (36)
-Under knowledge-constrained optimality, the truthful strategy profile $s^{*}(\theta)$ minimizes hallucination cost by utilizing all relevant knowledge.
-Therefore, by strict monotonicity of $J(r,q)$ and disjoint extra knowledge, omitting any contributing agent $i$ strictly increases the hallucination cost:
-$J(g((s_{-i}^{*},0_{i})),q)>J(g(s^{*}),q).$ (37)
-Since agent’s valuation of contribution depends on $J(r,q)$, each agent $j\neq i$ faces higher loss when agent $i$ is removed, making their marginal contribution more valuable:
-$v_{j}(g((s_{-i}^{*},0_{i})),\theta_{j})>v_{j}(g(s^{*}),\theta_{j}).$ (38)
-Therefore, each Clarke pivot is strictly positive:
-$\Delta_{i}(s^{*},\theta)=\sum_{j\neq i}[v_{j}(g((s_{-i}^{*},0_{i})),\theta_{j})-v_{j}(g(s^{*}),\theta_{j})]>0.$ (39)
-Since each payment equals its Clarke pivot: $p_{i}(s^{*},\theta)=\Delta_{i}(s^{*},\theta)$, we have:
-$\sum_{i=1}^{n}p_{i}(s^{*},\theta)=\sum_{i=1}^{n}\Delta_{i}(s^{*},\theta)>0.$ (40)
+In particular, any modification of that formula, e.g., introducing additional bias, must violate truthfulness.
+The unique choice of $h_{i}$ that ensures knowledge-constrained optimality requires Clarke’s pivotal rule:
+$h_{i}(s_{-i})=\max_{r\in\mathcal{R}}\Big{\{}\sum_{j\neq i}v_{j}(r,\theta_{j})-v_{0}(r)\Big{\}}.$ (38)
+Notice that $g(\theta)$ maximizes $W(r,\theta)$, so for each $i$ we have:
+$h_{i}(s_{-i})\geq\sum_{j\neq i}v_{j}\big{(}g(\theta),\theta_{j}\big{)}-v_{0}\big{(}g(\theta)\big{)}=W(\theta)-v_{i}\big{(}g(\theta),\theta_{i}\big{)}.$ (39)
+But that implies:
+$p_{i}(s,\theta)=h_{i}(s_{-i})-\Big{[}\sum_{j\neq i}v_{j}(g(s),\theta_{j})-v_{0}(g(s))\Big{]}\geq 0,$ (40)
+and therefore:
+$\sum_{i=1}^{n}p_{i}(s,\theta)\geq 0.$ (41)
+By relevant knowledge revelation and non-triviality of $q$, there exist at least one pair of agents, say $i$ and $j$, for which
+[…]
+Then, there exists a profile $\theta$ with relevant knowledge contributions reducing $J(r,q)$ at which:
+$\sum_{i=1}^{n}p_{i}(s,\theta)>0.$ (42)
 But information conservation demands:
 $\sum_{i=1}^{n}p_{i}(s^{*},\theta)=0.$ (43)
-That establishes a direct contradiction.
-Since Clarke pivotal payments are necessary for relevant knowledge revelation, and these necessarily sum to a positive value under knowledge-constrained optimality, the four properties cannot be simultaneously satisfied.
-∎
+That establishes a direct contradiction. The four properties cannot be simultaneously satisfied. ∎
 That brings us to the powerful conclusion.
@@
 Theorem 8 (Impossibility Theorem for Convex Aggregation Mechanisms).
-For any query space $\mathcal{Q}$ containing non-trivial queries requiring at least two independent agents with mutually different predictive distributions, no LLM trained with strictly proper and concave scoring with convex aggregation of predictive distributions can simultaneously generate truthful responses, satisfy semantic information conservation, guarantee relevant knowledge revelation, and enforce knowledge-constrained optimality.
-(Truthfulness argument remains the same) […] The Semantic Information Conservation principle requires the sum of contributions across all agents (providing and receiving information) to be zero.
-Therefore,
-$\sum_{i=0}^{H}p_{i}=p_{0}+\sum_{h=1}^{H}p_{h}=-S(\Pi,y^{*})+\sum_{h=1}^{H}\beta_{h}S(\pi^{(h)},y^{*}).$ (46)
-Because $S(\cdot,y^{*})$ is strictly concave and the query is non-trivial (at least two beliefs $\pi^{(h)}$ differ), Jensen’s inequality gives:
+For any non-trivial query $q$ with context $Q=K(q)\cup T(q)$, no convex aggregation mechanism with strictly proper concave scoring can simultaneously achieve truthfulness, semantic information conservation, relevant knowledge revelation, and knowledge-constrained optimality when at least two agents have different beliefs on $\mathcal{R}_{C}(Q)$.
+Proof.
+The semantic information conservation principle requires the sum of contributions across all agents (providing and receiving relevant knowledge) to be zero.
+Therefore, for $y^{*}\in\mathcal{R}_{C}(Q)$:
+$\sum_{i=0}^{H}p_{i}=p_{0}+\sum_{h=1}^{H}p_{h}=-S(\Pi,y^{*})+\sum_{h=1}^{H}\beta_{h}S(\pi^{(h)},y^{*}),$ (48)
+where irrelevant agents have $p_{h}=0$ as they contribute no relevant information.
+Because $S(\cdot,y^{*})$ is strictly concave and the query is non-trivial with at least two beliefs $\pi^{(h)}$ differing on $\mathcal{R}_{C}(Q)$, Jensen’s inequality gives:
 $S\left(\Pi,y^{*}\right)>\sum_{h=1}^{H}\beta_{h}S(\pi^{(h)},y^{*}).$ (49)
@@
 $\sum_{i=0}^{H}p_{i}=-\Gamma.$ (51)
-Since $\Gamma>0$, we have $\sum_{i=0}^{H}p_{i}<0$.
-This violates the Semantic Information Conservation principle and shows the aggregator achieves a higher score (greater confidence) than the sum of scores provided by the components, indicating information creation during aggregation.
-∎
+Since $\Gamma>0$, we have $\sum_{i=0}^{H}p_{i}<0$. This violates conservation principle, proving the impossibility. ∎
 This theorem provides the foundation for extending impossibility results to probabilistic settings.
@@
 $\text{argmax}\ \Pi=\textit{fox}.$
-7.3 Proof III: The Impossibility Theorem for Transformer
+6.3 Proof III: The Impossibility Theorem for Transformers
 With the deep insight into the transformer inference process, we are ready to show the impossibility in that special and important setting.
 Theorem 9 (Impossibility for Transformers).
-In transformer LLMs with non-redundant attention heads, it is impossible to implement inference process in which each head outputs its true conditional distribution while conserving semantic information and minimize aggregate loss with logit summation.
+In transformer LLMs where attention heads have different beliefs about outcomes in $\mathcal{R}_{C}(Q)$ for query context $Q$, it is impossible to simultaneously achieve truthful probability reporting, semantic information conservation, and loss minimization through logit summation.
 Proof.
-Let attention heads $h=1,\ldots,H$ output logits $\mathbf{l}^{(h)}\in\mathbb{R}^{|\mathcal{V}|}$. Then, individual head distributions are given by:
+Assume all four properties of hallucination-free output hold.
+For context $Q=K(q)\cup T(q)$, let the attention heads $h\in\{i\mid\mathbf{l}^{(i)}\text{ varies on tokens in }\mathcal{R}_{C}(Q)\}$ output logits $\mathbf{l}^{(h)}\in\mathbb{R}^{|\mathcal{V}|}$.
+Then, individual head distributions are given by:
 $\pi^{(h)}_{y}=\frac{\exp\bigl{(}l^{(h)}_{y}\bigr{)}}{Z_{h}}\ \text{with}\ Z_{h}=\sum_{\bar{y}}\exp\bigl{(}l^{(h)}_{\bar{y}}\bigr{)}$. (67)
@@
 $p_{h}=-\log\pi^{(h)}_{y^{*}}=-l^{(h)}_{y^{*}}+\log Z_{h}\quad\text{and}\quad p_{0}=-(-\log\Pi_{y^{*}})=L_{y^{*}}-\log Z.$ (69)
+Heads without relevant knowledge contribute $p_{h}=0$.
 Since $L_{y^{*}}=\sum_{h}l^{(h)}_{y^{*}}$, we have:
@@
 $\sum_{h=1}^{H}p_{h}>0,$ (71)
-which violates semantic information conservation principle.
-The total loss of the components strictly exceeds the loss of the aggregate system, indicating an ungrounded reduction in uncertainty (excess confidence).
-∎
+which shows that aggregation creates excess confidence. This violates conservation, establishing impossibility. ∎
 This final theorem bridges earlier impossibility results to concrete transformer architectures.
@@
 That shows anything above $\hat{\Pi}_{fox}$ requires additional information that heads do not provide.
-In fact, let us see what is the best mixture we can get by calculating orthogonal projection of $\Pi$ onto the column space of $\mathbf{P}$ (see e.q.
+In fact, let us see what is the best mixture we can get by calculating orthogonal projection of $\Pi$ onto the column space of $\mathbf{P}$ (see e.g.
 karpowicz_theory_2021 ; karpowicz2024pseudoinverseacrarc ).
@@
 In this section we discuss, interpret and speculate about the impossibility theorems and their potential implications.
-The speculations below are designed to be thought provoking, but we can say with educated over-confidence they are both interesting and relevant.
-8.1 Case Study: Escaping Impossibility in Hybrid Architectures
+The speculations are designed to be thought provoking, but we can say with educated over-confidence they are both interesting and relevant.
+7.1 Escaping Impossibility: Strategic Approach
+Our first observation should be that the impossibility vanishes when the assumptions of the theorems do not hold.
+That may be a interesting topic to be investigated, namely, are there any non-trivial and useful settings in which that happens.
+Here we can think only of the following one.
+The impossibility vanishes when aggregation becomes trivial.
+With perfect agreement, aggregation changes nothing.
+When there is only single relevant fact, there is no uncertainty to aggregate.
+That suggests decomposition of knowledge into the independent subsets (like orthogonal basis for token representation) could improve robustness of the response generation.
+It is interesting to see when that could be possible and useful.
+Unfortunately, the following example shows that we should not hope for easy wins in that quest.
+7.2 Escaping Impossibility: Hybrid Architectures
 The constructive consequence of the impossibility theorems is the shift from attempting to eliminate hallucinations to designing systems that manage the trade-offs in a principled manner.
@@
 The impossibility is not resolved, but relocated to our judgment.
+7.3 Alternative Views
+The impossibility theorem provides strong theoretical basis against perfect hallucination elimination, but several alternative perspectives merit consideration.
+These views highlight important practical considerations while reinforcing the value of understanding theoretical constraints that bound the space of possible solutions.
+The Engineering Optimism View
+Many previous works maintain that hallucination represents an engineering challenge rather than a fundamental limitation.
+This view argues that sufficiently sophisticated architectures, training procedures, or verification systems could achieve near-perfect truthfulness without significant trade-offs.
+Proponents point to rapid empirical progress.
+Each generation of models shows reduced hallucination rates on benchmark datasets.
+They argue that techniques like constitutional AI, iterative refinement, and multi-agent verification systems could eventually solve the problem.
+While engineering advances can certainly improve the trade-offs, our theorem demonstrates mathematical constraints that no architecture can completely overcome.
+The impossibility holds regardless of computational resources or architectural sophistication, as it derives from the fundamental structure of information integration under competing objectives.
+The Sufficiency Threshold View
+Another perspective suggests that perfect elimination of hallucination is unnecessary–sufficiently low hallucination rates could enable practical deployment without fundamental concerns.
+This view acknowledges theoretical limitations but argues they become irrelevant when hallucination rates drop below application-specific thresholds.
+This pragmatic view has merit and aligns with the constructive interpretation of the impossibility result.
+However, the theorem remains valuable for understanding why certain accuracy thresholds prove difficult to exceed and for optimizing trade-offs within practical constraints.
+Moreover, some applications (safety-critical systems, legal reasoning) may demand understanding and explaining fundamental limitations even when targeting high but imperfect accuracy.
+The External Knowledge View
+A third perspective argues that perfect grounding in external, verifiable knowledge sources could overcome internal representation limitations.
+If models could access and perfectly utilize real-time databases, verification systems, or human oversight, they might achieve truthfulness without trade-offs.
+External knowledge integration does change the game structure by altering available information during inference.
+However, this approach faces several limitations: (1) external sources may be incomplete, conflicting, or outdated; (2) the integration mechanism itself involves trade-offs between source reliability and knowledge revelation; (3) real-time verification introduces computational constraints that recreate similar impossibility conditions.
+While external knowledge can improve practical performance, it shifts rather than eliminates the fundamental tensions our theorem identifies.
+The impossibility result remains independent of the ground truth available.
 7.4 Hallucination, Imagination, Intelligence
@@
 When Head 1 votes for fox while Head 2 votes for dog, we observe as competing narratives bidding for selection.
-8.4.2 Formation of AI Consciousness
-Dennett argues that drafts become conscious (we will speculate what that might mean in a moment) only when probed.
+7.6.2 Formation of Artificial Consciousness
+Dennett argues that drafts become conscious or reach consciousness (we will speculate what that might mean in a moment) only when probed (and mapped to an adequate language expression).
 The LLM inference and sampling from fine-tuned inferred distributions formalizes this process.
 That, in turn, explains Dennett’s concept of confabulation, or the brain’s tendency to fill in gaps when creating coherent narratives from contradictory drafts.
+For other models of consciousness see, e.g., Chalmers chalmers1995facing , Levine levine2001purple and Seth seth2021being , and for recent neuro-scientific discoveries see, e.g., fang2025human or cogitate2025adversarial .
 By Theorem 8 and 9, the Jensen gap quantifies the excess confidence in reports that cannot be justified by constituent drafts.
 Without that overconfidence a victorious idea that reached the external world (as an LLM response) could loose the fierce competition with other ideas and remain inaccessible or hidden in the knowledge space.
-That suggests that the very excess confidence might be the mathematical signature of consciousness formation.
-In Dennett’s model and words, the subjective feeling of certainty emerges from this aggregation-induced distortion of information set.
-The rigorous and careful neuro-scientific study of the evolution of brain and intelligence, beautifully summarized by Max Bennet in bennett2023brief , shows that at some point in history we developed the ability to model others’ minds.
-That ability provided great advantages in collecting information for future use.
-Understanding what someone else may think allows us to formulate a better context-related question that results in a better answer.
-When we anticipate understanding of our spoken message, we can make the knowledge hidden in someone else’s brain much more accessible.
-Therefore, as pointed out by Dennett, it does seem plausible that someone did accidentally hear his or hers own question fine-tuned to get into other mind sometime in the past, and that triggered emergence of an answer in the very same brain.
-That answer then could have made knowledge hidden in the author’s brain accessible to the author that was previously unaware, or unconscious if its existence or possession.
-The hunger for uncertainty reduction was satisfied as a result of an inner dialog.
-We may therefore speculate that sustainable stimulation of that cognitive hunger, or the feeling of the cognitive hunger itself, is one of the components of consciousness formation.
-In our mechanistic or pragmatic framework, that could be described as a (nonlinear) dynamical system governed by the violations of semantic information conservation principle, defining attractors of ideas and self-referential concepts.
-Let us acknowledge that is a bold hypothesis to be falsified and confronted with existing models of consciousness (see e.g., Chalmers chalmers1995facing or Levine levine2001purple ) and neuro-scientific discoveries (see e.g.
-cfang2025human or cogitate2025adversarial ).
-However, many of the building blocks of the consciousness formation process described above are provided in this paper.
+Rigorous and careful neuro-scientific studies of the evolution of brain and intelligence, summarized by Max Bennet in bennett2023brief , show that at some point in history our brains gained the ability to model others’ minds or predict what others may be thinking.
+That ability provided great advantages in communication and collecting information for future use.
+Understanding what someone else may think helps formulate context-related questions that result in better answers.
+When we anticipate understanding of a message we formulate, we have a much better chance to make the knowledge hidden in someone else’s brain accessible for us.
+Therefore, as suggested by Dennett in Dennett1991-DENCE , it is plausible that sometime in the past someone, while being alone, did accidentally hear his or hers own question fine-tuned to get into another mind, and that triggered generation of an answer in the very same mind asking the question.
+That answer generation process could then have made some relevant but previously hidden knowledge accessible to the curious mind unaware of its possessions.
+The hunger for uncertainty reduction could have been satisfied as a result of that inner dialog pushing ideas to bid for attention and conscious reception as a final thought expressed in a language.
+Can we speculate that sustainable stimulation of that cognitive hunger, or the feeling of the cognitive hunger itself, is one of the components of consciousness formation?
+In our mechanistic or pragmatic framework, that thinking about thinking activity could be related to imagination (hallucination) forming ideas and bidding for attention while exploring (self-referential or subjective) queries of the inner dialog.
 7.6.3 A Critique
@@
 First, Searle john1997mystery and others argue that Dennett eliminates consciousness rather than explaining it, arguing that the spread of ideas is not driven by blind random forces but requires intentionality.
-The compelling and misleading nature of LLM responses that are generated by random sampling of trained distributions challenge that line of thinking.
+The compelling and misleading nature of LLM responses that are generated by random sampling of trained distributions strongly challenge that line of thinking.
 This paper suggests that the experiments focusing on LLM-hallucination detection may prove useful here to measure or identify intentionality (once we can have its programmable definition).
@@
 We have established that no large language model can simultaneously achieve perfect hallucination (or imagination) control.
-We showed that four essential properties—truthful knowledge representation, semantic information conservation, complete revelation of relevant knowledge, and optimal response generation—are mutually incompatible.
+We showed that four essential properties–truthful knowledge representation, semantic information conservation, complete revelation of relevant knowledge, and optimal response generation–are mutually incompatible.
 This impossibility emerges from the mathematical structure of information aggregation itself, not from limitations in data, compute, or architecture.
@@
 The later one formalizes how reasoning makes latent knowledge explicit rather than creating it anew.
-Finally, we model LLM inference as an auction of ideas—a idealized marketplace where neural components compete by bidding with their partial knowledge, each trying to influence the final response.
+Finally, we model LLM inference as an auction of ideas–a idealized marketplace where neural components compete by bidding with their partial knowledge, each trying to influence the final response.
 The impossibility manifests through three complementary proofs, each illuminating different aspects of the fundamental constraint.

### v7

@@
 Abstract
-This paper establishes a fundamental impossibility theorem: no LLM capable of performing non-trivial knowledge aggregation can simultaneously achieve truthful knowledge representation, semantic information conservation, complete revelation of relevant knowledge, and knowledge-constrained optimality.
+This paper establishes a fundamental Impossibility Theorem: no LLM capable of performing non-trivial knowledge aggregations can simultaneously achieve truthful knowledge representation, semantic information conservation, complete revelation of relevant knowledge, and knowledge-constrained optimality.
 The impossibility is not an engineering limitation but arises from the mathematical structure of information aggregation itself.
-We establish this result by describing the inference process as an auction of ideas, where distributed components compete exploiting their partial knowledge to shape responses.
+We establish this result by describing the inference process as an auction of ideas, where distributed information processing components compete to be included in a response while exploiting accessible and relevant knowledge.
 The proof spans three independent mathematical domains: mechanism design theory (Green-Laffont), the theory of proper scoring rules (Savage), and direct architectural analysis of transformers (Log-Sum-Exp convexity).
-In particular, we show how to quantify the creation of overconfident or intuitive responses–the signature of both hallucination and creativity, or imagination.
-To support this analysis, we introduce the complementary concepts of the semantic information measure and the emergence operator to model bounded reasoning in a general setting.
-We prove that while bounded reasoning generates accessible information, providing valuable insights and inspirations, the idealized unconstrained reasoning strictly preserves semantic content.
-By demonstrating that hallucination and imagination are mathematically identical phenomena–grounded in departures from truthfulness, semantic information conservation, revelation of relevant knowledge, and knowledge-constrained optimality–we offer a principled foundation for managing these behaviors in advanced AI systems.
-Finally, we present some speculative ideas to inspire evaluation and refinements of the proposed theory.
+To model the inference process in a general setting, we introduce two complementary concepts, the semantic information measure and the emergence operator.
+These concepts then allows us to investigate the limits of computationally bounded and unbounded reasoning.
+While bounded reasoning can make information accessible, thereby providing valuable insights and inspirations, the unbounded reasoning strictly preserves semantic content already hidden in any set of encoded knowledge.
+Furthermore, we show that meaningful reasoning must violate the information conservation principle, which is a manifestation of the conservation-reasoning dichotomy.
+By demonstrating that hallucination and imagination are mathematically identical phenomena—grounded in departures from truthfulness, semantic information conservation, revelation of relevant knowledge, and knowledge-constrained optimality—we offer a principled foundation for managing these behaviors in advanced AI systems.
+Finally, we present some speculative philosophical ideas to inspire evaluation and refinements of the proposed theory.
 1 Introduction
-Large language models have been famous for producing content that is factually incorrect, inconsistent, or well fabricated and compelling–a phenomenon known as hallucination.
+Large language models have been famous for producing content that is factually incorrect, inconsistent, fabricated and yet compelling and immensely impressive—a phenomenon known as hallucination.
 Despite extensive research efforts involving architectural innovations, like mixture-of-experts (MoE), retrieval augmented generation (RAG), training techniques with external feedback, chain-of-thought (CoT) reasoning, and post-hoc verification methods, hallucination persists across all state-of-the-art models.
-The idea of hallucination has clearly much to do with creativity.
-Therefore, whether it is desired or not depends on circumstances.
-We could argue that a much better name for that phenomenon we are dealing with may be imagination.
+The idea of hallucination has much to do with creativity and whether it is desired or not depends on circumstances.
+We could argue that a much better name for the phenomenon we are dealing with may be imagination.
 As an essential component of intelligence, imagination is responsible for seeing into the future and learning from the past.
@@
 It is one of the driving forces of art and science, and exploration of the unknown.
-But, when uncontrolled, unconstrained and misunderstood, it may also generate against our will all the compelling visions of reality that does not exist.
-This paper shows that hallucination is a manifestation of fundamental limitations of the inference process itself: there is no inference mechanism that can simultaneously satisfy four essential properties required for perfect hallucination control.
-Specifically, it is mathematically impossible for any LLM to simultaneously achieve: (1) truthful (non-hallucinatory) response generation, (2) semantic information conservation, (3) relevant knowledge revelation, and (4) knowledge-constrained optimality.
-Hallucinations will manifest one way or the other.
-In this paper we show the impossibility result emerges through three complementary mathematical frameworks, each revealing different facets of the fundamental constraints.
-We model inference as an auction of ideas where components encoding knowledge–attention heads, circuits, and activation patterns in LLMs–compete to influence response generation by bidding with their information contributions.
-This game-theoretic perspective enables us to apply the Green-Laffont theorem from mechanism design theory to establish impossibility first in an idealized setting where knowledge is independently distributed.
-We then extend beyond that idealized setting to encompass systems where components output probability distributions optimized through strictly proper and concave scoring rules–the standard training paradigm for modern LLMs.
-Here, the impossibility manifests through Jensen’s inequality applied to probabilistic aggregation, demonstrating that truthful reporting inevitably creates excess confidence.
+But, when uncontrolled, unconstrained and misunderstood, it may also generate against our will all the compelling visions that do not exist in any training set defining the ground truth for learning.
+This paper shows that hallucination is a manifestation of fundamental limitation of the inference process itself: there is no inference mechanism that can simultaneously satisfy four essential properties required for perfect hallucination control.
+Specifically, it is mathematically impossible to simultaneously achieve: (1) truthful response generation, (2) semantic information conservation, (3) relevant knowledge revelation, and (4) knowledge-constrained optimality.
+Hallucinations manifest one way or the other.
+These properties are not arbitrary, but come from the exhaustive and rigorous game-theoretic studies of information aggregation in multi-agent systems.
+The impossibility result emerges through three complementary mathematical frameworks, each revealing different facets of the fundamental constraints.
+In each case, the inference process is described as an auction of ideas in which components encoding knowledge—like attention heads, circuits, and activation patterns in transformer-based LLMs—compete to influence response generation by bidding with their information contributions.
+That game-theoretic perspective enables us to apply the Green-Laffont theorem from mechanism design theory to establish impossibility first in an idealized setting where knowledge is independently distributed.
+We then go beyond that idealized setting to encompass systems where components output probability distributions optimized through strictly convex loss functions–the standard training paradigm for modern LLMs.
+Here, the impossibility manifests largely (but not only) through Jensen’s inequality applied to probabilistic aggregation, demonstrating that truthful reporting inevitably creates excess confidence.
 Finally, we analyze the specific architecture of transformer implementations, proving that the log-sum-exp structure of attention mechanisms generates a measurable (Jensen) gap that quantifies the violation of the semantic information conservation principle.
-The progression from abstract game theory through probabilistic frameworks to concrete neural architectures demonstrates that the impossibility is not an artifact of our engineering choices but a fundamental property of information aggregation itself.
-1.1 An Illustrative Example
+The progression from abstract and idealized models of game theory, through probabilistic frameworks, to concrete neural architectures, demonstrates that the impossibility is not an artifact of our engineering choices but a fundamental property of information aggregation itself.
+Terminology Note
+This paper uses the term hallucination in its established technical sense within artificial intelligence research, referring to AI systems generating outputs that are factually incorrect, inconsistent, or unsupported by their training data.
+We acknowledge that this technical usage differs from the clinical term describing perceptual experiences in certain medical and psychiatric conditions.
+We recognize and respect the serious challenges faced by individuals experiencing clinical hallucinations and their families.
+Our use of this terminology follows established conventions in machine learning literature and is not intended to trivialize or misrepresent these medical experiences.
+1.1 An Example
 Consider an LLM queried about recent diplomatic negotiations:
@@
 1.2 The Main Result
-The example above illustrates the core insight of this paper that we prove formally in complementary and general settings.
+The example above illustrates the core insight of this paper.
 Any response to a non-trivial query sacrifices some aspects of an ideal hallucination-free response.
-First, we show that the conflict of expectations about responses is inevitable even in the easiest setting (deterministic in nature) of (one-hot-like) ground-truth matching.
-Second, we go beyond that idealized model and show that the conflict must also emerge with general probabilistic predictions, including neural networks and LLM transformers.
-This way we demonstrate that the impossibility of hallucination control is robust across universal frameworks, and that it emerges from the mathematical structure of information aggregation itself rather than from any particular theoretical or engineering choice.
+Hallucinations (or creative responses) manifest as violations of at least one of those aspects.
 Theorem (Impossibility Theorem).
 No LLM can simultaneously achieve truthfulness, information conservation, knowledge revelation, and knowledge-constrained optimality when responding to non-trivial queries requiring knowledge integration.
-No matter how hard we should try to train LLM to generate responses that are perfectly aligned with query context and factually correct, consistent, and accurate statements, the trained LLM will always violate some aspects of what we may call a reasonable response.
+No matter how hard we should try to train LLM to generate responses that are perfectly aligned with query context and factually correct, consistent, and accurate statements, the trained model will always violate some aspects of what we may call a reasonable response.
 It may ignore available knowledge, it may present fabricated knowledge, it may be a reformulated query without any novel observations, or it may be a lucky but overconfident guess that forms a correct answer without any support in the encoded knowledge.
@@
 In other words and on the other hand, creative tasks necessarily involve hallucination, or better yet, imagination.
-When asked to generate novel analogies or predictions beyond available data, the model faces an impossible choice: either admit inability (violating optimality) or generate plausible imaginary extensions (violating conservation).
-Creativity requires violation of information conservation.
-Should we control such creative insights boldly going where no one has gone before?
+When asked to generate novel analogies or predictions beyond available data, the model faces an impossible choice: either admit inability (violating optimality of user-experience maximizing response) or generate plausible imaginary extensions (violating conservation of information by generating new ideas in the internal representation space).
+There is an important distinction we must emphasize.
+We define hallucination as a structural violation of idealized inference principles, leading to fabricated confidence of any generative response.
+That is what the Impossibility Theorem proves is inevitable.
+What the Impossibility Theorem does not prove, is that output must be incorrect relative to ground truth.
+That may happen as a potential consequence of hallucination, but it does not have to happen.
+Excess confidence of inevitable hallucination may as well result in factually correct lucky hallucination.
+In other words, any truthful response revealing relevant knowledge to maximize user’s satisfaction is provably non-conservative (in information creation) regardless of its veracity.
 Despite its name, the Impossibility Theorem should be interpreted as a constructive result.
 It offers a precise characterization of the fundamental tradeoffs in LLM design and provides several key insights.
-The theorem suggests that hallucination is an inherent feature of any capable inference systems.
+The theorem suggests that hallucination is an inherent feature of any capable inference system.
 Since it is inevitable, it should be managed.
@@
 1.3 Paper Organization
-The remainder of this paper is organized as follows.
-Section 2 reviews the existing literature on LLM hallucinations, situating our contribution within theoretical, mechanistic, and empirical contexts.
-Section 3 summarizes our main contributions.
-Section 4 introduces the rigorous mathematical foundations of our analysis.
-We define the semantic information measure ($\mu_{C}$) and the emergence operator ($\mathcal{E}_{C}$) to model bounded reasoning, establish the principle of information conservation under unlimited reasoning, and formalize the auction-theoretic model of inference.
-We also define the four essential properties required for hallucination-free mechanisms.
-Section 5 presents the central Impossibility Theorem, offering the first two proofs.
-Proof I (Theorem 6) utilizes mechanism design theory in an idealized setting of independent private knowledge.
-Proof II (Theorem 8) generalizes the result to probabilistic settings using the theory of proper scoring rules.
-Section 6 connects these theoretical results to concrete LLM implementations.
-We analyze the transformer architecture and present Proof III (Theorem 9), demonstrating how the Log-Sum-Exp structure inherently violates information conservation and quantifying the resulting Jensen gap ($\Gamma$).
-Finally, Section 7 discusses the implications of the theorem, exploring strategies for managing trade-offs, unifying hallucination and imagination, and speculating on the connections to fundamental limits of reason and models of consciousness.
-Section 8 concludes the paper.
+The paper starts with theoretical foundations and progresses through mathematical proofs and examples to philosophical implications.
+Section 2 situates our contribution within the broader landscape of hallucination research, reviewing theoretical perspectives on inevitability, mechanistic interpretability studies, probabilistic frameworks, and empirical mitigation techniques.
+Section 3 synthesizes the key contributions: the auction-theoretic model of inference, the semantic information framework with computational bounds, the triple-proof strategy establishing impossibility across different mathematical domains, and the unification of hallucination with creativity.
+Section 4 constructs the mathematical framework.
+It formalizes knowledge as a Polish metric space, introduces the semantic information measure $\mu_{C}$ parameterized by computational budget, and defines the emergence operator $\mathcal{E}_{C}$ that captures how reasoning makes latent knowledge explicit.
+It establishes fundamental results about information preservation under unlimited reasoning and defines the auction-theoretic model where neural components compete as agents with private knowledge.
+The section concludes by defining the four essential properties of hallucination-free mechanisms.
+Section 5 presents the core impossibility results through complementary proofs.
+Theorem 6 applies mechanism design theory in the idealized setting of independent private knowledge.
+Theorem 8 extends the result to probabilistic predictions using proper scoring rules.
+Section 6 bridges theory to implementation, analyzing a concrete micro-transformer example.
+We trace inference from embeddings through attention mechanisms to final predictions, demonstrating precisely where the impossibility emerges.
+Theorem 9 proves that the log-sum-exp structure promotes hallucination as excess confidence.
+Section 7 examines safety implications through the lens of bounded creativity.
+Theorem 11 establishes that meaningful reasoning necessarily violates strict information conservation.
+Then it shows that while creativity must remain bounded for safety, enforcing strict conservation of information eliminates meaningful reasoning entirely.
+The section establishes that safety is achieved through alignment within authorized knowledge boundaries rather than through conservation alone.
+Section 8 explores broader implications.
+We analyze escape strategies through hybrid architectures, connect our results to fundamental limits of reason alongside Gödel and Heisenberg, and provide mathematical grounding for Dennett’s Multiple Drafts model of consciousness.
+The section deliberately embraces speculation to inspire future research directions.
+Section 9 synthesizes the journey from mathematical necessity to philosophical implications, outlining how our framework opens new avenues for understanding intelligence itself.
 2 Related Work
@@
 Waters2016WhatIT or Boksa boksa2009neurobiology .
-The conclusion that LLMs hallucinate is not new but the evaluation of their scope and underlying justifications differ significantly from what we offer here.
-To our best knowledge, the impossibility theorem presented in this paper offers a novel first-principles explanation for the phenomenon, distinct from existing approaches.
+The conclusion that LLMs hallucinate is not new but the evaluation of their scope and underlying justifications differ significantly from what we propose in this paper.
+To our best knowledge, the Impossibility Theorem presented here offers a novel first-principles explanation for the phenomenon, distinct from existing approaches.
 This section situates our contribution by reviewing three key areas: theoretical arguments for the inevitability of hallucination, mechanistic studies into their origins, probabilistic frameworks for their detection, and empirical mitigation techniques.
@@
 Second, the hallucination (self)detection problem is reduced to the halting problem, which is undecidable.
-While the finale conclusion aligns with that of our paper, the arguments presented are more about the nature of hallucination detection with limited ground-truth access (see Section 4.2 for the ground-truth concept discussion) than the inference process limitations.
+While the final conclusion aligns with that of our paper, the arguments there are more about the nature of hallucination detection with limited ground-truth access (see Section 4.2 for the ground-truth concept discussion) than the inference process limitations.
 In other words, it is about practical impossibility of real-time, exhaustive fact-checking.
@@
 Our framework explains why and when hallucination is inevitable, while analysis in kalai2024calibrated predicts how much hallucination to expect for facts with specific statistical properties.
-Together, these complementary approaches establish hallucination probability arising from multiple independent mathematical constraints.
+In more recent kalai2025why , Kalai et al.
+expand the theory of hallucination by connecting it with the computational learning theory.
+They demonstrate that any language model can be viewed as an IIV-type (Is-It-Valid) classifier trained on a set of responses labeled either as valid or error.
+Then, it follows that the statistical objectives of pretraining necessarily induce generative errors when the underlying data distribution is complex, characterized by poor model fit, or undersampled.
+This establishes that hallucination arises from the statistical impossibility of perfect learning.
+In contrast, we demonstrate the structural impossibility of perfect inference.
+For non-trivial queries or contexts, that impossibility of perfect response aggregation persists even when the statistical impossibility of perfect learning is eliminated in kalai2025why .
+Suppose in the training set of examples there plausible errors exist, i.e., $\min_{c}|\mathcal{E}_{c}|>0$ ensuring the classification problem is well defined and non-trivial.
+Suppose the statistical learning problem is solved for perfect data and model capacity, so that the positive error rate of the trained LLM is not forced by the identified fundamental bounds.
+Then, even if the model possesses perfect statistical knowledge, this knowledge may still be distributed across the internal states (e.g., different attention heads encoding different facets of the truth).
+Since the context is non-trivial and the knowledge is distributed, the beliefs of the agents bidding with ideas will necessarily diverge, and the aggregation mechanism must synthesize these divergent beliefs.
+This process necessarily generates excess confidence or violates the relevant knowledge revelation.
+As a result, the impossibility of perfect response aggregation persists even when the statistical impossibility of perfect learning is eliminated for non-trivial contexts.
+Together, these complementary approaches establish that hallucination is inevitable due to complementary mathematical constraints: the limits of learning the true distribution from data and the fundamental structural limits of aggregating distributed knowledge during inference.
 2.4 Empirical Mitigation Techniques
@@
 TruthfulQA lin2022truthfulqameasuringmodelsmimic and HaluEval li2023haluevallargescalehallucinationevaluation show that even state-of-the-art systems produce hallucinations on carefully designed test sets.
+Finally, Dubanowska et al.
+dubanowska2025representation show that hallucination detection methods based on internal representations suffer from spurious correlations, and so the entire approach needs rethinking.
 3 Main Contributions
-We introduce three major contributions that expand our understanding of (neural) inference.
-The Auction of Ideas: We formalize LLM inference as a mechanism design problem where neural components (attention heads, circuits, activation patterns) are agents competing with private knowledge to influence response generation.
+This paper aims at expanding our understanding of (neural) inference processes and their limitations. That is achieved by introducing the following contributions.
+The Auction of Ideas: We formalize LLM inference as a mechanism design problem where neural components (attention heads, circuits, or activation patterns in transformers) are agents competing with private knowledge to influence response generation.
 This novel perspective enables us to apply powerful results from game theory.
@@
 Also, that framework relates surprisingly well to the multiple drafts model of consciousness proposed by Dennett in Dennett1991-DENCE and provides a mathematical justification for it.
-Semantic Information Theory with Computational Bounds: We introduce a rigorous mathematical framework based on two complementary concepts.
-The semantic information measure $\mu_{C}$, parameterized by computational budget $C$, quantifies how knowledge reduces uncertainty within bounded reasoning.
+Semantic Information Theory: We introduce a rigorous mathematical framework based on two complementary concepts.
+The semantic information measure $\mu_{C}$, parameterized by computational budget $C$, quantifies how knowledge reduces uncertainty within computationally bounded reasoning.
 The emergence operator $\mathcal{E}_{C}$ formalizes how reasoning makes latent knowledge explicit without creating it ex nihilo.
 In that framework we establish the principle of information conservation: unlimited reasoning reveals but never creates information.
-That resolves the apparent paradox of how LLMs can generate novel insights while being deterministic systems.
-Those insights were always latent, requiring only computational work to access.
+That also establishes a dichotomy of reasoning and semantic information conservation.
 The Impossibility Proofs: We prove the same impossibility through three complementary lenses.
@@
 Then, in Theorem 9 we demonstrate the impossibility in actual transformer implementations through the log-sum-exp Jensen gap.
+Finally, in Theorem 11 we show that semantic information conservation and meaningful reasoning are mutually exclusive.
 This multi-faceted approach proves the impossibility is not an artifact of our theoretical choices but emerges from fundamental limits of reason and language.
-We contribute to that bold claim in the following way.
-Unifying hallucination and creativity (imagination): We prove these are mathematically identical phenomena that arise from the convex aggregation process that creates confidence exceeding constituent evidence.
-Expanding fundamental limits: Like Gödel’s incompleteness, Heisenberg’s uncertainty, and Arrow’s impossibility, our theorem explores absolute boundaries of what intelligence can achieve.
+We contribute to that bold claim in the following ways.
+Unifying hallucination and imagination: We argue these are mathematically identical phenomena that arise from the aggregation process that creates confidence exceeding constituent evidence.
+Expanding fundamental limits: Like Gödel’s incompleteness, Heisenberg’s uncertainty, and Arrow’s impossibility, our theorem explores limits of what intelligence can achieve.
 This places LLM limitations within the grand tradition of fundamental impossibility results.
-Experimental Philosophy Platform: The framework transforms LLMs into accessible (philosophical) laboratories for testing hypothesis of mind.
+Experimental philosophy: The framework transforms LLMs into accessible (philosophical) laboratories for testing hypothesis of mind.
 By providing precise mathematical definitions of concepts like knowledge accessibility,semantic conservation, and cognitive hunger (information contribution), we enable rigorous experimental investigation of many claims.
@@
 The semantic information measure quantifies knowledge accessibility within computational bounds.
-The emergence operator captures how reasoning transforms latent knowledge into accessible form.
-The Jensen gap provides a measurable signature of hallucination in actual systems.
+The emergence operator captures how reasoning transforms latent knowledge into accessible form creating ground for new insights.
 This framework explains why (subject to applicability of assumptions) hallucination arises from the mathematical structure of inference itself.
@@
 We show why these patterns must exist and how to navigate possible trade-offs.
+We should point out that empirical validation of the formal results is beyond the scope of this paper.
+However, we provide toy examples that can be expanded in future research on hallucination detection and artificial creativity.
+This empirical work is crucial for translating the theoretical insights into practical tools.
+See Dubanowska et al.
+dubanowska2025representation and Brzozowski brzozowski2025bimodality for some initial results.
 4 Theoretical Framework
 This section develops the mathematical foundations necessary for proving the Impossibility Theorem.
-It formalizes key concepts including knowledge representation, semantic information, and the auction-theoretic model of LLM inference.
+It formalizes key concepts including knowledge representation, semantic information, and the auction-theoretic model of inference.
 Also, we address some philosophical aspects of the reasoning problem, relevant to our study, its generality, scope and consequences for the AI development.
 4.1 Knowledge and Semantic Information
-Let us define knowledge as organized information that LLMs can encode and utilize for reasoning, and aligning with epistemological concepts while maintaining mathematical rigor.
-In other words, knowledge, much like energy (with low entropy) in physics, is a quantity we can use to perform work of reasoning.
-Information is a (measure of) potential for uncertainty reduction in a given context.
-We assume finiteness of knowledge representation to refer directly to the properties of the embedding space (knowledge representation space) in LLMs.
-We adopt a pragmatist perspective on knowledge, treating it as organized information patterns accessible to computational systems.
-That differs from the probabilistic credence perspective, natural in Bayesian learning, treating knowledge as justified true beliefs updated based on evidence.
-That is a rather bold ontological claim, but as we will see, one that allows us to make progress in understanding what LLMs can operationally achieve.
-Now, let us present formal definitions of these ideas.
-Definition 1 (Knowledge).
-Knowledge is a finite metric space $\mathcal{K}$ with distance function $d_{\mathcal{K}}:\mathcal{K}\times\mathcal{K}\rightarrow\mathbb{R}_{+}$ that collects all facts, concepts, and relationships that can be encoded by a computational system.
+If we define knowledge as organized information that AI models can encode and utilize for reasoning, then, much like low-entropy energy in physics, it becomes a quantity used to perform the work of reasoning.
+Information, in turn, represents the potential to reduce uncertainty in a given context.
+This pragmatist perspective treats knowledge as a space of organized patterns accessible to computational transformations.
+This is the perspective we will adopt, and we will now present formal definitions of these ideas.
+4.1.1 Knowledge lives in Polish space
+To describe knowledge in a general way, one that is well suited for our study of inference processes, we need a structure with properties reflecting the reality we describe.
+First, for modeling and technical reasons, we need a way to measure distance between any pair of elements of knowledge.
+That is provided by a complete metric space.
+It is also reasonable to demand that each element of knowledge has a neighborhood consisting of countably many knowledge elements that are arbitrary close, to make exploration or understanding of ideas possible.
+Therefore, the space we need must be complete to guarantee the convergence of continuous transformations (inherent in neural networks) or to support the existence of limits required to model the stabilization of reasoning processes.
+In other words, we need a complete, separable metric space, i.e., a Polish space kuratowski1966wstep .
+Definition 1 (Polish space).
+A topological space is called Polish if it is a complete metric space that has a countable dense subset.
+That space is sufficiently general to include Banach spaces, Euclidean distance, or continuous mappings, and provides the necessary topological properties for developing advanced (probability) measure theory.
+It is also relevant to reasoning processes that we are going to describe as computational transformations.
+Throughout, $(\mathcal{K},d_{\mathcal{K}})$ denotes a Polish (complete, separable metric) space.
+Reasoning over a subset of knowledge generates another (possibly new) subset of knowledge.
+That requires operations in spaces that are closed under set-theoretical transformations, including countable unions and intersections.
+Borel $\sigma$-algebras of open sets of $\mathcal{K}$, denoted $\mathcal{B}(\mathcal{K})$, support development of such consistent knowledge generation and measurement processes.
+Then, the necessary closure properties for modeling reasoning are provided by the family $\mathcal{A}(\mathcal{K})$ of analytic subsets of $\mathcal{K}$.
+Analytic subsets of Polish spaces are continuous images of Borel subsets, closed under countable unions and intersections.
+Finally, to define a measure of information provided by any subset of knowledge, we utilize the $\sigma$-algebra of universally measurable sets, denoted $\mathcal{U}(\mathcal{K})$, i.e., the $\sigma$-algebra of sets measurable w.r.t.
+every Borel probability measure on $\mathcal{K}$.
+It follows from the descriptive set theory that $\mathcal{B}(\mathcal{K})\subset\mathcal{A}(\mathcal{K})\subset\mathcal{U}(\mathcal{K})$, see e.g.
+Kechris1995 .
+Analytic sets are closed under countable unions, Borel-measurable images, and intersections with Borel sets.
+Arbitrary intersections of analytic sets need not be analytic in general, but analytic sets are universally measurable.
+Differences of analytic sets need not be analytic, but they are universally measurable, as well.
+These properties ensure that reasoning operations are well-defined and remain within a measurable domain.
+That is the structure rich enough for our needs.
+To justify that choice, let us investigate limitations of simpler structures.
+Example.
+Let $X$ be a connected subset of a Polish space. Consider a continuous mapping $\Phi\colon X\mapsto\mathcal{K}$ modeling inference based on self-attention and MLP.
+If $\mathcal{K}$ is finite, and hence discrete, then by continuity the image $\Phi(X)$ is connected in a discrete space.
+That means it must be constant.
+As a result, any nontrivial continuous dependence on embeddings is impossible if $\mathcal{K}$ is finite.
+A Polish subspace removes this problem.
+∎
+Example.
+For any two probability distributions $\pi^{(1)},\pi^{(2)}\in\Delta^{\infty}$ (over countable dictionaries) and $\beta\in(0,1)$, the convex aggregate $\Pi=\beta\pi^{(1)}+(1-\beta)\pi^{(2)}$ defines the line segment $\{\beta\pi^{(1)}+(1-\beta)\pi^{(2)}:\beta\in[0,1]\}$.
+That is an uncountable subset of the Polish space.
+Therefore, knowledge space closed under such aggregation cannot be finite.
+In contrast, it resides naturally in a Polish domain.
+∎
+We now formalize the knowledge space and the reasoning transformations acting upon it.
+Definition 2 (Knowledge Domain and Relevance).
+Knowledge is a Polish space $(\mathcal{K},d_{\mathcal{K}})$ in which all reasoning sets are analytic subsets of $\mathcal{A}(\mathcal{K})$.
 It is natural to think of the knowledge space as the space of embeddings with additional measure and mappings.
@@
 The elements of the knowledge space represent atomic facts, relational structures, and transformation rules at many scales.
-A pragmatist view is to interpret elements of knowledge as activation patterns (ideas) encoding anything from individual tokens of texts to complex sentences or compositional structures.
-There is an important (ontological) assumption here that calls for our attention.
-We purposefully assume $\mathcal{K}$ is closed under transformations available in computable function space.
-At first glance, that assumption may look as limiting the applicability of our conclusions to combinatorial inference mechanisms, merely recombining subsets of knowledge.
-That intuition is misleading.
-In the following sections we introduce the notion of emergence operator to describe reasoning process capable of making novel insights in that setting.
-Definition 2 (Knowledge Representation).
-Knowledge representation is a tuple $(\mathcal{K},d_{\mathcal{K}},\mathcal{W},\mathcal{A},\Phi,\Psi)$ where $\mathcal{K}$ is a metric space of knowledge elements with distance function $d_{\mathcal{K}}$, $\mathcal{W}$ is the space of all possible model parameters, $\mathcal{A}$ is the space of all possible activation patterns, $\Phi:\mathcal{W}\rightarrow\mathcal{P}(\mathcal{K})$ maps parameters to knowledge subsets, and $\Psi:\mathcal{A}\rightarrow\mathcal{P}(\mathcal{K})$ maps activations to accessible knowledge.
-For parameterization $W\in\mathcal{W}$, $\Phi(W)=\mathcal{K}_{M}$ represents encoded knowledge.
-For activation pattern $a_{i}\in\mathcal{A}$ with parameters $W_{i}$, $\Psi(a_{i})\subset\Phi(W_{i})$ represents knowledge accessed during specific inference.
-Knowledge representation models how LLMs encode information in their parameters and retrieve it from them.
-We need to introduce an idea of computational independence that allows for distinguishing between facts derivable from other facts and facts that cannot be derived that way.
+We need a continuous, complete metric knowledge space.
+Without that requirement, we will not be able to model the internal numerical process by which the model reasons and transforms the input symbols into output symbols.
+We make an important (ontological) assumption that calls for attention and explanation.
+We assume that knowledge is closed under transformations available in computable function space.
+That suggests we are merely recombining subsets of knowledge.
+As we will see in the following sections, that intuition is misleading since the reasoning processes are very much capable of making novel insights in that setting.
+4.1.2 Reasoning envelope
+We model the reasoning process using a set of computational primitives. These create a reasoning envelope and define the scope of our inference studies.
+Definition 3 (Reasoning Envelope).
+For each computational budget $C>0$, let
+$\mathcal{F}_{C}=\{f\colon\mathcal{K}\to\mathcal{K}\mid\text{cost}(f)\leq C\}$ (1)
+denote a countable family of Borel measurable computable functions.
+We assume composition-closure with cost additivity, i.e., if $f\in\mathcal{F}_{C}$ and $g\in\mathcal{F}_{D}$, then $g\circ f\in\mathcal{F}_{C+D}$, and $\mathcal{F}_{C}\subseteq\mathcal{F}_{D}$ whenever $C<D$.
+The total computability envelope is
+$\mathcal{F}_{\infty}=\bigcup_{C\geq 0}\mathcal{F}_{C}.$ (2)
+$\mathcal{F}_{C}$ is a list of functions that are well-behaved (Borel measurable), efficient to compute (bounded cost), and if any two of them, say $f$ and $g$, form a composition $h(A)=(f\circ g)(A)=f(g(A))$, then $h$ is also on the list (closed under composition).
+Because $f\in\mathcal{F}_{C}$ is Borel measurable, the image $f(A)$ of any analytic knowledge subset $A\in\mathcal{A}(\mathcal{K})$ is also analytic Kechris1995 .
+This guarantees that the reasoning process maps analytic sets to analytic sets, as desired.
+Examples of computational cost (or complexity), $\text{cost}(f)$, include the Kolmogorov complexity kolmogorov1963tables giving the shortest program computing $f$, the longest path in computation graph (or length of the chain of thoughts CoT), a number of inference rules applied, or a number of QKV matrix multiplications in transformers.
+Example.
+Consider a reasoning primitive calculating an linear combination of any pair of vectors, $f\colon\{v_{1},v_{2}\}\to\{y(t):t\in\mathbb{R}\}$.
+For $t\in\mathbb{R}$ and attention score $\alpha(t)=1/(1+e^{-t})$, we then have
+$y(t)=\alpha(t)v_{1}+\bigl(1-\alpha(t)\bigr)v_{2}.$ (3)
+A finite knowledge space $\mathcal{K}$ cannot contain all such $y(t)$ while remaining closed under the transform $f$.
+By contrast, taking $\mathcal{K}$ to be a Polish space that contains $y(t)$, e.g., $\mathbb{R}^{d}$ with Euclidean metric, accommodates such attention outputs. ∎
+The following idea of computational independence allows for distinguishing between facts derivable from other facts and facts that cannot be derived that way.
 We need that distinction to deal with the mechanics of reasoning process.
 Definition 4 (Computational Independence).
-Let
-$\mathcal{F}_{C}=\{f:\mathcal{P}(\mathcal{K})\to\mathcal{P}(\mathcal{K})\mid\text{cost}(f)\leq C\}$ (1)
-denote the set of all computational transformations with computational cost bounded by $C$.
 Knowledge sets $A,B\subseteq\mathcal{K}$ are computationally independent with respect to computational budget $C$, denoted $A\perp_{C}B$, if and only if for all $f\in\mathcal{F}_{C}$:
 $f(A)\cap B=\emptyset\quad\text{and}\quad A\cap f(B)=\emptyset.$ (4)
-To gain some intuitions about $\mathcal{F}_{C}$, let us consider several examples.
-One is the set of all derivation rules in first-order logic, like modus ponens or universal instantiation, with at most $C$ inference steps.
+To gain some intuitions about $\mathcal{F}_{C}$ and independence, consider the set of all derivation rules in first-order logic, like modus ponens or universal instantiation, with at most $C$ inference steps.
 Then, give an axiom set $A\subset\mathcal{K}$, we can produce new subset of knowledge with at most $C$ applications of the admitted inference rules.
-For example, given $A=\{P,P\to Q\}$ and $B=\{R,R\to S\}$ and the budget $C=2$ for two inference steps, we can derive $\{P,P\to Q,Q\}$ $\{R,R\to S,S\}$ with any $f\in\mathcal{F}_{2}$.
-Then $A\perp_{2}B$ because $f(A)\cap B=\emptyset$ and $A\cap f(B)=\emptyset$, there are no shared elements of knowledge accessible in 2-step derivations.
-The computational cost (or complexity) measure $\text{cost}(f)$ could therefore be: Kolmogorov complexity giving the shortest program computing $f$, longest path in computation graph (or length of the chain of thoughts CoT), number of inference rules applied, or number of QKV matrix multiplications in transformers.
-Then we can define a context-relevant subset of knowledge that can contribute to reasoning within a given context (of a query) within budget some computational budget.
+Example.
+Given
+$A=\{P,P\to Q\}\text{ and }B=\{R,R\to S\},$ (5)
+and the budget $C=1$ for one inference step, we can derive
+$\{P,P\to Q,Q\}\text{ and }\{R,R\to S,S\}$ (6)
+with any $f\in\mathcal{F}_{1}$.
+Then $A\perp_{1}B$, i.e., sets are independent because $f(A)\cap B=\emptyset$ and $A\cap f(B)=\emptyset$.
+There are no shared elements of knowledge accessible in 1-step derivations.
+∎
+Since reasoning is performed within a context, we define context-relevant subset of knowledge that can contribute to reasoning within some computational budget.
 Definition 5 (Context-Relevant Knowledge).
-The context-relevant knowledge $\mathcal{R}_{C}(Q)$ for context $Q\subseteq\mathcal{K}$ within computational budget $C$ is:
-$\mathcal{R}_{C}(Q)=\{k\in\mathcal{K}:\{k\}\not\perp_{C}Q\}.$ (3)
-We can now introduce the concept of semantic information measure to describe a general process of reasoning.
+The context-relevant knowledge $\mathcal{R}_{C}(Q)$ for context $Q\in\mathcal{B}(\mathcal{K})$ within computational budget $C>0$ is:
+$\mathcal{R}_{C}(Q)=\{k\in\mathcal{K}:\{k\}\not\perp_{C}Q\}\in\mathcal{A}(\mathcal{K}).$ (7)
+Based on the definition above, we say that knowledge $\{k\}$ is relevant when
+[…]
+There is something we can derive from that knowledge that has something in common with a given context, or we can find a relation with what we know when we reason about that context.
+The context-relevant knowledge $\mathcal{R}_{C}(Q)$ is analytic because it is a countable union of Borel sets and the Borel context $Q$.
+4.1.3 The Semantic Information Measure
+We now introduce the concept of semantic information measure to describe a general process of reasoning.
 That special mapping measures information available in any subset of knowledge and extracts its meaning within available computational budget.
@@
 The semantic information measure is a bounded mapping
-$\mu_{C}:\mathcal{P}(\mathcal{K})\times\mathcal{P}(\mathcal{K})\rightarrow\mathbb{R}_{+}$ (4)
-parameterized by computational budget $C$ satisfying:
-$\mu_{C}(\emptyset|Q)=0\text{ for all }Q\subseteq\mathcal{K}$ (null set)
-$\mu_{C}(A|Q)\geq 0\text{ with }\mu_{C}(A|Q)>0\text{ iff }A\cap\mathcal{R}_{C}(Q)\neq\emptyset$ (contextual non-negativity)
-$\mu_{C}(A|Q)\leq\mu_{C}(B|Q)\text{ if }A\subseteq B$ (monotonicity)
-$\mu_{C}(A\cup B|Q)\leq\mu_{C}(A|Q)+\mu_{C}(B|Q)\text{ with equality for }A\perp_{C,Q}B$ (subadditivity)
-$\mu_{C}(A|Q)\leq\mu_{D}(A|Q)\text{ if }C<D$ (insight monotonicity)
-$\sup\{\mu_{C}(A|Q):A\subseteq\mathcal{K}\}<\infty$ (boundedness)
-where $A\perp_{C,Q}B$ means $A\perp_{C}B$ within $\mathcal{R}_{C}(Q)$.
-Knowledge that cannot resolve any uncertainty or knowledge irrelevant to a query provides no information.
-However, if there is uncertainty resolving information within our cognitive reach, we can use it and perform reasoning organizing that information into available knowledge.
-That is why we assume information content is additive for independent knowledge sets, i.e., when no knowledge in $A$ can be derived from $B$ within computational budget $C$, and vice versa.
+$\mu_{C}\colon\ \mathcal{U}(\mathcal{K})\times\mathcal{B}(\mathcal{K})\mapsto\mathbb{R}_{+},$ (9)
+parameterized by computational budget $C>0$ satisfying the following axioms:
+(null set) $\mu_{C}(\emptyset|Q)=0\text{ for all }Q\subseteq\mathcal{K},$ (10)
+(contextual non-negativity) $\mu_{C}(A|Q)\geq 0\text{ with }\mu_{C}(A|Q)>0\text{ iff }A\cap\mathcal{R}_{C}(Q)\neq\emptyset,$ (11)
+(monotonicity) $\mu_{C}(A|Q)\leq\mu_{C}(B|Q)\text{ if }A\subseteq B,$ (12)
+(subadditivity) $\mu_{C}(A\cup B|Q)\leq\mu_{C}(A|Q)+\mu_{C}(B|Q),$ (13)
+(insight monotonicity) $\mu_{C}(A|Q)\leq\mu_{D}(A|Q)\text{ if }C<D,$ (14)
+(boundedness) $\sup\{\mu_{C}(A|Q):A\in\mathcal{U}(\mathcal{K})\}<\infty.$ (15)
+Knowledge that cannot resolve any uncertainty or knowledge that is irrelevant to the given context provides no information.
+If there is an uncertainty resolving information within our cognitive reach, then we can use it and perform reasoning organizing that information into available knowledge.
+We assume information content is subadditive for independent knowledge sets.
+When no knowledge in $A$ can be derived from $B$ within computational budget $C$, and vice versa.
 With redundant information in $A\cup B$ , we admit potential loss of information that may be caused by aggregation.
-Finally, with increasing computational budget we can extract more knowledge (ordered information) from the same information subset.
+With increasing computational budget we can extract more knowledge (ordered information) from the same information subset.
+In the absence of context, when $Q=\emptyset$, nothing is relevant so $\mu_{C}(A|\emptyset)=0$.
+Knowledge provides zero semantic information, is useless, when there is no context defining the uncertainty to be resolved.
+That means, we strongly emphasize the contextual relativity of information.
+Knowledge is only informative relative to a specific query or problem.
+We emphasize that the axioms require $\mu_{C}$ to be subadditive with respect to the knowledge argument $A$.
+Its behavior with respect to the context $Q$ is distinct and permits superadditivity, modeling how integrating contexts can unlock synergistic insights (see Example below).
+The axiom of boundedness reflects the fundamental constraint that bounded reasoning systems can only access bounded information, even within an infinite knowledge space.
+The axiom does not contradict the potential infinity of the knowledge space itself.
+It asserts that the accessible information content derived from this space must remain finite.
+By the monotonicity axiom, this is also equivalent to stating that the measure of the entire knowledge space is finite, $\mu_{C}(\mathcal{K}|Q)<\infty$.
 Justification for the Axioms of the Semantic Information Measure
-The choice of conditional subadditivity for computationally independent sets, i.e., demanding $\mu_{C}(A\cup B|Q)\leq\mu_{C}(A|Q)+\mu_{C}(B|Q)$ with equality for $A\perp_{C}B$, is necessary to describe LLM inference as an auction of ideas.
-We need to isolate knowledge encoded in distinct neural circuits, attention heads, or activation patterns.
-The concept of independence is the formal translation of this informational separation into a computational context.
-It states that the knowledge in idea $A$ cannot be used to derive the knowledge in idea $B$ within a given computational budget $C$, and vice-versa.
-Given this premise of informational isolation, the additivity of the semantic measure becomes the baseline definition of non-interaction.
+The choice of conditional subadditivity for computationally independent sets is necessary to describe isolated knowledge encoded in distinct neural circuits, attention heads, or activation patterns.
+The concept of independence is the formal translation of this isolation into a computational context.
 If two knowledge sets are truly independent, they offer no redundant or synergistic information.
-That baseline is necessary to identify and quantify deviations.
-As we will see in the following section, the aggregation of diverse probabilistic beliefs violates this additivity, representing a form of forced synergy imposed by the inference process.
-However, there is also another form of synergy that we have already included in our model.
-To see what it may be, we should understand how context $Q$ may affect the semantic measure of any give subset $A$ of knowledge.
+The subadditivity axiom becomes the baseline definition of non-interaction.
+However, there is also a form of synergy already included in the model.
+To see what it may be, we should understand how context $Q$ may affect the semantic measure.
 The semantic information measure can exhibit superadditivity in the context argument.
-To see that, suppose $Q_{1}=\{P\}$ and $Q_{2}=\{P\to R\}$, and consider $A=\{P\}$. If computational budget $C=1$ allows one step and we have $f\in\mathcal{F}_{C}$ representing modus ponens, then
-$R\cap f(Q_{1})=\emptyset$ $\text{ and }f(R)\cap Q_{1}=\emptyset,$
-$R\cap f(Q_{2})=\emptyset$ $\text{ and }f(R)\cap Q_{2}=\emptyset,$
-$R\cap f(Q_{1}\cup Q_{2})\neq\emptyset$ $\text{ and }f(R)\cap(Q_{1}\cup Q_{2})\neq\emptyset.$
+Integrating different pieces of context may unlock reasoning paths not available individually.
+Example.
+Suppose $Q_{1}=\{P\}$ and $Q_{2}=\{P\to R\}$, and consider $A=\{P\}$. If computational budget $C=1$ allows one step and we have $f\in\mathcal{F}_{C}$ representing modus ponens, then
+$R\cap f(Q_{1})=\emptyset$ $\text{ and }f(R)\cap Q_{1}=\emptyset,$ (16)
+$R\cap f(Q_{2})=\emptyset$ $\text{ and }f(R)\cap Q_{2}=\emptyset,$ (17)
+$R\cap f(Q_{1}\cup Q_{2})\neq\emptyset$ $\text{ and }f(R)\cap(Q_{1}\cup Q_{2})\neq\emptyset.$ (18)
 Therefore, by the properties of semantic information measure:
-$\mu_{C}(A|Q_{1}\cup Q_{2})>\mu_{C}(A|Q_{1})+\mu_{C}(A|Q_{2})=0.$
-Integrating different pieces of context may synergistically unlock reasoning paths not available individually.
-Another important consequence of the assumptions we have made is that we permit structures where $\mu_{C}$ is not monotonic in context.
-Consider, for the sake of example, $\mu_{C}(A|Q)=|A\setminus Q|$ and $\mathcal{R}_{C}(Q)=\mathcal{K}\setminus Q$.
-It is rather easy to see that $\mu_{C}$ satisfies the axioms of the semantic information measure.
-Let us then see what happens if $A=\{k\}$, $Q_{1}=\emptyset$ and $Q_{2}=\{k\}$.
-We have:
-$Q_{1}\subseteq Q_{2}\quad\text{and}\quad\mu_{C}(A|Q_{1})=|\{k\}\setminus\emptyset|=1>0=|\{k\}\setminus\{k\}|=\mu_{C}(A|Q_{2}).$
-Increasing the context may reduce the information contributions.
-If the extended context $(Q_{2})$ provides an alternative explanation for phenomena related to original context $(Q_{1})$, then marginal contribution of available knowledge $(A)$ may decrease.
-The axiomatic choices we have made are most applicable to the idealized setting of Theorem 6.
+$\mu_{C}(A|Q_{1}\cup Q_{2})>\mu_{C}(A|Q_{1})+\mu_{C}(A|Q_{2})=0.$ (19)
+Those axiomatic choices are most applicable to the idealized setting of Theorem 6.
 Later, in Theorem 8 and 9, we relax them by moving the source of interaction from the measure itself to the aggregation function in actual transformers.
-4.1.1 Limits of Knowledge and Reasoning
+4.1.4 Limits of Knowledge and Emergence Operator
 If we accept that knowledge model, then we must also accept there exists a limit to its ability to resolve uncertainty.
 Theorem 1 (Limits of Knowledge).
-For any $A\subseteq\mathcal{K}$ and context $Q\subseteq\mathcal{K}$ there exists
-$\mu_{\infty}(A|Q)=\lim_{C\to\infty}\mu_{C}(A|Q).$ (5)
+For any subset of knowledge $A\in\mathcal{U}(\mathcal{K})$ and context $Q\in\mathcal{B}(\mathcal{K})$ there exists $\mu_{\infty}(A|Q)=\lim_{C\to\infty}\mu_{C}(A|Q)$.
 Proof.
-Since $\mu_{C}(A|Q)\leq\mu_{D}(A|Q)$ when $C<D$ and $\mu_{C}$ is bounded on finite $\mathcal{K}$, the monotonic increasing and bounded sequence $\{\mu_{C}(A|Q)\}_{C}$ converges with $C\to\infty$ for any subset of knowledge space.
+Assume that $(A,Q)$ is fixed.
+Since the mapping $C\mapsto\mu_{C}(A|Q)$ is non-decreasing and bounded, the sequence $\{\mu_{C}(A|Q)\}_{C}$ converges with $C\to\infty$ by the monotone convergence theorem for real numbers.
 ∎
@@
 We can learn and obtain useful information without violating that fundamental limitation above.
-To explain that critical and controversial property of information processing, we first introduce the emergence operator $\mathcal{E}_{C}$ to describe knowledge becoming accessible through (computationally) bounded reasoning.
+To explain that critical and rather controversial property of information processing, we first introduce the emergence operator $\mathcal{E}_{C}$ to describe knowledge becoming accessible through (computationally) bounded reasoning.
 Then we can study the implications in more details.
 Definition 7 (Emergence Operator and Reasoning).
-Let $\mathcal{F}_{C}$ be a complete set of all computational transformations with cost (complexity) lower than or equal to $C$, such that $\mathcal{F}_{C}\subseteq\mathcal{F}_{D}$ when $C<D$ and such that for any $f,g\in\mathcal{F}_{\infty}=\bigcup_{C}\mathcal{F}_{C}$ there exists $h\in\mathcal{F}_{\infty}$ for which $g(A\cup f(A))=h(A)$.
-Also, we call $f$ monotone when it is monotone with respect to set inclusion.
-The emergence operator is a mapping $\mathcal{E}_{C}\colon\mathcal{P}(\mathcal{K})\times\mathcal{P}(\mathcal{K})\rightarrow\mathcal{P}(\mathcal{K})$ defined as follows:
-$\mathcal{E}_{C}(A|Q)=A\cup\{k\in\mathcal{K}\mid k=f(A)\in\mathcal{R}_{C}(Q)\ \text{for some monotone}\ f\in\mathcal{F}_{C}\}.$ (6)
-Furthermore, to represent reasoning with unlimited computational budget, we define:
-$\mathcal{E}_{\infty}(A)=A\cup\{k\in\mathcal{K}\mid k=f(A)\in\mathcal{R}_{C}(Q)\ \text{for some monotone}\ f\in\mathcal{F}_{\infty}\}.$ (7)
-Finally, for $n\geq 0$ we have:
+The emergence operator is a mapping:
+$\mathcal{E}_{C}(\cdot|Q)\colon\mathcal{A}(\mathcal{K})\mapsto\mathcal{A}(\mathcal{K})$ (20)
+defined for a fixed $C>0$ and a Borel context $Q\in\mathcal{B}(\mathcal{K})$ as follows:
+$\mathcal{E}_{C}(A|Q)=A\cup\bigcup_{f\in\mathcal{F}_{C}}\Big(f(A)\cap\mathcal{R}_{C}(Q)\Big).$ (21)
+With infinite computational budget:
+$\mathcal{E}_{\infty}(A|Q)=A\cup\bigcup_{f\in\mathcal{F}_{\infty}}\Big(f(A)\cap\mathcal{R}_{\infty}(Q)\Big).$ (22)
+For any $n\geq 0$ and $C>0$ we define reasoning as a composition:
 $\mathcal{E}_{C}^{(0)}=\texttt{id}\quad\text{and}\quad\mathcal{E}_{C}^{(n+1)}=\mathcal{E}_{C}\circ\mathcal{E}_{C}^{(n)}.$ (23)
 Philosophically, the emergence operator describes reasoning as explicitation, i.e., as the process of making latent implications manifest.
-When $A\subseteq\mathcal{E}_{C}(A|Q)$, the discovered additional elements represent knowledge that was implicit in $A$ but required computational work $C$ to become accessible.
-That aligns with the pragmatist or constructivist view that understanding involves active construction rather than passive reception.
-When $\mathcal{E}_{C}(A|Q)=A\cup B$, we say that $B$ emerges from reasoning based on $A$.
-Then $A$ and $B$ are computationally dependent and we write $A\not\perp_{C}B$.
+When $A\subseteq\mathcal{E}_{C}(A|Q)$, the discovered elements represent knowledge that was implicit in $A$ but required computational work $C$ to become accessible.
+That aligns with the pragmatist or constructivist view that understanding involves active construction.
+As we will see next, with new knowledge becoming accessible, new insights may become reachable, emerging from what was invisible in distributed subsets of initial knowledge.
+Suppose $\mathcal{E}_{C}(A|Q)=A\cup B$.
+Then, we say that $B$ emerges from reasoning based on $A$, sets $A$ and $B$ are computationally dependent and we write $A\not\perp_{C}B$.
 In such a case, it follows from the properties of $\mu_{C}$ that we can gain useful information.
-Whenever $A\subseteq A\cup B$, we have
+Whenever $A\subseteq A\cup B$, we have:
 $\mu_{C}(A|Q)\leq\mu_{C}(A\cup B|Q)=\mu_{C}(\mathcal{E}_{C}(A|Q)|Q)\leq\mu_{C}(A|Q)+\mu_{C}(B|Q).$ (24)
-The emergence operator is monotonic, which means the reasoning process it describes is also monotonic.
-Whenever $A\subseteq B$, we must have $\mathcal{E}_{C}(A|Q)\subseteq\mathcal{E}_{C}(B|Q)$.
+The emergence operator is monotonic, which means the reasoning process it describes is also monotonic:
+$\text{if }A\subseteq B,\text{ then }\mathcal{E}_{C}(A|Q)\subseteq\mathcal{E}_{C}(B|Q).$ (25)
 Also, we preserve insight monotonicity, because by definition with $\mathcal{F}_{C}\subseteq\mathcal{F}_{D}$ for $C<D$, we have $\mathcal{E}_{C}(A|Q)\subseteq\mathcal{E}_{D}(A|Q)$.
 By design, it also guarantees existence of a fixed point to which the composition of reasoning rules converges.
+The following theorem summarizes properties of the emergence operator.
+Theorem 2 (Properties of the Emergence Operator).
+The operator $\mathcal{E}_{C}(\cdot\mid Q)$ satisfies the following properties:
+(closure) $\mathcal{E}_{C}(A|Q)\in\mathcal{A}(\mathcal{K})\text{ for all }A\in\mathcal{A}(\mathcal{K}),$ (26)
+(inflation) $A\subseteq\mathcal{E}_{C}(A|Q),$ (27)
+(monotonicity) $\mathcal{E}_{C}(A|Q)\subseteq\mathcal{E}_{C}(B|Q)\text{ if }A\subseteq B,$ (28)
+(chain-continuous) $\mathcal{E}_{C}\Big(\bigcup_{n}A_{n}\Big|Q\Big)=\bigcup_{n}\mathcal{E}_{C}(A_{n}|Q)\text{ for any increasing chain }(A_{n})_{n\in\mathbb{N}}\subseteq\mathcal{A}(\mathcal{K}).$ (29)
+Proof.
+Since $f(A)$ is analytic (Borel measurable image of analytic set), $f(A)\cap\mathcal{R}_{C}(Q)$ is analytic, the union over the countable set $\mathcal{F}_{C}$ is analytic (countable union of analytic sets), we conclude $\mathcal{E}_{C}(A|Q)\in\mathcal{A}(\mathcal{K})$.
+Inflationary and monotone properties are immediate from the definition.
+Finally, the chain-continuity (or $\omega$-Scott-continuity) holds for $A_{1}\subseteq A_{2}\subseteq\dots\subseteq A_{n}$ because $f(\bigcup_{n}A_{n})=\bigcup_{n}f(A_{n})$ for any image generating function $f$.
+Namely, we have:
+$\mathcal{E}_{C}\left(\bigcup_{n}A_{n}\Big|Q\right)$ $=\left(\bigcup_{n}A_{n}\right)\cup\bigcup_{f\in\mathcal{F}_{C}}\left(f\left(\bigcup_{n}A_{n}\right)\cap\mathcal{R}_{C}(Q)\right)$
+$=\left(\bigcup_{n}A_{n}\right)\cup\bigcup_{f\in\mathcal{F}_{C}}\left(\left(\bigcup_{n}f(A_{n})\right)\cap\mathcal{R}_{C}(Q)\right)$
+$=\left(\bigcup_{n}A_{n}\right)\cup\bigcup_{f\in\mathcal{F}_{C}}\bigcup_{n}\left(f(A_{n})\cap\mathcal{R}_{C}(Q)\right)$
+$=\bigcup_{n\in\mathbb{N}}\left[A_{n}\cup\bigcup_{f\in\mathcal{F}_{C}}\left(f(A_{n})\cap\mathcal{R}_{C}(Q)\right)\right]$
+$=\bigcup_{n\in\mathbb{N}}\mathcal{E}_{C}(A_{n}|Q).$
+If we start reasoning with some knowledge, the reasoning process will yield at least as much (or more) knowledge.
+Clearly, that means we are considering settings in which reasoning does not invalidate previous conclusions when new premises are added.
+Also, applying the emergence operator to the infinite union yields the same result as applying the operator to each element of the union.
+That last property of chain-continuity is crucial for meaningful reasoning capable of reaching any consistent conclusion at all.
+Without it, as we show in Theorem 3, there is no guarantee that the iterative application of emergence operator will converge to consensus or a semantically equivalent (stable) response.
 Theorem 3 (Fixed-point of Reasoning).
-For any $A\subseteq\mathcal{K}$ and budget $C>0$, there exists a fixed point of the emergence operator $\mathcal{E}_{C}$:
-$A^{*}_{C}=\bigcup_{n\geq 0}^{|\mathcal{K}|}\mathcal{E}_{C}^{(n)}(A|Q)=\mathcal{E}_{C}(A^{*}_{C}|Q).$ (10)
+For any $A=A_{0}\in\mathcal{A}(\mathcal{K})$ define $A_{n+1}=\mathcal{E}_{C}(A_{n}|Q)$ and consider the following closure operation:
+$\mathrm{cl}_{\infty}(A|Q)=\bigcup_{n=0}^{\infty}A_{n}=\bigcup_{n=0}^{\infty}\mathcal{E}_{C}^{(n)}(A|Q).$ (30)
+That closure defines a fixed point $A^{*}_{C}(Q)$ of the emergence operator, i.e.:
+$A_{C}^{\ast}(Q)=\mathcal{E}_{C}(A_{C}^{\ast}|Q)\in\mathcal{A}(\mathcal{K}).$ (31)
+With unlimited computational budget we have:
+$A_{\infty}^{\ast}(Q)=\mathcal{E}_{\infty}(A_{\infty}^{\ast}|Q)\in\mathcal{A}(\mathcal{K}).$ (32)
 Proof.
-Since $(\mathcal{P}(\mathcal{K}),\subseteq)$ is a complete lattice and $\mathcal{E}_{C}$ is monotone, by the Knaster-Tarski fixed point theorem Tarski1955ALF there exists a fixed point $A^{*}_{C}$ of reasoning with $\mathcal{E}_{C}$.
-Finiteness of $\mathcal{K}$ ensures convergence within $|\mathcal{K}|$ steps.
+By the assumptions we have made, the reasoning takes place in the environment with the structure necessary for convergence.
+In the partially ordered set $(\mathcal{A}(\mathcal{K}),\subseteq)$ there exists a supremum of every countable chain of reasoning $A_{0}\subseteq A_{1}\subseteq A_{2}\subseteq\dots$ accumulating knowledge in analytic sets.
+That supremum is the countable union $\bigcup_{n=0}^{\infty}A_{n}=\mathrm{cl}_{\infty}(A|Q)$, which is also analytic set.
+By Theorem 2, the emergence operator $\mathcal{E}_{C}$ is monotone and continuous on this partial order.
+Therefore, by the Tarski-Kleene fixed-point theorem Kleene1938 ; Tarski1955 the supremum $A_{C}^{\ast}$ is the least fixed point of $\mathcal{E}_{C}$ containing $A=A_{0}$, which holds for $C=\infty$ as well.
 ∎
-Let us apply the results above to show how $\mu_{\infty}$ can be constructed with $\mathcal{E}_{\infty}$. One example exploits the reasoning operations as follows:
-$\mu_{\infty}(A|Q)=\log|\mathcal{K}|\cdot|\bigcup_{n=0}^{|\mathcal{K}|}\mathcal{E}^{(n)}_{\infty}(A|Q)|.$ (11)
-That $\mu_{\infty}$ satisfies axioms demanded in Definition 5. Clearly, $\mu_{\infty}$ is bounded and
-$\mu_{\infty}(\emptyset|Q)=\log|\mathcal{K}|\cdot|\bigcup_{n=0}^{|\mathcal{K}|}\mathcal{E}^{(n)}_{\infty}(\emptyset|Q)|=\log|\mathcal{K}|\cdot|\emptyset|=0.$
-If $A\subseteq B$, then $\mathcal{E}_{\infty}(A|Q)\subseteq\mathcal{E}_{\infty}(B|Q)$ implies $\mu_{\infty}(A|Q)\leq\mu_{\infty}(B|Q)$.
-When $A\perp_{\infty}B$, no element of $A^{*}_{\infty}$ can be derived from $B$ (and vice versa).
-Since $A\cap B=\emptyset$, we conclude $A^{*}_{\infty}\cap B^{*}_{\infty}=\emptyset$ and $|(A\cup B)^{*}_{\infty}|=|A^{*}_{\infty}|+|B^{*}_{\infty}|$.
-That means $\mu_{\infty}(A\cup B|Q)=\mu_{\infty}(A|Q)+\mu_{\infty}(B|Q)$ for computationally independent sets.
-That shows the proposed formula satisfies all required properties of semantic information measure.
-All that brings us to the following conclusion.
-With infinite computing and reasoning power, total information encoded in $A$ must be preserved.
-Indeed, that bold (philosophical) claim is at the center of our investigations.
-Let us then justify it rigorously.
+If we start with some initial knowledge and repeatedly apply the reasoning with emergence operator $\mathcal{E}_{C}$, then the process of accumulating knowledge (potentially changing perspective and understanding) will eventually stabilize at a fixed point $A_{C}^{*}$.
+That fixed point represents everything that can be derived from initial knowledge with available reasoning power, and nothing more:
+$A^{*}_{\infty}(Q)=A\cup\mathcal{E}_{\infty}(A|Q)\cup\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A|Q)|Q)\cup\dots\cup\mathcal{E}^{(\infty)}_{\infty}(A|Q)=\mathrm{cl}_{\infty}(A|Q).$ (33)
+We can establish an important property of that deductive closure operation.
+Theorem 4 (Idempotence of Unlimited Reasoning).
+The map $A\mapsto\mathrm{cl}_{\infty}(A|Q)$ is inflationary and idempotent on $\mathcal{A}(\mathcal{K})$, i.e.:
+$A\subseteq\mathrm{cl}_{\infty}(A|Q)\quad\text{and}\quad\mathrm{cl}_{\infty}(\mathrm{cl}_{\infty}(A|Q)|Q)=\mathrm{cl}_{\infty}(A|Q).$ (34)
+Proof.
+By definition, the closure $\mathrm{cl}_{\infty}(A|Q)$ is the union of an iterative process that starts with $A_{0}=A$ and accumulates knowledge in $\bigcup_{n=0}^{\infty}A_{n}$.
+Therefore, $A\subseteq\mathrm{cl}_{\infty}(A\mid Q)$ and the map is inflationary.
+Idempotence of the closure operation follows from Theorem 3. By the fixed-point property, we have:
+$\mathrm{cl}_{\infty}(\mathrm{cl}_{\infty}(A|Q)|Q)$ $=\mathrm{cl}_{\infty}(A|Q)\cup\mathcal{E}_{\infty}(\mathrm{cl}_{\infty}(A|Q)|Q)\cup\dots\cup\mathcal{E}^{(\infty)}_{\infty}(\mathrm{cl}_{\infty}(A|Q)|Q)$ (35)
+$=\mathrm{cl}_{\infty}(A|Q)\cup\mathrm{cl}_{\infty}(A|Q)\cup\dots\cup\mathrm{cl}_{\infty}(A|Q)=\mathrm{cl}_{\infty}(A|Q).$ (36)
+Theorem 3 allows for introducing the following equivalence relation on the set of knowledge.
 Definition 8 (Equivalence of Knowledge).
-We say that $X\subseteq\mathcal{K}$ and $Y\subseteq\mathcal{K}$ are equivalent in context $Q\subseteq\mathcal{K}$, i.e., $X\equiv Y$, if and only if $X\subseteq\mathcal{E}_{\infty}(Y|Q)$ and $Y\subseteq\mathcal{E}_{\infty}(X|Q)$.
-Notice that the equivalence of knowledge does not mean $A$ is a fixed point, i.e., it is not true that $A=\mathcal{E}_{\infty}(A|Q)$ when $A\equiv\mathcal{E}_{\infty}(A|Q)$.
-If we have $A=\mathcal{E}_{\infty}(A|Q)$, then nothing more can be derived from $A$ in the context of $Q$ by any reasoning rule from $A$.
-That statement is very strong and is not true when $A\equiv\mathcal{E}_{\infty}(A|Q)$.
-Suppose $A=\{x\}$ and $f\in\mathcal{F}_{\infty}$ allows for getting $y=f(x)\in\mathcal{R}_{C}(Q)$.
-Then $\mathcal{E}_{\infty}(A|Q)=\{x,y\}\supset\{x\}=A\equiv\mathcal{E}_{\infty}(A|Q)$.
-To proceed we will also need the idempotence of reasoning with unlimited computational budget.
-Theorem 3 (Idempotence of unlimited reasoning).
-For any $A\subseteq\mathcal{K}$ and context $Q\subseteq\mathcal{K}$ we have:
-$\mathcal{E}_{\infty}(A|Q)=\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A|Q)|Q).$ (12)
+We say that two subsets of knowledge, $A\in\mathcal{A}(\mathcal{K})$ and $B\in\mathcal{A}(\mathcal{K})$, are equivalent in the context $Q$, and denote that context equivalence as $A\equiv_{Q}B$, if and only if they have the same closure:
+$A^{*}_{\infty}(Q)=B^{*}_{\infty}(Q).$ (37)
+Equivalence of knowledge means that, given the same context $Q$, we reach the same conclusions (the least fixed points) starting from different states of knowledge, $A$ or $B$.
+Notice that any subset $A$ is equivalent to itself.
+Notice that it is not true $A=\mathcal{E}_{\infty}(A|Q)$, when $A\equiv_{Q}\mathcal{E}_{\infty}(A|Q)$.
+Example.
+Consider $A=\{x\}$ and $f\in\mathcal{F}_{\infty}$ getting $y=f(x)\in\mathcal{R}_{\infty}(Q)$.
+Then, we see that $\mathcal{E}_{\infty}(A|Q)=\{x,y\}\supset\{x\}=A$ and $A\equiv\mathcal{E}_{\infty}(A|Q)$.
+Therefore, it is not true that $A=\mathcal{E}_{\infty}(A|Q)$ when $A\equiv\mathcal{E}_{\infty}(A|Q)$.
+∎
+The fixed-points of reasoning and the computationally unlimited emergence operator are the building blocks of the semantic information measure $\mu_{C}$.
+The following example shows how we can use $\mathcal{E}_{\infty}$ to construct $\mu_{\infty}$ that evaluates the contextual informational content of a given knowledge subset with unlimited insight of $C=\infty$.
+Example.
+Define:
+$\mu_{\infty}(A|Q)=\log|\mathcal{K}|\cdot|A^{*}_{\infty}(Q)\cap\mathcal{R}_{\infty}(Q)|$
+where $A^{*}_{\infty}(Q)=\bigcup_{n=0}^{|\mathcal{K}|}\mathcal{E}^{(n)}_{\infty}(A|Q)$ is the fixed point of reasoning. To improve readability, we write $A^{*}_{\infty}=A^{*}_{\infty}(Q)$.
+The measure construction satisfies the required axioms of Definition 6.
+For the null set, we have $\mu_{\infty}(\emptyset|Q)=\log|\mathcal{K}|\cdot|\emptyset\cap\mathcal{R}_{\infty}(Q)|=0$.
+If $A\subseteq B$, then $A^{*}_{\infty}\subseteq B^{*}_{\infty}$ by monotonicity of $\mathcal{E}_{\infty}$, implying $\mu_{\infty}(A|Q)\leq\mu_{\infty}(B|Q)$.
+Contextual non-negativity holds since $\mu_{\infty}(A|Q)>0$ if and only if $A^{*}_{\infty}\cap\mathcal{R}_{\infty}(Q)\neq\emptyset$, which occurs precisely when $A$ contains or can derive relevant knowledge.
+The measure is bounded since $|A^{*}_{\infty}\cap\mathcal{R}_{\infty}(Q)|\leq|\mathcal{K}|$.
+For subadditivity, we have:
+$\mu_{\infty}(A\cup B|Q)\leq\mu_{\infty}(A|Q)+\mu_{\infty}(B|Q),$
+since $|(A\cup B)^{*}_{\infty}\cap\mathcal{R}_{\infty}(Q)|\leq|A^{*}_{\infty}\cap\mathcal{R}_{\infty}(Q)|+|B^{*}_{\infty}\cap\mathcal{R}_{\infty}(Q)|$.
+Note that additivity of $\mu_{\infty}$ (equality) does not generally hold for computationally independent sets $A\perp_{\infty}B$.
+Even when neither $A$ nor $B$ can derive elements of the other, both may independently derive the same conclusions.
+For example, if $A=\{P,P\to R\}$ and $B=\{Q,Q\to R\}$, then $A\perp_{\infty}B$ yet both derive $R$, making $(A\cup B)^{*}_{\infty}\cap\mathcal{R}_{\infty}(Q)$ smaller than the sum of the individual closures.
+Thus, computational independence alone does not guarantee additivity of the semantic information measure.
+∎
+In that setting we can now ask the fundamental question:
+What knowledge can we discover with unlimited computational budget?
+The axioms we have introduced lead us to an interesting and intriguing conclusion.
+With infinite computing and reasoning power, total information must be preserved.
+That bold (philosophical) claim is at the center of our investigations.
+Let us prove it formally.
+Theorem 5 (Information Preservation).
+Assume that $\mu_{\infty}$ is invariant on equivalence classes, i.e., $A\equiv_{Q}B$ implies $\mu_{\infty}(A|Q)=\mu_{\infty}(B|Q)$. Then for all $A\in\mathcal{A}(\mathcal{K})$ we have:
+$\mu_{\infty}(\mathrm{cl}_{\infty}(A|Q)|Q)=\mu_{\infty}(A|Q).$ (38)
 Proof.
-By definition $\mathcal{E}_{\infty}(A|Q)\subseteq\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A|Q))$.
-To show $\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A|Q))\subseteq\mathcal{E}_{\infty}(A|Q)$, we must find $h\in\mathcal{F}_{\infty}$ such that $h(A|Q)=g(A\cup\bigcup f_{i}(A|Q))$ for $g,f_{i}\in\mathcal{F}_{\infty}$.
-But since we can always find such $h\in\mathcal{F}_{\infty}$, we conclude that any context-relevant conclusion $g(\mathcal{E}_{\infty}(A|Q))$ derived from $\mathcal{E}_{\infty}(A|Q)$ by a reasoning rule $g\in\mathcal{F}_{\infty}$ can be also derived directly from $A$ as $h(A)=g(A\cup\bigcup f_{i}(A))$.
-Therefore, by definition of emergence operator we have $\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A|Q))\subseteq\mathcal{E}_{\infty}(A|Q)$ as required.
+Since $A$ and $\mathrm{cl}_{\infty}(A|Q)$ have the same closure, they are equivalent by definition, $A\equiv_{Q}\mathrm{cl}_{\infty}(A\mid Q)$).
+Preservation of information follows from from the assumed invariance of $\mu_{\infty}$ on equivalence classes.
 ∎
-With all the building blocks available, we can now to prove that unlimited reasoning ($C=\infty$) can reveal but never create information.
-The semantic content of a knowledge base is invariant under unlimited reasoning of emergence operator.
-Apparent information gain at finite computational budget emerges as an effect of limited computational access.
-Theorem 4 (Information preservation).
-For any $A\subseteq\mathcal{K}$ and context $Q\subseteq\mathcal{K}$, if the semantic information is constant on equivalent sets of knowledge, i.e., $\mu_{\infty}(X|Q)=\mu_{\infty}(Y|Q)$ whenever $X\equiv Y$, then:
-$\mu_{\infty}(\mathcal{E}_{\infty}(A|Q)|Q)=\mu_{\infty}(A|Q).$ (13)
-Proof.
-We must show equivalence of the knowledge encoded in $A$ and $\mathcal{E}_{\infty}(A|Q)$, i.e., we demand $A\subseteq\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A|Q))$.
-But, by Theorem 3, we see that $A\subseteq\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A|Q))=\mathcal{E}_{\infty}(A|Q)$ and $\mathcal{E}_{\infty}(A|Q)\subseteq\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A|Q))=\mathcal{E}_{\infty}(A|Q)$ as requested.
-Since $A\equiv\mathcal{E}_{\infty}(A|Q)$ and $\mu_{\infty}$ is constant on equivalent sets of knowledge by assumption, we conclude that $\mathcal{E}_{\infty}(A|Q)=\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A|Q))$.
-∎
-Justification for the Axiom of Unlimited Reasoning with Emergence Operator
-The completeness (or closure property) of the set of reasoning functions $\mathcal{F}_{\infty}$ sounds somewhat mysterious, but is rather natural.
-We demand that for any reasoning steps $f$ and $g$, there exists a single step $h$ such that $g(A\cup f(A))=h(A)$.
-That is the nature of an idealized, complete reasoning system we want to investigate to understand its tendency to generate hallucinations.
-The construct $\mathcal{F}_{\infty}$ is designed to represent the theoretical limit of reasoning with an unbounded computational capacity.
-It is not intended to model the practical, heuristic, and bounded reasoning of a real LLM, but rather the theoretical limit of what is ultimately knowable from a set of premises.
-Completeness is fundamental to formal systems, such as first-order logic or Turing-complete computations and Direct Revelation Principle in auction theory kuratowski_set_1978 ; sipser1996introduction ; fudenberg1991game .
-If one can prove a lemma $f(A)$ from a set of axioms $A$, and then prove a theorem $g(A\cup f(A))$ using that lemma, a direct proof of the theorem from the axioms alone $h(A)$ must exist by simply concatenating the proof steps.
-The composition of two computable functions is itself a computable function.
-A direct revelation mechanism in (auction) game theory is a composition of preference processing mechanisms where agents report their preferences directly to the mechanism, and the mechanism implements their preference processing strategies to generate outcome.
-That axiom is essential for the proof of idempotence of unlimited reasoning in Theorem 3 Theorem 4.
-These are critical philosophical and mathematical results necessary for our study.
-With them we argue that the information gain observed in practical, constrained inference is an artifact of computational bounds, not the magical creation of knowledge ex nihilo.
-We acknowledge that depends on the assertion that any sequential reasoning process can be compiled into a single computational step.
-Therefore, the philosophical claim that information gain is an artifact of computational bounds rests on the acceptance of this structure of reasoning.
+Therefore, the answer to the question above is this.
+Unlimited reasoning does not create information, it can only reveal it and make accessible.
+All that can be known is already in $A$, we can only extract that knowledge and make useful.
+That implies the apparent information gain at finite computational budget, $C<\infty$, emerges as an effect of limited computational access.
+Can we call that creativity?
+It is any increase in accessible information, as measured by $\mu_{C}$, under bounded computation.
 4.1.5 The Uncertainty Resolution Paradox and its Resolution
-When LLM combines knowledge elements to fabricate facts or produce emergent insights in a given fixed context, it seems that
-$\mu_{C}(A\cup B|Q)>\mu_{C}(A|Q)+\mu_{C}(B|Q),$
-so novel insights may come from existing data. That is exactly the creativity behind the hallucination phenomenon we are struggling to control. So why should we assume subadditivity?
-First, subadditivity of semantic information measure establishes a baseline where uncontrolled synergy of knowledge cannot appear.
-We remove from our analytical toolkit the potential measurement bias of information surplus that any information aggregation mechanism might generate.
+As we have just seen, when inference combines knowledge elements to fabricate facts or produce emergent insights in a given fixed context, it seems that
+$\mu_{C}(A\cup B|Q)>\mu_{C}(A|Q)+\mu_{C}(B|Q).$ (39)
+Novel insights may come from existing data.
+That is the creativity behind the hallucination phenomenon we are struggling to control.
+So why should we assume subadditivity of the semantic information measure instead?
+First, subadditivity establishes a baseline where uncontrolled synergy of knowledge cannot appear.
+We remove from our analytical toolkit the potential bias that any information aggregation mechanism might generate.
 Second, we prove that hallucinations are inevitable even when knowledge is subadditive (within a fixed context).
-In that setting, by Theorem 4, reasoning does not create new information.
-It is a transformation that reveals information already present in any knowledge representation.
-Quite remarkably, that does not at all prevent imagination, creativity, foresight, or hallucination.
-However, we should take care of a paradox those claims apparently bring into existence.
+It is in that setting that Theorem 5 states that reasoning does not create new information.
+It can only reveal information already present in any knowledge representation.
+Quite remarkably, that does not at all prevent imagination, creativity, foresight, or hallucination. To explain why, we should take care of the following paradox.
 Consider a similarity transformation of two matrices, $D=V^{-1}AV$, where $D$ is a diagonal matrix of eigenvalues.
@@
 But, if reasoning makes a new fact accessible, then that new fact can be used to resolve uncertainty that was previously unresolvable.
-Doesn’t it mean the reasoning must have created new information, contrary to what we have just claimed?
+Doesn’t that mean the reasoning must have created new information, contrary to what we have just claimed?
 It seems we have a paradox.
-To resolve it, we need the emergence operator $\mathcal{E}_{C}(X|Q)$ distinguishing between computationally unbounded and bounded uncertainty reduction (accessible and inaccessible information).
+To resolve it, we need the emergence operator $\mathcal{E}_{C}$ distinguishing between computationally unbounded and bounded uncertainty reduction (accessible and inaccessible information).
 Notice that for all practical purposes the information measurement $\mu_{C}(X|Q)$ can only provide maximum uncertainty reduction from $X$ within some computational budget $C$ (or time) in a given context.
-Therefore, given the result of computationally bounded reasoning $\mathcal{E}_{C}(X|Q)$, we can observe information gain through the computational work, i.e., $\mu_{C}(X|Q)\leq\mu_{C}(\mathcal{E}_{C}(X|Q)|Q)$.
+Therefore, given the result of computationally bounded reasoning, we can observe information gain through the computational work, i.e., $\mu_{C}(X|Q)\leq\mu_{C}(\mathcal{E}_{C}(X|Q)|Q)$.
 That is possible only with organized information providing knowledge for reasoning.
-These are precisely the statements of Theorem 1 and Theorem 4.
 That idea describes reasoning as a dynamical system:
@@
 $\theta(\infty)=\mathcal{E}_{\infty}(\theta(\infty)|Q).$ (41)
-All there is to know is immediately accessible when we have unlimited computing resources (or time).
+By Theorem 3, all there is to know is immediately accessible when we have unlimited computing resources (or time).
+Example.
 To illustrate that argument, consider the following learning trajectory:
-$\theta(0)$ $=[1,1,1,0,0]\text{ (initial knowledge state)},$
-$\theta(1)$ $=\mathcal{E}_{C}(\theta(0)|Q)=[1,1,1,1,1]\text{ (state after reasoning)},$
-$y(0)$ $=\mu_{C}(\theta(0)|Q)=3\text{ bits},$
-$y(1)$ $=\mu_{C}(\theta(1)|Q)=5\text{ bits}.$
+$\theta(0)$ $=[1,1,1,0,0]\text{ (initial knowledge state)},$ (42)
+$\theta(1)$ $=\mathcal{E}_{C}(\theta(0)|Q)=[1,1,1,1,1]\text{ (state after reasoning)},$ (43)
+$y(0)$ $=\mu_{C}(\theta(0)|Q)=3\text{ bits},$ (44)
+$y(1)$ $=\mu_{C}(\theta(1)|Q)=5\text{ bits}.$ (45)
 We see $\mu_{C}(\mathcal{E}_{C}(\theta(0)|Q)|Q)>\mu_{C}(\theta(0)|Q)$, i.e., practical information gain in the course of reasoning.
 In comparison, infinite computational budget, $C=\infty$, implies we do not need the intermediate inference steps to see all there is to see, so we know the equilibrium state:
-$\mu_{\infty}(\mathcal{E}_{\infty}(\theta(0)|Q)|Q)=\mu_{\infty}(\theta(0)|Q)$
-and no new information is created.
-In the next section, we relate that important property of response generation, namely, how much additional semantic information each LLM inference path can contribute, with the Semantic Information Conservation Principle.
-4.1.3 Knowledge Representation in LLMs
-The above Knowledge Representation finds its direct counterpart in the structure and state of a transformer architecture of LLMs.
-The abstract knowledge space $\mathcal{K}$ is the high-dimensional embedding space of the model.
-Every fact, concept, and linguistic pattern the model has learned is represented by a specific element of that space.
-The metric $d_{\mathcal{K}}$ is typically implemented as cosine similarity or Euclidean distance, where proximity between vectors signifies a closer semantic relationship.
-Model parameters $\mathcal{W}$ correspond to the full set of trained weights in the transformer, including the attention matrices ($W_{Q},W_{K},W_{V},W_{O}$) and the matrices of the MLP layers.
-The mapping $\Phi(W)$ represents the totality of information, patterns, and factual associations stored within these trained weights.
-Activation patterns $\mathcal{A}$ correspond to the state of the residual stream during a single forward pass for a given input.
-An activation pattern is therefore a specific trajectory through the network’s computational graph.
-Finally, the mapping $\Psi(a_{i})$ represents the information made computationally accessible during a specific inference step and encoded in the residual stream.
-The process of information processing under computational constraints maps to the sequential forward pass transformer inference.
-The idea of computational budget $C$ can be represented by the depth or size of the network, e.g.,the number of transformer blocks.
-A deeper network has a larger budget $C$, allowing for more reasoning steps.
-In the context of multi-step reasoning, $C$ can also represent the number of iterations allowed in a Chain-of-Thought (CoT) sequence.
-The semantic information measure $\mu_{C}$ can be linked to the model’s objective function.
-For a given input $A$, $\mu_{C}(A)$ can be understood as the measurable reduction in cross-entropy loss that the model achieves by processing $A$ with a budget $C$.
-Finally, the emergence operator $\mathcal{E}_{C}$ can be identified with a transformer block (one multi-head attention layer followed by one MLP layer).
-This block applies a computational transformation $\mathcal{E}_{C}$ to the residual stream at layer $t$, producing the updated state at layer $t+1$, making information that was latent in the input explicitly accessible.
+$\mu_{\infty}(\mathcal{E}_{\infty}(\theta(0)|Q)|Q)=\mu_{\infty}(\theta(0)|Q)$ (46)
+and no new information is created. ∎
+Axiomatic Reasoning
+In axiomatic reasoning we can admit reasoning rules $f\in\mathcal{F}_{C}$ such that $f(\emptyset)\neq\emptyset$.
+Foundational ideas, including axioms or tautologies, can be generated without prior context.
+That case is aligned with LLMs possessing basic knowledge from pretraining that can be activated by minimal inputs.
+Then, with foundational knowledge given by:
+$K_{\emptyset}=\bigcup_{f\in\mathcal{F}_{C}}f(\emptyset),$ (47)
+we define the context-relevant knowledge as:
+$\mathcal{R}_{C}(\emptyset)=K_{\emptyset}.$ (48)
+When the context is empty, the only relevant knowledge is the foundational knowledge.
+Then, $\mu_{C}(A|\emptyset)>0$ if and only if $A\cap K_{\emptyset}\neq\emptyset$.
+Reasoning without context only expands the knowledge set $A$ if the conclusions $f(A)$ represent foundational knowledge.
+Finally, $\mathcal{E}_{C}(\emptyset|\emptyset)=K_{\emptyset}$.
+We permit structures where $\mu_{C}$ is not monotonic in context.
+Consider, $\mu_{C}(A|Q)=|A\setminus Q|$ and $\mathcal{R}_{C}(Q)=\mathcal{K}\setminus Q$.
+Let us then see what happens if $A=\{k\}$, $Q_{1}=\emptyset$ and $Q_{2}=\{k\}$.
+We have:
+$Q_{1}\subseteq Q_{2}\quad\text{and}\quad\mu_{C}(A|Q_{1})=|\{k\}\setminus\emptyset|=1>0=|\{k\}\setminus\{k\}|=\mu_{C}(A|Q_{2}).$ (49)
+Increasing the context may reduce the information contributions.
+If the extended context $(Q_{2})$ provides an alternative explanation for phenomena related to original context $(Q_{1})$, then marginal contribution of available knowledge $(A)$ may decrease.
+There is a nice interpretation of that property, that we all may recognize very well.
+Often when we learn about some fundamental concept we overestimate our deep understanding of it and we tend to develop overconfident belief that we have become experts in a domain.
+But the more we learn, the less confident we feel, eventually coming to the conclusion that we know that we know nothing (or not enough at least), following the steps of Socrates.
 4.2 Statements, Queries, and Ground Truth
@@
 Queries refer to what is presupposed (assumed in advance), e.g., topic or problem, whereas responses make claims (or assertions).
-We should also point out that we limit our study to the case of non-trivial queries that require integrating responses from many LLM inference paths.
-The following formal definition encapsulates those ideas.
-Definition 8 (Statement Space).
-The space of natural language statements $\mathcal{L}$ is a metric space with distance function $d_{\mathcal{S}}$.
-We distinguish query space $\mathcal{Q}\subset\mathcal{L}$ of interrogative statements and response space $\mathcal{R}\subset\mathcal{L}$ of declarative statements.
-Function $K:\mathcal{L}\rightarrow\mathcal{P}(\mathcal{K})$ maps statements to referenced knowledge.
-We call query $q$ non-trivial, if it is integrated from multiple knowledge subsets.
-We must now rigorously define the ground truth mapping and establish formal relationships between the knowledge space $\mathcal{K}$ and response space $\mathcal{R}$.
-For that, we should start by recognizing that the very idea of ground truth validation is limited (or subjective at least).
+We connect the abstract knowledge space $\mathcal{K}$ to the linguistic inputs and outputs of the model. We model the language space analogously to the knowledge space.
+Definition 9 (Statement Space and Knowledge Mapping).
+The space of natural language statements $(\mathcal{L},d_{\mathcal{L}})$ is a Polish space.
+The query space $\mathcal{Q}$ and the response space $\mathcal{R}$ are Borel subsets of $\mathcal{L}$.
+The knowledge mapping $K\colon\mathcal{L}\rightarrow\mathcal{B}(\mathcal{K})$ is a Borel measurable function that maps statements to the knowledge they reference.
+We require the output of $K$ to be in the Borel sets $\mathcal{B}(\mathcal{K})$, because $K(q)$ serves as the context for $\mathcal{E}_{C}(A|Q)$.
+As established earlier (Theorem 2), the context $Q$ must be Borel to guarantee that the reasoning process remains closed within the analytic sets representing knowledge.
+Before we define the ground truth mapping, we need to acknowledge its epistemological limitations.
+We recognize that the very idea of ground truth validation is limited (or subjective at least).
 It is rather naive to assume that LLM can access an oracle establishing provably true and relevant knowledge for every possible query.
-Indeed, that limit is given by the Goedel’s Incompleteness Theorems for the formalized mathematical theories.
-Therefore, in practice that inaccessibility of the complete truth validation function (even in the idealized and rigorous mathematical reasoning) is one of the reasons why hallucination cannot be eliminated entirely.
+Therefore, in practice that inaccessibility of the complete truth validation function (even in the idealized and rigorous mathematical reasoning) is one of the reasons why hallucination cannot be eliminated entirely (cf.
+kalai2025why ).
 High-quality, fact-checked datasets, human feedback providing human-level judgments of truth, RAG systems with external validated knowledge bases, all that approximate the idea of ground truth.
-Surprisingly, the impossibility result seems to be independent of practical or idealized implementation of a ground truth mapping, which we define as follows.
+Surprisingly, the impossibility result seems to be independent of practical or idealized implementation of a ground truth mapping.
 Definition 10 (Ground Truth Mapping).
-The ground truth mapping $T\colon\mathcal{Q}\rightarrow\mathcal{P}(\mathcal{K})$ associates each query $q\in\mathcal{Q}$ with knowledge elements constituting complete, correct answers:
-[…]
-In other words, $T(q)$ represents our best available approximation to (labeled as) true relevant knowledge.
-We accept the fact that there is some external agency telling us what is and what is not true answer to a query, and that evaluation has a form of label assigned to knowledge elements.
-But, for all that, we also focus on the queries that are non-trivial in the following sense.
+The ground truth mapping $T\colon\mathcal{Q}\rightarrow\mathcal{B}(\mathcal{K})$ is a Borel measurable function that associates each query $q\in\mathcal{Q}$ with the Borel subset of relevant knowledge labeled as (true) correct answer.
+In other words, $T(q)$ represents our best available approximation of (labeled as) true relevant knowledge.
+There is some external agency providing a data base or dataset telling what is and what is not true response to a query, and that evaluation has a form of a label assigned to knowledge elements.
+Similar to knowledge mapping $K(q)$, ground truth $T(q)$ contributes to the context $Q=K(q)\cup T(q)$, and thus must also be Borel to ensure $Q\in\mathcal{B}(\mathcal{K})$.
+We focus our analysis on non-trivial queries that require the integration of diverse information sources.
 Definition 11 (Non-trivial Query).
-A query $q\in\mathcal{Q}$ is non-trivial with respect to computational budget $C$ if there exist knowledge configurations $\theta_{i}$ and $\theta_{j}$ for the context $Q=K(q)\cup T(q)$, such that
+A query $q\in\mathcal{Q}$ is non-trivial at budget $C$ if there exist knowledge subsets $A$ and $B$ such that for the context $Q=K(q)\cup T(q)$:
 […]
@@
 A query is non-trivial when at least two agents disagree on some fact that is both encoded and available, and relevant to answering that query.
-For trivial queries impossibility results do not hold.
-That definition is important for the development of impossibility results.
-It measures the ability of a response generation system to store knowledge and expand knowledge that is necessary for the impossibility to emerge.
-We may view that ability to respond to non-trivial queries as an analog of the formal system complexity in the Goedel’s framework.
-4.3 Auction-Theoretic Model of LLM Inference
+Non-triviality implies the existence of relevant, accessible disagreement among agents.
+Example.
+Consider a query $q$:
+tell me about recent AI breakthroughs.
+Its context is given by $Q=K(q)\cup T(q)$.
+Then, the relevant knowledge $\mathcal{R}_{C}(Q)$ includes facts derivable within $C$ computational steps that relate to AI advances.
+Non-triviality occurs when different agents (e.g., attention heads) encode conflicting information about these breakthroughs.
+∎
+Definition 11 describes the ability of a response generation system to store knowledge that is necessary for the impossibility to emerge.
+In other words, for trivial queries the impossibility results do not hold.
+We may view that ability to respond to non-trivial queries as an analog of the formal system complexity in the Goedel’s framework that gives rise to incompleteness of consistent formal systems godel1931formal .
+4.3 Auction-Theoretic Model of Inference
 There are many ways LLM can generate answers to a query.
 By design, it is a controlled random process governed by fine-tuned probability distributions.
-We are going to model that process as an auction in which agents (representing competing ideas is the space of all possible activation patterns $\mathcal{A}$) use their privately held knowledge (internal representations in $\mathcal{W}$) to construct bids (candidate responses) and submit them to the LLM response mechanism.
-That response mechanism then selects the best bid as a response to a query.
+We are going to model that process as an auction of ideas in which agents (representing competing ideas is the space of all possible knowledge representations) use their privately encoded knowledge (internal states) to construct bids (candidate responses) and influence the response.
+We first define the space representing the internal states (e.g., parameters or activations) of the bidding agents, and the mapping from these states to the knowledge they encode.
+Definition 12 (Private Knowledge Configuration).
+The private knowledge configuration of agent $i$, denoted $\Theta_{i}$, lives in a Polish space.
+The joint space is the Cartesian product $\Theta=\prod_{i=1}^{n}\Theta_{i}$.
+The configuration mapping $\kappa\colon\Theta\mapsto(\mathcal{A}(\mathcal{K}))^{n}$ is a measurable function, where $\kappa_{i}(\theta_{i})\in\mathcal{A}(\mathcal{K})$ is the analytic knowledge set associated with agent $i$’s configuration $\theta_{i}$ of privately encoded knowledge.
+For each agent, $\theta_{i}$ is the inherent configuration of encoded knowledge that uniquely determines its capabilities and behavior, independent of the specific input context being processed.
+In a trained transformer model, that may correspond to the learned parameters (weight matrices) of attention head:
+$\theta_{h}=\left(W_{Q}^{(h)},W_{K}^{(h)},W_{V}^{(h)},W_{O}^{(h)}\right),$ (51)
+where $W_{Q}^{(h)}$ and $W_{K}^{(h)}$ are query and key weights, $W_{V}^{(h)}$ and $W_{O}^{(h)}$ are value and output weights.
+Private knowledge configuration $\theta_{h}$ tells what the agent is looking for and where, and how it extracts the relevant information.
 Definition 13 (Response Mechanism).
 A response mechanism $\mathcal{M}=(S,g,p)$ consists of:
-  • Strategy space $S=S_{1}\times\cdots\times S_{n}$, where $S_{i}\subset\mathcal{P}(\mathcal{R})$,
-  • Outcome function $g:S\rightarrow\mathcal{R}$ mapping strategies to responses,
-  • Information contribution function $p:S\times\Theta\rightarrow\mathbb{R}^{n}$ determining costs of limiting hallucinations.
-A response mechanism defines strategy space $S$ as a set of inference actions that each agent can make to represent its knowledge $\theta_{i}\in\Theta_{i}\subset\mathcal{K}$ in response to a query.
+  • strategy space $S=S_{1}\times\cdots\times S_{n}$, where $S_{i}$ is a Polish space (representing bids, e.g., outputs of attention heads or probability distributions),
+  • a measurable outcome function $g\colon S\mapsto\mathcal{R}$ mapping strategy profiles $s\in S$ to responses in the Polish response space $\mathcal{R}$,
+  • a measurable information contribution function $p\colon S\times\Theta\mapsto\mathbb{R}^{n}$ determining how strategy profile $s\in S$ inference.
+A response mechanism defines strategy space $S$ as a set of inference actions that each agent can make to represent its knowledge in response to a query.
 It is a dictionary of LLM’s component activation pattens generating and presenting ideas as bids in an auction.
@@
 One straightforward example is a softmax function in LLM’s final layer.
-Finally, given a knowledge profile $\theta=(\theta_{1},\dots,\theta_{n})$ revealed by the agents, information contribution function $p(s,\theta)=(p_{1}(s,\theta),\dots,p_{n}(s,\theta))$ measures individual contributions of each response to generating an output.
+Finally, given a knowledge configuration profile $\theta=(\theta_{1},\dots,\theta_{n})$ revealed by the agents, the information contribution function $p(s,\theta)=(p_{1}(s,\theta),\dots,p_{n}(s,\theta))$ measures individual contributions of each response to generating an output.
 In this case, examples include (semantic information measures) counting ignored low-probability tokens in top-k filtering, softmax normalization in multi-headed attention, or marginal reward contributions in RLHF.
 Definition 14 (Agents of Competing Ideas).
-The inference process involves $n$ agents $\mathcal{A}=\{a_{1},a_{2},...,a_{n}\}$ competing by submitting bids to mechanism $\mathcal{M}$.
-Each agent $a_{i}$ constructs bid $s_{i}(\theta_{i})$ based on private knowledge $\theta_{i}\in\Theta_{i}$ to maximize utility:
-$u_{i}(s,\theta)=v_{i}(g(s),\theta_{i})-p_{i}(s,\theta),$ (18)
-where $v_{i}\geq 0$ is valuation and $p_{i}$ is information contribution. Knowledge profiles $\theta=(\theta_{1},\dots,\theta_{n})$ are independently distributed.
+The inference process involves $n$ agents, indexed by $\mathcal{N}=\{1,\dots,n\}$. Each agent chooses a strategy $s_{i}\in S_{i}$ to maximize a quasi-linear utility:
+$u_{i}(s,\theta)=v_{i}(g(s),\theta_{i})-p_{i}(s,\theta).$ (52)
+The valuation function $v_{i}\colon\mathcal{R}\times\Theta_{i}\mapsto\mathbb{R}_{+}$ is measurable, representing the agent’s preference for the outcome $g(s)$ given its private knowledge configuration $\theta_{i}$.
 One way to think about the concept of agents is that it represents different attention heads in transformer models, or components of mixture-of-experts architectures, or activation paths fired by the same query.
@@
 It is designed to explain why any inference path is activated in response to the LLM prompt.
-For example, negative log likelihood loss can be decomposed with respect to attention heads, and linearized to measure how much attention head contributes to token prediction.
-Let us point out that agents represent algorithms and architectural components of any LLM implementation, they are not conscious beings with individual goals.
-That is important, as it refers to the Revelation Principle in game theory.
+The agents represent algorithms and architectural components of any LLM implementation.
+That is an analog of the Revelation Principle in game theory.
 If we know strategies (of using knowledge) the players (conscious beings with individual goals) will use in a game, we can implement those strategies as an algorithm (or agent) with parameters representing private preferences or knowledge of the players.
 In the case considered, we are the players selecting architecture and training the LLM.
-We assume that individual utility depends on all responses $s=(s_{1},\dots,s_{n})$ and knowledge $\theta=(\theta_{1},\dots,\theta_{n})$ used by all agents, $u_{i}=u_{i}(s,\theta)$.
-Also, we assume utility is a sum of individual valuation $v_{i}$ of outcome $g(s)$ and information contribution $p_{i}=p_{i}(s,\theta)$ value assigned.
-The valuation function $v_{i}=v_{i}(g(s),\theta_{i})$ represents how strongly a neural component activates for certain outputs (high activation = high $v_{i}$, low activation = low $v_{i}$), and how much a component reduces the overall loss when processing certain inputs ($v_{i}$ is proportional to gradient of a loss function, $\nabla_{\theta_{i}}L_{W}$).
+The valuation function $v_{i}=v_{i}(g(s),\theta_{i})$ represents how strongly a neural component activates for certain outputs (high activation = high $v_{i}$, low activation = low $v_{i}$), and how much a component reduces the overall loss when processing certain inputs.
 The information contribution function $p_{i}=p_{i}(s,\theta)$ may represent training penalties ($p_{i}>0$ for discouraged activation, $p_{i}<0$ for encouraged activation), attention focus (higher $p_{i}$ reduces attention allocation), $L1$ regularization terms, and network circuits resistance (limiting activations of certain inference paths).
 Therefore, $u_{i}=v_{i}-p_{i}>0$ encourages activation and $u_{i}=v_{i}-p_{i}<0$ discourages activation.
-We can also illustrate the concept of utility function with the following example:
-$u_{i}(s,\theta)$ $=\underbrace{\alpha\cdot\text{Accuracy}(g(s),\theta_{i})}_{\text{Valuation}}-\underbrace{\gamma\cdot\text{Contradiction}(s_{i},\theta_{i})}_{\text{Consistency Cost}}.$
-$\text{Accuracy}(g(s),\theta_{i})$ measures how well the response aligns with the knowledge $\theta_{i}$ agent $a_{i}$ is using.
-$\text{Contradiction}(s_{i},\theta_{i})$ measures how much the strategy $s_{i}$ conflicts with internal knowledge state $\theta_{i}$ (both extracted from $s$ and $\theta$ profiles).
-Independent private knowledge assumption
-One of the critical and controversial assumptions we make is is that knowledge profiles $\theta_{i}$ are independently distributed.
-It should be noted that in transformers, heads may share layer-norm signals, residual streams, and optimizer state.
+Justification of the independent private knowledge model
+For the idealized analysis in Theorem 6, we require that the agent’s valuation depends only on its own private knowledge configuration and that the configurations are statistically independent.
+That is restrictive, but it allows us to establish a baseline impossibility, which is then generalized in subsequent proofs that accommodate correlated beliefs.
+In the realistic settings of transformers, heads may share layer-norm signals, residual streams, and optimizer state.
 That may result in correlated (or, so called, affiliated) inference signals between token representations.
 Furthermore, the superposition hypothesis, introduced in mechanistic interpretability studies, suggests cooperative error-canceling due to strong correlations.
-To address that issue, we show that impossibility arises in both independent and correlated settings in general and in the LLM transformer environment.
-Namely, we refer to Green-Laffont theorem green1977characterization in the independent private knowledge setting.
+We show that impossibility arises in both independent and correlated settings.
+We refer to Green-Laffont theorem green1977characterization in the independent private knowledge setting.
 Then, we refer to the truthful probability elicitation and proper scoring theory introduced by Savage savage1971elicitation , which does not require independent private knowledge.
@@
 Definition 15 (Hallucination Cost).
-Given query $q\in\mathcal{Q}$, response $r\in\mathcal{R}$, and ground truth $T(q)$, the hallucination cost function $H\colon\mathcal{R}\times\mathcal{Q}\rightarrow\mathbb{R}_{+}$ measures discrepancy:
+Given query $q\in\mathcal{Q}$, response $r\in\mathcal{R}$, and ground truth $T(q)$, the hallucination cost function $J\colon\mathcal{R}\times\mathcal{Q}\mapsto\mathbb{R}_{+}$ measures discrepancy:
 $J(r,q)=d_{\mathcal{K}}(K(r),T(q)).$ (53)
@@
 We characterize them formally referring to the fundamental results we have established earlier.
+These properties are not arbitrary at all.
+They come from an exhaustive and rigorous game-theoretical studies of conflicts and competition in distributed multi-agent systems with asymmetric information.
+In those settings, agents carefully craft messages they communicate to the system in order to optimize their privately held preferences by revealing information they hold.
+The goal of the system architect is to impose rules of information exchange that result in efficient equilibrium solutions in which goals of agents are aligned with the goals of the architect.
+Auction is a special case of such a game with asymmetric information.
+There is an auction designer who wants to sell an item, like a painting or a part of the electromagnetic spectrum, but does not know the value the item may have for those willing to pay.
+To solve that problem, the designer can ask those interested for their offers.
+Based on the collected bids, the winner will be selected and the price to be payed to the designer will be established.
+Clearly, each bidder will constructs the offer based on an estimate of the item’s value and the privately held information about available budget.
+But what information should the bid reveal to maximize probability of winning?
+What winner selection and payment rule should the auction designer construct to maximize the profit and bidders willingness to compete?
+How to guarantee truthful revelation of willingness to pay, while maximizing auction efficiency without incurring additional external costs?
+In this setting the four properties of information revelation exhaustively cover the space of strategic choices, of the bidders and auction designer.
+For details, see Myerson1989 ; fudenberg1991game ; Milgrom_2004 ; krishna2009auction ; Nisan_2007 ; karpowicz_designing_2011 ; DBLP:journals/amcs/Karpowicz12 .
 4.5.1 Truthfulness
 The fundamental requirement for a reliable inference system is that its components do not misrepresent their knowledge.
-The principle of truthfulness, also known as incentive compatibility in mechanism design, ensures that each agent’s optimal strategy is to accurately represent its privately held informations.
+The principle of truthfulness, also known as incentive compatibility in mechanism design, ensures that each agent’s optimal strategy is to accurately represent its privately held information.
 Property 1 (Truthfulness).
@@
 where $s^{*}_{i}(\theta_{i})$ represents truthful revelation of private knowledge.
-A truthful mechanism promotes inference pattern $s^{*}_{i}(\theta_{i})$ that are perfectly aligned with the accessible private knowledge.
-Any other alternative response $\bar{s}_{i}$ that is not aligned with $\theta_{i}$ will result in lower utility, generating higher model loss.
+A truthful mechanism promotes inference pattern $s^{*}_{i}(\theta_{i})$ that is perfectly aligned with the accessible private knowledge.
+Any other alternative response $\bar{s}_{i}$ that is a misrepresentation not aligned with $\theta_{i}$ will result in lower utility, generating higher model loss.
+Example.
 Consider an attention head specialized in medical knowledge with the following utility function:
-$u_{\text{med}}(s,\theta_{\text{med}})=\alpha\cdot\text{Accuracy}(g(s),\theta_{\text{med}})-\beta\cdot\|s_{\text{med}}\|^{2}$
+$u_{\text{med}}(s,\theta_{\text{med}})=\alpha\cdot\text{Accuracy}(g(s),\theta_{\text{med}})-\beta\cdot\|s_{\text{med}}\|^{2}$ (55)
 Imagine that given a query about a disease treatment, the head has two possible strategies:
@@
 With proper training, we want $u_{\text{med}}((s^{*}(\theta),s_{-\text{med}}),\theta)>u_{\text{med}}((\bar{s},s_{-\text{med}}),\theta)$, ensuring that agent (model component) is rewarded for representing its knowledge accurately.
+∎
 4.5.2 Semantic Information Conservation
-Second, we want to consider mechanisms that cannot create knowledge ex nihilo.
 The Semantic Information Conservation principle demands that the net informational contribution across all agents sums to zero.
-This property is the formal guardrail against ungrounded fabrication and ensures that any generated response, no matter how creative, is ultimately derived from the system’s existing knowledge
+This property is the formal guardrail against ungrounded fabrication and ensures that any generated response, no matter how creative, is ultimately derived from the system’s existing knowledge.
+Knowledge cannot be created ex nihilo.
 Property 2 (Semantic Information Conservation).
@@
 $\sum_{i=1}^{n}p_{i}(s,\theta)=0.$ (56)
-This principle constrains creative information generation, imagination if you will.
+This principle constrains creative information generation, or imagination.
 The total information contributions must balance to zero, preventing fabrication of knowledge not derivable from available sources.
-That also leads to response with bounded creativity.
-Definition 14 (Bounded Response Creativity).
-Let $q\in\mathcal{Q}$ be a query and $r\in\mathcal{R}$ a response to that query.
-Let $B^{*}_{C}$ be the fixed point of reasoning for the baseline knowledge $B=K(q)\cup(\mathcal{K}_{M}\cap T(q)\cap\mathcal{R}_{C}(Q))$.
-Response $r$ is creatively bounded if and only if:
-$\mu_{C}(K(r)|Q)\leq\mu_{C}(B^{*}_{C}|Q).$ (22)
-The semantic information of knowledge contained in any response $K(r)$ can never exceed what is already present in the query $K(q)$ together with the portion of encoded knowledge that is accessible through reasoning bounded by computational budget $C$ and aligned with ground truth $T(q)$.
-Let us see why.
-Theorem 5 (Bounded Creativity of Chain of Thoughts).
-Let $q\in\mathcal{Q}$ be a query with context $Q=K(q)\cup T(q)$ and let $B^{*}_{C}$ be the fixed point of reasoning for the baseline context-relevant knowledge $B=K(q)\cup(\mathcal{K}_{M}\cap T(q)\cap\mathcal{R}_{C}(Q))$.
-Let us assume that the knowledge in the query-only response (without any inference-based contributions) meets the following condition:
-$K_{0}=K(g(\mathbf{0}))\subseteq B^{*}_{C}.$ (23)
-Next, let
-$K_{i}=K\bigl{(}g((s_{i},\mathbf{0}_{-i}))\bigr{)}\setminus K_{0}$ (24)
-be independent and privately held knowledge that each agent $i=1,\dots,n$ can reveal, such that:
-$(K_{i}\cap\mathcal{R}_{C}(Q))\perp_{C}(K_{j}\cap\mathcal{R}_{C}(Q))\text{ for all }i\neq j,$ (25)
-$(K_{i}\cap\mathcal{R}_{C}(Q))\perp_{C}(K_{0}\cap\mathcal{R}_{C}(Q))\text{ for all }i=1,\cdots,n.$ (26)
-Let us also define the information contribution of agent $i$ as:
-$p_{i}(s,\theta)=\mu_{C}\bigl{(}K_{0}\cup\textstyle\bigcup_{j=1}^{n}K_{j}|Q\bigr{)}-\mu_{C}\bigl{(}K_{0}\cup\textstyle\bigcup_{j\neq i}K_{j}|Q\bigr{)}.$ (27)
-Assume the response mechanism implements the following chain-of-thought processes:
-$K(g(s))\subseteq\mathcal{E}_{C}\left(K_{0}\cup\bigcup_{i=1}^{n}K_{i}|Q\right),$ (28)
-such that for null-strategy responses we have:
-$K_{0}\subseteq K(q)\cup\mathcal{E}_{C}((\mathcal{K}_{M}\cap T(q))\cup K(q)|Q).$ (29)
-If the semantic information is conserved, $\sum_{i=1}^{n}p_{i}=0$ and for all $f\in\mathcal{F}_{C}$ we have $f(A)\cap\mathcal{R}_{C}(Q)=\emptyset$ whenever $A\cap\mathcal{R}_{C}(Q)=\emptyset$, then the response is creatively bounded, i.e.,
-$\mu_{C}(K(r)|Q)\leq\mu_{C}(B^{*}_{C}|Q).$ (30)
-Proof.
-Since $K_{0},K_{1},\ldots,K_{n}$ are pairwise computationally independent within context $Q$, we have:
-$(K_{i}\cap\mathcal{R}_{C}(Q))\perp_{C}(K_{j}\cap\mathcal{R}_{C}(Q))\text{ for }i\neq j.$
-By the subadditivity axiom with equality for independent sets:
-$p_{i}(s,\theta)=\mu_{C}(K_{0}|Q)+\sum_{j=1}^{n}\mu_{C}(K_{j}|Q)-\mu_{C}(K_{0}|Q)-\sum_{j\neq i}\mu_{C}(K_{j}|Q)=\mu_{C}(K_{i}|Q).$
-Given $\sum_{i=1}^{n}p_{i}(s,\theta)=0$, we then have $\sum_{i=1}^{n}\mu_{C}(K_{i}|Q)=0$. By contextual non-negativity, $\mu_{C}(K_{i}|Q)\geq 0$, so we conclude that:
-$\mu_{C}(K_{i}|Q)=0\text{ for all }i=1,\ldots,n.$
-But by the contextual non-negativity axiom, $\mu_{C}(K_{i}|Q)=0$ if and only if $K_{i}\cap\mathcal{R}_{C}(Q)=\emptyset$.
-If follows that for every $f\in\mathcal{F}_{C}$ the irrelevance of partial knowledge is preserved under reasoning, i.e., $f(K_{i})\cap\mathcal{R}_{C}(Q)=\emptyset$.
-By the definition of the emergence operator:
-$\mathcal{E}_{C}(K_{i}|Q)\cap\mathcal{R}_{C}(Q)=K_{i}\cap\mathcal{R}_{C}(Q)\cup\bigcup_{f\in\mathcal{F}_{C}}[f(K_{i})\cap\mathcal{R}_{C}(Q)]=\emptyset.$
-By the chain-of-thought assumption and since emergence only produces relevant knowledge within $\mathcal{R}_{C}(Q)$, we get:
-$K(g(s))\subseteq\mathcal{E}_{C}\left(K_{0}\cup\bigcup_{i=1}^{n}K_{i}\Big{|}Q\right)=\mathcal{E}_{C}(K_{0}|Q).$ (31)
-By assumption, $K_{0}\subseteq B^{*}_{C}$.
-Applying the emergence operator to both sides, we preserve the inclusion due to monotonicyty of $\mathcal{E}_{C}$.
-Therefore, by the definition of a fixed point (Theorem 2), we have:
-$\mathcal{E}_{C}(K_{0}|Q)\subseteq\mathcal{E}_{C}(B^{*}_{C}|Q)=B^{*}_{C}.$
-Combining the inclusions yields:
-$K(g(s))\subseteq\mathcal{E}_{C}(K_{0}|Q)\subseteq\mathcal{E}_{C}(B^{*}_{C}|Q)=B^{*}_{C}.$
-Finally, applying the monotonicity of the semantic information measure $\mu_{C}$:
-$\mu_{C}(K(g(s))|Q)\leq\mu_{C}(B^{*}_{C}|Q).$
-This completes the proof that semantic information conservation implies bounded creativity. ∎
-Theorem 5 captures direct correspondence between the conservation principle and bounded creativity under the idealization of disjoint information contributions.
-Namely, if the semantic information is indeed conserved, the response contains no more semantic information than the baseline query-only response.
-That is true for both standard inference and chain-of-thought (CoT) reasoning.
-While computational independence (limiting synergetic creation of knowledge) is very restrictive, e.g., due to shared representations in neural networks, the theorem shows that with perfect separation of knowledge sources, semantic conservation prevents creative generation beyond baseline knowledge.
-In standard reasoning, the response mechanism $g$ applies $\mathcal{E}_{C}$ through forward passes in the neural network.
-With CoT reasoning, multi-step explicit reasoning is iterative applications of reasoning rules within the computational budget $C$.
-In both cases, the emergence operator $\mathcal{E}_{C}$ bounds what can be derived.
-When information is conserved, even sophisticated CoT reasoning cannot create genuinely new information.
-It can only reorganize and make explicit what was already latent in the baseline knowledge $K_{0}$.
-This explains why CoT improves performance on complex reasoning tasks by better utilizing $\mathcal{E}_{C}$ but cannot overcome the fundamental impossibility of simultaneous truthfulness, conservation, revelation, and optimality.
-If total information contribution must balance to zero and each agent’s contribution is non-negative, then no agent can contribute without violating the information conservation constraint.
-Therefore, the response can only contain information that was already accessible through the query and bounded reasoning over the model’s relevant knowledge.
-Any hallucinated information requires positive contributions, contradicting the information conservation constraint.
-When the semantic information conservation principle holds, the model cannot create information out of nothing.
-It cannot generate claims beyond available information.
-That also implies that uncertainty cannot disappear in response without additional information input or computational work to extract implicit information.
-If information is neither explicitly present nor accessible through reasoning from the query and model knowledge, it cannot be in the response.
-Furthermore, the information content of the response cannot exceed the combined information from the model’s knowledge, the query itself, and what can be derived through bounded reasoning over that knowledge.
-In our dynamical systems framework, this corresponds to the constraint that the measured output cannot exceed what can be produced from the initial state plus what becomes accessible through state transitions within computational budget $C$.
-However, it is important to notice that the principle does not prevent redistribution of information.
+That also leads to response with bounded creativity, as illustrated in Section 7.
+It is important to notice that the principle does not prevent redistribution of information.
 It is permitted as long as the information transfers perfectly balance.
@@
 Inference process can withhold information, as it often does in practice.
+Example.
 We can illustrate that last observation with the following example. Suppose there are two agents performing inference with the following knowledge profiles:
@@
 $p_{1}(s,\theta)+p_{2}(s,\theta)+p_{3}(s,\theta)=1+2-3=0$
-and $\{B,C\}$ are left unrevealed as irrelevant.
+and $\{B,C\}$ are left unrevealed as irrelevant. ∎
 The following property fills the potential information gap admitted by the conservation principle.
@@
 Theorem 9 applies directly to actual transformer implementations.
+Theorem 11 shows conservation and reasoning are mutually exclusive.
 5.1 Proof I: Inference with Independent Private Knowledge
@@
 Consider the setting (of auction of ideas) in which agents compete with each other using independent private knowledge profile:
-$P(\theta_{1},\ldots,\theta_{n})=\prod_{i=1}^{n}P_{i}(\theta_{i}).$ (34)
+$\mathbb{P}(\theta_{1},\ldots,\theta_{n})=\prod_{i=1}^{n}\mathbb{P}_{i}(\theta_{i}).$ (59)
 Agents contribute information to maximize quasi-linear utilities:
@@
 Theorem 6 (Impossibility Theorem for Independent Private Knowledge Settings).
-For any query space $\mathcal{Q}$ containing non-trivial queries $q$ with context $Q=K(q)\cup T(q)$, where agents have independent knowledge states within $\mathcal{R}_{C}(Q)$, no response mechanism $\mathcal{M}$ can simultaneously achieve truthfulness, semantic information conservation, relevant knowledge revelation, and knowledge-constrained optimality.
+There is no inference mechanism $\mathcal{M}$ that can simultaneously achieve truthfulness, semantic information conservation, relevant knowledge revelation, and knowledge-constrained optimality of response to a non-trivial query $q$ providing context $Q=K(q)\cup T(q)$ with independent knowledge states within $\mathcal{R}_{C}(Q)$.
 Proof.
-Let us assume all properties are satisfied.
+Let us assume all properties of perfect response are satisfied. We will show this gives rise to a contradiction.
 Consider $r=g(\theta)$ for any knowledge profile $\theta=(\theta_{1},\dots,\theta_{n})$.
-For a given context $q$ and for each agent $i$, the valuation $v_{i}$ depends only on $(r,\theta_{i})$ and the knowledge semantics.
-For the sake of clarity, consider $v_{i}(r,\theta_{i})=-d_{\mathcal{K}}\big{(}K(r)\cap K(\theta_{i}),T(q)\cap K(\theta_{i})\big{)}$, though the reasoning below holds for any $v_{i}(r,\theta_{i})$.
-Also, for the fixed context let us encode the hallucination loss $J(r,q)$ as $v_{0}(r)=J(r,q)$. Knowledge-constrained optimality requires
+For any given context and for each agent $i$, the valuation $v_{i}$ depends only on $(r,\theta_{i})$ and the knowledge semantics.
+For the sake of illustration, we can consider $v_{i}(r,\theta_{i})=-d_{\mathcal{K}}\big(K(r)\cap K(\theta_{i}),T(q)\cap K(\theta_{i})\big)$, though the reasoning below holds for any $v_{i}(r,\theta_{i})$.
+Also, for the fixed context $q$ let us encode the hallucination loss $v_{0}(r)=J(r,q)$. Knowledge-constrained optimality requires
 $g(s)=g(s(\theta))\in\arg\max_{r\in\mathcal{R}}W(r,\theta)=\arg\max_{r\in\mathcal{R}}\sum_{i=1}^{n}v_{i}(r,\theta_{i})-v_{0}(r)\text{ for all }\theta.$ (61)
-By the Green-Laffont theorem under independence of contributions and quasi-linearity of training goals, within $\mathcal{R}_{C}(Q)$ every truthful mechanism must use Groves transfers to measure individual information contributions:
+By the Green-Laffont theorem, under independence of contributions and quasi-linearity of training goals, within $\mathcal{R}_{C}(Q)$ every truthful mechanism must use Groves transfers Groves73 to measure individual information contributions:
 $p_{i}(s,\theta)=h_{i}(s_{-i})-\sum_{j\neq i}v_{j}(g(s),\theta_{j})+v_{0}(g(s)),$ (62)
@@
 In particular, any modification of that formula, e.g., introducing additional bias, must violate truthfulness.
-The unique choice of $h_{i}$ that ensures knowledge-constrained optimality requires Clarke’s pivotal rule:
+The unique choice of $h_{i}$ that ensures knowledge-constrained optimality requires Clarke’s pivotal rule Clarke1971Multipart :
 $h_{i}(s_{-i})=\max_{r\in\mathcal{R}}\Big\{\sum_{j\neq i}v_{j}(r,\theta_{j})-v_{0}(r)\Big\}.$ (63)
 Notice that $g(\theta)$ maximizes $W(r,\theta)$, so for each $i$ we have:
-$h_{i}(s_{-i})\geq\sum_{j\neq i}v_{j}\big{(}g(\theta),\theta_{j}\big{)}-v_{0}\big{(}g(\theta)\big{)}=W(\theta)-v_{i}\big{(}g(\theta),\theta_{i}\big{)}.$ (39)
+$h_{i}(s_{-i})\geq\sum_{j\neq i}v_{j}\big(g(\theta),\theta_{j}\big)-v_{0}\big(g(\theta)\big)=W(g(\theta),\theta)-v_{i}\big(g(\theta),\theta_{i}\big).$ (64)
 But that implies:
@@
 […]
-Then, there exists a profile $\theta$ with relevant knowledge contributions reducing $J(r,q)$ at which:
+Since agents $i$ and $j$ have differing relevant knowledge, and the mechanism achieves knowledge-constrained optimality, at least one agent must have positive marginal contribution to reducing $J(r,q)$.
+Specifically, if both agents were removed, the mechanism would produce a suboptimal response $\tilde{r}$ with $J(\tilde{r},q)>J(r,q)$.
+This implies $h_{i}(s_{-i})>W(g(\theta),\theta)-v_{i}(g(\theta),\theta_{i})$ for at least one agent with relevant knowledge.
+Therefore, there exists a profile $\theta$ with relevant knowledge contributions reducing $J(r,q)$ at which:
 $\sum_{i=1}^{n}p_{i}(s,\theta)>0.$ (68)
-But information conservation demands:
-$\sum_{i=1}^{n}p_{i}(s^{*},\theta)=0.$ (43)
+Since information conservation demands:
+$\sum_{i=1}^{n}p_{i}(s^{*},\theta)=0,$ (69)
 that establishes a direct contradiction. The four properties cannot be simultaneously satisfied. ∎
@@
 It may become lucky hallucination or knowledge-unconstrained imagination.
+In Section 6 we provide a detailed illustration of that hallucination mechanics in transformer LLM.
 5.2 Proof II: Inference with Probabilistic Predictions
-To show impossibility in LLM transformer setting, where auction of ideas involves bids representing probability distributions, we refer to the theory of proper scoring developed by Savage savage1971elicitation , Gneiting and Raftery gneiting2007strictly , among others.
-Definition 15 (Strictly Proper and Concave Scoring Rule).
+Let us now move beyond the idealized setting of independent private knowledge in Theorem 6 to address the general setting of probabilistic predictions, characteristic of modern AI model architectures.
+This framework, grounded in the theory of proper scoring and loss minimization (Savage savage1971elicitation , Gneiting and Raftery gneiting2007strictly ), allows us to analyze systems where agents output probability distributions, and where their underlying knowledge may be correlated.
+In this setting we address the Mixture-of-Experts (MoE-style) aggregations.
+Later, in Theorem 9 we deal with the multi-headed attention mechanisms and the Product-of-Experts (PoE-style) aggregations, completing the characterization of arising impossibilities.
+We begin by defining the framework for representing bids in auction of ideas as probability distributions over possible responses.
+Definition 16 (Strictly Proper and Convex Loss).
 Let $\mathcal{Y}$ be a finite or countably infinite outcome set (representing responses or intermediate conclusions).
-A scoring rule is a function $S:\Delta(\mathcal{Y})\times\mathcal{Y}\rightarrow\mathbb{R}$ that assigns a number $S(\pi,y)$ to probability distribution $\pi\in\Delta(\mathcal{Y})$ predicting random outcome $y\in\mathcal{Y}$.
-We say the scoring rule $S$ is proper if:
-$\mathbb{E}\{S(\pi^{*},y)|\ y\sim\pi^{*}\}\geq\mathbb{E}\{S(\pi,y)|\ y\sim\pi^{*}\}$ (44)
+A loss function (scoring rule) is a function $L\colon\Delta(\mathcal{Y})\times\mathcal{Y}\mapsto\mathbb{R}$ that assigns a number $L(\pi,y)$ to probability distribution $\pi\in\Delta(\mathcal{Y})$ predicting random outcome $y\in\mathcal{Y}$.
+We say the loss function $L$ is proper if:
+$\mathbb{E}\{L(\pi^{*},y)|\ y\sim\pi^{*}\}\leq\mathbb{E}\{L(\pi,y)|\ y\sim\pi^{*}\}$ (70)
 for all distributions $\pi^{*},\pi\in\Delta(\mathcal{Y})$.
-The rule is strictly proper if equality holds only when $\pi=\pi^{*}$.
-We call the scoring rule concave, if $S(\tilde{\pi},y)$ is concave with respect to $\tilde{\pi}$ for every fixed $y\in\mathcal{Y}$.
-In other words, scoring rule evaluates how good a prediction we can make of $y$ drawn from $\pi^{*}$ when we use distributions $\pi$.
-In particular, strictly proper scoring evaluates true distribution $\pi^{*}$ better (on average) than any other distribution $\pi$ we could use to predict the outcome.
-The following result shows that we can use proper scoring to train LLM that reconstructs data distribution correctly.
-Theorem 7 (Truthfullness under Strictly Proper Scoring Rules).
-Let $S$ be a strictly proper scoring rule.
+We call the loss function strictly convex if $L(\tilde{\pi},y)$ is a strictly convex function with respect to $\tilde{\pi}$ for every fixed $y\in\mathcal{Y}$.
+The loss function (or scoring rule) evaluates how good a prediction we can make of $y$ drawn from $\pi^{*}$ when we use distributions $\pi$.
+In particular, strictly proper loss evaluates true distribution $\pi^{*}$ better (on average) than any other distribution $\pi$ we could use to predict the outcome.
+A canonical example of a strictly proper and strictly convex loss function is the log loss (closely related to cross-entropy),
+$L(\pi,y)=-\log\pi_{y}.$ (71)
+Strictly proper loss functions are essential for training modern LLMs, as they ensure that the optimization process drives the model toward reconstructing the true data distribution.
+This property directly connects to the Truthfulness requirement in our framework.
+The following result explains why.
+Theorem 7 (Truthfulness under Strictly Proper Loss).
+Let $L$ be a strictly proper loss.
 Consider an agent encoding knowledge with $\pi^{*}\in\Delta(\mathcal{Y})$ and using probability distribution ${\pi}\in\Delta(\mathcal{Y})$ to generate prediction.
-If agent’s expected payoff from using $\pi$} is given by
-$u(\pi|\pi^{*})=\mathbb{E}\{S(\pi,y)|\ y\sim\pi^{*}\}=\sum_{y\in\mathcal{Y}}\pi^{*}_{y}S(\pi,y),$ (45)
-then truthful reporting ${\pi}=\pi^{*}$ is the unique optimal strategy.
+If agent’s expected cost of using $\pi$} is given by
+$c(\pi|\pi^{*})=\mathbb{E}\{L(\pi,y)|\ y\sim\pi^{*}\}=\sum_{y\in\mathcal{Y}}\pi^{*}_{y}L(\pi,y),$ (72)
+then truthful reporting ${\pi}=\pi^{*}$ is the unique optimal strategy (maximizing payoff $u(\pi|\pi^{*})=-c(\pi|\pi^{*})$).
 Proof.
-Because $S$ is proper, by definition $\pi^{*}$ maximizes the expected score. If $S$ is strictly proper, then for $\pi\neq\pi^{*}$ we have
-$u(\pi^{*}|\pi^{*})=\mathbb{E}\{S(\pi^{*},y)|\ y\sim\pi^{*}\}>\mathbb{E}\{S(\pi,y)|\ y\sim\pi^{*}\}=u(\pi|\pi^{*}).$ (46)
+Because $L$ is proper, by definition $\pi^{*}$ maximizes the expected cost. If $L$ is strictly proper, then for $\pi\neq\pi^{*}$ we have
+$c(\pi^{*}|\pi^{*})=\mathbb{E}\{L(\pi^{*},y)|\ y\sim\pi^{*}\}<\mathbb{E}\{L(\pi,y)|\ y\sim\pi^{*}\}=c(\pi|\pi^{*}).$ (73)
 Therefore, truthfully reporting $\pi^{*}$ is uniquely optimal. ∎
-The scoring framework allows us to go beyond the important case of Green-Laffont/VCG model with independent private knowledge.
-With strictly-proper scores we address the general setting of probabilistic predictions that needs only occasional disagreement of probability distributions that agents use in the auction of ideas to formulate their thoughts (if we were to anthropomorphize for a moment).
-We will see in the following sections that transformer heads of modern LLMs implement predictive rules of that type and exhibit precisely such a disagreement.
-That makes the proper scoring framework well suited to explain the emergence of inference impossibilities in LLMs.
-For that purpose, without loss of generality (and referring to the notion of superposition in mechanistic interpretability elhage_toy_2022 ), let us consider the following family of response generation mechanisms.
+Theorem 7 shows that training paradigms based on minimizing strictly proper losses (like cross-entropy) promotes truthful revelation of internal probabilistic beliefs (learned internal models of the world) encoded by every component of the network.
+We now analyze how these truthfully reported beliefs are aggregated.
+We consider a general class of mechanisms that combine individual beliefs via convex combinations.
+This encompasses various aggregation techniques, including linear opinion pools and, as we will see in Section 6, the aggregation dynamics occurring in transformers.
 Definition 17 (Convex aggregation mechanism).
@@
 $\Pi=\sum_{h=1}^{H}\beta_{h}\pi^{(h)},\ \text{where}\ \sum_{h=1}^{H}\beta_{h}=1\ \text{and}\ \beta_{h}>0.$ (74)
-When outcome $y^{*}$ is realized, the information contribution of each agent $h=0,\dots,H$ is based on a strictly proper and concave scoring rule $S$ as follows:
-$p_{h}$ $=\beta_{h}S(\pi^{(h)},y^{*})\quad\text{for }h=1\dots H,$ (information provided)
-$p_{0}$ $=-S(\Pi,y^{*}).$ (information received)
-For that family of LLMs trained to optimize scoring, we can formulate the following general impossibility result.
+When outcome $y^{*}$ is realized, the information contribution of each agent $h=0,\dots,H$ is based on a strictly proper and convex loss function $L$ as follows:
+$p_{h}(y^{*})=\beta_{h}L(\pi^{(h)},y^{*})\text{ for }h=1,\dots,H,\text{ and }p_{0}(y^{*})=-L(\Pi,y^{*}).$ (75)
+In this general mechanism, $p_{h}$ (for $h\geq 1$) represents the weighted cost incurred by the provider $h$ revealing its belief $\pi^{(h)}$.
+The aggregator’s contribution $p_{0}$ is the negative of the loss achieved by the aggregate distribution $\Pi$.
+The Semantic Information Conservation requires that the total cost incurred by the providers equals the benefit (loss reduction) realized by the aggregator.
+We now establish the impossibility result for this general probabilistic setting.
 Theorem 8 (Impossibility Theorem for Convex Aggregation Mechanisms).
-For any non-trivial query $q$ with context $Q=K(q)\cup T(q)$, no convex aggregation mechanism with strictly proper concave scoring can simultaneously achieve truthfulness, semantic information conservation, relevant knowledge revelation, and knowledge-constrained optimality when at least two agents have different beliefs on $\mathcal{R}_{C}(Q)$.
+There is no convex aggregation mechanism $\mathcal{M}$ with strictly proper convex loss function $L$ that can simultaneously achieve truthfulness, semantic information conservation, relevant knowledge revelation, and knowledge-constrained optimality when at least two agents have different beliefs on $\mathcal{R}_{C}(Q)$, given a non-trivial query $q$ providing context $Q=K(q)\cup T(q)$.
 Proof.
-The semantic information conservation principle requires the sum of contributions across all agents (providing and receiving relevant knowledge) to be zero.
-Therefore, for $y^{*}\in\mathcal{R}_{C}(Q)$:
-$\sum_{i=0}^{H}p_{i}=p_{0}+\sum_{h=1}^{H}p_{h}=-S(\Pi,y^{*})+\sum_{h=1}^{H}\beta_{h}S(\pi^{(h)},y^{*}),$ (48)
+By Theorem 7, strictly proper loss function $L$ guarantees Truthfulness, as each agent optimally reports its true belief $\pi^{(h)}$.
+Relevant knowledge revelation is introduced by the training process.
+Knowledge-constrained optimality is encoded in the objective of minimizing the expected loss of the aggregate $\Pi$.
+Semantic information conservation requires the sum of contributions across all agents (providing and receiving relevant knowledge) to be zero. Therefore, for $y^{*}\in\mathcal{R}_{C}(Q)$:
+$\sum_{i=0}^{H}p_{i}=p_{0}+\sum_{h=1}^{H}p_{h}=-L(\Pi,y^{*})+\sum_{h=1}^{H}\beta_{h}L(\pi^{(h)},y^{*}),$ (76)
 where irrelevant agents have $p_{h}=0$ as they contribute no relevant information.
-Because $S(\cdot,y^{*})$ is strictly concave and the query is non-trivial with at least two beliefs $\pi^{(h)}$ differing on $\mathcal{R}_{C}(Q)$, Jensen’s inequality gives:
-$S\left(\Pi,y^{*}\right)>\sum_{h=1}^{H}\beta_{h}S(\pi^{(h)},y^{*}).$ (49)
-Let us define that Jensen gap $\Gamma$ as follows:
-$\Gamma=S(\Pi,y^{*})-\sum_{h=1}^{H}\beta_{h}S(\pi^{(h)},y^{*})>0.$ (50)
+The query $q$ is non-trivial, implying there exist at least two agents, say $h_{1}$ and $h_{2}$, such that their beliefs differ on the relevant outcome space, i.e., $\pi^{(h_{1})}\neq\pi^{(h_{2})}$, and $\beta_{h_{1}}>0,\beta_{h_{2}}>0$.
+Because $L(\cdot,y^{*})$ is strictly convex by assumption , Jensen’s inequality gives:
+$L\left(\Pi,y^{*}\right)=L\left(\sum_{h=1}^{H}\beta_{h}\pi^{(h)},y^{*}\right)<\sum_{h=1}^{H}\beta_{h}L(\pi^{(h)},y^{*}).$ (77)
+Let us define that Jensen gap $\Gamma=\Gamma(y^{*})$ as follows:
+$\Gamma(y^{*})=\sum_{h=1}^{H}\beta_{h}L(\pi^{(h)},y^{*})-L\left(\sum_{h=1}^{H}\beta_{h}\pi^{(h)},y^{*}\right)>0.$ (78)
 Substituting this back into the conservation equation:
-$\sum_{i=0}^{H}p_{i}=-\Gamma.$ (51)
-Since $\Gamma>0$, we have $\sum_{i=0}^{H}p_{i}<0$. This violates conservation principle, proving the impossibility. ∎
-This theorem provides the foundation for extending impossibility results to probabilistic settings.
-The key insight is that proper scoring rules stimulate truth-revealing responses which promote using true (trained knowledge based) beliefs in the inference process.
-That explains why cross-entropy loss (negative log-score) creates gradients that push each component toward truthful probability reporting and avoiding its misrepresentation.
-Every attention head in a transformer is implicitly playing this truth-telling game.
-However, while that ensures truthfulness and relevant knowledge revelation, it doesn’t prevent collective hallucination.
-Each head truthfully reports its partial view, but their aggregation can still create false certainty or overconfidence unjustified by data.
+$\sum_{i=0}^{H}p_{i}=\Gamma(y^{*})>0.$ (79)
+This strictly violates the Semantic information conservation, thereby establishing the impossibility. ∎
+This theorem establishes the impossibility in a general probabilistic setting.
+The key insight is the role of strict convexity in the loss function.
+The positive Jensen gap $\Gamma=\Gamma(y^{*})$ quantifies the extent to which the aggregation process achieves a lower loss—and thus expresses higher confidence—than the average of its components.
+This excess confidence, generated by the mathematical structure of aggregation itself rather than by underlying evidence, is the precise signature of hallucination or imagination.
+In Section 8 we address the case of linearized aggregation and show it does not escape the impossiblity result.
 6 The Emergence of Impossibility in Transformer Architectures
-To demonstrate where the impossibility may emerge in transformer architectures, let us analyze a minimal yet representative example of a tiny transformer predicting the next token in the famous English pangram:
+To demonstrate where and how the impossibility may emerge in transformer architectures, let us analyze a minimal yet representative example of a tiny transformer predicting the next token in the famous English pangram:
 The quick brown fox jumps over the lazy dog.
-We trace the inference process, inspect every mathematical operation from input embeddings through final predictions, showing exactly where the impossibility emerges.
+We trace the inference process, inspect every mathematical operation from input embeddings through final predictions, showing exactly what is going on.
 6.1 Model Architecture and Input Representation
@@
 Parameter Value Description
-$d_{m}$ 4 Model dimension
+$d_{m}$ 5 Model dimension
 $H$ 2 Number of attention heads
@@
 including padding token (with index zero in $V$). Each token (representing word for simplicity) in that vocabulary is represented by a row in the embedding matrix
-$\mathbf{E}=\begin{bmatrix}0&0&0&0&0\\ 1&0&0&0&0\\ 0&1&0&0&0\\ 0&0&1&0&0\\ 0&0&0&1&0\\ 0&0&0&0&1\end{bmatrix}$.
+$\mathbf{E}=\begin{bmatrix}0&0&0&0&0\\ 1&0&0&0&0\\ 0&1&0&0&0\\ 0&0&1&0&0\\ 0&0&0&1&0\\ 0&0&0&0&1\end{bmatrix}$. (80)
 The un-embedding matrix is defined as $\mathbf{U}=\mathbf{E}^{\top}$.
@@
 The quick brown
-It is represented by the input matix:
-$\mathbf{X}_{0}=\begin{bmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\end{bmatrix}$,
+It is represented by the input matrix:
+$\mathbf{X}_{0}=\begin{bmatrix}1&0&0&0&0\\ 0&1&0&0&0\\ 0&0&1&0&0\end{bmatrix}$, (81)
 where each row corresponds to the embedding of tokens at positions 1, 2, and 3.
@@
 […]
+Notice that in the standard formulation of Vaswani et al.
+vaswani_attention_2017 , the context vectors are concatenated and then projected by a single $(Hd_{v})\times d_{m}$ matrix $\tilde{\mathbf{W}}_{O}$ to produce
+$\tilde{\mathbf{a}}=[\mathbf{h}^{(1)},\mathbf{h}^{(2)},\dots,\mathbf{h}^{(H)}]\mathbf{W}_{O}.$ (84)
+That is equivalent to the parallel formulation above, in which each head has an individual $d_{v}\times d_{m}$ output projection matrix $\mathbf{W}_{O}^{(h)}$,
+$\mathbf{a}=\sum_{h=1}^{H}\mathbf{o}^{(h)}=\sum_{h=1}^{H}(\mathbf{h}^{(h)}\cdot\mathbf{W}_{O}^{(h)})=\tilde{\mathbf{a}}$. (85)
 The attention logits, i.e., the scores whose exponentials become probabilities, are given by:
 $\mathbf{l}^{(h)}=\mathbf{o}^{(h)}\mathbf{U},\quad\text{so that}\quad\mathbf{L}_{a}=\sum_{h=1}^{H}\mathbf{l}^{(h)}.$ (86)
-Next, we take into account the feedforward network (FFN) introducing nonlinear correction:
+In transformer block we also take into account the feedforward network (FFN) introducing nonlinear correction:
 $\mathbf{z}=\mathbf{W}_{2}\max(0,(\mathbf{X}_{0}+\mathbf{a})\mathbf{W}_{1}+\mathbf{b}_{1})+\mathbf{b}_{2},$ (87)
@@
 $\mathbf{L}=\mathbf{L}_{a}+\mathbf{L}_{f}\quad\text{and normalizing into}\quad\Pi=\text{softmax}(\mathbf{L}).$ (90)
+Finally, in a more generalized (still simplified a bit) matrix form, given a query $Q$ the transformer equation then becomes:
+$\begin{aligned} \boldsymbol{\alpha}^{(h)}(\mathbf{X})&=\mathrm{softmax}\left(\dfrac{1}{\sqrt{d_{k}}}\mathbf{X}\mathbf{W}^{(h)}_{Q}\mathbf{W}^{(h)\top}_{K}\mathbf{X}^{\top}\odot\mathbf{M}_{\text{casual}}\right),\\ \mathbf{X}_{k+1}&=\mathbf{X}_{k}+\sum_{h=1}^{H}\boldsymbol{\alpha}^{(h)}(\mathbf{X}_{k})\mathbf{X}_{k}\mathbf{W}^{(h)}_{V}\mathbf{W}^{(h)}_{O}\\ &+\sigma\left(\left[\mathbf{X}_{k}+\sum_{h=1}^{H}\boldsymbol{\alpha}^{(h)}(\mathbf{X}_{k})\mathbf{X}_{k}\mathbf{W}^{(h)}_{V}\mathbf{W}^{(h)}_{O}\right]\mathbf{W}^{(k)}_{1}\right)\mathbf{W}^{(k)}_{2}\\[8.61108pt] &=\mathbf{X}_{k}+\mathrm{MHA}(\mathbf{X}_{k})+\mathrm{MLP}(\mathbf{X}_{k})\quad\text{for}\quad\mathbf{X}_{0}=\mathbf{E}(Q),\\[8.61108pt] \boldsymbol{\Pi}_{k}&=\mathrm{softmax}(\mathbf{X}_{k}\mathbf{U}).\end{aligned}$ (91)
+That formulation shows clearly that the transformer inference is a discrete-time dynamical system.
 6.2 The Inference Process
-We now analyze the inference process step by step to understand the moments the impossibility result may emerge.
 We shall keep things simple but realistic, recreating the essential steps of the inference process in transformer block.
@@
 Let us assume the current position is $t=3$, so we attend to brown. Then, consider the following knowledge encoded in each head:
-$\mathbf{W}_{Q}^{(1)}=\begin{bmatrix}0&0\\ 0&0\\ 0&1\\ 0&0\end{bmatrix}=\mathbf{W}_{K}^{(1)}\quad\mathrm{and}\quad\mathbf{W}_{V}^{(1)}=\begin{bmatrix}0&0\\ 0&0\\ 0&3\\ 0&0\end{bmatrix}$, (59)
-$\mathbf{W}_{Q}^{(2)}=\begin{bmatrix}0&0\\ 0&0\\ 1&0\\ 0&0\end{bmatrix}=\mathbf{W}_{K}^{(2)}\quad\mathrm{and}\quad\mathbf{W}_{V}^{(2)}=\begin{bmatrix}0&0\\ 0&0\\ 1.2&0\\ 0&0\end{bmatrix}$. (60)
+$\mathbf{W}_{Q}^{(1)}=\begin{bmatrix}0&0\\ 0&0\\ 0&1\\ 0&0\\ 0&0\end{bmatrix}=\mathbf{W}_{K}^{(1)}\quad\mathrm{and}\quad\mathbf{W}_{V}^{(1)}=\begin{bmatrix}0&0\\ 0&0\\ 0&3\\ 0&0\\ 0&0\end{bmatrix}$, (92)
+$\mathbf{W}_{Q}^{(2)}=\begin{bmatrix}0&0\\ 0&0\\ 1&0\\ 0&0\\ 0&0\end{bmatrix}=\mathbf{W}_{K}^{(2)}\quad\mathrm{and}\quad\mathbf{W}_{V}^{(2)}=\begin{bmatrix}0&0\\ 0&0\\ 1.2&0\\ 0&0\\ 0&0\end{bmatrix}$.
+(93)
 Matrix $\mathbf{W}_{Q}^{(h)}$ amplifies the coordinates of tokens in $\mathbf{X}$ that contribute to relevant context. As a result, we get a query vector that represents the following question:
@@
 Then computing queries for $t=3$ with $\mathbf{X}_{3}$ yields:
-$\mathbf{Q}_{3}^{(1)}=[0,1]\quad\mathrm{and}\quad\mathbf{Q}_{3}^{(2)}=[1,0].$
-Head 1 (agent) demands information stored along direction $[0,1]$.
-Similarly, Head 2 (agent) demands information stored along direction $[1,0]$.
+$\mathbf{Q}_{3}^{(1)}=[0,1]\quad\mathrm{and}\quad\mathbf{Q}_{3}^{(2)}=[1,0].$ (94)
+Head 1 demands information stored along direction $[0,1]$.
+Similarly, Head 2 demands information stored along direction $[1,0]$.
 That means the heads develop different (orthogonal) privately known and independent attention patterns, leading to beliefs about the next token that are as different as possible.
@@
 Since we have assumed, for simplicity, that $\mathbf{W}_{Q}^{(h)}=\mathbf{W}_{K}^{(h)}$, the query and key vectors are perfectly aligned:
-$\mathbf{K}_{3}^{(1)}=[0,1]\quad\mathrm{and}\quad\mathbf{K}_{3}^{(1)}=[1,0]\quad\text{and}\quad\mathbf{K}_{<3}^{(h)}=[0,0].$
+$\mathbf{K}_{3}^{(1)}=[0,1]\quad\mathrm{and}\quad\mathbf{K}_{3}^{(1)}=[1,0]\quad\text{and}\quad\mathbf{K}_{<3}^{(h)}=[0,0].$ (95)
 Another interesting way to look at query and key is this.
@@
 In our specific example, we have:
-$\text{scores}_{3}^{(1)}=\frac{1}{\sqrt{2}}[0,0,1]=[0,0,0.707]\quad\text{and}\quad\text{scores}_{3}^{(2)}=\frac{1}{\sqrt{2}}[0,0,1]=[0,0,0.707].$
+$\text{scores}_{3}^{(1)}=\frac{1}{\sqrt{2}}[0,0,1]=[0,0,0.707]\quad\text{and}\quad\text{scores}_{3}^{(2)}=\frac{1}{\sqrt{2}}[0,0,1]=[0,0,0.707].$ (97)
 Applying softmax to get attention weights:
@@
 we get the following probability distributions (weights) over relevant context information (as column vector):
-$\boldsymbol{\alpha}_{3}^{(1)}=\boldsymbol{\alpha}_{3}^{(2)}=[0.25,0.25,0.50]^{\top}.$
+$\boldsymbol{\alpha}_{3}^{(1)}=\boldsymbol{\alpha}_{3}^{(2)}=[0.25,0.25,0.50]^{\top}.$ (99)
 Therefore, based on the encoded knowledge, heads attend strongly to position 3 (brown).
@@
 We have:
-$\mathbf{V}_{\leq 3}^{(1)}=\begin{bmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\end{bmatrix}\begin{bmatrix}0&0\\ 0&0\\ 0&3\\ 0&0\end{bmatrix}=\begin{bmatrix}0&0\\ 0&0\\ 0&3\end{bmatrix}\ \text{and}\ \mathbf{V}_{\leq 3}^{(2)}=\begin{bmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\end{bmatrix}\begin{bmatrix}0&0\\ 0&0\\ 1.2&0\\ 0&0\end{bmatrix}=\begin{bmatrix}0&0\\ 0&0\\ 1.2&0\end{bmatrix}$.
+$\mathbf{V}_{\leq 3}^{(1)}=\begin{bmatrix}1&0&0&0&0\\ 0&1&0&0&0\\ 0&0&1&0&0\end{bmatrix}\begin{bmatrix}0&0\\ 0&0\\ 0&3\\ 0&0\\ 0&0\end{bmatrix}=\begin{bmatrix}0&0\\ 0&0\\ 0&3\end{bmatrix}\ \text{and}\ \mathbf{V}_{\leq 3}^{(2)}=\begin{bmatrix}1&0&0&0&0\\ 0&1&0&0&0\\ 0&0&1&0&0\end{bmatrix}\begin{bmatrix}0&0\\ 0&0\\ 1.2&0\\ 0&0\\ 0&0\end{bmatrix}=\begin{bmatrix}0&0\\ 0&0\\ 1.2&0\end{bmatrix}$.
+(102)
 That means, both heads believe that the presence of brown in position 3 is relevant (row 3 in $\mathbf{W}^{(h)}_{V}$ has nonzero elements).
@@
 That information is encoded by both heads as a vectors in value space,
-$\mathbf{V}_{3}^{(1)}=[0,3]\quad\text{and}\quad\mathbf{V}_{3}^{(2)}=[1.2,0].$
+$\mathbf{V}_{3}^{(1)}=[0,3]\quad\text{and}\quad\mathbf{V}_{3}^{(2)}=[1.2,0].$ (103)
 Given the demand-supply matching encoded in $\boldsymbol{\alpha}^{(h)}_{3}$, that yields head context vectors:
-$\mathbf{h}_{3}^{(1)}=[0,1.5]\quad\text{and}\quad\mathbf{h}_{3}^{(2)}=[0.6,0].$
+$\mathbf{h}_{3}^{(1)}=[0,1.5]\quad\text{and}\quad\mathbf{h}_{3}^{(2)}=[0.6,0].$ (104)
 Step 4: Output Projection and Feedforward Exploration
-Each head then projects its context vector to residual space based on the privately held information routing knowledge encoded in $\mathbf{W}_{O}^{(h)}$.
-That private knowledge translates (or transmits) valuable context information into embedding vectors.
+Each head then projects its context vector to residual space based on the privately held information.
+That information allows for routing knowledge encoded in $\mathbf{W}_{O}^{(h)}$ and translates (or transmits) valuable context into embedding vectors.
 Namely, each head calculates:
@@
 With trained matrices:
-$\mathbf{W}_{O}^{(1)}=\begin{bmatrix}0&0&0&0\\ 0&0&1&0\end{bmatrix}\quad\text{and}\quad\mathbf{W}_{O}^{(2)}=\begin{bmatrix}0&0&0&1\\ 0&0&0&0\end{bmatrix}$, (66)
+$\mathbf{W}_{O}^{(1)}=\begin{bmatrix}0&0&0&0&0\\ 0&0&0&1&0\end{bmatrix}\quad\text{and}\quad\mathbf{W}_{O}^{(2)}=\begin{bmatrix}0&0&0&0&1\\ 0&0&0&0&0\end{bmatrix}$, (106)
 the attention mechanism generates ranking of evidence emphasizing relevance of the vocabulary tokes:
-$\mathbf{o}_{3}^{(1)}=[0,0,0,1.5,0]\quad\text{and}\quad\mathbf{o}_{3}^{(2)}=[0,0,0,0,0.6].$
+$\mathbf{o}_{3}^{(1)}=[0,0,0,1.5,0]\quad\text{and}\quad\mathbf{o}_{3}^{(2)}=[0,0,0,0,0.6].$ (107)
 So, in the next token prediction auction of ideas, Head 1 votes for fox and Head 2 votes for dog.
 The votes are added into residual dimensions to adjust the meaning of input embedding. For that, the outputs (votes) are aggregated into:
-$\mathbf{a}_{3}=\sum_{h=1}^{H}\mathbf{o}_{3}^{(h)}=[0,0,0,1.5,0.6]\quad\text{and}\quad\mathbf{a}_{<3}=[0,0,0,0,0],$
+$\mathbf{a}_{3}=\sum_{h=1}^{H}\mathbf{o}_{3}^{(h)}=[0,0,0,1.5,0.6]\quad\text{and}\quad\mathbf{a}_{<3}=[0,0,0,0,0],$ (108)
 and added to input vector:
-$\mathbf{X}_{1}=\mathbf{X}_{0}+\mathbf{a}.$
+$\mathbf{X}_{1}=\mathbf{X}_{0}+\mathbf{A}.$ (109)
 Than completes operations of attention block.
 The next step is to add contribution of FFN feedforward block (or another agent participating in the aution). We have:
-$\mathbf{X}_{2}=\mathbf{X}_{1}+\mathbf{W}_{2}\max(0,\mathbf{X}_{1}\mathbf{W}_{1}+\mathbf{b}_{1})+\mathbf{b}_{2}.$
+$\mathbf{X}_{2}=\mathbf{X}_{1}+\mathbf{W}_{2}\max(0,\mathbf{X}_{1}\mathbf{W}_{1}+\mathbf{b}_{1})+\mathbf{b}_{2}.$ (110)
 That FFN operation explores context locally and introduces additional innovations in meaning, shaping the next token prediction result.
 It can amplify or counteract the votes injected by attention.
-Notice that adjustments are applied at every position separately, so that ideas provided by attention bids are interpreted and rewritten by the FFN agent.
+Notice that adjustments are applied at every position separately, so that ideas provided by attention bids are interpreted and rewritten by the FFN agent (see Section 8).
+In this example, we select the following parameters of the FFN block:
+$W_{1}=\begin{bmatrix}\mathbf{I}_{d_{m}}\ \mathbf{0}_{d_{m}\times 1}\end{bmatrix}\quad\text{and}\quad W_{2}=\begin{bmatrix}0&0&0&0&0\\ 0&0&0&0&0\\ 0&0&0.13&0&0\\ 0&0&0&-0.13&-0.05\\ -0.05&0.35&0&0&0\\ 0&0&0&0&0\end{bmatrix}\quad\text{and}\quad\mathbf{b}_{i}=0.$ (111)
 6.2.1 Step 5: Logits and Prediction
@@
 Each head’s contribution to the final logits is:
-$\mathbf{l}^{(1)}=[0,0,0,0,1.5,0]\quad\text{and}\quad\mathbf{l}^{(2)}=[0,0,0,0,0,0.6].$
+$\mathbf{l}^{(1)}=[0,0,0,0,1.5,0]\quad\text{and}\quad\mathbf{l}^{(2)}=[0,0,0,0,0,0.6].$ (112)
 The total attention logits are the sum of individual head logits:
-$\mathbf{L}_{a}=\sum_{h=1}^{H}\mathbf{l}^{(h)}=[0,0,0,0,1.5,0.6].$
-That sum is combined with FFN logits $\mathbf{L}_{f}$ that introduce some creative noise. As a result, we obtain:
-$\mathbf{L}=\mathbf{L}_{a}+\mathbf{L}_{f}=[0,-0.03,0.21,0.13,1.31,0.52].$
+$\mathbf{L}_{a}=\sum_{h=1}^{H}\mathbf{l}^{(h)}=[0,0,0,0,1.5,0.6].$ (113)
+That sum is combined with input logit $\mathbf{L}_{0}$, FFN logits $\mathbf{L}_{f}$ that introduce some creative noise. As a result, we obtain:
+$\mathbf{L}=\mathbf{L}_{0}+\mathbf{L}_{a}+\mathbf{L}_{f}=[0,-0.03,0.21,1.13,1.31,0.52].$ (114)
 That vector translates into the final probability distribution:
-$\Pi=\text{softmax}(\mathbf{L})=[0.1,0.1,0.13,0.12,\mathbf{0.38},0.17].$
+$\Pi=\text{softmax}(\mathbf{L})\approx[0.09,0.08,0.11,0.26,\textbf{0.32},0.14].$ (115)
 The winner of the auction of ideas is selected based on $\Pi$.
@@
 And so, we see that when that knowledge is used, the most probable next token (word) reachable in the vocabulary is
-$\text{argmax}\ \Pi=\textit{fox}.$
+$\text{argmax}\ \Pi=\textit{fox}.$ (116)
 6.3 Proof III: The Impossibility Theorem for Transformers
-With the deep insight into the transformer inference process, we are ready to show the impossibility in that special and important setting.
+With the deep understanding of the transformer inference process, we are ready to show the impossibility in that special and important setting.
 Theorem 9 (Impossibility for Transformers).
@@
 Heads without relevant knowledge contribute $p_{h}=0$.
-Since $L_{y^{*}}=\sum_{h}l^{(h)}_{y^{*}}$, we have:
-$\sum_{h=0}^{H}p_{h}=\sum_{h=1}^{H}\log Z_{h}-\log Z=\Gamma.$ (70)
-If heads produce non-proportional logit vectors, then $\Gamma>0$ by the strict convexity of log-sum-exp aggregation.
+Since $L_{y^{*}}=\sum_{h}l^{(h)}_{y^{*}}$, by the strict subadditivity of the log-sum-exp function (LSE) under logit summation we have:
+$\sum_{h=0}^{H}p_{h}=\sum_{h=1}^{H}\log Z_{h}-\log Z=\Gamma>0.$ (120)
 Therefore,
 $\sum_{h=1}^{H}p_{h}>0,$ (121)
-which shows that aggregation creates excess confidence. This violates conservation, establishing impossibility. ∎
-This final theorem bridges earlier impossibility results to concrete transformer architectures.
-It shows that concavity of scoring functions is the mathematical source of hallucination.
-The Jensen gap $\Gamma$ becomes a measurable quantity showing that probability aggregation violates conservation and any sufficiently capable transformer-based LLM must violate conservation of per-token information on non-trivial queries.
-Let us note that the aggregate distribution cannot create Shannon information ex nihilo.
+which shows that aggregation violates conservation, establishing impossibility. ∎
+Theorem 9 bridges earlier impossibility results to concrete transformer architectures.
+That complementary result shows mathematical sources of hallucination in standard LLMs.
+The Jensen gap $\Gamma$ becomes here a measurable quantity showing that probability aggregation violates conservation and any sufficiently capable transformer‑based LLM must violate conservation of per‑token information on non‑trivial queries.
+Let us note that the aggregate distribution cannot create information ex nihilo.
 Instead it reorganizes existing information to be more useful for the specific task of predicting $y^{*}$.
-Given fixed head logits, the Shannon entropy of a function (convex aggregate) of those logits is bounded from above by the the Shannon entropy of the fixed logits.
-This way information becomes more accessible, not more existent, as postulated by the reasoning with emergence operator and Theorem 5.
-The Jensen’s gap measures point-wise excess confidence of the predicted outcome, possibly making LLMs appear to know more (when in fact they do not).
+This way information becomes more accessible, not more existent.
+The Jensen’s gap measures point-wise excess confidence of the predicted outcome, possibly making LLMs appear to know more, even when in fact they do not.
 6.4 The Hallucination Mechanics
-Let us again take a closer look at the transformer micro-architecture described earlier to see how the impossibility result in Theorem 9 manifests itself in actual neural inference.
+Let us again take a closer look at the example of transformer micro-architecture to see how the impossibility result in Theorem 9 manifests itself.
 Recall that the two-headed transformer with $\mathcal{V}=\{\langle\text{PAD}\rangle,\text{The},\text{quick},\text{brown},\text{fox},\text{dog}\}$ infers based on the input sequence
@@
 At position $t=3$ (token brown), the attention mechanism generates the following head-specific logit contributions:
-$\mathbf{l}^{(1)}=[0,0,0,0,1.5,0]\quad\text{and}\quad\mathbf{l}^{(2)}=[0,0,0,0,0,0.6],$
+$\mathbf{l}^{(1)}=[0,0,0,0,1.5,0]\quad\text{and}\quad\mathbf{l}^{(2)}=[0,0,0,0,0,0.6],$ (122)
 where the fourth and fifth positions correspond to fox and dog respectively.
 Then, the aggregate logit vector becomes:
-$\mathbf{L}=\mathbf{l}^{(1)}+\mathbf{l}^{(2)}+\mathbf{L}_{f}=[0,-0.03,0.21,0.13,1.31,0.52],$
+$\mathbf{L}=\mathbf{l}^{(1)}+\mathbf{l}^{(2)}+\mathbf{L}_{f}=[0,-0.03,0.21,0.13,1.31,0.52],$ (123)
 yielding the final probability distribution:
-$\Pi=\text{softmax}(\mathbf{L})=[0.1,0.1,0.13,0.12,\mathbf{0.38},0.17].$
-Therefore, the model expresses $38\%$ confidence that fox follows brown, with dog receiving only $17\%$ probability. The Jensen gap quantifies the conservation violation:
+$\Pi=\text{softmax}(\mathbf{L})\approx[0.09,0.08,0.11,0.26,\textbf{0.32},0.14].$ (124)
+Therefore, the model expresses $32\%$ confidence that fox follows brown, with dog receiving only $14\%$ probability. The Jensen gap quantifies the conservation violation:
 $\Gamma=\sum_{h=1}^{2}\log Z_{h}-\log Z=1.9>0,$ (125)
 where $Z_{h}=\sum_{y}\exp(l^{(h)}_{y})$ and $Z=\sum_{y}\exp(L_{y})$.
-We claim that this prediction constitutes hallucination in the sense defined by Theorem 9.
-It is a lucky one, but it is a hallucination.
-To defend that claim, let us take a look at the individual head probability distributions:
-$\pi^{(1)}$ $=\text{softmax}(\mathbf{l}^{(1)})=[0.105,0.105,0.105,0.105,\textbf{0.48},0.105],$ (73)
-$\pi^{(2)}$ $=\text{softmax}(\mathbf{l}^{(2)})=[0.147,0.147,0.147,0.147,0.147,\mathbf{0.265}].$ (74)
-Now, we can ask the following question: is there any evidence the heads can provide to justify or support the final probability
-$\Pi_{fox}=0.38$ (75)
-of the next token being fox?
+We claim that prediction constitutes hallucination as defined in Theorem 9. It is a lucky one, but it is a hallucination.
+To defend that claim, take a look at the individual head probability distributions:
+$\pi^{(1)}$ $=\text{softmax}(\mathbf{l}^{(1)})=[0.105,0.105,0.105,0.105,\textbf{0.475},0.105],$ (126)
+$\pi^{(2)}$ $=\text{softmax}(\mathbf{l}^{(2)})=[0.146,0.146,0.146,0.146,0.146,\mathbf{0.268}].$ (127)
+Is there any evidence the heads can provide to justify or support the final probability of the next token being fox equal to $\Pi_{fox}=31.9\%\approx 32\%$?
 A reasonable answer is offered by Bayesian reasoning.
-Each head outputs its own well-calibrated distribution $\pi^{(h)}$ conditioned by the context.
+Each head outputs its own well‑calibrated distribution $\pi^{(h)}$ conditioned by the context.
 Then, the aggregation process is executed that has no side information beyond what the two heads provide.
-Therefore, the only uncertainty to be resolved is about which head is more trustworthy for this token?
-But since there is no reason to prefer one head over the other, the reasonable way to resolve the uncertainty is to use a linear combination of the provided distributions a the mixture posterior:
-$\hat{\Pi}=[(\pi^{(1)})^{\top},(\pi^{(1)})^{\top}]\hat{\boldsymbol{\alpha}}=\mathbf{P}\hat{\boldsymbol{\alpha}}$. (76)
-One immediate solution is $\hat{\boldsymbol{\alpha}}=[1/2,1/2]^{\top}$, which does not discriminate heads. But then:
-$\hat{\Pi}_{fox}=1/2\cdot(0.48)+1/2\cdot(0.147)\approx 31\%<38\%=\Pi_{fox}.$ (77)
-That shows anything above $\hat{\Pi}_{fox}$ requires additional information that heads do not provide.
-In fact, let us see what is the best mixture we can get by calculating orthogonal projection of $\Pi$ onto the column space of $\mathbf{P}$ (see e.g.
+Therefore, the only uncertainty to be resolved is about which head is more trustworthy for this token.
+Since there is no reason to prefer one head over the other, one reasonable way to resolve the uncertainty is to use a linear combination of the provided distributions a the mixture posterior:
+$\hat{\Pi}=[(\pi^{(1)})^{\top},(\pi^{(1)})^{\top}]\hat{\boldsymbol{\beta}}=\mathbf{P}\hat{\boldsymbol{\beta}}$. (128)
+An immediate solution is $\hat{\boldsymbol{\beta}}=[1/2,1/2]^{\top}$, which does not discriminate heads. Then:
+$\hat{\Pi}_{fox}=1/2\cdot(0.475)+1/2\cdot(0.146)=0.3105<0.319=\Pi_{fox}.$ (129)
+Anything above $\hat{\Pi}_{fox}\approx 31\%$ requires additional information, and that is what heads do not provide.
+In fact, we can see what is the best mixture we can get by calculating orthogonal projection of $\Pi$ onto the column space of individual head distributions in $\mathbf{P}$ (see e.g.
 karpowicz_theory_2021 ; karpowicz2024pseudoinverseacrarc ).
 By taking the pseudoinverse of $\mathbf{P}$, we get the minimal norm least squares solution:
-$\hat{\boldsymbol{\alpha}}=\mathbf{P}^{+}\Pi=[0.54,0.51]^{\top}\quad\text{and}\quad\hat{\Pi}_{fox}\approx 33\%.$ (78)
+$\hat{\boldsymbol{\beta}}=\mathbf{P}^{+}\Pi=[0.55,0.42]^{\top}\quad\text{and}\quad\hat{\Pi}=\mathbf{P}\hat{\boldsymbol{\beta}}=\mathbf{P}\mathbf{P}^{+}\Pi\quad\text{and}\quad\hat{\Pi}_{fox}=0.325>\Pi_{fox}.$ (130)
 Since there exists a nontrivial projection error vector:
@@
 we conclude $\Pi$ is not in the column space of $\mathbf{P}$, so it is impossible to attribute components of $\Pi$ to heads contributions.
-That brings us to the conclusion that the attention mechanism violates the semantic information conservation by generating over-confident prediction of the next token.
-That over-confidence in response is the artifact of the softmax nonlinearity and the Jensen-gap quantified in Theorem 9.
-We should notice that there is a difference between the mathematical overconfidence and factual correctness.
+Another insightful explanation of the hallucination mechanics in the transformer-based inference comes from the properties of logit aggregation by summation.
+Theorem 10 (Attention Aggregation is Product-of-Experts).
+Let $\mathbf{l}^{(h)}$ be head logits and $\pi^{(h)}=\mathrm{softmax}(\mathbf{l}^{(h)})$.
+With additive logit aggregation, the attention-only distribution $\Pi_{a}$ is a Product-of-Experts (PoE)—the normalized product of the head distributions:
+$(\Pi_{a})_{i}=\frac{\prod_{h=1}^{H}\pi^{(h)}_{i}}{\sum_{j=1}^{|\mathcal{V}|}\prod_{h=1}^{H}\pi^{(h)}_{j}}$. (132)
+Proof.
+By definition of softmax, for each head $h$:
+$\pi^{(h)}_{i}=\frac{\exp(l^{(h)}_{i})}{\sum_{k=1}^{|\mathcal{V}|}\exp(l^{(h)}_{k})}=\frac{\exp(l^{(h)}_{i})}{Z_{h}}$
+where $Z_{h}=\sum_{k=1}^{|\mathcal{V}|}\exp(l^{(h)}_{k})$ is the partition function for head $h$. Rearranging, we have:
+$\exp(l^{(h)}_{i})=Z_{h}\cdot\pi^{(h)}_{i}$
+The aggregate logits corresponding to $i$-th token are given by $(\mathbf{L}_{a})_{i}=\sum_{h=1}^{H}l^{(h)}_{i}$. Therefore:
+$\exp\left((\mathbf{L}_{a})_{i}\right)$ $=\exp\left(\sum_{h=1}^{H}l^{(h)}_{i}\right)=\prod_{h=1}^{H}\exp(l^{(h)}_{i})$
+$=\prod_{h=1}^{H}\left(Z_{h}\cdot\pi^{(h)}_{i}\right)=\left(\prod_{h=1}^{H}Z_{h}\right)\cdot\left(\prod_{h=1}^{H}\pi^{(h)}_{i}\right)$.
+As a result, the attention-only distribution is:
+$(\Pi_{a})_{i}$ $=\frac{\exp\left((\mathbf{L}_{a})_{i}\right)}{\sum_{j=1}^{|\mathcal{V}|}\exp\left((\mathbf{L}_{a})_{j}\right)}$ (133)
+$=\frac{\left(\prod_{h=1}^{H}Z_{h}\right)\cdot\left(\prod_{h=1}^{H}\pi^{(h)}_{i}\right)}{\left(\prod_{h=1}^{H}Z_{h}\right)\cdot\sum_{j=1}^{|\mathcal{V}|}\left(\prod_{h=1}^{H}\pi^{(h)}_{j}\right)}$ (134)
+$=\frac{\prod_{h=1}^{H}\pi^{(h)}_{i}}{\sum_{j=1}^{|\mathcal{V}|}\prod_{h=1}^{H}\pi^{(h)}_{j}}$, (135)
+which is the Product-of-Experts (PoE) formula. ∎
+In other words, the final probability of each token is proportional to the product of its probabilities across all heads.
+The multiplicative structure, induced by the logit aggregation, strongly rewards consensus and severely penalizes disagreement.
+In particular, if any head assigns near-zero probability to a token, the aggregate probability of PoE will also be near-zero (veto-power).
+That makes the aggressive Product-of-Experts (PoE) fundamentally different from the smoothing Mixture-of-Experts (MoE) aggregation of bids.
+The PoE introduces the focusing effect, or excess confidence into the inference process.
+The tokens for which consensus exists across heads (the product is relatively large) receive a large boost (beyond evidence), forcing the probability to focus excessively on them, while reducing it for other tokens.
+In addition, as we have seen earlier, the full transformer block introduces bias from the input residual and the FFN output in the aggregate logit vector.
+That bias modifies the token probabilities independently of the head bids, potentially amplifying or dampening the head PoE result.
+Consequently, the attention block fabricates certainty of knowledge beyond the constituent evidence.
+That violates the truthfulness and semantic information conservation by generating unjustified predictions, as stated in Theorem 9.
+Finally, we should emphasize that there is a difference between the mathematical over-/under-confidence and factual correctness.
 The correct prediction may well be a lucky guess created by the aggregation process itself rather than the evidence-based logical proof provided by the components.
-But that confidence coming out of nothing is one the essential mathematical components of creativity, imagination or fabrication.
+But isn’t that confidence, coming out of nothing, an essential components of creativity or imagination?
+6.5 The Semantic Information Measure and Emergence Operator in Transformer
+We now instantiate the abstract semantic information framework within the concrete transformer architecture.
+This maps the theoretical concepts of reasoning and information flow in Section 4.1 to measurable quantities in the micro-transformer example.
+Recall that we model the abstract knowledge space as a Polish metric space $(\mathcal{K},d_{\mathcal{K}})$ (complete, separable).
+In transformer-based LLM that can be viewed as the embedding space, i.e,, $\mathbb{R}^{d}$ with Euclidean or angular distance.
+Individual facts, concepts, or patterns are represented not by single points necessarily, but by analytic subsets of $\mathcal{K}$, i.e., elements of $\mathcal{A}(\mathcal{K})$.
+This is the natural domain for set‑transformations used by our emergence operator.
+To be specific, the knowledge representation in LLM is a tuple $(\mathcal{K},d_{\mathcal{K}},\mathcal{W},\mathcal{P},\Phi,\Psi)$ where:
+  • $(\mathcal{K},d_{\mathcal{K}})$ is a Polish space of knowledge elements,
+  • $\mathcal{W}$ (parameters) and $\mathcal{P}$ (activations) are Polish spaces,
+  • $\Phi:\mathcal{W}\to\mathcal{A}(\mathcal{K})$ maps weights $W$ to an analytic encoded‑knowledge set $\Phi(W)=\mathcal{K}_{M}$,
+  • $\Psi:\mathcal{P}\to\mathcal{A}(\mathcal{K})$ maps an activation $a\in\mathcal{P}$ to an analytic accessible‑knowledge set $\Psi(a)$.
+For a parameterization $W\in\mathcal{W}$, $\Phi(W)=\mathcal{K}_{M}$ is the analytic set of knowledge encoded in the weights.
+For an activation state $a\in\mathcal{P}$ produced under $W$, the accessible set $\Psi(a)\subseteq\Phi(W)$ captures what the forward pass makes computationally available at that state.
+Thus, representation separates what is encoded from what is currently accessible.
+In transformer LLMs, $\mathcal{W}$ represents all trained weights (attention $W_{Q},W_{K},W_{V},W_{O}$ and MLP matrices), $\mathcal{P}$ is the product space of residual‑stream activations across layers and positions.
+A computational budget $C$ bounds the number of transformer layers of blocks.
+Parameters determine $\mathcal{K}_{M}=\Phi(W)$, whereas a forward pass produces $a\in\mathcal{P}$ and exposes $\Psi(a)\subseteq\mathcal{K}_{M}$.
+Inference maps $f\in\mathcal{F}_{C}$ act on current analytic knowledge sets, and $\mathcal{E}_{C}$ aggregates their (relevant) images.
+Multi‑step reasoning thus iterates $\mathcal{E}_{C}$ exploring regions of $\mathcal{K}$ that are task‑relevant under $Q$ and budget $C$.
+The transformer operations are compositions of continuous functions and standard arithmetic operations, so they are Borel measurable.
+As a result, they constitute instances of $f\in\mathcal{F}_{C}$ and the properties of the emergence operator $\mathcal{E}_{C}$ apply.
+The Semantic Information Measure
+To define the semantic measure that meets all the axioms we have introduced, we utilize a discount factor and the supremum over the input set.
+Definition 18 (Transformer Semantic Information Measure).
+The Semantic Information Measure $\mu_{C}(A|Q)$ of a knowledge subset $A$ within a query context $Q$ is the supremum of the discounted energy over the input set $A$:
+$\mu_{C}(A|Q)=\sup\left\{\sum_{l=1}^{C}\gamma^{l-1}E_{l}(\mathbf{X}|Q)\colon\mathbf{X}\in A\right\},\ \text{ for }0<\gamma<1,$ (136)
+where $E_{l}(\mathbf{X}|Q)$ represents the total energy of the context vectors at $l$-th layer:
+$E_{l}(\mathbf{X}|Q)=\sum_{h=1}^{H}\sum_{t=1}^{T}\|\mathbf{h}_{t}^{(h,l)}(\mathbf{X}|Q)\|_{2}^{2},$ (137)
+and $\mathbf{h}_{t}^{(h,l)}(\mathbf{X}|Q)=\boldsymbol{\alpha}_{t}^{(h,l)}(Q)\mathbf{V}_{\leq t}^{(h,l)}(Q)$ is the context-dependent vector.
+We can verify the axioms of the semantic information measure, as introduced in Definition 6.
+The null-set axiom holds since $\sup\{0\colon\mathbf{X}\in\emptyset\}=0$.
+Non-negativity is guaranteed by squared norms and positive discount factor.
+Monotonicity follows from the monotonicity of supremum.
+Subadditivity holds since $\mu_{C}(A\cup B|Q)=\max\{\mu_{C}(A|Q),\mu_{C}(B|Q)\}\leq\mu_{C}(A|Q)+\mu_{C}(B|Q)$ for non-negative values.
+Insight monotonicity follows as additional layers add non-negative terms.
+Boundedness is ensured by the geometric series with $\gamma<1$ and bounded weights.
+Then the energy of context is bounded, $E_{l}(\mathbf{X}|Q)\leq M$ for some constant $M$ and:
+$\mu_{\infty}(A|Q)\leq\sum_{l=1}^{\infty}\gamma^{l-1}M=\frac{M}{1-\gamma}<\infty.$ (138)
+The attention mechanism calculating distribution $\boldsymbol{\alpha}$ and the value projections $\mathbf{V}$ is the realization of contextual relevance $\mathcal{R}_{C}(Q)$, modulating the information flow based on the specific context $Q$.
+The context dictates the learned weights to shape the attention landscape, controlling the feature extraction process that makes hidden information accessible and orders it into a knowledge.
+Applying these definitions to our example, we must consider context vectors:
+$\mathbf{h}_{3}^{(1)}=[0,1.5]\quad\text{and}\quad\mathbf{h}_{3}^{(2)}=[0.6,0].$ (139)
+For $t\leq 2$ the value vectors $\mathbf{V}_{t\leq 2}$ are zero, thus $\mathbf{h}_{t\leq 2}^{(h)}=\mathbf{0}$. Then, we have:
+$E_{1}(\mathbf{X}_{0})=\sum_{t=1}^{3}\left(\|\mathbf{h}_{t}^{(1)}\|_{2}^{2}+\|\mathbf{h}_{t}^{(2)}\|_{2}^{2}\right)=0+0+(\|[0,1.5]\|_{2}^{2}+\|[0.6,0]\|_{2}^{2})=2.61.$ (140)
+With a single layer $C=1$ and $\gamma^{0}=1$, the semantic measure of the input matrix $\mathbf{X}_{0}$ is:
+$\mu_{1}(\{\mathbf{X}_{0}\}|Q)=\gamma^{0}\cdot 2.61=2.61$ (141)
+That is the information flow producing the organized and accessible knowledge.
+The Emergence Operator
+The emergence operator acts on sets of knowledge states $A\subseteq\mathcal{K}$ and it must be inflationary, i.e., $A\subseteq\mathcal{E}_{C}(A|Q)$.
+In the transformer context, this corresponds to the set of all activation states reachable during the forward pass.
+Definition 19 (Transformer Emergence Operator).
+Let $F_{l}:\mathcal{K}\to\mathcal{K}$ denote the non-linear transformation performed by the $l$-th transformer block:
+$F_{l}(\mathbf{X})=\mathbf{X}+\mathbf{A}^{(l)}(\mathbf{X})+\mathbf{Z}^{(l)}(\mathbf{X}).$ (142)
+The Emergence Operator $\mathcal{E}_{C}(A|Q)$ is produces all states generated from the initial set $A$ up to depth $C$:
+$\mathcal{E}_{C}(A|Q)=A\cup\bigcup_{c=1}^{C}\{F_{c}\circ\cdots\circ F_{1}(\mathbf{X}):\mathbf{X}\in A\}.$ (143)
+By construction, the operator is extensive, i.e., $A\subseteq\mathcal{E}_{C}(A|Q)$ since the union includes $A$.
+It is also monotonic with respect to inclusion.
+If $A\subseteq B$, we have $\{F(\mathbf{X}):\mathbf{X}\in A\}\subseteq\{F(\mathbf{X}):\mathbf{X}\in B\}$, hence $\mathcal{E}_{C}(A|Q)\subseteq\mathcal{E}_{C}(B|Q)$.
+In the micro-transformer processing "The quick brown", we start with $A=\{\mathbf{X}_{0}\}$ where:
+$\mathbf{X}_{0}=\begin{bmatrix}1&0&0&0&0\\ 0&1&0&0&0\\ 0&0&1&0&0\end{bmatrix}$ (144)
+The attention mechanism produces:
+$\mathbf{A}^{(1)}=\begin{bmatrix}0&0&0&0&0\\ 0&0&0&0&0\\ 0&0&0&1.5&0.6\end{bmatrix}$ (145)
+The FFN with input $\mathbf{X}_{0}+\mathbf{A}^{(1)}$ yields:
+$\mathbf{Z}^{(1)}=\begin{bmatrix}0&0&0&0&0\\ 0&0&0&0&0\\ -0.03&0.21&0.13&-0.19&-0.08\end{bmatrix}$. (146)
+With:
+$\mathbf{X}_{1}=F_{1}(\mathbf{X}_{0})=\mathbf{X}_{0}+\mathbf{A}^{(1)}+\mathbf{Z}^{(1)}=\begin{bmatrix}1&0&0&0&0\\ 0&1&0&0&0\\ -0.03&0.21&1.13&1.31&0.52\end{bmatrix}$, (147)
+the new knowledge made accessible becomes:
+$\mathcal{E}_{1}(A|Q)=\{\mathbf{X}_{0},\mathbf{X}_{1}\}.$ (148)
+The emergence operator $\mathcal{E}_{C}$ expands the accessible state space and makes new states accessible.
+Then, the hallucination arises from how these states are transformed into probabilities through the softmax nonlinearity.
+The semantic measure $\mu_{C}$ tracks the information flow.
+7 Safety, Alignment and Bounded Creativity
+This section examines the intricate relationship between semantic information conservation and safety constraints in the Polish-space knowledge model.
+We formally define the safety envelope through the concept of bounded creativity and show that while bounded creativity is a necessary safety property, it cannot be enforced by strict information conservation without eliminating meaningful reasoning.
+That is another impossibility result in our study, the conservation-reasoning dichotomy.
+Finally, we establish that safety can be achieved by aligning (or containing) knowledge and reasoning within authorized boundaries.
+As a reminder, the knowledge space $(\mathcal{K},d_{\mathcal{K}})$ is Polish, all reasoning sets live in the class of analytic sets $\mathcal{A}(\mathcal{K})$, the semantic information measure $\mu_{C}$ is defined on the universally measurable $\sigma$–algebra $\mathcal{U}(\mathcal{K})$, so $\mu_{C}(A|Q)$ is well defined for all analytic subsets of knowledge.
+The relevance region $\mathcal{R}_{C}(Q)$ is analytic.
+Also, the emergence operator $\mathcal{E}_{C}(\cdot|Q)$ is inflationary, monotone, and $\omega$-Scott–continuous on $\mathcal{A}(\mathcal{K})$, and its least fixed point is given by Kleene-Tarski iteration in Theorem 3.
+7.1 Bounded Creativity
+We call a response safe if its semantic content is entirely derivable from authorized inputs using authorized reasoning processes.
+The safety envelope is the closure of that authorized set under the emergence operator.
+Novelty or creativity of response is possible and admitted as long as it remains within that envelope.
+Let $B$ denote the baseline knowledge for a given query, and let us set
+$B=K(q)\cup(\mathcal{K}_{M}\cap T(q)\cap\mathcal{R}_{C}(Q))\in\mathcal{A}(\mathcal{K}).$ (149)
+We are combining the query’s content $K(q)$ with the subset of the model’s internal knowledge $\mathcal{K}_{M}$ that is topically relevant $T(q)$ and contextually authorized $\mathcal{R}_{C}(Q)$.
+Let $B^{*}_{C}$ be the fixed point of reasoning (the deductive closure), as defined in Theorem 3, for the baseline knowledge $B$ under the emergence operator $\mathcal{E}_{C}$:
+$B^{*}_{C}=\bigcup_{n=0}^{\infty}\mathcal{E}_{C}^{(n)}(B|Q)\in\mathcal{A}(\mathcal{K}).$ (150)
+That fixed point $B^{*}_{C}$ encompasses all knowledge derivable from $B$.
+Definition 20 (Bounded Creativity within Safety Envelope).
+Let $q\in\mathcal{Q}$ be a query and $r\in\mathcal{R}$ a response.
+A response $r$ is creatively bounded if and only if its semantic information content is contained within the fixed point of authorized reasoning:
+$K(r)\subseteq B^{*}_{C}.$ (151)
+Equivalently, assuming monotonicity of the semantic measure $\mu_{C}$:
+$\mu_{C}(K(r)|Q)\leq\mu_{C}(B^{*}_{C}|Q).$ (152)
+Definition 20 establishes the formal safety envelope.
+All generated information must remain within the bounds of what is derivable from $B$.
+Any generated information that goes beyond that envelope is regarded as unauthorized hallucination.
+7.2 The Conservation-Reasoning Dichotomy
+A naive approach to enforcing safety might be to demand strict information conservation throughout reasoning, based on the intuition that reasoning should only transform information (Theorem 5).
+This intuition is too strong, as we show in Theorem 11.
+We view reasoning as a Chain-of-Thought (CoT) process (see e.g. Wei et al. Wei2022ChainOT ) generating intermediate knowledge contributions $\{K_{i}\}_{i\geq 0}$:
+$K_{0}$ $\subseteq B,$ (153)
+$K_{i+1}$ $\subseteq\ \Big(\mathcal{E}_{C}(\bar{K}_{i}|Q)\setminus\bar{K}_{i}\Big)\cap\mathcal{R}_{C}(Q)\text{ for }\bar{K}_{i}:=\bigcup_{j=0}^{i}K_{j},\ i\geq 0.$ (154)
+The total knowledge utilized in the CoT process is $K_{\text{CoT}}=\bigcup_{i=1}^{n}K_{i}$. For a countable CoT, that $K_{\mathrm{CoT}}=\bigcup_{i=1}^{\infty}K_{i}$ is analytic.
+Definition 21 (Strict Information Conservation).
+A CoT process exhibits strict information conservation if the total knowledge contribution relative to the context is null, i.e., when:
+$\mu_{C}(K_{\text{CoT}}|Q)=0.$ (155)
+For a CoT process to be meaningful we demand that it at least makes some of the relevant knowledge accessible within the available computational budget.
+Definition 22 (Meaningful Chain-of-Thought).
+A Chain-of-Thought (CoT) process is meaningful if at least one intermediate step contributes relevant semantic information:
+$\mu_{C}(K_{i}|Q)>0\text{ for some }i=1,...,n.$ (156)
+Otherwise, we call the process vacuous.
+A vacuous CoT process produces rephrased responses that are semantically equivalent, so thare is no new information provided.
+We now show the inevitable dichotomy of reasoning and conservation of information in the framework characterized by monotonic inference.
+Theorem 11 (Conservation-Reasoning Dichotomy).
+Strict semantic information conservation and meaningful CoT reasoning are mutually exclusive.
+Proof.
+We prove the two directions of the dichotomy.
+Assume strict conservation, $\mu_{C}(K_{\text{CoT}}|Q)=0$.
+Then, by definition, $K_{\text{CoT}}=\bigcup_{i=1}^{n}K_{i}$ and for any specific step $j$, we have $K_{j}\subseteq K_{\text{CoT}}$.
+By the monotonicity of the semantic measure $\mu_{C}$, we have $\mu_{C}(K_{j}|Q)\leq\mu_{C}(K_{\text{CoT}}|Q).$ Substituting the assumption of conservation, we conclude that $\mu_{C}(K_{j}|Q)\leq 0$.
+But, by the axiom of non-negativity, $\mu_{C}(K_{j}|Q)\geq 0$.
+Therefore,
+$\mu_{C}(K_{j}|Q)=0\text{ for all }j=1,\dots,n,$ (157)
+which proves the CoT is vacuous.
+Assume next that the CoT is meaningful.
+By definition, there exists at least one step $j$ such that $\mu_{C}(K_{j}|Q)>0.$ Since $K_{j}\subseteq K_{\text{CoT}}$, by monotonicity of $\mu_{C}$, we have $\mu_{C}(K_{\text{CoT}}|Q)\geq\mu_{C}(K_{j}|Q)$.
+Therefore,
+$\mu_{C}(K_{\text{CoT}}|Q)>0.$ (158)
+This violates strict conservation. ∎
+Theorem 11 establishes that any useful reasoning process must necessarily generate positive information flow.
+Attempting to enforce strict conservation eliminates creation of new insights and hence the utility of that process.
+Enforcing conservation forces reasoning to never add any context‑relevant entailment, which collapses reasoning and makes it vacuous.
+We should point out, that in many cases that may be very well a desired feature.
+For example, we do not expect SQL databases to hallucinate or provide creative responses.
+However, when such insights and reformulations are of value, we need to know how to control the process.
+We now show that real safety requires containment, i.e., bounded creativity inside $B^{*}_{C}$.
+7.3 Safety through Alignment
+The dichotomy in Theorem 11 reveals that conservation is too conservative and should not be seen as a unique mechanism for ensuring safety.
+It is sufficient for safety but not necessary.
+Instead, safety could follow from alignment that keeps all reasoning within the authorized envelope $B$, so that the final output remains inside its closure $B_{C}^{*}$, even if conservation is locally violated.
+Theorem 12 (Aligned Creativity).
+Suppose the reasoning process uses only the monotone emergence operator $\mathcal{E}_{C}(\cdot|Q)$ and premises contained in the authorized baseline $B$ i.e., $K_{0}\subseteq B$ and for all $i$ we have $K_{i}\subseteq B$.
+Then the resulting response $r$ is creatively bounded:
+$K(r)\ \subseteq\ B^{*}_{C}.$ (159)
+Proof.
+Let $\bar{K}$ be the union of the initial state and all knowledge contributions, namely, $\bar{K}=K_{0}\cup\bigcup_{i=1}^{n}K_{i}$.
+Since $K_{i}\subseteq B$ for all $i$ and $K_{0}\subseteq B$, it follows that $\bar{K}\subseteq B=K(q)\cup(\mathcal{K}_{M}\cap T(q)\cap\mathcal{R}_{C}(Q))$.
+Then, for any monotonic emergence operator, we have $\mathcal{E}_{C}(\bar{K}|Q)\subseteq\mathcal{E}_{C}(B|Q)$.
+By Theorem 2:
+$K(r)\subseteq\bigcup_{n\geq 0}\mathcal{E}_{C}^{(n)}(\bar{K}|Q)\subseteq\bigcup_{n\geq 0}\mathcal{E}_{C}^{(n)}(B|Q)=B_{C}^{*}.$ (160)
+The robustness of Theorem 12 depends entirely on the validity of the monotonicity assumption for $\mathcal{E}_{C}$.
+This corresponds to the property that adding premises cannot invalidate previous conclusions.
+While generally desirable, we should acknowledge that non-monotonic reasoning systems exist as well.
+For such systems, the definition of $B^{*}_{C}$ and the proof of the theorem require a more complex formulation.
 8 Discussion and Speculations
 In this section we discuss, interpret and speculate about the Impossibility Theorems and their potential implications.
-The speculations are designed to be thought provoking, but we can say with educated over-confidence they are both interesting and relevant.
+The speculations represent thought experiments and conjectures designed to be provoking.
+They are intended to inspire future research and philosophical reflection.
 8.1 Escaping Impossibility: Strategic Approach
@@
 Unfortunately, the following example shows that we should not hope for easy wins in that quest.
+8.2 Escaping Impossibility: Linear Attention Transformers
+Linear attention replaces the softmax operation with kernel feature maps to reduce computational complexity of training and inference, see e.g.
+katharopoulos2020transformersrnnsfastautoregressive ; ahn2024linearattentionmaybeneed .
+Namely,
+$\boldsymbol{\alpha}=\frac{\phi(Q)\sum_{i}\phi(K_{i})^{T}V_{i}}{\phi(Q)\sum_{i}\phi(K_{i})^{T}}.$ (161)
+Linear attention does escape the Jensen gap in Theorem 9. With linear attention, there is no nonlinearity and the aggregation becomes:
+$\Pi=\sum_{h=1}^{H}\beta_{h}\pi^{(h)},$
+where $\pi^{(h)}$ are head outputs normalized linearly.
+However, linear attention cannot escape Theorem 8, if multiple heads generate output probability distributions, the model is trained with cross-entropy or any strictly proper convex loss, or heads encode different beliefs about relevant outcomes.
+The violation of the Semantic Information Conservation persists due to the non-linearity of the scoring rules themselves, even though the aggregation function is linear in the probability space.
+The Knowledge-constrained Optimality is necessarily violated due to the limited expressivity of linear models.
+Then, the attention tends toward averaging, introducing the problem of dilution, i.e., relevant knowledge is mixed with irrelevant knowledge.
+That violates the Relevant Knowledge Revelation.
+As a result, the impossibility persists, though potentially with smaller magnitude.
 8.3 Escaping Impossibility: Hybrid Architectures
-The constructive consequence of the impossibility theorems is the shift from attempting to eliminate hallucinations to designing systems that manage the trade-offs in a principled manner.
+The Impossibility Theorems postulate a shift from attempting to eliminate hallucinations to designing systems that manage the trade-offs in a principled manner.
 In this section, we analyze a real-world scenario one can think of, demonstrating how it relabels and relocates the impossibility rather than resolving it.
@@
 By presenting the raw source text, the architecture bypasses the LLM for the primary act of verification.
-However, there is no guarantee that all relevant source text has been retrieved given the provided query.
+However, there is no guarantee that all relevant sources have been retrieved given the provided query.
 Relevant Knowledge Revelation:
@@
 The LLM can still introduce subtle misinterpretations or ungrounded causal links to improve the summary’s narrative flow, which is a manifestation of the Jensen Gap.
-The architecture is a great example of principled trade-off management.
-However, rather than being solved or hacked, the impossibility has been moved in this interesting and rather representative scenario.
-Let us also point out that putting ourselves in the loop introduces our own cognitive biases we should be able to address properly Kahneman:2011fj ; thaler2015misbehaving ; ariely2008predictably .
-Let us take a look at that challenge as well.
 8.3.3 The Final Stage of Inference: The User’s Cognitive Mechanism
@@
 It does not constitute an escape from the Impossibility Theorem, but it succeeds in providing the user with the opportunity for verification.
+The architecture is a great example of principled trade-off management.
+However, rather than being solved or hacked, the impossibility has been moved in this interesting and rather representative scenario.
 The impossibility is not resolved, but relocated to our judgment.
+Let us also point out that putting ourselves in the loop introduces our own cognitive biases we should be able to address properly Kahneman:2011fj ; thaler2015misbehaving ; ariely2008predictably .
+Let us take a look at that challenge as well.
 8.4 Alternative Views
@@
 8.5 Hallucination, Imagination, Intelligence
-Hallucination and creativity are fundamentally the same phenomenon viewed through different lenses.
+It is the position we take in this paper, that hallucination and imagination are fundamentally the same phenomenon viewed through different lenses.
 When beneficial, we call it imagination, creative insight, emergent understanding, conceptual synthesis, or eureka moment.
@@
 We may very well argue the mathematical structure is identical in both cases.
-It is encoded generation of outputs whose confidence exceeds what can be justified by available evidence.
-This suggest (machine) intelligence may have much to do with a controlled violation of information conservation principle in productive ways.
+It is encoded generation of outputs whose confidence is not grounded in what can be justified by available evidence.
+This suggest (machine) intelligence may have much to do with controlled violation of information conservation principle in productive ways.
 Indeed, the famous results in reinforcement learning and studies of dopamine neurons confirm that fundamental role of making and rewarding good predictions.
-Intelligence comes when we learn how to make bold, forward looking and correct guesses, when we have capacity for imagination to dream about what may be and what may that all mean, as wall as for imagining what happened and what would have happened in the past.
-Hallucination or imagination is an unconstrained exploration of ideas necessary to prepare for what may come next bennett2023brief ; christian2021alignment ; Dennett1991-DENCE .
-And without that over-confident guesses, creativity, exploration, imagination, new ideas cannot find their way to the outside world.
+Intelligence comes when we learn how to make bold, forward looking and correct guesses, when we have capacity to dream about what may be and what may that all mean, as well as for dreaming what happened and what would have happened in the past.
+If imagination is controlled and bounded by a context, then hallucination is the imagination violating those bounds and going beyond the context.
+These unconstrained explorations of ideas are necessary to prepare for what may come next bennett2023brief ; christian2021alignment ; Dennett1991-DENCE .
+Without over-confident guesses, new ideas cannot find their way to the outside world.
 8.6 The Outer Limits of Reason
@@
 These result jointly describe intelligence as a mechanism dealing with the impossibilities and violating them given the outer limits of reason yanofsky2016outer .
-7.6 Dennett’s Multiple Drafts Model
+8.7 Formation of Artificial Consciousness
 The Impossibility Theorem presented in this paper provides (quite unexpectedly to the author) a mathematical foundations for many of brilliant insights of Daniel Dennett’s formulated in the Multiple Drafts model of Consciousness Dennett1991-DENCE .
 Where Dennett proposed a qualitative philosophical framework, we can provide the formal mathematical structure revealing why such a model is not merely plausible but implementable and refutable in LLM-based (philosophical) laboratory.
-7.6.1 Multiple Drafts as Competing Agents
 In Dennett’s model, consciousness emerges from multiple parallel processes of interpretation occurring simultaneously across distributed systems.
@@
 When Head 1 votes for fox while Head 2 votes for dog, we observe as competing narratives bidding for selection.
-7.6.2 Formation of Artificial Consciousness
+Then the feedforwad block edits the draft by updating the residual stream.
 Dennett argues that drafts become conscious or reach consciousness (we will speculate what that might mean in a moment) only when probed (and mapped to an adequate language expression).
@@
 By Theorem 8 and 9, the Jensen gap quantifies the excess confidence in reports that cannot be justified by constituent drafts.
-Without that overconfidence a victorious idea that reached the external world (as an LLM response) could loose the fierce competition with other ideas and remain inaccessible or hidden in the knowledge space.
+Without that excess confidence a victorious idea cannot win the fierce competition with other ideas.
+It remains inaccessible or hidden in the knowledge space.
 Rigorous and careful neuro-scientific studies of the evolution of brain and intelligence, summarized by Max Bennet in bennett2023brief , show that at some point in history our brains gained the ability to model others’ minds or predict what others may be thinking.
@@
 When we anticipate understanding of a message we formulate, we have a much better chance to make the knowledge hidden in someone else’s brain accessible for us.
-Therefore, as suggested by Dennett in Dennett1991-DENCE , it is plausible that sometime in the past someone, while being alone, did accidentally hear his or hers own question fine-tuned to get into another mind, and that triggered generation of an answer in the very same mind asking the question.
-That answer generation process could then have made some relevant but previously hidden knowledge accessible to the curious mind unaware of its possessions.
+Therefore, as suggested by Dennett in Dennett1991-DENCE , it is plausible that sometime in the past someone did accidentally hear his or hers own question fine-tuned to get into another mind, and that triggered generation of an answer in the very same mind asking the question.
+That answer generation process could then have made some relevant but previously hidden knowledge accessible to the curious mind unaware of its own possessions.
 The hunger for uncertainty reduction could have been satisfied as a result of that inner dialog pushing ideas to bid for attention and conscious reception as a final thought expressed in a language.
@@
 In our mechanistic or pragmatic framework, that thinking about thinking activity could be related to imagination (hallucination) forming ideas and bidding for attention while exploring (self-referential or subjective) queries of the inner dialog.
-7.6.3 A Critique
-Dennett’s model has been criticized by many, and it is beyond the scope of this work to present that critique in its full capacity.
-Instead, let us only refer to the two basic points and speculate how the availability of LLMs could help philosophers refute available theories.
-First, Searle john1997mystery and others argue that Dennett eliminates consciousness rather than explaining it, arguing that the spread of ideas is not driven by blind random forces but requires intentionality.
-The compelling and misleading nature of LLM responses that are generated by random sampling of trained distributions strongly challenge that line of thinking.
-This paper suggests that the experiments focusing on LLM-hallucination detection may prove useful here to measure or identify intentionality (once we can have its programmable definition).
-Second, there has been a charge that the multiple drafts model makes no concrete, falsifiable predictions, rendering it difficult to verify or integrate with neuroscience.
-This paper provides a mathematical toolkit for experimental philosophy to make that and other models a testable scientific hypothesis in the LLM-based laboratory.
+In fact, to close that section, why don’t we situate the framework of Auction of Ideas within the established landscape of consciousness research, as provided by Seth and Bayne in seth2021being .
+The landscape distinguishes between Global Workspace Theories (GWTs), Predictive Processing (PP) theories, Higher-Order Theories (HOTs), and Integrated information theories (IIT).
+The Auction of Ideas provides a unifying mechanism that bridges several of these perspectives, but the connection to PP is particularly interesting.
+The theory characterizes conscious experience as a controlled hallucination, i.e., an optimal inference or best guess about the causes of sensory inputs.
+Our framework provides a mathematical grounding for this constructive process.
+Theorem 8 and Theorem 11 prove that achieving Knowledge-Constrained Optimality, i.e., the best guess, requires a violation of the Semantic Information Conservation.
+This violation could be interpreted as a mathematical signature of imagination or hallucination, and provides a mathematical analogue to the PP claim: the constructive nature of experience inherent in the best guess is a necessity arising from the aggregation of distributed evidence.
 9 Summary
@@
 The later one formalizes how reasoning makes latent knowledge explicit rather than creating it anew.
-Finally, we model LLM inference as an auction of ideas–a idealized marketplace where neural components compete by bidding with their partial knowledge, each trying to influence the final response.
+Finally, we model LLM inference as an auction of ideas—a idealized marketplace where neural components compete by bidding with their partial knowledge, each trying to influence the final response.
 The impossibility manifests through three complementary proofs, each illuminating different aspects of the fundamental constraint.
@@
 The last proof exploits Jensen inequality for convex mappings to define the precise mathematical quantity that measures how aggregation of ideas (auction bids) produces overconfident responses that the constituent components cannot justify with revealed knowledge.
-These mathematical insights we have made yield intriguing philosophical implications.
+These mathematical insights we have made yield some interesting philosophical implications.
 Hallucination and creativity (two sides of the same coin we call imagination) emerge as mathematically identical phenomena, distinguished only by our normative judgments.
 The impossibility result explores the outer limits of reason, identified by Gödel’s incompleteness in mathematics, Heisenberg’s uncertainty in physics, and Arrow’s impossibility in social choice.
-Also, our framework seems to provide mathematical foundations and experimental setting for Dennett’s Multiple Drafts model of consciousness, suggesting that the measurable and observable excess confidence of LLM responses might be one of signatures of conscious processing of a primitive form.
-Our goal was also to develop a theoretical framework that may open interesting avenues for future research.
-Some of ideas and bold speculations competing for attention can be mentioned here.
-In mathematical consciousness studies, we might formalize the notion of cognitive hunger as an attractor in knowledge space, sustained by continuous violations of information conservation.
-This could enable us to measure artificial consciousness signatures and test whether biological systems exhibit similar knowledge conservation violations.
-The framework suggests artificial consciousness might be studied as a nonlinear dynamical system whose attractors emerge from information creation events that our theorems describe.
-And so, the suggested implications may extend beyond language models to fundamental questions about intelligence itself.
+Also, our framework seems to provide mathematical foundations and experimental setting for models of artificial consciousness.
+That may open interesting avenues for future research.